A Bibliography of The Journal of the American Statistical Association

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

20 May 2021 Version 2.11

Title word cross-reference

#4059 [CMS87].

.471 [CB94]. .632+ [ET97]. 2^{31} − 1 [FM82]. 2 × 2 [Agr98, BS97c, BB98, MPS98, SW98b, SW98c, VHE91, AG83, GR81, Gui93, MPG85, MPG86]. 2 × 2 × 2 [CGM83]. [p(1 − p)]^{−1/6} [Geb71]. A [Jac89]. AR(1) [CR89]. AR[k] [Dha93]. ARIMA(p, 1, q) [SD85]. B(x; 2/3, 2/3) [Geb71]. C_{p} [RS94b]. chi^{2} [BN74]. χ_{r}^{2} [JH82]. chi_{1}^{2} [BN74]. D [AC95c, Stu82]. ED50 [JJW84]. F [LRR91, Dye82, HC84, WHH88]. G [MG84]. I [DL88, SY84]. II [BM81, BM83, Bha85, Esc86, LaR86, MB82, SM84b]. III [Bof85]. ∫_{−∞}^{+∞} f(t)exp(−t^{2})dt [CS90b]. k [Boo86, BC82b, EL90, Flu84, GF96a, HN87, MW81, MW82, Pen97, SS87a, Wei81]. K2J [YF95]. k × l [Qui98a]. L [ER86, KP87, PS82]. L2 × K [YF90]. L^{1} [dHY93]. L_{1} [EM92, WS94]. L_{p} [Spo82]. log [HB86, Koe86]. M [Akr91b, Ber91a, KP90, HRR81, MZ89]. N [Hal94, Rel72, BD81b, jW97, CLNL87, CL85, FD85, KT94a]. N, p [Hal94]. P [BB94, Bor87, LS97, Daw95, SMP95, TW89, Wee91, BS87, SO83]. P (X_{1} + X_{2} + ⋯ + X_{k} < z) [YM85, YM86]. P − P [Hol95]. r [Zac86, JJW84]. r × c [MP83, PH81]. S [RV90c, CRH94]. t [Ben83, Hjo94, Joh78, KT88, LLT89, PMF88]. v [FOS95].

-Dimensional [Daw95]. -estimates [MZ89]. -estimation [KP87]. -estimator [JJW84]. -estimators [CRH94, HRR81]. -linear [HB86, Koe86]. -ness [Hjo94]. -optimal [Stu82]. -optimality [Jac89]. -Optimum [AC95c]. -Ratio [Pen97]. -Sample [EL90, BC82b, HN87, MW81, MW82, SS87a]. -spectral [MG84]. -spherical [FOS95]. -statistics [PS82]. -System [RV90c]. -values [TW89].

1 [GHMW94]. 1800- [Sol95]. 1980 [EKT89]. 1990s [Ano91].

21st [Ket97]. 2nd [Tuk47b]. 2SAE [Kar81].

32-bit-word [Che71].

62009a [Bor87].

71 [CMS87]. 73 [FH81c]. 74 [Krz81]. 75 [JTB81, RC82, WM81]. 76 [BC82a, DD82, FH81a, KP82, Law82, MW82, MT82, Pra82, Sch82, de 83a]. 77 [HG83]. 78 [NS86, Rin85].

80 [BK86, MPG86, YM86]. 80k [FH81c]. 81 [Mac87, Sim87a]. 81a [Krz81]. 81e [WM81]. 81f [JTB81]. 82 [Bor87, Cla89, Kot89]. 82c [FH81a]. 82d [MW82, RC82]. 82h [KP82, Law82, MT82, Sch82]. 83 [LB94c]. 83a [de 83a]. 83b [BC82a]. 83e [DD82]. 83h [Pra82]. 84 [Fab94, Whi92]. 84d [HG83]. 84i [NS86]. 85a [Rin85]. 86 [MS95]. 87b [YM86]. 87c [BK86]. 87j [Mac87]. 88 [Wee94]. 88h [Kot89]. 88k [Cla89]. 89 [Sta94]. 89a [Bor87].

90 [SG95a]. 92 [Wes98]. 92e [Fab94, Whi92]. 92i [MS95]. 93 [MS98a]. 94e [Wee94]. 96e [SG95a]. 98c [Wes98].

A-Bomb [PSVS92]. A. [Oak85]. Aalen [HM91]. Ability [Tuk47b, Tuk47c]. Absolute [RC93, Sch87a]. Accelerated [CM97a, Yam92, Sch88, SS82a]. Accelerated-Time [CM97a]. Acceleration [JJ93]. Acceptability [BRB76, BS97c]. Acceptance [LS91b, SS80, Tuk50a, KP81b, KP82]. acceptance-rejection [KP81b, KP82]. Acceptance/rejection [SS80]. Accidents [Wei92]. account [SO83]. Accounting [MR94, Roc95]. accounts [KY86]. Accuracy [BRB76, Che90, CCL90, DC77, Haw93, SM92, Wam70, Hod87, JJS83, Tuk79b]. Accurate [SW97, TK86]. Achieve [CN94]. Achieving [HOLDG98, Min98]. Acid [Haa90]. Acknowledgement [MS95]. Action [APR92, Gai96, LS91b].

Activity [GJDT98]. Adapting [DJ95, HT97a, HT97b, Por97]. Adaptive [AMCS98, BF93a, BWC98, BGH93, CKM97, GRS98, GR96b, Gu90, Hog74, HOS94b, KP92, LT94a, LS91a, LW97, Pri94, Ruk84, Ruk93, SS96a, Sch95a, TFAH94, Tho90b, CS87, Fan92, HB88, Par93a]. Addendum [Wes98]. Additive [Bac89, BR94a, Cho92, DS97b, HM91, KO95, KPR91, LRM96, LCWH97, Niu96, OR98]. additivity [MJ82]. Address [Ket97, Moo98]. Adequacy [CW97, HW92a, Jen86a]. Adjoint [Dua90]. Adjust [IRC98]. Adjusted [CW94a, NC99, ZD97]. Adjusting [EKT89, RG94, SO83]. Adjustment [HB99, Sch91a, SK91, SR95, HT82, Kar87, O’G82, Sch81, Sch82, Wat87]. Adjustments [FMSP90]. Admissible [Sob91, MV85]. Adolescence [MSMP92]. Adrian [DH83]. Adult [HHH92]. Advances [Hoa73]. Advantages [MH99]. Adversaries [LS91b]. Adverse [GLB98, Kad90, YKB97]. Advertising [MSL93]. Affine [HMO97a, HMO97b]. Affine-Invariant [HMO97a, HMO97b]. Africa [GJDT98]. After [OMF^{+}93, Rub96a, XXP93, FH78, FH81c]. Against [AP96, CS92a, CF95, Hay90, LGWC93, Shi91, SS95a, Wan96, dCP99a, BHS81, HM88a, Lee87, ZC88]. Age [AK92, Cha91, HR90, NSS97, RS90, YP92, SS82b]. Ages [Par93b]. Aggregate [BP86, KK84b]. Aggregated [MST97]. Aggregates [PH90a]. aging [HPP86]. Agreed [LS91b]. agreements [SG83]. Agricultural [CLY92]. AHEAD [JS97b]. AIDS [Bac90, DXP92, DA96, EFS92, GJDT98, Har90b, HMC^{+}93, JMV94, TCS94, TDW95, XXP93, ZDL96, RG94]. Akaike [Nef82]. algebraic [SB81]. ALGOL [Gol63]. Algorithm [Cho93, GN82, GR92a, JJ93, KT95, LK92b, Luc93b, Sch96, Sho93, WT90b, WZ97a, XLR91, CC74, Dyk83, JK89, LLS87, MP83, PH81, Rel72, Spi82, TZ81]. Algorithms [Joh96b, Joh98b, McC97, WT90b, LB94c, LB94d, Lin74, SS87b]. aliasing [HW88]. Aligning [RN97]. Alignment [LNL95, LNL99]. All-or-None [Bak98]. Allocation [KM91a, Ode91, Pag90, Sha85, SN90, WSLP90, BM85, Sha83]. Almost [Eas94]. Also [Goo96]. Alternative [EH91, Fay96a, Hay90, MSMP92, Mun96, NG94, SW95, Sim74, PH86, SM84b]. Alternatives [AC95b, CF95, Fol96, GR92b, JSKK98, LH91, LGWC93, RC93, SS95a, dCP99a, BD88, BS86a, HN87, KR84, MW81, MW82, Nai87c, Stu56, ZC88]. Always [RRZ94]. Ambiguous [SM90b]. Amer [BK86, Bor87, BC82a, CMS87, Cla89, DD82, Fab94, FH81a, FH81c, HG83, JTB81, Kot89, KP82, Krz81, Law82, LB94c, MW82, Mac87, MS98a, MPG86, MT82, MS95, NS86, Pra82, Rin85, RC82, Sch82, Sim87a, SG95a, Sta94, Wee94, Wes98, WM81, Whi92, YM86, de 83a]. American [BB90, Pea91]. Among [Moo98, SMF94, KD87, Thu85, Uus85]. Amount [Nyc91, Hjo94]. Analgesic [DS97a, GB97, LL97a, SBD97a, SBD97b]. Analog [BW94, SM81]. analogs [Rot83, Yan84]. analogy [SW83]. Analyses [HAS97, PSVS92, WC92, Ole85, Sim74]. Analysis [Ada98, Ala84, AC93, AASM93, AC95a, ASL91, AL95, AD90, BRR92, BPS96, Bak98, BH93, BHB96, BC96, BS93, BY95, BS91, Bri92, BR98a, BR98b, BD91b, BH97, Car94, CWW95, CJ96, CM97b, CM91a, CNWW98, Con89, Con92, CH92, CMMY91, DwSH94, DG99, DAO94, DR91, DLB91, DS97a, DH99, DM96, DPWH99, DD90, dE97, EP94, EK93, EGGS97, FZ98, FSE98, Fay96a, Flu85, FR99, FP95, For92, FB83, Fun95, GSTM92, GBJ96, GB97, GWG98, GN82, GH90, Goo96, GRB^{+}95, Gro90, GL96, Hab82, HW95a, HTB94, HM97, Hei95, HT97c, HH91, IL94, JT97, Joh90, Joh94, Joh96a, Kad90, Kne98, KC96, Lan98, Lav91a, LWY93, LB94b, Lin91, LW92, LL97a, LCG91, LCWH97, Man90, MM95, MSK96, Mur95, MSL93, Niu96, OMF^{+}93, OV95, PB98, PM97, PS96]. Analysis [Pri96, PHL98, QTC96, Qui98a, RS98, RSY93, RG93, RRZ95, Roc95, RT96b, Ryu94, STR97, SG96a, SB87, SBD97a, SBD97b, SJ92, SW90, Sil97, SD97, Sol91, Sol95, SEH95, SHH78, SRSD94, SL98b, Sto91, SK93b, SY95, Tar99, TCS94, Tuk47b, Tuk47c, Ver93, Wak96, Wan92, WSJ92, WWZtO97, WLGC98, WW98, Weg90, Wes97a, WM91, Yam92, YK96, YJW95, YB97, Zuc98, dCP99b, Ame74, BOP88, BM85, BD88, BB89, BH88, BD81a, Boi86, BC82a, Bro83, CGG85, CM84a, Chr87, Cla87b, Cla89, Clo82, CG84, DWL87, Dwy67, Efr88, Fab91, Fat83, FH86, Fri89, Gas85, GH86, GH88, Goo91, GH87b, GY93, HW88, HR84, HVS84, Hol87, IL91, IHC84, IS87, JG83, JK89, KL85, Kee82, LS86, LV88, MB88b, NS83, NS86, Ols74, PH74, Pud82, RW85]. analysis [Rav89, Sim87b, SHS86, Sto89, Tak83, Tsi82, TH82, TV85, WLW89, Won82a, Wor34, Wor74, ZB86]. analysis-of-variance [BD88]. Analytical [Hod94, dHY93]. Analyzing [BO95, DB97, GH94, Gra92, Mar92, WKL95]. Ancestral [GT95]. Ancillarity [God82a, He90]. ancillary [Bue82]. and/or [Gui88]. Anderson [SS87a]. Andrews [Hoa73]. animal [MV88, MC84]. Annealing [GT95]. ANOCOVA [BS86a]. Anomalous [Ben38]. ANOVA [Boi86, KR86, Tak83]. Answer [BZ99]. Answered [Bin98, Bre98, GKL98a, GKL98b, Jud98, Kad98, San98]. Anti [MSL93]. Anti-Drinking [MSL93]. Any [Eas94]. apparent [Efr86b]. applicability [Hur85b]. Application [AME96, ALJ99, AK92, BSB98, BF92, BPS95, CH95a, CG97, CS96b, Cra94, CS90b, DwSH94, DDV99, Fin90, GT98, GRB^{+}95, GR92a, Haa95, HW92a, HR90, HDG97, HMC^{+}93, JMV94, Joh96a, KS96, LR96, Li92, LTW96, LS91b, MW98a, MW98b, MSK96, MST97, O’N94, O’S90, OO97, PJT96, PH90a, RW98, SC96a, Sas97, SR95, SH90, SEH95, The93, WM90, WHG98, Yam92, YK96, Ye98b, ZDL96, Abr88, CT98, GG87, LLS87, Ole85, Sco85, SS88, VW87, Wij85]. Applications [AML95, AHB97, AGM98, BPS96, Bak98, Bin98, BS97b, BO95, BWC98, Bre98, CCE85, CNWW98, CMMY91, DN98, DG92, FG94, Gal94, GJ94, GKL98a, GKL98b, GJDT98, GT95, Gho92, GL92, Gra92, Hig98, HLC92, Hou95, IRC98, Jud98, Kad98, LZ98, Liu94, LWT94, LKC97, ME96, MLRR92, Niu96, NGG97, QH98, Riv98, SW95, San98, Sil97, SG96b, Sto91, TDW95, WEST91, AG85, Boi86, BJM81, CS87, DH93, Dwy67, ER86, Hog74, Hwa86, JTB80, JTB81, KSC89, Las94, McC82, Mor83, TB81, Yan84, YD84, Tuk52].

Applied [MR97a, MR97b, dCP99b, Ran84]. Apportionment [Sch91a]. Appraisal [Lon67]. Approach [Amm93, BR94a, Bro97, CP91, CPS92, CKLH92, CM97b, DF98, DG92, EH91, Fon92, FK94, Fyg97, GMP97, GNK96, GJP96, Hal73, HW95a, HL98, HZ98, HL94, HDG97, IL94, KA91, Kol96, KS97, KP81a, LaR91, LK97, Lav91a, LT98, Lju93, MM93, MV91, NCDF96, NGH^{+}95, O’S90, Pai97, Pre96, Raz90, RCL96, SH92, Shu97, SK97, SG97, Tho91b, WW91a, WLGC98, WZ97b, WK98, YL97, BD88, BJM81, CT98, CS87, Fin83, Gil84, Gil86, GM86b, HW94, HT82, JSG88, KT94a, Mor86, OS87, Oma81, Rav89, Roy86, Sin86, Tam88b, TJW88, WA83, ZK91]. Approaches [CG90, DN98, GS90a, SR94]. Appropriateness [Ueb93]. Approximate [Atw84, Bly86, BC93, CS90a, Che90, Cla87a, Hsi97, KT94b, KT94c, PFV89, Swe84, TL97]. Approximately [Agr98, BB98, EMM95, MPS98, Sev94, SW98b, SW98c, Wie98]. Approximating [GW91, HL97]. Approximation [BS98, GJ99, Jen91, Lie94, Pes93, Shu97, Geb71, Mar84, MT81, MT82]. Approximations [BDM97, CCL90, DKRW97, Erk94, FG94, GR96a, Gla89, Haw93, JS94, Joe95, KH84, Nau82, SGP98, WK88, WBB93, dHY93, DT87, DKMO89, ER86, Hod87, PMF88, Sha89a, Sha89b, TK86, TKK89]. arbitrary [Wes87a]. Archimedean [GR93]. Area [FG91, GNSC98, HH91, LR95a, MSML97, PR90b, Rag93]. Areas [WC91]. ARIMA [CNS90, HT82, KA86, Krä86, Mar86, Sol84, TR97]. ARIMA-model-based [HT82]. Arising [EFS92, GMD95, ZDL96]. arm [Fox74]. ARMA [HM89, AP96, BY86, GL92, GW86, HM88b, Mau95, NP95, Roc92, TT84]. ARMAX [Sto86]. Arms [RG94]. Army [Hod94]. Array [Tan93]. Array-Based [Tan93]. Arrays [WW91a]. Arrest [CNWW98]. Artificial [DM90, HD97]. Arts [Moo98]. Ascertained [Wan92]. Asked [Bin98, Bre98, GKL98a, GKL98b, Jud98, Kad98, San98]. Aspects [HH91, GKR88, Hub83]. Assays [HZS^{+}90, Min93, ZDL96]. Assessed [CG92]. Assessing [CCL90, CW97, Gle92, GCW96, HW90a, HL94, Hod87, Kra90, LA93, QN98, Qui98a, SS91b, Vel98]. Assessment [GJR98, KM91b, MO98, ZT97]. assessments [Agn85]. assessors [VM83]. Assignment [CH91, Gav78, Kie78, Mal78, SM78a, SM78b, Sim78, HP88]. Assoc [BK86, Bor87, BC82a, CMS87, Cla89, DD82, Fab94, FH81a, FH81c, HG83, JTB81, Kot89, KP82, Krz81, Law82, LB94c, MW82, Mac87, MS98a, MPG86, MT82, MS95, NS86, Pra82, Rin85, RC82, Sch82, Sim87a, SG95a, Sta94, Wee94, Wes98, WM81, Whi92, YM86, de 83a]. associated [Boo82, YD84]. Association [Alb97, ALHH92, Goo96, Hab95, KA95, Lau92, ACK87, Clo82, Gil86, GH88, GKR88, KMT99]. Associations [SK93a]. Assumption [CB97, PHS94, Rom90, CM85, KL85, Kru88]. Assumptions [RT96a, FJV89]. Assurance [Riv98]. Astronomical [VB97]. Astronomy [AML95, WM90]. Asymmetric [Efr92, AP87, SG83, Zel86]. Asymmetrical [WW91a, Joh78]. Asymmetry [KP97, Boo82]. Asymptotic [Cha93, CD96, Cro85, Cro74, DKRW97, Kun80, PP88, SL87, Sha86, Wei81, Whi89, WCH98, XLR91, Ber88b, Fab94, Gro81, Whi92]. Asymptotically [LaR86, MZ89, CC74, IHC84, KR86, Sin86, WL84]. Asymptotics [CKLS96, Ste95, SCW96, SCW97, Bha85]. Athletes [MH94b]. Atlantic [NSS97]. attained [But84]. Attainment [AK92]. Attainments [WHGA96]. Attenuation [SKC94, FH78, FH81c]. Attraction [HW92b, Wan95]. attributes [Tam81]. Auditing [CS96b]. Augmentation [WT90b, TW87]. author [Bue82, Daw82, Efr87, Efr94, Efr96, Fab91, Gew82, God82b, Goo91, Hol86, Kit87, Las94, Li91, Sha82]. authors [BS87, BH88, BF85, CB87, DG93, DH83, HHM88, HL81, LL87, LF82, LV88, MB88a, PT82, RHS93, SCM91, TW87, WMR91]. Autocorrelated [BD91b, HZ90, Kob85, ZB86]. autocorrelation [Ali89, BHS81, DP89, KD87, TT84]. Autocorrelations [CH95b, RT96a, Chi74]. Automated [EL97, Joh96a, OR98]. Automatic [Cho93, GKK91, OYR86]. automatically [HHM88]. Autoregression [DM95, TS90]. Autoregressions [SG97]. Autoregressive [AR90, Ahn97, BDR92, CT93, LRM96, LL97b, Qui98b, SL93a, Sch96, SS90b, Ter94, YR95, Bha81, BHS81, BA83, Chi74, FH81a, FH81b, Har84, KY85, Mac86, Mac87, Spl83, Sti87, TT84, Tsa89]. Auxiliary [Hig98, PS99, WPD98, Isa83, Smo82, Wri83]. Average [AI95, BDR92, HMT92, LL97b, SL93a, TR97, YR95, HS89a, Spl83]. Averaged [CHR94]. Averages [RP96, Mac86, Mac87]. Averaging [RMH97, RB90]. AZT [RG94].

B [Daw82, Efr94, Sim87a, HS98c]. B-Spline [HS98c]. backward [BB88, Sti88]. Bad [GS97]. Bahadur [Rog88, Ruk93]. Balaena [RZ98]. Balance [GJDT98]. Balanced [Ber88a, CD91, Hem79, KW96b, Pen97, RS96, SCC98, Bro83, CS89, HS89c, MS88a, PH86]. Ban [PB92]. bandits [Fox74]. Bands [BD93, ES93, FS95a, HMY97, Owe95, BK89, HB88, Kan88, KSY85, Nai87a, Pie87, TT86, Uus83]. Bandwidth [BGH93, FJ90, HMT92, JMS96, OP96, OR98, PM90, RSW95, Sch95a, WW93, jW97, Sta89b]. Bandwidths [Rup97]. Barndorff [Bue82]. Barndorff-Nielsen [Bue82]. Bartlett [Riv86]. Base [Pea91, SS90a, WW91b]. Baseball [Alb94, AASM93]. Based [BAP95, BPL97, BY95, BW94, BO97, BF93a, CM93, CS93, CH95b, DM90, DM93, DM96, Efr92, Fan96, GS90a, HYP97, HMY97, Hol99, Hös94a, Hua97, Hun94a, Joe95, Joh90, Kai94, KBB93, KS94, LRR91, LQT97, Lin91, LS93, LT96b, LS97, MSH90, MSH93, McW90, Mee90, Mur95, OP96, Pap96, Pes93, PMW93, SDW98, SJ92, SG97, Tab90, Tan93, WH90, WZ97b, VS93, Atw84, Bha85, BS86b, BHT88, CSS89, CS87, DR98, DvdM81, HM88a, HM83, LL87, LQ85, LL81, MH82, MC84, OW82, Ran89, Ros84a, Sch86b, SW82, Sta89a, Ste81, Tar87, TH82, Wat87, CD96, HT82]. Baseline [Peñ98]. Basis [Kra90]. Basu [Leh81]. Bathtub [Gla80]. Batting [CB94]. Bayes [AC95a, ALM97, BP96, BM99, BPR95, BD91b, CG90, CH83, CS96a, CZ77, CLWZ95, DL81, Del89, DKRW97, DH83, Efr96, Gel96, Geo96, GM86a, GL87, Gho92, Gol83a, Goo92, Hsi97, KR95, LL87, LCG^{+}92, LR97c, Lou84, MV85, Mor83, Mor96, NS93, NM83a, NT92, O’H87, Pen97, Pet92, Rag93, Rei96, ST96, Sob91, Sob93, Tom94, VW95, Vid98, WF82, VS93]. Bayes-like [WF82]. Bayes/Non [Goo92]. Bayes/Non-Bayes [Goo92]. Bayesian [AG83, AG85, Alb88b, AC93, Alb97, AC97b, AG96, AP89a, BH93, Ber82, BD88, BB89, BS93, BH88, Ber91b, BWC98, Buc95b, CO90, CS91, CMW97, CKLH92, CB87, Che94a, CKM97, CGM98a, CGM98b, CS86, CRS^{+}91, CNWW98, CV95, CCC96, CMMY91, DKW97, Daw82, DKS97, DSM98, DJK87, EGMP93, EGG95, EW95, EGGS97, FOS97, FS98b, Fon92, FR99, FS94, GHRPS90, GSTM92, GNK96, Goe83, GB98, Hal73, HW95a, Hay94, Hei95, HH91, IL91, Jay79, JW88, Joh96a, KR96, KW95, KP92, KKT98, KLW94, KY96, Lav91a, LWW91, Lav91c, LRM96, Len88, LHT89, LB94a, Liu94, LNL95, Liu96, MS85a, McC98, MV91, MB88a, MR97a, MG97a, MR97b, NSS97, NCC96, Nyc88, PJT96, PSS93, PS96, Per97, PS94b, Qui98a, RMH97, Rig97, RW97, Rue95, SR94, SN96, Sev94, Sha89b]. Bayesian [SW80, Sha83, SD97, SS82b, SK97, Smo82, SM92, SKF98, SY95, SHGL96, SCCF93, Tam88b, TW98a, TW98b, Tom96, TS96, VM83, Var98, VK92, Vil90, Wak96, WB96, Wei97, Wei95, WHM85, Wes95, Won87, WK98, Zel86, Zha98b]. Bayesians [BD91a]. Be [PH90b, SN96, BF83, BHT88]. Bearings [ASL96]. Bearings-Only [ASL96]. Behavior [AL94, Cha93, GS93, Hal94, Man90, Rom90, Fox74, MY95, Nai84, SM84a, Sch84b]. Behavioral [EGG95]. Behrens [JW88]. being [NM83a]. belief [Gol86]. Beliefs [SHBF94]. Believe [CB94]. Benchmarking [CCD97, HT87]. Berger [Bor87, MB88a]. Berkeley [Met50]. Bernoulli [MO85, Tam85]. Best [AP87, HV86, Man90, Rol94, SH81, SC96a, dCP99a, But84, CC74, EH83, EH84, Esc86, Hsu82, HE83, Sam85b, Tam88b, Wil84]. Best-Case [Man90]. Best-Choice [SC96a, Sam85b, Tam88b]. best-fitting [But84]. best-linear-unbiased [Esc86]. bet [SG83]. Beta [CW91, Geb71, SS80]. beta-integral [Geb71]. Better [Cha91, Efr87, RS90, SN96, Ruk87, Tam85, Wil84]. Between [BDR92, GR97a, GR97b, HH90, HHS97, HPS98, HH94, HHH92, Krz79, Krz81, LS91b, Llo90, Rig97, SS90a, BJB95, CR82, Gew82, Gew84, HS89c, Kob86, dAJ86]. Between-groups [Krz79, Krz81]. Beyond [LT94b, Tuk85]. Bias [CDW93, Isa74, JS97a, LT97, LB96, MW98a, MW98b, O’N94, Pea61, PMW93, RW98, Rup97, SS90b, WLGC98, CSC86, MZ89]. Biased [BBH96, GS91b, ST87, Efr86b]. biasing [Oma81]. Bickel [BC82a, Hoa73]. bidec [KT88]. Bifurcation [KR93]. Binary [Agr97, AC93, BO95, CLWZ95, Coo96, FSE98, FLZ96, Fyg97, GM96, HL98, IRC98, LLR95, LR97a, MSML97, McC94, NCC96, OK96, PHL98, QN98, SB95, WK98, ZDL96, Zha98a, AP83, APS81, BS84b, Kee82, LZ89, Man88, Min87, Ros89, TH82, Wu85]. binary-type [APS81]. Binned [Min98]. Binomial [BS83a, CB94, Che90, CV95, Dea92, Hal94, Pes93, SK90, VM96, AG85, BD81b, Bly86, CL85, De 83b, Fis79, Geb71, HPE81, KT94a, LL86, MT49, Rel72]. bioassay [MS88b, Sch84a]. bioequivalence [BCH95]. Biological [GMP97, LNL99]. Biomedical [GS90b]. Birth [Hof92, Sti86]. bispectral [HW88]. bit [Che71]. Bivariate [Ame74, ALHH92, BHNO92, CR92, CM91a, Die82, ES99, FL95, Gür96, Hay94, LK95, Oak89, ON89, Pru93, Ryu93, SV93, SK97, WJ93, WY96, WKL95, BB74, Cha84, GR93, KR82, MO85, Sar87]. Blind [LC95b]. Block [JMS97, HS89c, Hsu82, Jac89, KR84, YBN85]. Blocking [BS91]. blockmodels [WW87a]. blocks [CS89, IHC84, SCM91, Vos86]. Blood [HL94, MR97a, MR97b, OO97]. Bodies [PC96, Riv98]. Body [DMLC97]. Bomb [PSVS92]. Bone [DwSH94]. Book [Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano99a, Ano99b, Met50, Gei48]. Bootstrap [Ber90b, BBH94, Bre92, Bur94, CHP96, DM95, DT87, Efr90, ET97, FJ90, FH89, FH90, GMCM96, HH90, JB92, LT96b, LS97, MBL98, Mar90, NG94, Pap96, PC96, PR94, SS96b, Sha96, Sti85, SW97, SGP98, TS90, Zel93, ZD97, dHY93, Ber88b, BY88, BK89, Efr87, Efr94, LL87, Rom88, Sch85]. Bootstrap-Adjusted [ZD97]. Bootstrap-Based [Pap96]. Bootstrapping [Akr86, DC94, FP84, HB88, SW91a, Sim86]. Borges [Geb71]. Born [HHH92]. Borrowed [Pri96]. Both [Mad59, WR95, CM86, KT94a, Tuk79b]. Bought [GS93]. Bound [BF90, Sat95, BE85, Kim86, Oma81]. Boundary [CK92, Erk94, HW92a, Hsi97, Cha88b]. Bounded [HR94, MH94a, Ric97, Sam85a, Tab90, AP87, CH93, DH93, Hoe63, Hub83, KW82, KSC89, MH90, Rup85, dJdWW88]. Bounded-Influence [HR94, Tab90, Sam85a, Hub83, KW82, KSC89, MH90, Rup85]. bounded-influence-regression [dJdWW88]. Bounds [BP97, CS96b, Kar81, Spu99, WK92, Del89, DH93, Gla89, HHO88, KK84a, Kob86, Lea81, Nai83, Nai87b, YM85, YM86]. Bowhead [Buc95a, RGZ95a, RGZ95b, RZ98, Sch95b, TL95, Wol95]. Box [Jen91, BDM97, ST99, Spi82, Tay85]. Box-and-Cox [Spi82]. Box-Type [BDM97]. Brain [RTF92]. Brains [Met50]. branching [SW87]. Brazil [Sas97]. Breakdown [Alb94, CMNS99, EM92, HM97, HSP90, Hös94a, MH94a, MY91, SW95, SR92, Str93, CH93, Sim89, YZ88]. breakdown-point [YZ88]. Breast [GWG90, Gra92, LZ98]. Breast-Cancer [GWG90]. Breiger [HL81]. Breiman [LV88]. Bridge [MS96]. Brief [Goo92, JMS96]. Broken [GMP97]. Brown [Rot83]. Brown-Forsythe [Rot83]. Brownian [Ste94]. Brownlee [TW46b]. Burden [ZDL96]. Burst [EP94]. Bursts [LMR96]. Business [The93].

C [Bow87, Met50, Sim87a, Tuk49b]. C. [God82b, MB88a]. C.-M. [God82b]. Cahoon [Bow87]. Calculating [GS90a]. Calculation [SP91, TW87]. Calculations [MLRR92]. Calibrated [PB92, Daw82]. Calibrating [BR95, Ber90a, Loh87, Oak85]. Calibration [CDW93, DS92, JK96, MZ97, Thé99, VM83, ZD97]. call [Ros84a]. Callbacks [PMW93]. Camouflaged [OO97]. Campaign [MSL93]. Campbell [Mar80]. Can [PH90b, BHT88, Hjo94]. Canadian [SR95]. Cancer [BC90, Bak98, BHB96, BSB98, GWG90, Gra92, LZ98, MPCB95, DH83].

Canonical [BY95, GH86, GF96a, Gil86, GH88]. Capital [Gro90]. Capture [CM86, Pol91, KMT99, MSW^{+}83]. Capture-recapture [CM86, KMT99, MSW^{+}83]. Capturing [Soo94]. carcinogenesis [MC84]. Carcinogenicity [HZS^{+}90, MV88]. Card [RCH94]. Cards [McP42]. Care [CS91, DG99, NGG97]. Carlo [CL95, JTB81, Ros95b, BBGL97, BS96b, CPS92, CC96b, GT95, GRS98, GJR98, Hig98, Hur85a, JTB80, Joh96b, Joh98b, LC98a, LC98b, MU49, MP95, NG94, Ode91, OB93, Ros95c, SMP95, Sha89b, TV85, WT90b]. CART [CGM98a, CGM98b, DSM98, KKT98, Zha98b]. Case [AGM98, Bak98, BHB96, Bin98, BWC98, Bre98, Bre96, CGL93, CWW95, CNWW98, DN98, DC94, DMLC97, Efr97, FS95b, Fyg97, GKL98a, GKL98b, GJDT98, GWG90, Hei91, Hod94, IRC98, Jud98, Kad98, LZ98, Man90, Mei94, Per97, QH98, RCL96, RSV95, San98, Sev91, SHBF94, TFAH94, Akr88, Gui83, Hoe84, Pic87, SM84b, Wil84]. Case- [FS95b]. Case-Control [BHB96, Bre96, CGL93, CWW95, RCL96]. Case-Deletion [Per97]. Cassel [God82b]. casual [Kru88]. Catch [Mil92]. Categorical [BZ92, Con92, Cro95, DS97a, dE97, EGGS97, Fyg97, GB97, GH94, GH95, HZ98, LA94, Lan96, LL97a, PB94, SBD97a, SBD97b, Con89, DJK87]. Categories [AMP90, Clo82]. Categorization [Roj96]. Category [GS93]. Causal [AI95, AIR96, GM96, Hec96, Mof96, RG96, Ros96, SG97, Hol86]. Causality [BDR92]. Cause [AKM94, GFT^{+}93]. Cause-Specific [AKM94, GFT^{+}93]. Caused [BL93]. CD4 [LCG^{+}92, TDW95]. CDF [WBB93]. Cell [Bri92, Cox95, HS94, LCG^{+}92, PB98, BOP88, Hab88]. Censored [AML95, AB97, CW94a, Da̧b90, DS97a, DG92, FG94, FSE98, GB97, GC91, GWG90, GL96, HW95a, Har92, HP92, JMV94, KS97, LLE93, LL97a, LYY99, LGWC93, LWT94, Pru93, SDW98, SBD97a, SBD97b, WH90, WT90a, WZ97a, WY96, Yan97, Yan99, AA87, Alb88a, BR83, Bha85, BC82b, CC74, DJK87, Esc86, GY93, Gui88, KW85, LaR86, MB82, Pad86, SM83, SM84a, Tsi82, VJ88, WG83, Wei84, WP85, Zac86, ZB86]. censored-data [VJ88]. Censoring [Har90a, HR97, SW97, BM81, BM83, SY84]. Census [CRS^{+}91, EKT89, GBN96, Joh90, RZ98, Sch91a, SS95b, CT98, Tuk85, Wol86]. centenary [Sti88]. Central [WK94, Mar81, ZC88]. central-mixture [ZC88]. Century [Ket97]. Certain [BHNO92, Bur94, Dha93, SM84a]. Chain [BBGL97, BM95, CC96b, FK94, GT95, GRS98, GJR98, Hig98, Joh96b, Joh98b, Lou96, MTY95, Ros95c, RP96, Sam90, Ros95b]. Chains [GJR98]. Chairman [Tuk50a]. Change [AL95, BH93, For94, Wan98a, HRR81, KC93, SW86, WG83]. change-of-variance [HRR81]. Changepoint [Bro97, CH97b]. Changepoints [CG97]. Changes [DDV99, IT94, JD91, KM96a, LD91]. Changing [Nor90]. Chaotic [Ber91b, Lel94, MEGN92]. Characteristic [Llo98, YD84]. Characteristics [GBN96, SLS97, SRR95]. Characterization [MLMT94, SS88, Vos93, JG83]. Characterizations [Chr91, Gla80]. characterize [Sto82]. Characterizing [Hof92]. Chart [RP96]. Charts [Liu95, LT96b]. Checking [Pap96]. Checks [SGP98]. Chemicals [HZS^{+}90].

Chemotherapy [Min93]. Chi [HP92, Lee87, Nai87c, Ryu94, Wil50, Ame74, Fay85, Hab88, KOS85, LaR86, Ole85]. Chi-square [Wil50, Ame74]. Chi-Squared [HP92, Lee87, Ryu94, Fay85, Hab88, KOS85, LaR86, Ole85]. Chi-squared-type [Nai87c]. Child [GR92a, Sas97]. Children [SB90, WHGA96]. Chilean [EV92]. Choice [BD97, BGH93, FJ90, HMT92, SC96a, Sch95a, Sev96, Wil50, Yam90, Bro81, Car82, GC82, Hur85b, Par93a, SA86, Sam85b, Sil84, Tam88b]. Choosing [DC94, KP91, Nyc91, Oma81]. chosen [HHM88]. Chow [Thu85]. Christ [Sti86]. Christian [BH88]. Churchman [Tuk49b]. Circadian [PF97]. City [DXP92]. Claims [Coo91, GHCF94]. Class [ES99, FKYZ99, For92, Jen91, Joh90, Lin91, LCG91, OKS97, Xio95, DM88b, HPP86, LZ89, NS83, NS86, OW82, Ste81, Tsi82]. Classes [BL92, CL92b, GMWZ99, HS98a, Wil50, KOS85, Wal86]. Classical [Bra90, FOS97, GK93, LB94a]. Classification [CD89, FS94, KP92, LT96a, Pre96, Sch84b, SR95, Sta91, Zha98a, Isa74, Kor81, LV88, Ruk84, Sin86]. classifications [Agr83, Clo82, Gro81]. Classified [Goo96, Goo91]. Classifiers [Moj99]. Clevenson [Tsu84, Tsu86]. Clevenson-Zidek-type [Tsu84, Tsu86]. Clinical [BO95, CCC96, DLB91, EF91, EFS92, GFT^{+}93, GM96, LD92, LB94a, Roc95, TFAH94]. Clonogenic [Min93]. closed [Sch87a]. Closeness [MKSB90, GS89, RHS93]. closure [Tuk50a]. Cloud [CS93]. Clumps [HL97]. Cluster [DD90, FH82, Nau65, NM98, RSY93, Tam95, Tho90b, BM85, MS85a, Roy86, TR87]. Clustered [CR92, GR92a, HS98b, HZ96, CM86, Ros89]. Clustering [BR92, DR98, FL95, KST98, LJK93, MNS99, RMS90, Won82b]. Clusters [NW97, Har81, PFV89, Won82b]. Clutter [BR98c, DR98]. Coale [YP92]. Coale- [YP92]. Coarse [HR90]. Coastal [PM97]. Cochran [CMT55, Riv86]. Cod [NSS97]. Codes [CRS^{+}91]. Coefficient [CT93, CTC96, SW98a, WCH98, Cha84, GH87a, Mui85, Sen86, SG82]. Coefficients [CW94a, Pag90, RRZ94, Tho90a, Hod88, Kob86, Loh87, NM83b]. Coherent [Mei94, Ger82]. coincidences [DM89]. Cointegration [Ahn97, BY95]. Collaborators [Ano95a, Ano97a, Ano98c]. Collapsed [Liu94]. collecting [APS81]. College [MH94b]. Collinearity [FM92, SB97]. Column [Ye98b]. Combination [SGC95]. Combinations [SL93b, DH93, Flu85]. Combined [AA87, Alb88a, BM81]. Combining [CV95, DKRW97, DPWH99, Efr96, FK99, Gel96, Geo86, Geo96, GS91b, HM85, JK96, LT96a, LB94b, LL86, MH99, Moj99, Mor96, Rei96, Sch86a, ST96, ZM93, DH83, Gas85, Kin85, LQ85, LL81]. Comment [Agr98, Bin96, Bin98, Bjö78, BB98, Bor87, BC82a, Bre98, Bro92, Buc95a, BK99, Chr99, Cre97, DH91, DSM98, DS97a, Elt96, Fab94, FZ98, Gav78, Gel96, GB97, Geo96, GMS97, Han99, He99, Hec96, Joh98a, Jud96, Jud98, Kad98, Kie78, Kne98, KKT98, Kot89, Lan98, Li98b, LL97a, Mal78, Mar80, McC98, MT98, MPS98, Mof96, Mor96, Por97, RS98, Rei96, RG96, RW98, Ros96, San98, Sch95b, ST96, SC99, Sim78, Sim87a, SHL98, SF97, Ste99, TV85, TL95, Var98, WW98, WPD98, Wol95, Zha98b, Efr94, JV85, LF82, LV88]. Comments [EKWS91, Tuk79a, Tuk85, BS87, BH88, Bow87, Bue82, CB87, Daw82, DH83, Efr87, God82b, Goo91, HW94, HHM88, HL81, Kit87, Las94, MB88a, Oak85, PT82, SCM91]. Common [Agr98, BB98, DD90, Flu84, GF96a, HPS98, MPS98, RN97, RV98, SW98b, SW98c, VHE91, GG87, Mat83, MPG85, MPG86, Noe87, SW88]. Comparability [GH95]. Comparative [CC96b, HH91, Ols74, WM91, BT81]. Compare [Llo98]. Comparing [AKM94, Bac90, Boo86, Bra90, CS92a, DDV99, Efr84, Hol95, Jac93, MH99, Nai83, PW96, SRR95, Sob93, SK90, TL97, CMST88, Jac89]. Comparison [AD90, Bur94, CGG85, CG92, CHW92, FT94, HA94, HA95, IHC84, JD91, JS97a, Kob85, Kul95, LTW96, Lin74, MV88, MKSB90, MRW87, PM90, Sev91, SN90, SKF98, SO94, Tin65, WJ93, WPK99, CS85, CR82, Krz79, Krz81, Las94, LD91, Wes87a, WPSL84]. Comparisons [GB98, Kun90, LTW96, ME96, NM83b, PH86, Pen97, SMF94, BB88, Bof85, EH83, EH84, Fli84, Hay89, HVS84, HE83, Rea84, Rin83b, SI85, TC72, TR87, Uus85, Wit88]. Compatible [AP89b]. Competing [FW96, FG99, HW95b, Kra90, Mei94]. Competitions [Est92]. Complementary [Cox95]. complete [IHC84, KR84]. Completely [Pai97]. Completeness [Kre86, Leh81]. Complex [ALM97, BP94b, GK93, Riz92, Sit92, SR96, SRSD94, BS85, Fay85, RW88]. Compliance [Bak98, BP97, CCC96, EF91, GM96, OMF^{+}93, SCCF93]. Component [Sch91b, SHGL96, Tar99, Wes95, Ger82, HV86, Pud82, SHS86, Wes87a]. Components [Ahn97, BR94a, CD91, EL90, GP92, HA94, HS94, LB96, McC94, Roe94, She93, Vel98, Wee91, AH81, Bro81, ELR87, Flu84, HA95, Khu81, KR86, Krz79, Krz81, Loh86, MS88a, Pri74, Wes87b]. Composite [CS96a, DR91, HL98, KL97, LKM98, SS99, Wel90b]. Composition [NSS97, Hil70, SS82b]. Compositional [BR94a]. Compound [Col87, CW94c, WR94, AY82]. Compromise [Eas91, Goo92]. Computable [Wes87b, WR94]. Computation [Dha93, Hab95, KY96, MG97a, Str93, VHE91, WZ97a]. Computational [Alb88b, BRB76, DC77, HH91]. Computations [Ber77, BS98, Efr90, Hur85a, Liu94, WK94, WT90a, LLS87]. Computed [SC94, Wei97]. Computer [CMMY91, KR76, Lon67, Wam70, Ye98b, Fis79, KP81b, KP82, Mar84, Tei65]. Computerized [BWC98]. Computers [FMR62, Che71]. Computing [DKRW97, ES90a, HMP87, Hir92, LK92b, MBR91, MPG85, MPG86, VW95, WK92, Lin74]. Concatenated [AMCS98]. Concave [GMPS99]. Concavity [Pra81, Pra82]. Concept [Soo94, PT82]. Conception [HW90b]. concerning [BS84a, Ber89]. Conclusion [KSS96]. Concomitant [DM88b]. Concomitant-variable [DM88b]. Concurrent [KL90]. Condition [Ber77, Cas85, HW90a]. Conditional [BBGL97, FK99, FP95, For94, GJ99, GFT^{+}93, GH94, HWY99, HK92, Joe95, Joh92, KA95, KT94b, LL97b, MBR91, Mar64, PC97, Ros84b, Sch96, Sev95, Soo92, SC97c, Tsa87, Val87a, Val93, WM98, Xio95, AP89b, Coh81, CGM83, Con89, FW88, Gew84, HHO88, KT94c, SC85]. Conditionally [AC97b, CMW97, KSC89, Riz92]. Conditioning [Kot89, Mil92, Rob87]. Conditions [Ros95c, Bie83, Ros95b, SL87]. cone [DL88]. Cones [BM96]. Confidence [BS96a, BHS98, Ber90b, BB94, BK89, BS96b, BD93, Bur94, CG90, CJ95a, CJ95b, CJ96, CH95b, CS96b, CW90, CHP96, ES93, FS95a, FHJW96, Ger82, Gri92, Har90a, HH94, HMY97, Hol99, HB99, KSY85, LH91, LL81, Llo90, Mar90, MZ97, Mee90, MB82, Mur95, Owe95, Pan97, PC96, Pes93, Roe90, Sat95, SG95b, Spu99, SW97, TT86, Tho90a, TF90, VHE91, WCH98, ZD97, Ber88a, BS83a, Bly86, BCH95, CH83, Cla87a, DT87, Efr87, FMD82, FH89, FH90, FS87, HB88, IS87, Joh78, Kan88, Khu81, Kim86, LL87, Loh87, Mal82, MSW^{+}83, MPG85, MPG86, Mey87, Mor86, Nai83, Nai87a, Nai87b, Nyc88, Oma81, PFV89, Pie87, Ric82, Sch85, SG95a, SI85, Uus83, Uus85, Wee93, Wee94, YM85, YM86]. Confidentiality [Fel72]. configural [Mor86]. Confirmation [Fun93]. Conflict [Luc93a]. confounding [SB81]. Congenital [SHBF94]. Congressional [Sch91a]. Congruential [FM82]. Conjecture [SMF94]. Conjugate [CV92, GN82, GPS95, JJ93]. Conjugate-Gradient [GN82]. Consequences [GK90, Whi81a]. conservative [Ben83]. conservativeness [Wei81]. Considerations [PSVS92, VM83]. Consistency [GS98a, Har81, Tar99, Bie83, Sha89a]. Consistent [FG95, Fel74, Jia98, LB93, Riz92, TT84, ZL95, Wit88]. constancy [Nyb89]. constant [Nai83]. constant-width [Nai83]. Constrain [BHS98]. Constrained [CC96a, CW94c, GSTM92, Gey91, Gho92, KP81a, MM93, RMS90, Cha84, Sch84a, VW87]. Constraint [RS93]. Constraints [AG96, HT97c, RG93, Wes97b, DL88, HB86, Nai87b, Wes98, Wol87]. Constructed [LK95]. Constructing [FHJW96, VY80, Ric82]. Construction [Hay94, Jac93, JMS97, MST97, WW91a, Wel90a, CJ88, SB81, Vos86, Wat87]. Consumer [GS93, Zie90]. contagious [Hol87]. Contain [BS96a]. Containing [Pan97]. Containment [GLB98]. Contaminated [Gle93]. Content [CM91b]. Contents [Gag43]. Context [DF98, BM85, Gut82]. Contextual [Pre96]. Contextually [WM91]. Contingency [AMP90, Alb97, BS97a, BS97c, CS92b, DD90, Fri94, Goo94, KP97, KA95, LW91, Qui98a, RG93, YF90, ACK87, AG83, BOP88, CG84, Coh81, CGM83, CG82, Dal82, Gas85, GH86, Gil86, GH88, GKR88, Gro81, GSS89, Koe86, MP83, MPG85, MPG86, Nai87c, PH81, Sar89, Sim74]. Contingent [CH91, SG83]. Continuous [CR92, FL95, GJR98, KC96, BB74, Poo74, Sar87, Sto88, TH82]. continuous-time [Sto88]. Continuum [BS94]. contour [Har87]. Contoured [CN94]. Contrast [Wes97b, JK89]. Contrasts [Mar92, Spu99, MRW87, Wes98]. Contributions [BR94a, SRR95]. Control [Ano90, BC90, BHB96, BKP90, Bre96, CGL93, CWW95, CNF95, CS92a, FIB90, FS95b, GJ94, GHMW94, GS94, HT90, KA91, Kra90, Liu95, LT96b, LCW98, RCL96, RP96, SMP95, SN90, TL97, CMST88, Fli84, Jac89, MRW87, Spu88]. controllable [HVY83]. Controlled [GM96, LB94b, VY80]. Controlling [Owe94]. Controls [IRC98, Jac93, HY86]. Convergence [BPR99, Cha91, CC96b, DG92, GJR98, Joh96b, Joh98b, Min98, Ros95c, ZM95, CH88, Ros95b]. Convex [FHTW97, Gey91, Har87]. convolution [Dye82, Spr83]. Coordinate [Lea81]. Coordinate-free [Lea81]. Coordinates [Weg90]. copulas [GR93]. Correct [WPD98, FM83, Kim86]. Correcting [MW98a, MW98b, PMW93, RW98, Ye98a, FH78, FH81c]. Correction [BK86, CMS87, Cla89, FH90, FH81a, FH81c, HM89, HA95, JTB81, KT94c, Krz81, LB96, LB94c, Mac87, Mar90, MPG86, NS86, Ros95b, SKC94, SG95a, Sta94, SCW97, Wee94, WM81, Whi81b, YM86]. Corrections [Rin85]. Correlated [Alt90, BR94b, DLW95, KPR91, Llo90, ML94, QH98, Wan98b, ZT97, Hay89]. Correlation [AP96, BY95, BJB95, CTC96, DBB^{+}94, Dwy45, GS98a, OV95, Sch94a, Sch91b, WEST91, Bin82, BF85, GH87a, Gil86, KY86, Mui85, PH74, SG82, VJ88, dAJ86]. Correlations [Ber91a, BG91a, Gri92, Owe94, Wes97b, Wes98]. Correspondence [Goo96, RG93, GH87b]. Corrigenda [EH84, HG83, de 83a]. Corrigendum [DD82, KP82, Law82, MW82, MT82, Pra82, RC82, Sch82]. Corrupted [OP96]. Coseismic [KM96a]. Cost [Bak98, GJ94, GLB98, DD81, DD82, Sam85b]. Cost-Effective [GJ94]. Cost-Effectiveness [Bak98]. costs [FE88]. Count [McC98, MR97a, MR97b, PB98, TW98a, TW98b, Var98, Sim87b]. Counting [GM86b, CSC86]. Countries [GMWZ99]. Counts [CL95, DB97, Gro90, PB98, TDW95, BOP88, Hab88]. County [CCG^{+}97a, CCG^{+}97b, Cre97, GMS97, SF97]. Coupled [Joh96b]. Coupling [Joh98b]. Coupling-Regeneration [Joh98b]. Covariables [RCL96]. Covariance [AS90b, CD91, CD95, Che70, CLT96, GS91a, Hou95, JB92, Rig97, Roc92, SG92, Sie93, Ver93, WHG98, XLR91, YB97, BK84, CGG85, HVS84, Kar81, Lee88, Mar64, MB88b, Rei82, SB87, TB86]. covariances [Mar64]. Covariate [Fon92, FS95b, HHS97, HAS97, JT97, LY93, PF91, SH90, WWZtO97, WK88]. Covariate-Matched [HHS97]. Covariates [BL93, BZ92, BD93, CGL93, CN94, FSE98, GWG90, HDG97, HH96, IRC98, Neu98, Roc95, WB96, Zuc98, vR98, Kar87]. Coverage [CL92b, GWG98, Mar90, CT98, Wol86]. coverages [Yan84]. Cox [Bue82, BD93, Bur94, DwSH94, Gil84, LW89, Lin91, LY93, Peñ98, Sas93, ST99, SH90, Spi82, SC85, Tay85]. Craig [Sti86]. Creation [DLW95]. Credence [O’H90]. Crime [Sta91]. Criteria [BS91, Jen91, MT95a, ZM95, Zha92]. Criterion [Bly93, HZ90, KW95, MNS99, Nef82, RHS93]. Critical [JB92]. Criticism [CPP96]. Critique [SC80]. Cross [ET97, Goo96, HY98, KAT91, NWGP84, PC84, RFB97, SBS94, Sha93, Tho91a, Wan91, Agr83, Clo82, Efr83, Goo91, GA82, Hur85b, KY86, Par93a, ST87, Sil84, Sto82]. cross-classifications [Agr83, Clo82]. Cross-Classified [Goo96, Goo91]. cross-correlation [KY86]. cross-section [GA82, Sto82]. Cross-Sectional [Wan91]. Cross-Validated [Tho91a, NWGP84]. Cross-Validation [ET97, HY98, KAT91, RFB97, SBS94, Sha93, PC84, Efr83, Hur85b, ST87, Sil84]. cross-validatory [Par93a]. Crossed [BR98a, BR98b, FZ98, Kne98, Lan98, RS98, WW98]. crossings [Cha88b]. crossover [Gui93]. Crystal [SW90]. Cumulant [PSS93]. Cumulative [AKM94, BK99, BS98, Chr99, Han99, HLC92, IT94, LKCH99a, LKCH99b, SC99, Ste99, SW97]. Cumulatives [RV90c]. Current [BF92, RT96b, SK93b, vR98, NS93]. Curtailment [LYY99]. curvature [CW85]. curvatures [Cla87b, Cla89]. Curve [AC95a, AGM94, FHMP96, GK95, MP95, OO97, RS90, RS93, Wan98a, Efr88, HRR81, Mül87, Rei82]. Curves [BR92, BR98a, BR98b, CM97a, CW90, DBB^{+}94, DLW95, FZ98, FL98, HYP97, Kne98, KA91, Kul95, KW97, Lan98, Lee96, Llo98, Pru93, RS98, WW98, Abr88, CC83, HS89b, Lee88, SW87]. Customer [BZ99]. Cut [dE97]. Cut-Points [dE97]. Cycle [ZWWB96]. Cyclic [CH95b]. Cyclical [Wes95].

D [Gol63, Hoa73, Sim87a]. D. [Bue82, Daw82]. Daily [Gro90, NCDF96]. Daniel [DH83, Gol63]. Danish [HHH92]. Darkly [Bar91]. Darling [SS87a]. Data [AML95, AB97, AC93, Alb94, Alt90, AC95b, ASL91, ASL96, AD90, AS90b, BPS96, Bak98, BF92, Bin96, BW92, BO97, BO95, BZ99, Bri92, Car94, CL92a, CH95a, CH97a, CS93, Cha96, CTC96, CCD97, CW94a, Che94b, Coo91, CCC96, Cro95, DwSH94, DH91, DAO94, Daw95, DS97a, DM97, DLW95, DG92, DA96, dE97, Efr98, EP94, Elt96, EV92, Eub97, FL98, FSE98, Fay96a, Fay96b, FHJW96, FHTW97, For92, For94, GFT^{+}93, GHRPS90, GSTM92, GB97, GC91, GH94, GJP96, GWG90, Goo96, GRB^{+}95, GK93, Gra92, Gra95, GL96, GR92a, GS95, Gür96, GT91, HS98b, HW95a, Har92, HL98, HW90b, HYP97, HR90, Hei91, Hem79, HLC92, Hir92, Hod94, HP92, HOS95, HL96, HW95b, Hur85b, Ibr90, IRC98, JD91, JMV94, Joh96a, Jud96, JS97b, KL97]. Data [KL99, KM91b, KSW92, KP92, KLW94, KC96, LT97, LI98a, Lav91b, LR96, Led94, LWY93, LLE93, LS95, LRR91, Li92, Li95, LQT97, LL97a, LYY99, Lit93, LS93, LGWC93, Liu96, LS97, LWT94, MZ94, Man90, Mar92, McC94, McC98, MFS84, Min98, ML94, MR97a, Moy99, MSK96, MR97b, MST97, NP95, NT95, Niu96, O’S93, OP96, OK96, Pai97, PM90, PB94, PB98, PF91, PC97, PSVS92, PML98, Pre96, Pru93, PHL98, QN98, Rao96, RRZ95, RR95, RW96, RT96b, RB90, Rub96b, Sas97, SDW98, Sch99, Sch96, SGC95, Seg92, SLS97, SS96b, SCC98, SS99, SBD97a, SBD97b, SC94, SJ92, Shu97, SG96b, SR95, SD97, Sit92, SEH95, SM92, SHH78, SM90b, SL98a, SL98b, SKF98, SC97c, SK93b, TCS94, TW98a, TW98b, TDW95, Var96, Var98, Vel95, VFR97, WH90]. Data [Wan91, Wan92, WSJ92, WWZtO97, Weg90, WT90b, WT90a, WSLP90, Wei97, WZ97a, WY96, WW91b, WHG98, WKL95, WCH98, XXP93, YKB97, Ye98a, ZDL96, ZLRS98, vR98, AP83, Ala84, AA87, Alb88a, AG85, AP89a, BR83, Bha85, BS84b, BC82b, Bro83, BP86, CR84, CC74, CR89, Chr87, DJK87, Efr94, Esc86, FH89, FH90, Goo91, GA82, GY93, Gui88, HW86, HVY83, HR84, HC84, Hwa86, KL85, Kee82, KW85, KK84b, KT94a, KA86, Law87, LD91, LB94c, LB94d, LS87, Min87, O’N84, Pad86, RW88, Ros89, SM83, SS86, Sim87b, Spe85, SG82, Sto86, Sto82, TW87, Tuk79a, Var88, VJ88, Wan89, WP85, WLW89, Wit88, Wol86, Wu85, ZB86]. Data-Dependent [NP95, WSLP90]. Data-Driven [Efr98, KL97, KL99, Led94, PM90, Hur85b]. Dave [CB94]. David [Bue82, HHM88, Tuk49a]. Dawid [Oak85]. Dealing [PF91]. Death [Hof92, SKF98, Sti86]. Death-Penalty [SKF98]. Decaying [Ber91a, BG91a]. Deciding [HPS98]. Decision [HH94, LB94a, SM90b, Jay79, Sha89b, WF82]. Decision-Theoretic [LB94a]. Decisions [Kad90, SA92]. Decisive [GKB98]. Decomposition [BC96, Har85, KO96, Tak83]. Decompositions [Amm93]. Deconvolution [CT97, EL97, LC95b, MO98, OO97, WW93]. deconvolving [CH88]. decreasing [SM83]. Default [BM99]. Defects [Sho93]. Definite [JS94]. Definiteness [GL92]. Definition [SW95]. Deformable [AGP91]. deformation [Sto88]. Degrees [OR92, PWM92]. Delaney [Bro92]. Delay [DXP92]. Delayed [BBH96]. Delays [Har90b, Wan92]. Deletion [Per97, Vos86]. Demand [HK93, HK92]. Deming [Man94]. Demographic [Pol91]. Demography [DKW97]. Densities [Fre94, GR96a, GS90a, GR96b, Joh92, LK95, RV90c, SBS94, Len88, TK86]. Density [BF93a, BF90, CCG^{+}97a, CCG^{+}97b, Che94a, Cho93, CB90, Cre97, Efr97, Efr98, EL97, EW95, FJ90, Gou97, GMS97, HYP97, Ize91, JMS96, JS97a, Min98, OP96, PO94, Roe90, RW97, Rup97, SS96a, SF97, Ter90, VW95, WJ93, WY96, jW97, YM99, CH88, DH86, DM88c, FSS84, Har84, Har81, Har87, Jon89, LF82, OS87, ST87, SW83, TS85, WMR91, Won82b, Won83, Yak85]. Density-Based [BF93a]. Department [Hwa86]. Dependence [Goo94, HZ98, LR97b, QN98, RS94a, SV93, SJ92, BS84b, Gew82, Gew84, GJ84, HM88a, Har84, JH82, Joe89]. Dependent [AME96, Ada98, ALHH92, CLWZ95, FT94, FSE98, GV96, HDG97, HS94, LT96b, NP95, Pep91, SC97a, SC97b, SSK90, WSLP90, vR98, Agn85, Wei81, Zeh88, Sch83b]. Depends [Ye93]. Deposition [Haa90, Haa95, Oeh93]. Depressive [TFAH94]. Depth [CRS99, He99, Koe99, LS93, LS97, LS99, MS99, OH99, PM99, RH99, RHV99]. Derivation [GH95]. Derivative [HMT92]. Derivatives [jW97, Dwy67, HS89a, McC82]. derived [ELR87, Rot83, Sch81, Sch82]. Design [CM93, CLY92, CC96a, CM91b, CMMY91, EK93, Fan92, HT97a, HT97b, Min93, MP95, Pen97, Por97, RG95, STR97, Tom96, AP83, BHS81, BS85, BHT88, CSS89, IF82, O’G82, SHS86, Tam85]. Design-adaptive [Fan92]. design-based [BHT88]. Designed [FR90, Gav78, IL94, Kie78, Mal78, SM78a, SM78b, Sim78]. Designs [Akr90, Akr91a, AA94, AAB97, AB97, AC95c, BS91, BPS95, BDM97, CC96a, CW94c, DL92, DPWH99, HYP97, Jac93, JMS97, JR96, KW96b, Kus98, LB94a, MS92, PR93, SR96, TA90, Tho91b, Tom94, VK92, WZ97b, Wie98, CJ88, CMST88, CS89, HS89c, Hsu82, Jac89, KR84, KR88, Min87, SB81, Stu82, Vos86, Vos93, WG88, Wit88, Won84, Wu85, YBN85, dAJ86]. Destination [Var96]. destructively [dAJ86]. Detect [RMS90]. Detecting [DR98, KR82, KR88]. Detection [Atk94, BSPK98, BC90, FNP95, IT94, KM96a, LPT91, LZ98, Sho93, JG83, KP83, Whi81a]. Determination [Fin90, JMS97, SG95b, Noe87, SG95a]. Determine [BHS98, Est92]. Determining [Fer98, Roe94, Sch94b, VB97, CC74, Cre88, NP83]. Deterministic [Buc95a, CMMY91, RGZ95a, RGZ95b, Sch95b, TL95, Wol95, Nyb86]. Deterrent [Gro90]. Developed [PH90b]. Development [FMR62, KS96]. Developmental [BO95, SGC95]. Developments [Ize91, Tsu86]. Deviance [CJ96, Sim89]. Deviation [Jen91, RC93]. Deviation-Type [Jen91]. Deviations [SB90]. Device [Pet92]. Diagnosing [Joh98b, dP98]. Diagnosis [Bac90, XXP93]. Diagnostic [CPP96, Eas94, LTW96, Llo98, QH98, SS86, SL93b, Wei95, SC85]. Diagnostics [AC97b, BAP95, BF97, BVT96, CP91, CJ95a, Cho92, CCL90, CHW92, CC96b, FMSP90, FM92, Fun95, GS92, HLC92, LR95b, LR92, MSH90, PY99, SC92a, Soo92, SS91b, Tho91a, Vel95, Pol84]. dialing [Pot87]. dichotomies [SO83]. Dichotomous [For94, Isa74]. Dickey [VW95]. Dictionaries [Hau92]. Differ [LW91]. Difference [HH90, HHS97, Llo90, MS98b, Pes93, PH90a, Hay86, MS98a, WP85, Wor34]. Differences [YK96, MC84]. differencing [Sol84]. Different [SKF98]. Differential [RG94, RV90c]. Differentiation [AM92]. Diffusions [GT98]. digit [Pot87]. Digital [FMR62]. Diltiazem [OMF^{+}93]. Dilution [ZDL96]. Dimension [CL92a, Fer98, Li92, WR94, Li91]. Dimensional [Daw95, DDD94, GS98a, O’S90, Abr88, Dal82, NWGP84]. Dimensionality [Bre92, Sch94b]. Dimensionless [Cho93]. dimensions [Har87]. diminishing [BJM81]. Dipoles [RTF92]. Direct [ALJ99, HH96, BJM81]. directed [HL81, WW87a, Won87]. Directional [FHJW96, PML98, Pre96, FH89, FH90]. Directions [Coo98a, Coo98b, Li92, Li98b]. directly [Eri81]. Dirichlet [AG83, Esc94, GB98, MG97b]. Discharge [KR96]. Disclosure [BKP90, DDD94]. discontinuous [Gle85]. discovering [CF87]. Discovery [Bil97]. Discrete [BPS96, BRMZR97, BD97, CR92, FW96, FL95, HH96, Tel67, BS83b, Pic87]. Discretization [GJR98, Moj99]. Discretized [DR96, Jon89, O’S91]. discriminability [GC82]. Discriminant [BC96, BS97b, Fun95, HTB94, HM97, TH82, CM84a, Fat83, FH86, Fri89, LV88, Rav89, SU74]. Discrimination [BK85b, FBF97, KST98, KM91a, Li96, O’N94, PR93, Rig97]. Discussion [CRS99, Gel94, Koe99, LS99, MS99, OH99, PM99, BF85, DG93, Efr96, Fab91, Gew82, HR84, Hol86, Hub83, LL87, Li91, Mor83, RHS93, Sha82, TW87, VSK85, WMR91, WHM85]. Disease [LZ98, LTP94, SK93a, SH90, SHBF94, WCXG97, WHG98, YL97]. Diseases [BBH96]. Disorder [TFAH94]. Dispersion [GH95, LB96, Hab82, Mat83]. Displays [Fri94, Goo91]. Distance [CCB96, EL90, Har92, Yam90, Boo81, Dha93, ELR87, Rom88, Sim87b, TB86, WPSL84]. Distances [BPL97]. Distinguishability [SEH95]. distributed [CM85, Ger82, LaR86]. Distribution [BS93, BK99, BS98, CH97a, Chi74, Chr99, CT97, FS98a, Fin90, FP95, FK94, Gav78, Gel94, GRB^{+}95, Hal94, HWY99, Han99, HRW90, HL97, Kai94, Kie78, KS93, LKCH99a, LKCH99b, LMR96, LK92b, LRR91, Lie94, MZ96, Mal78, ME96, McW90, Mer94a, Mer94b, ML94, MYZ67, MSK96, Nau65, OV95, Owe95, Ryu93, SM78a, SM78b, SC99, SAK90, Sim78, SP91, Ste99, Wan95, WP85, YM85, YM86, AY82, Ben83, BK89, BS86b, CL85, CH83, Cro85, De 83b, Esc86, FE88, Fis79, Fli84, Geb71, Gla89, IHC84, IS83, KR84, KM96b, LLT89, Len88, MO85, MT81, MT82, OW82, Poo74, Ran89, Rin83a, Rin85, Riv86, RS82, Sam81, Sar87, SZJ87, SS88, Stu56, Tad80, Tei65, VJ88, Wal86, WL84, Zac86]. Distribution-Free [Kai94, McW90, WP85, YM85, YM86, Fli84, IHC84, KR84, OW82, Ran89, Rin83a, Rin85, Stu56, WL84]. Distributional [LK91, Zha92]. Distributions [ALHH92, BPS96, BM92, Buc95b, Che95, CF95, Cra94, DXP92, DJ98, EGGS97, GR81, GBJ96, Hay94, Hir92, JS94, KO95, KW96a, KR93, LA94, LM93, NS83, NS86, NCDF96, RS94a, Sat95, SV93, Ste90, VM96, WBB93, dHY93, AG83, AP89b, BM81, CKC83, Dye82, FOS95, GW84, GJ84, Gle85, Has69, Hil70, HMP87, HL81, HPP86, Jen86b, JTB80, JTB81, Kun80, MO85, MO88, MB82, Nau82, NWGP84, PT83, SW89, SM83, SS82a, SB87, Suz85, TC72, TW87, Tay85, WLW89, Wes87a, Zel87, vd87]. Disturbance [Sri72, Kob86]. Disturbance-Variance [Sri72]. Disturbances [Sri72]. divergence [Rea84]. Diversity [PT82]. DNA [Bal95, BGH97, DRR92, KM91b]. Do [CB94, CH94, Oak85, Tuk60]. Dogma [Moy99]. Doksum [BC82a]. Domain [HW92b, HZ90, Pap96, RTF92, SC92a, Ste95, Wan95, Tan83, ZH87]. domination [VM83]. Donald [Efr94]. Donoho [MY95]. Don’t [RSV95, SG83]. Dose [CNF95, HYP97, HB99, Lee96, OK96, PSVS92, RG94, CC83]. Dose- [HYP97]. Dose-Related [OK96]. Dose-Response [Lee96, PSVS92, CC83]. Double [Buo90, Dor94, Efr86a, NW97]. Double-Scan [NW97]. Doubly [AML95, JMV94]. Down [RCH94, Fin93]. Draft [Ang91]. Drinking [MSL93]. Driven [Efr98, KL97, KL99, Led94, PM90, Hur85b]. Driving [MSL93]. Drop [Lit95]. Drop-Out [Lit95]. Dropout [PHL98]. Dropouts [HMC^{+}93]. Drug [GJ94, YK96]. Drug-Use [YK96]. Dual [BM96, SR96, NM83a]. Duality [DL88]. Duration [FW96, Ryu94]. Durbin [SAK90]. Durbin-Watson [SAK90]. During [MSMP92]. Dynamic [ALJ99, BBGL97, Bol95, CMW97, Eas94, Fah92, FT94, Gri94, LC98a, LC98b, MFS84, Now92, OKS97, PSS99, SK91, SS91a, WHM85, Wes95, AH81]. Dynamical [Gri90]. Dynamics [Buc95a, RGZ95a, RGZ95b, Sch95b, TL95, Wol95].

Each [BD91a, PB98]. Early [BC90, Hur85a, LZ98, NGC^{+}95, SB90]. Earnings [RBD93]. Earthquakes [PS96]. Eaton [DH93]. ecological [BP86]. Economic [JS97b, HT87, Nef82]. ed [Tuk47b]. Edge [CGGM98a, CGGM98b, HW91, Joh98a, MT98, SHL98]. Editing [LS87]. Edition [Tuk47c]. Editor [Ano99d, Ano99e, LHA94, LHAL95]. Editorial [Ano95a, Ano97a, Ano98c]. Editors [Ano99c, CCG90, LHA96, LCK97, LWG93]. Edmund [Met50]. Educational [AK92, ZT97]. Edwards [Man94]. EEG [WPK99]. Effect [AK92, BW94, Gro90, HC96, MSMP92, MW98a, MW98b, OMF^{+}93, Raz90, RCH94, RW98, Ros93a, WPD98, YKB97, Yas95, BK85a, BK86, Ber82, Bin82, CWR88, Hod88, SHS86, WHH88]. Effective [GJ94, RSW95]. Effectiveness [Bak98, SG96b]. Effects [AL95, AGM98, AI95, AIR96, BP97, Bha81, BD91b, Car94, CLM93, CM91b, Cro95, FW96, GHCF94, HW90a, HW91, Har90a, Hec96, HDG97, Hol95, JTC99, LB94b, MM95, MH99, Mof96, MPCB95, RG96, Ros96, Sch91a, SLS97, ST99, SH90, VL96, WKY93, Wei92, WM91, Ye98a, ZDL96, ZWWB96, Bow87, CR89, CW85, Dav82, Eri81, Fat83, KH84, KR88, LB94c, LB94d, MS88a, Mis85, Rei82, Rei84, YD84, ZK91]. Efficacies [TA89]. Efficacy [QH98]. Efficiencies [Mac86, Mac87, TA90, Kob85, Stu56]. Efficiency [BL93, DB97, DP89, HHO88, Har84, JH82, LLR95, Neu98, RR95, Ruk93, STR97, Ski91, Tar87, Wel90b, Bin82, CD89, Gas85, Gro81, HD88, KD87, Rog88, Sha89a, Sim89, Wei81]. Efficient [Efr98, Efr90, ES97, KW82, Kus98, PS94a, Rob94, RK96, Sär96, SP91, Wu85, ZT97, vR98, CH93, JJW84, Sil84, SS87b, Spi82, Vos86, YZ88]. Effort [Ode91]. Eigenvalue [BC96]. Eisenhart [SM84b]. Elasticities [BG91b]. Elderly [CM91b]. Electoral [GK90]. Electorates [Gel94, Mer94a, Mer94b]. Electrical [RTF92]. Electricity [HK93, HK92]. Electron [QTC96]. Electronic [Lon67]. Elemental [Haw93]. Elementary [DG92]. elicitation [KW88]. eliminant [Pud82]. elimination [BB88, Mar88]. ELISA [GHMW94]. Elliptical [KBB93, SB87]. Elliptically [CN94]. Elongation [Gle93]. Emanuel [Tuk79a]. Emission [Efr99, MO98, Wei97, WEST91, VSK85]. Empirical [AC95a, ALM97, BD91b, CG90, CH83, CLWZ95, Efr96, FP95, Gel96, Geo96, GM86a, LL87, MS96, Mor96, NS93, Rag93, Rei96, Rup97, ST96, Sob93, SHBF94, Yan99, VS93, BK82, DL81, FH82, FP84, GL87, Las94, Lou84, Mor83, RC81]. Empirical-Bias [Rup97]. Employer [Kad90]. Employment [Yam92, GG87]. Endogenous [BJB95]. Endpoint [Pep91]. Endpoints [LTM92]. Ends [BHS98]. Energy [Hwa86]. Engineering [Tuk52]. Enhancing [Wal93]. Enlistment [Ang91]. Enriching [Wal93]. entered [BF83]. Entropy [DvdM81, GJP96, Joe89]. Entropy-based [DvdM81]. Entry [AK92]. Environmental [Niu96]. Epidemic [GJDT98, XXP93]. Epidemiologic [BSB98]. Epidemiology [Bre96]. Equal [Rig97, Ros93b, GG87]. equality [BM81, Kob86, Sin86, Wol87]. Equation [Pai97, Bin82, BF83, FP84]. Equations [BG91b, CRW98, HS98b, MBL98, MN93, O’N94, RV90c, Sri72, Kun80].

Equivalence [CC96a, CW94c, DM88c, GS94, Mun96, Wan97, BB88, Esc86, PP88]. Equivalent [PWM92]. Era [Ang91, Tei65]. Ericson [God82b]. Errata [LLT54, MS98a]. Erratic [Hal94]. Error [AS90a, BD97, Bre92, Buo90, Buo96, BS96c, CL92a, CKLS96, CS94, CNS90, Efr97, EKWS91, FS95b, GMP97, HJ92, HDG97, HTP95, JTC99, LR95a, Mad59, MW98a, MW98b, MT95a, MP96, MS91b, MS91c, O’N94, PR90b, PH90b, RW98, SR95, SB95, SC95, TDW95, WLGC98, WJ92, WW91b, AH81, Bha81, Efr83, Efr86b, HV86, Har85, Hay86, Kor81, Law81, Law82, Sam85a, Sch86a, Tar86, WK88, Wol86]. Errors [Alt90, BR94b, BH98, CGL93, CH97b, FS95a, Fel74, HOS95, HZ90, LS95, Lin85, Now92, PSVS92, RBD93, RCL96, Sch94a, SK91, SS90a, SS95b, Wan98b, Wie98, BK82, Bof85, CGG85, CR89, CSC86, Hwa86, KH84, Kob85, KD87, Sch86a, Tsa84]. Errors-in- [Now92]. Errors-in-Variables [CH97b, LS95, PSVS92, CGG85, Hwa86, Sch86a]. Erscheinungen [Fur46, Gei48]. Essay [Joh96a]. Establishing [DMLC97]. Estadistica [Tuk47a]. Estimate [MZ96, BN74, CC74, Dua83, Eri81, Fer82, GA82, Hur85b, Sta89a]. Estimated [FS98a, GL92, KK84b]. Estimates [BS97c, BH94, CH91, CW94a, Efr92, EKWS91, FKYZ99, Hab95, HT90, Hoa73, Hös94a, LT96a, MSMP92, MY91, MS91a, MS91b, MS91c, RN97, Ros93a, Sev96, SRC92, SR95, WC91, Atw84, BY86, GL83, Hec88, Jon89, KR86, Kot89, Lea84, MZ89, MS88b, MS95, Rob87, SM84b, Sta89b, TS85, TT84, Wes87b, Wes87a, YZ88]. Estimating [Bac90, Bak98, Bal95, BR83, BB89, BBH96, BP94b, BD81b, BBC98, BPN87, BS86b, BF85, BF90, BF93b, BR98c, CRW98, Cha88b, CL92b, CG92, DXP92, DA96, Efr83, Esc94, Est92, FG91, Fre94, GG87, GK90, GKB98, GR92a, GT91, Haa90, HS98a, HS98b, HWY99, HPS98, HW90b, HRW90, HW95b, HOS94b, JR85, Joh92, JK93, KS93, Lel94, LR97c, LW90, Lou84, MEGN92, Mil92, MPCB95, Mui85, NNR95, NCDF96, O’N94, Pag90, Pai97, Pud82, RZ98, Raz90, RS90, RCH94, Riv98, RS93, RS82, Sch99, SMP95, ST82, Spr83, Sti87, Sto88, Tuk85, Var96, WPD98, WS89, XXP93, Yan97, YL97, ZDL96, Zac86, CWR88, CC83, MPS88, RC74, Sha85, SS88, TC72, WP85, dAJ86]. Estimation [AC97a, AR90, Akr91b, Alb97, AF97, Amr95, ASL96, AH81, AK86, AGM98, AI95, AGM94, AS90b, ALM97, BRR92, BM85, BHB96, BVT96, BJ74, BM95, BP94b, Bin96, BG91b, BJB95, BM70, BS96c, CD91, CD95, CK92, CS90a, CHR94, CB94, CDW93, CL95, Che94a, CS96a, Che94b, CLY92, Cho93, CLM93, CB90, CKC83, CS94, CCB96, DD88, DW87, DD90, Dor94, Efr97, Efr98, EL97, Elt96, EW95, Fah92, FJ90, Fay96b, FHTW97, Fon92, For94, FR90, Gho92, GNK96, GNSC98, GMPS99, GH95, God82b, GV96, GC95, GM94, GS93, GMCM96, Gou97, Gri94, GL96, Gür96, HMT92, Har87, HJ92, Has69, He90, Hoo93, HMC^{+}93, HH96, HL96, HR97, HM91, HH91, Ize91, JMV94, JMS96, Jud96, KP91, KT95, KPR91, KP90, KA86, KA91, KS97]. Estimation [KS94, LR95a, LK97, LT97, LS95, Lee96, Leh83, Li95, LCG91, LM93, Mau95, McC94, MSH93, MV91, MN93, Mei94, Min98, Mis85, MP96, MS94, Muk96, MRvdV97, NSS97, Neu98, O’S93, OP96, OR96, OKS97, PS94a, PO94, PC96, PSS99, Pol91, Poo74, PR90b, Rag93, RS96, Rao96, RTF92, Rei84, Ric97, RH96, RRZ94, Roe90, RW97, Rol94, RK96, Rub96b, RV98, Rup97, SS96a, SR94, SW89, SG92, SS94, SS99, She93, SC94, SY84, Sit97, Ski91, SR96, SS96c, Sri72, SRSD94, SB95, SW98a, Sto86, SW91a, SI90, Tan83, TFP90, Ter94, Ter90, Val93, Vel95, VFR97, WJ93, Wan91, WR95, WWZtO97, Wan98a, WC99, WJ92, WY96, WM90, WR94, jW97, YM99, YR95, Zie90, de 81, de 83a, vR98, AG83, AP87, AP89a, Bet89, BS83b]. estimation [BHT88, Cha86, CT98, Che70, CF87, DC87, DB83, DM88c, DL74, Dyk82, DL88, FD85, Fin83, FSS84, Fry84, FH78, FH81c, GM86a, GL87, GSS89, HM88b, HM89, HW86, Isa83, IS83, JSG88, KP87, KW82, Kre86, KSC89, Lee88, Mat83, MV85, NS93, NWGP84, OS87, Par93a, Pri74, Roy86, RS86, Ruk86, RC80, RC82, SH81, Sam93a, SM83, Sch84a, ST87, Sen74, SW80, Sha83, SW87, Sim87b, SW83, SS82b, Smo82, Spi82, Spl83, SG82, Sta94, Ste87, SZJ87, Suz85, Tak81, Tam88a, TB86, TH87, Val85, VW87, WMR91, WG83, Wij85, Won83, WPSL84, Yak85, Zel86, vd87]. estimations [MSW^{+}83]. Estimator [AML95, CKLS96, Cha91, Dor94, ES97, Fel74, GS91a, GS91b, GR92b, LK92a, LR97c, Sev91, Sev96, Sit97, SM92, SO94, Tab90, WZ97a, YF90, dHY93, Akr86, Boo82, Cha84, CHL82, CH93, Cro85, DW87, DH86, FM83, GW86, Har84, JJW84, Kar81, MY95, MG84, Pad86, RC81, Rup85, YF95, Yan85]. Estimators [AM92, Ber91a, BPR99, BGH93, CZ77, DS92, DS91, Eas91, FHMP96, FS95b, GS91b, HW91, Hal94, Har92, Jia98, JDP94, JS97a, KBB93, KCHH93, KAT91, LR95a, Lin91, MKSB90, PR90b, Riz92, Ryu94, SW95, Sär96, Sas93, SM90a, SO94, SS90b, Thé99, Tin65, WR94, Zha99, VS93, BLE88, Bha85, Bie83, Boo81, BT81, CL85, Cas85, CRH94, DWL87, DB83, Dha93, Esc86, Geo86, Gol83a, GHP87, HRR81, HV86, Hoe84, HT81, KH84, Kob85, Kun80, Law81, Law82, Loh86, Moo81, Nai84, O’H87, PMF88, RC81, Ruk87, SM84a, SL87, Sha89a, SW83, Spo82, Tsu84, Tsu86, Val87a, Wol84]. Ethnic [YK96]. Ethnicity [Joh90]. Euclidean [Vil90]. Evaluating [BC90, BSB98, MSL93, QH98, SG96b]. Evaluation [CS92b, CS90b, FM82, HZS^{+}90, Joh90, LB94a, Luc93b, SAK90, SS95b, SHBF94, Ter94, WPK99, WC91, GC82]. Event [DB97, HL96, YK96, KP83]. Events [DR96, GKB98, HMC^{+}93, JTC99, NW97, Pep91, SH92, Sol95]. Evidence [BGH97, CH94, GCW96, LS91b, BS87, Bor87, CB87, KM83, Dia78]. Evolution [PS96]. Exact [AMP90, Agr98, BOP92, BB98, Buo90, CS89, Gle85, Gui88, Gui93, Hir92, KA95, Koo97, MS92, Mau95, MPS98, SMP95, ST92, SI85, Spu99, SK90, SW98b, SW98c, Uus83, Uus85, VHE91, WS96, HMP87, Hjo94, Mar84, MP83, MPG85, MPG86, MPS88, PH81, Wol87]. exact-approximation [Mar84]. Examination [BHB96, CW91]. Examinations [LZ98]. Example [Sta91]. Examples [GFT^{+}93, ZT97, Sim89]. Excess [FP95, MS91a, MS95]. Exchangeable [BO95, Gol86, OK96, QN98]. Exemplified [Gey91, Roe90]. Exercise [Hod94]. Exhibitions [CH94]. exist [Oak85]. Existence [LK92a, KR86]. exogenous [Spl83]. expansions [Kun80]. Expect [BD91a]. Expectations [DM97, Erk94, Joe95, WK92, TKK89]. Expected [CP91, Est92, HB86, Hab88]. Expenditure [SM90b, Zie90]. Experiment [FR90]. Experimental [AP83, EGMP93, EK93, Hoo89, Min93, PR93, Tom96, Tuk49b, dAJ86, Tuk54]. Experimentation [TW46b]. Experiments [CM93, CV95, CMMY91, DL92, Gav78, HW95a, HPS98, HH91, IL94, Kie78, Mal78, MP95, Roc92, SM78a, SM78b, SGC95, Sim78, ST92, SN90, Tom94, Ye98b, BS85, FB83, HM85, MV88, MC84]. Expert [CS95, GMD95, Ueb93, LS86]. Experts [PJT96, Agn85]. Explained [DBB^{+}94]. Explanatory [BJB95, EF91, Jon96]. Exploration [WHGA96]. Exploratory [Fri87, Sol91, Sol95, Gil86]. Exploring [Alb94, HZ98, Sam93a]. Explosive [Sol91]. Exponent [MEGN92]. Exponential [Bir54, CV92, CCL90, GW88, GPS95, Hay94, KT94b, LCG91, LM93, PSS93, Ryu93, Sla94, dCP99b, BLE88, BK85a, BK86, BM81, BB74, CKC83, Efr86a, FMD82, Has69, HL81, KT94c, MB82, MS85b, RS82, Sar87, SW80, Tad80, Tak81, TKK89, Zac86]. Exponentiality [dCP99a]. Exponentially [RP96, Ger82]. Exponentials [CB90, HRW90]. Exposure [BRR92, BL93, CCG^{+}97a, CCG^{+}97b, Cre97, GMS97, SK93a, SF97, SCCF93]. Exposure-Disease [SK93a]. Extended [Fah92, HS98b, Yan97, TT84]. extending [Hur85b]. Extension [LW90, Ryu93, Zuc98, BB74, SZJ87]. Extensions [Thé99]. Extra [NM98]. Extraction [CCD97, KA91, Mar86, WA83]. Extrapolation [CS94, DH91, Lav91b, SC95, Tuk52, BY88]. Extreme [HW92b, OV95, Zel93, Flu85].

F [DH83, Hoa73, Tuk49a, RSY93]. Fabrication [HOLDG98]. Face [SJ90]. Face-to-Face [SJ90]. Faces [PS94b, CM84b]. Facilitating [RT92]. Facilities [KR96]. Factor [Akr90, BP96, BD92, DL92, DH99, GL92, MS96, FB83]. Factorial [AA94, AAB97, AB97, BDM97, Tom94, Tuk47b, Tuk47c, CJ88, SB81, Vos86, Vos93]. Factors [BSB98, BM99, BPR95, DKRW97, GF96b, Hsi97, HTP95, KR95, LR97c, Pet92, VW95, Vid98, Del89, SCM91]. Failure [AMCS98, CKLH92, GFT^{+}93, GS95, HR97, LWY93, LW92, SDW98, SEH95, SL98a, Yam92, Gla80, Sch88, SW83, WLW89]. failure-rate [SW83]. Failure-Time [Yam92, HR97, Sch88]. Failures [The93, Suz85]. Fairness [SKF98]. False [BR95]. False-Match [BR95]. Families [CV92, GNK96, GPS95, KT94b, MO88, Sla94, BLE88, EH87, Efr86a, KT94c, Loh85, SW80]. Family [AGM98, CW94b, CCL90, GW88, Gle93, LK95, MSMP92, PSS93, WHG98, CKC83, CC83, Has69, HL81, JTB80, JTB81, MO85]. familywise [Hay86]. Far [HHM92, HHM88]. Fast [Atk94, FR99, GKK91, LB93, Luc93b, PY99, SAK90, VHE91, EH87, Sil84, Spi82, Spl83]. fat [FS98b]. Fatigue [CH95a]. fault [LS86]. faults [JR85]. favorable [Eth82]. FDG [OO97]. Feasible [MS94, SM92]. Features [AF97, BR98c, DR98, VY80, HVY83]. Fecundability [HW90b]. Federal [HT90]. Feedback [RH96, Gew82, Gew84]. Feedforward [Whi89, Fab94, Whi92]. Fertility [SA92, YP92, ZWWB96]. Few [HA94, HA95]. Fields [DKS97, KO96, HW94, Pic87]. Fienberg [HL81]. Filter [GM94, SW91a, Zeh88]. Filtered [DJ98]. Filtered-Variate [DJ98]. Filtering [Fah92, Koo97, Li96, NNR95, PS99]. Filters [PS99, Wat87]. finding [PH81]. Finetti [BM92]. Finetti-type [BM92]. Fingerprints [DRR92]. Finite [ALM97, BBH94, CDW93, Cra94, CCB96, Fel74, Fox74, HS98a, Lou96, MV91, SW95, SG96a, WS96, Wri83, AK86, Car82, Cha86, DD81, DD82, GM86a, Has69, Lin85, MS85a, Mey87, NS93, RC81, Smo82, Tam88a, Val85]. finite-population [RC81]. Finite-Sample [SW95, SG96a]. Finitely [KO95]. Firm [HT97c]. First [DM95, HW90b, Sch96, Vos86, BHS81, HM88a]. First-Order [DM95, Sch96, Vos86, BHS81, HM88a]. Fish [HAS97, Pol91, SS82b]. Fisher [Goo96, Hay86, JW88, Leh93, MP83, Par96]. Fishing [Mil92]. Fit [AP99, BOP92, BPL97, CW94b, Chr91, ES90b, Eub97, HP92, Hua97, KL96, LaR91, Lan96, Led94, Lin91, MNS99, OR92, Pap96, Peñ98, SW91b, Wel90b, Wil50, Akr88, Boo81, CM85, DKMO89, Gle85, KOS85, Koe86, Mas51, O’N84, Ole85, Rea84, Sim86, TR87, Wei84, Zel87]. Fits [CHW92]. fitted [TV85]. Fitting [CH97a, HYP97, Mad59, MS96, MP95, OR98, Sch96, WCR94, But84, CR84, Mül87]. Fixed [Bre92, Qui98a, Ste95, KH84, MS88a, Wor74]. Fixed-Domain [Ste95]. Flat [LMR96]. Flexible [GKK91, Gra92, HTB94]. Floe [BR92]. Flow [SR95]. Fluctuations [PM97]. Follow [MZ94, Suz85]. Follow-Up [MZ94]. follow-ups [Suz85]. Following [FR90, CS87]. follows [Krä86]. Food [BD97, PH90b]. Football [CH94, GS98b]. Force [SG96b, SR95]. Forecast [CNS90, SS90a, Krz87]. Forecasting [CMW97, DR91, HK93, KL90, LCMA92, MZTT98, TB85, WHM85]. Forecasts [AS90a, LT94b, SMD93, Sti87]. Foregone [KSS96]. Forensic [Bal95, DRR92]. Forensics [Gei92]. Forest [RC94]. Form [CS90b, JMV94, Hil74, Sch87a]. Formal [KW96a]. Forms [JS94, Lie94, HVY83, HW75]. Formulation [Soo92]. Forsythe [Rot83]. Four [GNK96, HZS^{+}90, Uus85]. Four-Person [GNK96]. Fourier [FM81, LK91, O’S91, Sto91]. Fraction [Yam92]. fractional [WA67]. Fractionally [LL97b, PH90a]. Fractionated [HW95a]. Fractions [KR93, SS99]. frailties [Oak89]. Frailty [LK97, Sas97]. Frame [FG91, Ski91, SR96]. Frank [HL81]. Free [Kai94, McW90, RMS90, CJ88, Fli84, IHC84, KR84, Lea81, OW82, Ran89, Rin83a, Rin85, Stu56, WL84, WP85, YM85, YM86, YBN85]. Freedom [OR92, PWM92]. Freeze [SC96a]. Frequencies [GC95, SHBF94, HB86]. Frequency [Coo91, HZ90, Pap96, RTF92, SC92a, Sco85, Hab88, Hil74, IS87, KK84b, Tan83, ZH87]. frequency-domain [ZH87]. Frequentist [CS91, HH91, SR94, CB87]. frequentistic [Fin83]. Frequently [GS93]. Freshman [MH94b]. Friedman [Wit88, JH82, LV88]. Friedman-type [Wit88]. Friends [Yam90]. Frontier [GMPS99, PS94a]. Full [MS96, SC96a]. Full-Information [MS96, SC96a]. Fully [AA94, OR98, Qui98b, TKK89, Wei97]. Function [Bjø96, BBC98, BD93, Cha91, CZ77, CT97, FHTW97, GS93, HWY99, HMC^{+}93, KY99, LaR91, Owe95, Pap96, Raz90, SW91b, WC99, YL97, BK82, BK89, Bie83, BS86b, Cla87a, DC87, FW88, GA82, KY86, LF82, LP86, Mui85, Pad86, Poo74, SW83, Sta89a, TT84, Yan85, de 81, de 83a]. Function-Indexed [KY99]. Functional [CT93, GRB^{+}95, HHS97, Kun80, Sam93a]. Functional-Coefficient [CT93]. Functionals [Che94b, LC95a]. Functioning [CM91b]. Functions [AKM94, BPS96, BP94b, BS97b, BK99, BS98, CW94b, CM98, Chr99, CV92, CMMY91, DS91, FW96, FHW95, FS98a, Fre94, GMPS99, GH95, Han99, HMY97, Hou95, HL96, KS93, LKCH99a, LKCH99b, Lel94, LTW96, LW90, LWT94, MS94, Muk96, NNR95, OB93, Rue95, SC99, Sla94, Ste99, SW97, SG97, WK94, Yan97, Yan99, AP87, BK85b, De 83b, Dic83, Dyk82, FD85, Khu81, KSY85, Nai87a, OYR86, PS82, SU74, SS88, TKK89, Val87b, Zel86]. Furlan [Gei48]. Further [Tsu86]. Future [DXP92, Pol91, Hog74].

G [Sta94, Tuk47b]. Galaxies [Roe90]. Games [EGMP93]. Gamma [EP94, GR81, SL80]. Gamma-Ray [EP94]. Gauss [HM85, Kar85, KP81a]. Gaussian [MT82, BC96, Ber91a, CMW97, DKS97, FB83, Gu90, IS83, Kit87, MT81, SW91a]. Gaussians [MZ96]. Gear [Mil92]. Geisser [BH88]. Gender [YK96]. Gene [LR96, Liu94]. General [Ano98c, Ano98d, BL92, Buo96, DDD94, ER86, GR96a, Gol83a, Goo96, GW91, HJ92, HR94, KP81a, LYY99, Luc93b, MY91, MLRR92, SC94, Vel98, Wes97b, Fin83, Gui83, HVY83, KSC89, MB88b, Pol84, Raf87, Ric82, Ste81, SC85, Tsi82, Wes98]. Generalizable [Soo92]. Generalization [Bjø96, EM90, Mee90, MSK96, VW95, Zeh88]. Generalizations [Spu88, WM91, Pot87]. Generalized [BCJ96, BH98, BC93, BS96c, CFGW97, CH97a, CCD98, CRH94, DG99, ES97, Fah92, FHW95, FM92, GS99, GNSC98, GN82, HS98b, HMM98, HLC92, HN92, Ibr90, JGC95, Jia98, LR95b, LB96, LM93, MO98, McC97, MP96, MG97b, Neu98, Pai97, RSY93, RS94a, SMF94, Stu88, SW91b, Tho90a, TW89, Val87b, WKY93, WLGC98, Wee91, Wee93, Wee94, YF90, ZK91, Alb88b, Bie83, CR82, DD88, DWL87, Dye82, Efr86a, FHO87, IL91, KSC89, Law81, Law82, Lea84, LV88, OYR86, PP88, PS86, Pud82, Rei84, WHM85]. generally [Hay89]. Generated [LJK93, MO85]. Generating [BPS96, HVY83, Mar84, Rel72, SL80]. Generation [AP99, KR76, KP81b, KP82, SS80]. generator [Che71]. Generators [FM82, Col87]. Genest [BH88]. Genetic [HZS^{+}90]. Geocode [GBN96]. Geographic [SS96c]. Geographically [MST97]. Geometric [Cha96, Riv98, GH87b, HM83]. Geometrical [HW91]. Geometry [CS93, Daw95]. George [PT82]. Georgia [RC94]. German [Fur46, Gei48]. Giant [Met50]. Gibbs [KT94c, Cha93, Chi95, GHRPS90, GSTM92, GS99, HC96, KT94b, Liu94, LNL95, RT92, ZK91, ZM95]. Gini [Sen86]. Give [Sat95]. Glass [Bar91]. Global [Lea84, WKL95]. Globally [FKYZ99]. GM [SRC92]. go [Tuk60]. Godfrey [Tuk47c]. Good [GS97, PT82]. Goodman [GR81]. Goodness [AP99, BOP92, BPL97, CW94b, ES90b, Eub97, HP92, Hua97, KL96, Koe86, LaR91, Lan96, LC93, Lin91, OR92, Peñ98, Wel90b, Wil50, Zel87, Akr88, Boo81, CM85, Gle85, KOS85, Mas51, O’N84, Ole85, Rea84, Sim86, TR87, Wei84]. Goodness-of-Fit [AP99, BOP92, BPL97, HP92, Hua97, LaR91, Lan96, Lin91, OR92, Peñ98, Wel90b, Koe86, Zel87, Akr88, Gle85, KOS85, O’N84, Ole85, Rea84, Sim86, TR87]. Gradient [GN82, JJ93]. Grading [Joh96a, NGH^{+}95]. Graduation [MSMP92, MH94b]. Granularity [HD88]. Graphical [Eas94, GC91, Lau92, MR94, Roe94, Cre88, Goo91]. Graphics [Coo96, CW97]. graphs [FH82, FS86, HL81, WW87a, Won87]. Gray [GC95]. Griddy [RT92]. Griddy-Gibbs [RT92]. Gross [SR95]. Ground [Niu96]. Ground-Level [Niu96]. Group [BC90, HOS94b, JT97, LB94a, PSS99, Rom90, Ryu94, STR97, HPE81, Mis85]. Group- [STR97]. Group-Sequential [JT97]. Group-Testing [HOS94b, HPE81]. Grouped [HLC92, VFR97, Ala84, SG82]. Grouping [Gil86, Vin69]. Groups [GF96a, Gra95, PW96, SRR95, Flu84, Krz79, Krz81, TV85]. Growth [AC95a, CH92, Abr88, BPN87, Lee88, Rei82]. Guatemala [GR92a]. Guide [Gol63]. Guided [ES90a]. Guidelines [HTP95]. Gumbel [Wan95].

H [Fab94, Hoa73, LV88, Tuk47b, Tuk47c]. Haberman [HL81]. Haenszel [WW87b, YF95]. Haenzel [YF90]. Half [RS96]. Half-Sample [RS96]. Hampel [Hoa73]. Handcock [HW94]. Handling [Kot94, RSV95]. Harmoniegesetz [Fur46, Gei48]. Harmony [Fur46, Gei48]. Harris [CCG^{+}97a, CCG^{+}97b, Cre97, GMS97, SF97]. Hartley [Riv86]. Hausman [Thu85]. Having [CV92, Agr83, Clo82]. Hazard [AME96, AKM94, BAP95, FW96, GS93, GMCM96, KST95, KS97, Peñ98, Ryu94, SW97, Yan99, SM83]. Hazard-Based [BAP95]. Hazards [FG99, GR92a, HN92, Peñ98, SDW98, Sch99, LW89, Wei84]. Head [Wei92]. Headache [DPWH99]. Health [CM91b, DG99, GWG98, MSML97]. Heaping [HR90]. Heart [SHBF94]. Hellinger [CCB96, Sim87b, Sim89, TB86]. Helmet [Wei92]. HELP [DH99]. Hematopoiesis [NGC^{+}95]. Hemisphere [Sol91]. Heptathlon [DAO94]. here [Tuk60]. Hessian [Coo98a, Coo98b, Li92, Li98b]. Heterogeneity [GS93, KR96, OK96, VL96, KMT99]. Heterogeneous [DPWH99, PSS99]. Heteroscedastic [AC95c, DM96, FS95a, Hoo93, MS98b, WCR94, Wie98, CR82, FH86, Kar81, MS98a, Tsa87]. Heteroscedasticity [LL97b, Roc92]. Heterosexual [GJDT98]. Hidden [Whi89, Fab94, Whi92]. hidden-layer [Fab94, Whi92]. Hierarchical [AC97b, BZ99, CS96a, CM97c, CNWW98, DG99, HK92, HC96, Joh92, Kor81, LCG^{+}92, MR97a, MR97b, PJT96, PS94b, Sil97, Sta91, WCXG97, Wes97a, WM91, Alb88b, Goe83, HPE81]. Hierarchy [SKF98]. High [CMNS99, GS98a, HM97, Hös94a, LT94b, MSMP92, MH94a, RG94, Str93, WR94, YZ88, Abr88, CH93, Har81, Won82b]. High- [RG94]. High-Breakdown [CMNS99, HM97]. high-density [Har81, Won82b]. high-dimensional [Abr88]. Higher [JS97a, Min98, Mar81]. Higher-Order [JS97a, Min98]. Highly [LWY93, Sam90, Ger82]. Hildreth [Cro85]. Hildreth-Houck [Cro85]. Hinkley [Bue82]. Hispanic [Joh90]. Histogram [DJ98, OP96]. Histograms [SC96b]. Historic [EV92]. Historical [GT91, IRC98, VB97, HY86]. history [Sti86, Tei65]. Hitting [Alb94, AASM93]. HIV [DLD91, DA96, GJ94, GHMW94, LI98a, LCG^{+}92, SJ92]. HIV-1 [GHMW94]. HMOs [GLB98]. Hoaglin [Sim87a]. Hodges [Boo82, Ros93a]. Hodges-Lehmann [Boo82, Ros93a]. holding [FE88]. Homicide [Gro90]. Homogeneity [CF95, JGC95, JB92, MM93, Mit91, NC99, Qui98a, Shi91, YF90, ZC88]. Homophily [Yam90]. homoscedasticity [RA89]. Horizon [SC96a, Fox74]. Horvitz [SO94]. Hospital [SRR95]. Hospitalization [LI98a]. Hot [JK93]. Houck [Cro85]. Hourly [HK93, RBD93]. Hours [RBD93]. Household [AL94, SA92]. Housing [EKT89]. HRS [JS97b]. Huber [Hoa73]. Hultquist [SB81]. Human [RTF92, Tuk47b, Tuk47c]. humans [DH83]. Hybrid [LW97, WZ97a, FMD82, Won82b]. Hybridization [ST92]. Hypercube [Owe94]. Hypercubes [Tan93, Ye98b]. Hyperdimensional [Weg90]. hypergeometric [Dic83, Wal86]. Hypotheses [AA94, AAB97, BPR95, Bre90, GCW96, Hem79, KW95, LK91, LTM92, Leh93, MBR98, Rob94, Ruk93, Sil92, BS84a, Ber89, Del89, GV86, HT81, Loh85, Rog88, SS82a, TW89, WR84]. Hypothesis [AME96, BM99, BS97c, CGM83, GK93, HRW90, KL97, KK84b, SSC94, SD85, SC97a, SC97b, Tom96, Tro95, Wel90b, BS87, Bor87]. Hypothesis-Testing [Wel90b].

I. [PT82, TV85]. Iain [HHM88]. Ice [BR92]. identically [CM85]. Identifiability [GS99, Now92, Pru93, Hod88]. Identification [AIR96, Bal95, BR92, BWC98, HS93, Hec96, Man88, Mof96, RG96, RW96, Ros96, TA94a, DG93, Sto86, TA94b]. Identified [LTM92]. Identifying [HYP97, LT98, PB87, Won82b]. Identities [Hou95]. IGF [HDG97]. IGF-I [HDG97]. Iglewicz [Sim87a]. ignorable [RRS98, RS86]. ignored [KMT99]. II [BHS81]. iid [BK82, SM92]. Illustrated [DH91, Lav91b]. Illustration [GHRPS90]. Image [AGP91, BO97, CGGM98a, CGGM98b, FR99, GC95, Joh98a, Kol96, MT98, O’S90, O’S93, QTC96, SHL98]. Images [BR92, CHR94, FS94, Joh94, Pre96]. Imaginary [Pet92]. Imaging [Rue95]. Imbedding [Lou96]. Immigration [GMWZ99]. Immunoassay [ZD97]. Immunologic [Hei91]. Impact [BHB96, BSB98, Kad90]. Imperfect [BP97, BW94, PHS94, QH98, WS89]. Implementation [MNS99, WT90b].

Implication [STR97]. Implications [BAP95, Cla96]. implicit [FJV89]. implied [HHO88]. Importance [Che94a, GR96b, Gle92, LCW98, MPS88, OB93, Per97, Zha96]. Importance-Weighted [Che94a]. impossible [Gut82]. Improper [GS99, HC96]. Improved [CS92a, DH93, FHJW96, GSS89, KA95, Loh86, Ruk86, Zha99]. Improvement [NGH^{+}95, Efr83]. Improvements [ET97, YB97]. Improving [BS97a, CJ95b, JS97b, Pea91, WC91]. Impulse [SG97]. Imputation [Bin98, Bre98, CRS^{+}91, GKL98a, GKL98b, HR90, HL94, Jud98, Kad98, Rub96a, San98, SCC98, SS99, Efr94, LS87, RS86]. Imputations [KLW94, LC95b]. Imputed [Bin96, Elt96, Fay96a, Fay96b, Jud96, LRR91, Rao96, Rub96b, SDW98, SS96b]. imputing [CS86]. inadequate [Leh83]. Inadmissibility [Moo81]. Incidence [AKM94, BBH96, DXP92, Har90b, MST97, MC84]. Including [LCG91, MLRR92]. Income [DDV99, DM97, GNK96]. Incomplete [AC95b, Car94, FLZ96, Geb71, Ibr90, KW96b, LY93, Lit93, Liu96, CS89, MJ82, Wei83, WL84, WLW89]. inconsistent [Fer82]. Incorporating [CS95, HY86, JT97]. incorrect [Mat83]. Increase [RZ98]. Increasing [GS95]. Incubation [Bac90, DA96, HMC^{+}93]. Independence [BBGL97, CB97, GR97a, GR97b, Goo94, KL99, OK96, Coh81, CGM83, Gro81, GSS89, KY86, Kru88, Sar89]. Independent [AC97b, GF96a, JK96, LT96b, ZM93, Krä86, SS85]. independently [HVY83]. Index [Ano97e, BVT96, CFGW97, DR91, DMLC97, GRB^{+}95, HH96, LS93]. Index-Mortality [DMLC97]. Indexed [KY99]. indexes [Sen86]. Indicator [Jon96]. Indicators [EH91]. Indices [CG92, FH82]. Indirect [OP96]. Individual [GBN96]. Inducão [Tuk47a]. induced [Oak89]. Industrial [TW46a, TW46b]. Industry [CRS^{+}91]. Inequalities [Fin90, Hou95, Ber89, GJ84, Hoe63]. Inequality [Ede90, ME96, VW87, Wij85, Wol87]. Inequality-constrained [VW87]. Infarction [OMF^{+}93]. Infection [Bac90, DLD91, DXP92, LCG^{+}92]. Infectivity [SJ92]. Inference [AMP90, Agr98, Ahn97, AG96, AY82, BM96, BB98, BC93, Buc95a, Buc95b, BA83, CS91, CS95, Cha90, CM93, CS93, CSS89, Coh81, DF98, DM95, DRR92, DKR95, EW95, FIB90, FOS97, Fra91, GHRPS90, GT95, GM96, HR90, Hin71, Jay79, KBB93, KA95, KT94b, LWW91, LR96, LWT94, MSML97, MH82, McC98, MPS98, NCC96, O’H90, PJT96, Pep91, PSS93, Pic87, PWM92, Qui98b, Raf87, RGZ95a, RGZ95b, RW85, RT96a, SDW98, Sch95b, Sev94, Sev95, SN90, SW98b, SW98c, TW98a, TW98b, TL95, Var98, Vil90, WH90, WSLP90, WS96, Wes95, Wol95, YP99, BS84b, Cha88a, CS87, FOS95, FM81, FJV89, FH82, GR93, Gui93, Hol86, Hoo89, Hsu82, Jen86a, KT94c, LHT89, LW89, MS85a, MPS88, Mor83, RW88, Sch83a, Swe84, Tuk49b]. Inferences [CH97b, Gus96, LQT97, MT95b, RK90, SK93a, SRC92, HM83]. Infinite [Cha90]. Infinitesimal [WZ97b]. Influence [BF97, BL92, CP91, DWL87, DS91, GMWZ99, HR94, LA93, LKC97, MH94a, Ric97, Tab90, Tho90a, Tho91a, Wei95, CH93, FJV89, Hub83, KW82, KSC89, MH90, Rup85, Sam85a, dJdWW88]. Influential [Coo79, JG83]. Information [CB94, Cla96, DPWH99, Goe83, GJP96, JT97, MS96, Par96, SC96a, SK93a, Sev95, Soo94, SEH95, VK92, WPD98, WC92, CS89, Isa83, Leh83, Sch86a, SG83, Smo82, Wri83]. Informative [FS94, GBJ96, PHL98, VS93]. Ingenieure [TW52]. Inhomogeneities [KM91b]. Initial [Koo97]. Injuries [Wei92]. Inland [KSJ94]. inner [Mey87]. Innovations [DM95]. Insights [GS93]. inspection [Lin85]. Instrumental [AK92, AIR96, BJB95, BS96c, Hec96, Mof96, RG96, Ros96, SB95, BT81, DL74]. Instruments [BJB95, ZT97]. Insurance [GWG98]. Intake [NCDF96]. Intangible [Soo94]. Integer [Vin69, Lin85, LR87]. integral [Geb71]. Integrals [CS90b, Wij85]. Integrated [Cha90, LL97b, SL93a, PP88, Tar86]. Integration [MS98b, MG97a, OB93, MS98a]. Intensities [Var96]. Intensity [AI95, CZ77, CHP96, DM88c]. Intent [Roc95]. Intent-to-Treat [Roc95]. Intentions [Man90]. Interaction [MM95, BN74, Boi86, Dal82, MJ82, TA89]. Interactional [SJ90]. interactions [Fab91]. Interblock [TA90, CS89]. intercepted [Var88]. Interchangeability [ES99, Wei83]. Interdependence [VB97, Fur46, Gei48]. Interdependenz [Fur46, Gei48]. Interdirection [JR96]. interdirections [Ran89]. Interest [GT98, Ye93]. Intermittently [FSE98]. Interpoint [BPL97]. interpolated [Jon89]. Interpolation [GM94, HT97a, HT97b, HOS95, Por97, Tuk52, KA86, Sch88, dJ89]. Interpretation [Coo94, Gar94, MSH93, GH87b, HM83, Leh81]. interpretations [Dic83]. Interpreting [SCM91]. interpretive [GKR88]. intersection [BS86a]. Interval [HR97, LH91, MV91, Mun96, Pes93, SDW98, Del89, FMD82, MPG85, MPG86, RS86]. Intervals [BS96a, BHS98, BS96b, Bur94, CG90, Che90, CJ95a, CJ95b, CJ96, CH95b, DH99, Gri92, Har90a, HB99, HD97, Llo90, Mar90, Mee90, Mur95, Sat95, SG95b, TS90, TF90, ZD97, Atw84, BS83a, DT87, Efr87, FS87, Gui83, Joh78, Khu81, LL87, LT88, Mal82, MB82, Mey87, Mor86, Nyc88, PFV89, Ric82, Sch85, SG95a, SI85, Sti85, Uus83, Uus85, Wee93, Wee94]. Intervention [MSL93]. Interview [MSML97]. interviewer [BS85]. Interviews [SJ90]. Intrablock [BS91]. Intrasubject [Wan97]. Intrinsic [BP96, MBR98]. Invariance [Rom90]. Invariant [HMO97a, HMO97b, Rol94, SAK90, Tak81, dCP99b, HV86, Moo81]. Inventory [LLT54]. Inverse [Fer98, Sch94b, BK84, FB83, IS83, Li91]. inversions [Ros42]. inverting [Tri84]. Investigating [Daw95, HS89a]. Investigation [MS96, Fur46, Gei48, Wol84]. Investment [HT97c]. Involving [CL95, PB98, PSS93, Sie93, ST99]. IQ [MW98a, MW98b, RW98, WPD98]. Iraqi [DKW97]. irreconcilability [BS87, Bor87]. Irrelevant [LT97]. Isotonic [Gey91, Bac89]. Issues [EFS92, Gei92, GS93, NGG97]. Item [FMR62, LCG91, MS96]. Iterated [YM99]. Iteration [Mar90]. Iterative [Bjö78, DL74, Hoo93, O’S90, RC74]. Iteratively [DS91].

J [BK86, Bor87, BC82a, CMS87, Cla89, DD82, Fab94, FH81a, FH81c, Hoa73, HL81, HG83, JTB81, Kot89, KP82, Krz81, Law82, LB94c, MW82, Mac87, MS98a, MPG86, MT82, MS95, NS86, Oak85, PT82, Pra82, Rin85, RC82, Sch82, Sim87a, SG95a, Sta94, Tuk47a, Wee94, Wes98, WM81, Whi92, YM86, de 83a]. J. [Bor87]. Jack [BH88]. Jackknife [SC95, Sha89a]. jackknifed [Fay85]. Jackknifing [PS82, Sim86]. James [GS91b, MB88a]. Japan [Yam92]. JASA [Sti88]. Jeffreys [IL91]. Jerome [LV88]. John [JV85, Sti86]. Johnstone [HHM88, TV85]. Joint [BS94, CLM93, JS94, LA94, Rog88]. jointly [MB82]. Joseph [Daw82]. Judging [Jen86a]. Judgment [BW94]. Jurors [SKF98].

Kadane [Daw82]. Kalbfleisch [BH88]. Kalman [Fah92, GM94, Koo97, SW91a, Zeh88]. KAP [GJDT98]. Kaplan [Akr86, CHL82, Efr88]. Kass [DH83]. Kendall [LQ85]. Kernel [Alt90, Bly93, BGH93, FHW95, Gou97, HW91, HW86, HW92a, JDP94, JS97a, MBR91, MS88b, Sch95a, SM90a, SW83, WJ93, jW97, WCH98, Bie83, Har84, Jon89, Mül87, Pad86, Sta89a, Sta89b, YM99]. Kernel- [HW91]. Kernel-Smoothed [MBR91]. Kernel-Type [JDP94, Pad86]. Kingston [Tuk47a]. Kinsey [CMT53, CMT55]. knockout [Isr81]. Know [RSV95]. knowing [Bha81]. Knowledge [CS95, Pea91]. Known [LW91, FE88, vd87]. Kolmogorov [Gle85, Gui88, Mas51]. Kolmogorov-type [Gui88]. Krasker [Rup85]. Krewski [DH83]. Kriging [Las94]. Kruskal [GR81, Wei81]. Kruskal-Wallis [Wei81]. Kurdish [DKW97].

L [HL81, MB88a, WM90]. L. [Gei48]. L.A.M.S.T. [BN74]. Labeling [GRB^{+}95, HIT86, Sim87a]. Labor [SG96b]. Labour [SR95]. Lack [Chr91, MNS99, SW91b]. Lack-of-Fit [Chr91, SW91b]. Lag [CHR94, NP95]. lagged [Eri81]. lags [TA94b]. Laplace [Cra94, Erk94, LR97c, Shu97, Sti86, TKK89]. Laplace-Metropolis [LR97c]. Laplacian [O’S90, O’S91]. Large [HW92a, Jen91, LRR91, NM98, OR92, PY99, RB90, BOP88, BE85, CLNL87, DL74, EH87, Rel72, SM84a, VJ88, WM80, WM81, Zel87]. Large-Sample [LRR91, VJ88, WM80, WM81]. larger [HG81]. Largest [Wan95, Dav82]. Last [Sch99]. Last-Move [Sch99]. Latency [DXP92]. Latent [BRMZR97, BZ99, CR92, CG84, For92, For94, Joh90, LR97a, LCG91, WPK99, ZT97, DM88b, Mis85, Pud82]. Latent-Class [Joh90, DM88b]. Latin [Owe94, Tan93, Ye98b]. Law [Ben38, Hil70, Hil74, HW75, Fur46, Gei48]. Lawrence [Bow87]. Laws [Roj96, Sch84b]. Layer [Whi89, Fab94, Whi92]. Layout [PB98, HC84]. layouts [Rin83a, Rin85]. LDU [CD96]. LDU-Based [CD96]. Lead [MW98a, MW98b, RW98, WPD98]. Leading [DR91]. League [CH94, GS98b]. Learning [EGMP93, Whi89, Fab94, Whi92]. Least [AI95, Bjö78, CD91, CW94a, CHW92, DS91, DS97b, ES90a, Gar94, GF96b, Hoo93, HM91, Krä86, Lon67, RSY93, Rou84, RSW95, Sch87a, Wam70, CR82, CGG85, CWR88, Chr87, DWL87, DP89, Dyk83, Hay86, Lea84, PP88, RC80, RC82, SS86, SS87b]. Least-absolute-value [Sch87a]. Leaves [Ros93b]. Left [GL96, Gui88]. Left- [GL96]. left-truncated [Gui88]. Legislative [GK90]. Lehmann [Boo82, Mun96, Ros93a]. Lemma [Li92]. Length [ME96, SS90a]. Leo [LV88]. Leonard [LF82]. Less [Bar91, BS96a]. Lessons [JS97b]. Letters [Ano99d, Ano99e]. Level [KP91, KM96a, Niu96, PM97, Tom94, CJ88, Ros89, TR87]. Levels [CO90, Fin90, GC95, LRR91, SKF98, YKB97, PH81]. Leverage [EM92, Fun93, Rv90a, Rv90b, SC92b]. Liberal [Moo98, Riv86]. lieu [APS81]. Life [BM92, Zuc98, FMD82, Ger82, HPP86, Kar87, MS85b, O’N84, SS82a, Tak81]. life-testing [MS85b]. Lifetime [CH95a, Suz85]. Lifetimes [HHH92]. like [WF82]. Likelihood [AC97a, AF97, AS90b, Ber91b, Bjø96, Bre90, CM93, CJ96, Chi95, CM98, DF98, DKR95, Eas91, Efr92, FHW95, For94, Fra91, GW88, Gey91, Hab95, HL98, HMY97, Hun94a, Jen91, KBB93, KT95, KAT91, KS94, LK97, LT97, Lav91c, LR96, LK92a, Li95, LQT97, Luc93b, MBL98, Mau95, McC94, McC97, Mur95, MRvdV97, Owe95, PO94, RH96, Sas93, Sev91, She93, SW86, SW91a, TL97, VM96, Wan96, WM98, WZ97a, Yas95, Atw84, Bha85, BS83b, CR82, CS90a, Cha84, Con89, De 83b, Dyk82, DL88, FD85, Fer82, Fra84, Fry84, GH86, HB86, Jun96, KW85, KC93, KM83, LLS87, Lin85, Pra81, Pra82, SM83, SM84a, Sch84a, SL87, SS94, Spr83, Sta89a, Sta94, TC72, TH87, TZ81, WR84, Wij85, Won82a, Won83, WPSL84]. Likelihood-Based [Hun94a, Sta89a]. Likelihoods [Efr96, Gel96, Geo96, Mor96, Rei96, ST96, TR95, SC85]. Likely [BS96a]. limacon [BCH95]. Limit [WBB93]. Limitations [Akr91a]. Limiting [KSJ94, LS97, MBR98]. Limits [LPT91, SW97, VHE91, Bly86, Cla87a, Ger82]. Linder [TW52]. Lindley [Daw82, Sha82]. Line [Nau65, JV85, TV85]. Linear [AC95c, BF97, BCJ96, BH98, BC93, Buo96, BS96c, CFGW97, CH97b, CTC96, CM84a, CB97, CZ77, Coo79, CCD98, DG99, DM93, DM96, Efr99, EMM95, ES90b, Fah92, FG94, FG95, FHW95, FS95a, FK99, For92, Fun95, GS99, GNSC98, GLG92, GH94, HS93, HMM98, HJ92, HM97, HLC92, HC96, Hof92, Hoo93, Hös94a, Ibr90, JGC95, JMS97, Jia98, Jon96, KL96, KS96, KT95, KP81a, LR95b, LWY93, LS95, LW92, LB96, Liu96, LJK93, Luc93b, MH94a, MKSB90, Mat83, McC97, MT95b, MG97b, MLRR92, Neu98, Now92, Per97, PML98, RMH97, Ric97, Riv98, RFB97, SC92a, Sch88, SB97, Sha93, SU74, SK91, SS91a, Sil92, SRC92, SHH78, SW91b, SL93b, Tho90a, Vel98, VL96, WKY93, WLGC98, Wes95, WM91, WEST91, YJ98, ZL95, ZM93, Alb88b, AP87, BLE88]. linear [BK82, BS84a, Ber89, Boo86, BM70, CR82, CC74, Cha84, Che70, CSS89, DP89, DH93, DL88, Efr84, Efr86a, Esc86, FW88, Fin83, Flu85, Gew82, Gew84, Gol83a, HB86, HVY83, HM83, IL91, KH84, KK84a, Kan88, KC93, Kob85, Kob86, Koe86, KP87, Kor81, KD87, Kun80, KSC89, LQ85, LP86, LT88, LB94c, LB94d, MH90, MB88a, NM83b, Nyb86, O’G82, O’H87, OYR86, Ole85, PS86, Pol84, Rav89, Rei84, RC81, RC80, RC82, SH81, Sam85b, Sch87b, TT86, Uus83, WHM85, WM80, WM81, Wol87, ZK91].

Linearization [CJ95a, CJ95b]. linearly [Moo81]. Lines [APR92, ES90a, Mad59, Spu99, Sch87a, SB81, TV85]. Link [CL92a, CW94b, CM98, McC98, TW98a, TW98b, Var96, Var98]. Linkage [BR95, Har81]. Linked [BD91b]. Linking [NFL92]. Liouville [RS94a]. Liquidity [HT97c]. Literacy [Wal93]. Little [Bre92]. Liver [SH90]. Local [AC97a, CRW98, CNF95, DBB^{+}94, FG94, FHW95, FHMP96, GR96b, Gus96, Haa95, LNL95, MS98b, RSW95, Rup97, SG96a, Sta89b, TH87, YJ98, HB88, MS98a, Mül87, Sam85a]. Localized [AM92]. Locally [BGH93, FKYZ99, GV86, SS96a, vR98, BM83]. Locating [CG97]. Location [Ber91a, Eas91, Hoa73, O’H90, PR90a, WR94, Boo81, CC83, Die82, Esc86, MH82, Mor86, Par93a, Swe84, TB86, Tay85, WP85]. location-scale [CC83, Swe84]. Log [FK99, GH94, KY99, Yan97, DL88, FW88, FHO87, Kor81, Ole85, Pra81, Pra82]. Log-Linear [FK99, GH94, DL88, FW88, Kor81, Ole85]. Log-Rank [KY99, Yan97, FHO87]. Logarithmic [CLT96]. Logical [Wes97b, Wes98]. Logistic [AL94, BL93, CWW95, CRS^{+}91, CS90b, Efr88, FLZ96, For92, Gey91, KR96, KM91a, O’N94, ST99, CD89, CC74, HMP87, Jen86a, Jen86b, Len88, LZ89, Stu88, Whi81b, Whi81c]. Logistic-Normal [CS90b]. Logit [MH94b, Soo92]. Lognormal [Haa90, BM70, Ruk86, SS85]. Long [LR97b, Luc93b, Ben83, Riv86]. Long-Range [LR97b]. long-tailed [Ben83, Riv86]. Longest [Lou96]. Longitudinal [BF97, CCC96, DS97a, DLW95, GB97, HZ98, KC96, LCG^{+}92, LL97a, PC97, PHL98, Sch96, Seg92, SBD97a, SBD97b, SL98b, SC97c, TCS94, TDW95, WCH98, ZLRS98, CR89]. Longleaf [RC94]. Look [APR92, BRB76, Shu97, Sti88]. Lorelogram [HZ98]. Lorenz [KM96b]. Loss [BL93, Rue95, AP87, Zel86]. Lost [DB97]. Lottery [Ang91, JK93]. Low [LT94b, RG94]. Low-Dose [RG94]. Lower [BF90, CS96b, Del89, Sat95, Kim86, WF82]. Lupus [AME96]. Lyapunov [MEGN92].

M [DH83, MY91]. M. [God82b]. Machines [Met50]. macrofunctions [Sto82]. magnitude [AK86]. Main [MM95]. making [WF82]. Mallow [RS94b]. Mallows [dJdWW88, MB88a]. Mallows-type [dJdWW88]. Malpractice [Coo91, GHCF94]. Man [WT90b]. Managerial [GS93]. Mann [BW94, Mee90]. Mantel [WW87b, YF90, YF95]. Mantel-Haenzel [YF90]. Many [MBR91, BF83, CM84b]. Mapping [WCXG97]. Maps [MST97]. Margin [Qui98a]. Marginal [Che94a, Chi95, Cla87b, Cla89, GR96a, GS90a, HZ96, Joh92, LK97, LA94, ML94, SL98a, Coh81, CGM83, GA82, KR82, LHT89, TK86, WLW89]. Marginals [Gus96, LWW91, LK95]. Margins [LW91]. Marital [YP92]. Mark [Oak85]. Markers [LTW96, SL93b]. Market [CW91]. Marketing [ALJ99, Eri81]. Markov [Ros95b, BBGL97, BM95, CLWZ95, CC96b, DwSH94, FS86, FK94, GT95, GRS98, GJR98, HM85, Hig98, Joh96b, Joh98b, KL85, Kar85, KK84b, KP81a, Lou96, MTY95, Pic87, Ros95c, RP96, Sam90, Sch83b, Yak85].

Markov-dependent [Sch83b, CLWZ95]. Markovian [Krz87, LNL99]. Marriage [BB90]. Marrow [DwSH94]. MARS [LS91a]. Marshall [Ryu93]. Martingale [Bro97, Gil84]. Masking [BK85a, BK86]. Mass [DMLC97, MS91a, MS95]. Match [BR95]. Matched [Da̧b90, FS95b, HHS97, Tam85, WG88]. matched-pairs [WG88]. Mathematical [BR92, Dia78, McP42]. Mating [Shu97]. Matrices [CD91, CD95, GL92, JB92, Rig97, XLR91, Kar81, Wor74]. Matrix [CLT96, CD96, GS91a, HW90a, LKC97, Sch91b, ZT97, Boi86, BK84, CLNL87, Che70, Dwy67, Mar64, Mat83, McC82, WA67]. Matrix-Logarithmic [CLT96]. Matter [CH94]. Maturation [AHB97]. Max [KSW92, MZ89]. Maximal [Ter90]. maximin [Gas85]. Maximized [BB94]. Maximizing [VK92]. Maximum [AF97, AS90b, BS83b, CJ96, Dyk82, Eas91, Efr92, FD85, For94, Fry84, GW88, Gey91, GJP96, HB86, Hab95, KBB93, KW85, KT95, KAT91, KS94, LLS87, LT97, LK92a, Mau95, McC94, McC97, MNS99, MRvdV97, Nau65, RH96, Sas93, SM83, Sev91, She93, SW91a, WZ97a, Atw84, Bha85, CR82, Cha84, DL88, Fer82, GH86, Hay86, Lin85, SM84a, Sch84a, SL87, Spr83, TC72, TZ81, Wij85, Won82a, WPSL84]. Maxwell [Bro92]. May [GS95]. McCracken [Gol63]. Mean [AS90b, Che95, Che94b, CHW92, HJ92, HL96, LR95a, Law81, Law82, PSS99, PR90b, RV98, SMF94, Sev91, SW91b, SY95, Tar86, The93, WM98, YB97, Zuc98, Bha81, CH83, CC74, HV86, Kar87, NS93, Sam85a, SW86, Tak81, Uus85]. Means [ALM97, BHS98, BM96, Cro97, Esc94, FNP95, Fon92, HH90, HHS97, Kun90, Lau92, MM93, PC97, RV98, SP91, SC97c, BD88, BB89, Ber89, BA83, Coh81, Cre82, DB83, GL87, HG81, Hay89, HT81, Lin74, Mal82, Mar64, SI85, Tsu86, YZ88, dJdWW88]. Measure [EL90, Hua97, MKSB90, ELR87, SS85]. Measured [HTP95, TDW95]. Measurement [Buo90, Buo96, BS96c, CL92a, CKLS96, CS94, Efr97, FS95b, GMP97, Gew82, Gle92, HDG97, JTC99, Joh90, MW98a, MW98b, RW98, Sho93, SK91, SB95, SC95, WLGC98, PT82, Rei82, Sch86a, WK88]. Measurement-Error [MW98a, MW98b, RW98]. Measurements [Agr97, Con92, HZ96, JD91, Kus98, LD92, LY93, LT96b, OR96, Con89, HW86, LD91]. Measures [Akr91a, AA94, AHB97, ALHH92, BL92, CPP96, Car94, DBB^{+}94, Gew84, Goo91, JR96, KW96b, Lit95, Lju93, MLRR92, OMF^{+}93, Roc92, Sev96, Tay85, TA90, Tho91b, Abr88, Goe83, Joe89, KR88, LLS87, LB94c, LB94d]. measures-data [LB94c, LB94d]. Measuring [For94, HHH92, Ye98a]. Mechanism [Lit95]. Mechanisms [SKF98]. Median [BD93, CHW92, ES90a, GS89, GNK96, RC93, SW98a, WH90, Yan99, YJW95, AP87, BC82b, Jun96, Rou84, SS87b, ZH87]. median-unbiased [AP87]. Medians [BHNO92, Gui83]. Medicaid [GLB98]. Medical [Coo91, GHCF94, NGG97, Ueb93]. Medicare [KR96]. Medicine [Gei92]. Medium [LT94b]. Mediziner [TW52]. Meetings [Hur85a]. Meier [Akr86, CHL82, Efr88]. Melanoma [BHB96, BSB98]. memoriam [Man94]. Memory [CKLH92]. Men [MPCB95]. Meta [DPWH99, LB94b, Sil97]. Meta- [Sil97]. Meta-Analysis [DPWH99, LB94b]. meteorological [HW94].

Method [Akr90, BR95, BO97, CCD97, CV95, Cra94, Dwy45, Eas94, Efr92, ET97, GKK91, Goo96, Gou97, GW91, HW91, HRW90, Hol95, KP91, LB93, Lou96, OR98, Pap96, PF91, Pes93, Rag93, RH96, RB90, SC97a, SC97b, SS97, SMP95, Tro95, VY80, WS94, YL97, YP92, Zuc98, CMS76, CMS87, Dua83, DL74, HS89a, Jay79, KY86, KP81b, KP82, Loh85, Mar84, MU49, NP83, Pud82, Ric82, Sch81, Sch82, Sil84, Spl83, TC72, Won82b, YF95]. Methoden [TW52]. Methodological [WHGA96]. Methodology [SHBF94, Tuk79b, BD88]. Methods [AB97, AGM94, Atk94, BOP88, BOP92, BBGL97, Bir54, BBH94, Bre92, BPS95, CG92, DM93, DM89, Efr96, Fer98, FHJW96, Gel96, GC91, Geo96, Goo96, GLT79, Gra92, Haa90, HT97a, HT97b, HWY99, HA94, Hig98, Jon96, KO96, LKM98, LK91, LT96b, LC98a, LC98b, MSH90, Mor96, NP95, NGG97, O’S91, Por97, Rei96, Seg92, ST96, She97, SW90, SG96b, SC80, SHH78, SHGL96, Yas95, Zie90, AG83, AG85, Alb88b, Ame74, CC83, Con89, DJK87, DH83, FM81, HB86, HA95, Hur85b, JV85, KR88, Kin85, Law87, LL81, Lou84, Mül87, NWGP84, O’G82, RW85, Ros89, SS80, SL80, Stu56, TV85, WP85, Tuk49a]. metrical [Fur46, Gei48]. metrische [Fur46, Gei48]. Metropolis [LR97c]. Microdata [BKP90]. Microscopy [QTC96]. Migraine [DPWH99]. Min [KSW92, MZ89]. min-max [MZ89]. Minimal [AP99, PHS94]. Minimax [Amr95, Hec88, Hub83, Nai87a, Sam81, Sob91, Wie98, vd87, Cas85, Geo86, Sam85a, Sch87b, Wat87]. minimization [YZ88]. Minimum [Bet89, Boo81, CCB96, Har92, Ryu94, Sie93, Sim87b, Ste87, TB86, Ame74, BN74, Dha93, IS83, WPSL84, de 81, de 83a]. Mining [Ye98a]. Minorization [Ros95b, Ros95c]. MINQUE [Pri74, Wes87b]. MINQUE-type [Wes87b]. Miracles [Kru88]. misallocation [CM84a]. Misclassification [AHB97]. Misclassified [GWG90]. Mismeasured [PF91]. Misregistered [CHR94]. Missing [AS90b, Che94b, CLWZ95, Efr94, Jon96, KLW94, Lit92, Pai97, RRZ95, RR95, SLS97, SKF98, WSJ92, WWZtO97, CS86, KA86, SS86, Sto86, Wit88]. Missing-Data [SKF98]. Missingness [CCC96]. Misspecification [CHW92]. Misspecified [GLG92, KY85, Whi81a]. Misspecifying [BL93]. misweighted [Cre82]. Mitofsky [Pot87]. Mitofsky-Waksberg [Pot87]. Mixed [BH98, BC93, BD92, BD91b, GS91a, HDG97, HC96, Jia98, Lau92, LB96, MS92, McC97, MPCB95, MG97b, Ric97, VL96, WLGC98, Wee91, ZDL96, ZLRS98, Bro83, KH84, LB94c, LB94d, SH81, ST82, SB81, SM84b]. Mixed-Effects [HDG97, MPCB95, VL96, ZDL96, LB94c, LB94d]. Mixing [CDP96, Gou97, HRW90, LK92b, MZ96, vd87]. Mixture [Cra94, CCB96, HRW90, LMR96, LR92, Lit93, MR97a, MR97b, O’S93, Roe94, RCL96, Wes97a, WC92, TC72, WPSL84, ZC88]. Mixtures [BR94a, EW95, Joh92, LB93, MZ96, PJT96, Pri94, RW97, AG83, Has69]. Mixtures-of-Experts [PJT96]. MLE [LK92b]. Mobile [Sam90]. Mode [Erk94, Fah92, Hsi97, RC74]. Model [Agr97, AP99, AMCS98, AC97b, ALJ99, AL95, BAP95, BP96, Bol95, BO95, BZ99, BD92, Buc95a, Buo96, BD93, Bur94, CM97a, CKLH92, CPP96, CW91, CGM98a, CGM98b, CLT96, CLM93, CDP96, CLWZ95, CHW92, Cro95, DwSH94, DR98, DSM98, DM96, Eas94, ES97, ES90b, FG99, FBF97, Gel94, GHCF94, GLG92, GS98b, GS93, GR92a, HJ92, HLC92, HN92, Hös94a, HR97, HM91, Joh94, Kit98, KKT98, KP81a, LRM96, LK92a, Lin91, LCG91, MR94, Mer94a, Mer94b, MR97a, MBR98, MPCB95, MR97b, MRvdV97, Pap96, PS94a, Peñ98, Per97, Pie87, PJ94, PH90a, PHS94, PR93, PHL98, QH98, RGZ95a, RGZ95b, RMH97, Ric97, RFB97, RT96b, RP96, Ryu94, SW95, Sam90, Sas93, Sas97, SDW98, Sch95b, SB97, SLS97, Sha93, Sha96, SK91, Sil92, ST99, SH90]. Model [SA92, Ste94, SW98a, SGP98, SW91b, TCS94, TL95, VL96, WKY93, WR95, WJ92, Wol95, WM91, WEST91, WCH98, Yam92, YP99, YR95, Yas95, Ye98a, Zha92, Zha98b, ZWWB96, Alb88b, BJ74, Bin82, BS83b, BHT88, BA83, CR82, CSS89, CS87, DD88, Dal82, Dha93, EH87, Fat83, FW88, FJV89, Fin83, FB83, Fry84, Gil84, GC82, GHP87, Har84, Hec88, HM83, HT82, Hjo94, IF82, Kar81, KD87, LP86, LT88, LW89, MJ82, MH82, MSW^{+}83, NM83b, PH86, Raf87, Rei84, RC80, RC82, Sch86a, Sch88, SM84b, Spl83, TT86, VSK85, Wan89, Wat87, Wei84, WM80, WM81, Wol87, YD84, dJ89]. Model-based [DR98, BHT88, Wat87]. Modeling [AME96, AHB97, AL94, BB90, BD97, BS84b, CPS92, CH95a, CNF95, CM97c, CCC96, DLD91, FOS95, Fyg97, GMD95, GBJ96, GS90b, GRB^{+}95, Gri90, HMM98, HYP97, Hei91, HOLDG98, LA94, LMR96, LCMA92, LS91a, LR97b, Lit95, MH94b, ML94, NGC^{+}95, NT95, OO97, PC97, PS94b, Pol91, SA92, SC97c, TB81, Tsa98, TDW95, Ueb93, WB96, Agr83, FS98b, HW94, Kit87, LLT89, Tsa89, WLW89, Tuk79a]. Models [AC97a, AR90, Ahn97, AC97b, AS90a, AI95, AC95c, BF97, BR94b, BCJ96, BL93, BBGL97, BH98, BDR92, Bre90, BC93, BR98a, BR98b, BD91b, BH97, BS96c, CD91, CMW97, CS90a, CKLS96, CFGW97, CH95a, CR92, CL95, CH97b, CT93, CWY97, CR89, Cho92, CB97, CM91a, Coo91, CCL90, CS94, CW97, CCD98, Cox95, CS90b, Cro97, CNS90, CCB96, DG99, DM93, DR96, Dea92, DB97, DH99, Fah92, FT94, FW96, FG95, FHW95, FZ98, FOS97, FL95, FLZ96, GW88, GHRPS90, GS99, GNSC98, GL92, GH94, Gri94, Gu90, Hab95, HS93, HKS98, HMM98, HW92a, HZ96, HK92, HR94, HDG97, HC96, Hoo93, HH96, HW95b, HS94, Hun94a, Ibr90, JGC95, Jia98, JTC99, Joh92, KSJ94, KP97, KL96, KS96, KPR91, Kne98, Koo97, LK97, LR95b, Lan96, LCG^{+}92]. Models [Lan98, Lau92, LMR96, LS95, LR97a, LB94b, LT98, LB96, LCG91, LR92, LL97b, Lit82, LW91, Lit93, LNL95, Lju93, LJK93, MR94, MO98, MH94a, MS92, Mau95, McC97, MS96, MG97b, MLRR92, NM98, Neu98, NP95, Niu96, Now92, OR98, OKS97, OR92, PB94, PJT96, Pet92, PWM92, PML98, Qui98b, RMH97, RS98, RG93, Riv98, RRZ95, RR95, RK96, SL93a, Sch96, SGC95, SS94, She93, SS91a, Sil97, SHH78, SL98a, SS91b, Sta91, SKF98, SB95, Ste90, SW91a, SHGL96, TR97, Ter94, Tho90a, Tsa98, Vil90, Wak96, WB96, WLGC98, WW98, Wan98b, Wee91, WJ92, WCR94, Wes95, Wes97a, Whi89, WHG98, Wie98, Yam92, YJW95, ZDL96, ZLRS98, ZL95, ZM93, dCP99b, Abr88, ACK87, AH81, Bac89, BK82, BK85a, BT81, BM70, BY86, CR84, Che70, Cla87b]. models [Cla89, Clo82, CM86, DM88b, DP89, DL88, EH87, Efr84, Fab94, FV88, GH88, Goe83, Goo91, HB86, HM85, Hwa86, IL91, Jun96, KH84, KK84b, Khu81, KR86, Koe86, KP87, KA86, Kor81, KY85, KSC89, LZ89, LR87, LB94c, LB94d, Man88, MH90, Mat83, MS88a, Nef82, O’H87, OYR86, Oak89, PC84, PS86, Pol84, Rei82, Ruk86, SD85, SH81, Sen74, Sha86, Sol84, Spi82, Sta89a, Sta94, Sto86, Stu88, Swe84, TA89, TT84, Tsa84, Tsa87, Var88, WHM85, Whi81a, Whi92, Wol86, Won87, ZK91, Zeh88]. moderate [Dav82, GR81]. Modification [Zuc98]. modifications [Gol83a]. Modified [Joh78, Qui98b, Shu97, HPE81, SW82, Won83]. Modular [TFP90]. Modulus [FM82]. Molecular [SRSD94]. Moment [CTC96, LRR91, CKC83, HHO88]. Moment-Based [LRR91]. Moments [AK92, HRW90, HN92, HL97, LB93, SC94, Mar81, TC72, TK86]. Monetary [PH90a]. Monitoring [GS90b, HOS95, Yas95]. Monotone [BS97b, GMPS99, HS98c, Lee96, MS94, VS93, DR83]. Monotonic [AHB97, SS97]. monotonicity [Fin93]. Monte [JTB81, NG94, Ros95b, BBGL97, BS96b, CPS92, CL95, CC96b, GT95, GRS98, GJR98, Hig98, Hur85a, JTB80, Joh96b, Joh98b, LC98a, LC98b, MU49, MP95, Ode91, OB93, Ros95c, SMP95, Sha89b, TV85, WT90b]. Morphology [BR92, FR99]. Mortality [AS90a, DMLC97, GWG90, HLC92, LCMA92, OMF^{+}93, SK91]. Mosaic [Fri94]. most [BM83]. Motion [PC96, Riv98, Ste94]. Motorcycle [Wei92]. Move [Sch99]. mover [Fry84]. mover-stayer [Fry84]. Moving [BDR92, Haa90, LL97b, RP96, SL93a, TR97, Tel67, YR95, Spl83]. Moving-Average [LL97b]. MR [BK86, Bor87, BC82a, CMS87, Cla89, DD82, Fab94, FH81a, FH81c, HG83, JTB81, Kot89, KP82, Krz81, Law82, MW82, Mac87, MT82, MS95, NS86, Pra82, Rin85, RC82, Sch82, SG95a, Wee94, Wes98, WM81, Whi92, YM86, de 83a]. MSE [Fel74]. much [Ruk87]. Multi [Fri94]. Multi-Way [Fri94]. Multiargument [MS94]. Multichannel [BO97, O’S93]. Multiclinic [BPS95]. multicomponent [BJ74]. Multidimension [CS93]. Multidimensional [BPL97, DL92, Lav91a, CG84, Yan84]. multifactor [JK89]. Multilevel [MH94b]. Multimodal [OB93, CKC83]. Multimodality [MS91a, MS95]. Multinomial [CMW97, Eub97, GJP96, He90, HLC92, NM98, OR92, RMS90, SG95b, Ala84, CD89, DKMO89, FS87, Hab82, Hoe84, Lee87, SG95a, Zel87]. multinomials [CG82, Sim86]. Multiparameter [AC97a, GV86]. Multiparty [Gel94, Mer94a, Mer94b]. multiperiod [HHO88]. Multiple [Agn85, Atk94, BRMZR97, BPR95, Bin98, Bof85, BS91, Bre98, BS94, CRS^{+}91, CC96a, Dic83, DT92, EH83, EH84, Fin90, GKL98a, GKL98b, GB98, GCW96, HW90a, HS93, HR90, HL94, Hol99, Jon96, Jud98, Kad98, Kun90, LTM92, LLR95, LR97a, LB96, LNL95, MZ97, ME96, MP96, PB98, Pen97, Pep91, RS86, Rub96a, SC97a, San98, SC97b, SMF94, SLS97, Ski91, SL93b, Tom96, Tro95, Wes97b, Wes98, Yam90, Zha98a, Ali89, BB88, BF85, CS85, CT98, CSS89, Chi74, DG93, Fin93, Fli84, Gew82, HS89a, HE83, IS87, Mui85, MRW87, Nai87a, Nai87b, Rin83b, Tam81, TB81, Uus85, Wal86, Wit88, Wol87]. Multiple- [HL94]. multiple-design [CSS89]. multiple-system [CT98].

Multiplicative [FM82, Hwa86, MJ82]. Multiplicity [HB99, CSC86]. Multiply [LRR91, CC74]. Multiprocess [Bol95]. Multirater [Joh96a]. MultiSample [JSKK98]. Multistage [FV88, SCC98, MS85a]. Multistate [Sch99]. Multivariate [Agr97, AR90, AA94, BL92, BDR92, BF93a, Buo90, BS98, CD91, Cha96, CM97b, DM93, EM90, Fah92, FOS97, FLZ96, Fol96, GWG98, GR96b, Gle92, GR92a, HMO97a, HMO97b, Hou95, Joe95, JSKK98, KST98, KL96, KP90, LA94, Lan96, LK92a, LS91a, LT98, LW92, LB93, Lit93, LS93, Liu95, Liu96, MZ97, ME96, MFS84, ML94, MLRR92, Now92, OV95, PR90a, PML98, Rei82, Rig97, RR95, RW96, Ros89, Rv90a, Rv90b, RP96, SBS94, Sch86b, Sie93, Sob91, SL98a, Tan94, The93, Tho91b, Tsa98, Vel95, WHG98, WR94, BS83b, BS84b, BA83, CH83, CSS89, Dav82, Dwy67, GH88, GJ84, HVY83, Joe89, JW88, KM96b, LZ89, MY95, MO88, NS83, NS86, Ols74, Ran89, Rei84, SS85, SW89, Spu88, SW86, SZJ87, SM81, TB86]. multivariate [VW87, WL84, WLW89, Wri83]. Multivariate-t [ME96]. multiway [Clo82]. Multiyear [CLY92]. Myocardial [OMF^{+}93]. mysticetus [RZ98].

N [Tuk49a, Tuk52]. Nadir [DMLC97]. Names [NFL92]. National [CM91b, CH94, GS98b, HZS^{+}90, MSML97, Sta91, ZT97]. Natural [GPS95, KO95, Sla94]. Naturwissenschafter [TW52]. Near [Cha90, Chr91, MNS99]. Near-Integrated [Cha90]. Nearest [BR98c, Sch86b]. Nearly [YBN85, BHT88, Cha88a, Vos86]. need [Isr81]. Needed [LS91b, SM90b]. Negative [Mar88, GA82, SM84b]. Neighbor [BR98c]. Neighborhoods [Pre96]. neighbors [Sch86b]. Nephritis [AME96]. Nerve [Bri92]. ness [Hjo94]. Nested [BR98a, BR98b, FZ98, KW95, Kne98, Lan98, Mar92, NG94, RS98, Sas97, WW98, WJ92, Yas95, MB88b, WR84]. nesting [Ros89]. Net [PJ94]. Network [BH97, Cox95, McC98, NT92, TW98a, TW98b, Var96, Var98, Whi89, Fab94, MP83, Whi92]. Networks [HD97, SI90]. Neumann [Bar82]. Neural [HD97, PJ94]. Neuroanatomy [DLB91]. Neurological [Wes97a]. Never [GKB98]. Newton [LB94c, LB94d]. Neyman [Fan96, Led94, Leh93]. Nielsen [Bue82]. Nino [Sol95]. No [BZ99, Raz90, YKB97, BK86, BN74, Bor87, BC82a, CMS87, Cla89, DD82, Fab94, FH81a, FH81c, HG83, JTB81, Kot89, KP82, Krz81, Law82, LB94c, MW82, Mac87, MS98a, MPG86, MT82, MS95, NS86, Pra82, Rin85, RC82, Sch82, Sim87a, SG95a, Sta94, TA89, Wee94, Wes98, WM81, Whi92, YM86, de 83a]. No-Observed-Adverse-Effect [YKB97]. Noisy [FHTW97, GC95, Joh94, Pre96]. Nomination [KS93, BS86b]. Non [DF98, HH91, RRS98, SM92, Gu90, Kit87]. Non-Bayes [Goo92]. Non-Bayesian [HH91]. Non-Gaussian [Gu90, Kit87]. Non-ignorable [RRS98]. Non-iid [SM92]. Non-parametric [DF98]. Non-response [RRS98]. Nonadditivity [GH90]. nonanticipative [Gut82]. Nonavailability [PMW93]. Noncentral [WK94]. Noncentrality [Ole85]. Nondetects [LPT91]. None [Bak98]. Nonequivariant [BS96a]. Nonhomogeneity [Pri96]. Nonhomogeneous [KY96]. Nonignorable [Con92, CCC96, PB94]. Noninformative [CW93, DG95, MV91]. Nonlinear [CPS92, CKLS96, CM93, CJ95a, CJ95b, CW90, CCL90, GW88, GR96b, GP92, Hei91, HDG97, LS95, LS91a, MPCB95, Niu96, OKS97, SA92, SC92b, SR92, Str93, TA94a, Val85, Vid98, BT81, Cla87a, Cla87b, Cla89, Eub86, Kar85, RW85, Shi84, TA94b, WA83, Whi81a]. Nonlinearity [LR97b]. Nonlinearly [EL97]. nonmetric [Rav89]. Nonmonotone [SV93]. Nonnegative [Rol94, WM98, KR86]. Nonnegligible [SS99]. nonnested [Efr84]. Nonnormal [CPS92, WBB93, MT81, MT82]. Nonnormalities [LM93]. nonnormality [Dav82, WM80, WM81]. Nonnull [CM91a]. Nonorthogonal [GN82]. Nonparametric [AA94, AAB97, AB97, AF97, AC95a, AD90, BRR92, BZ92, BW92, BPS95, BDM97, CS85, CDW93, CCD97, Che94b, Cho92, CF87, DN98, EL97, ES90b, ES93, Fan92, FG95, FHMP96, GR97a, GR97b, GMCM96, Gou97, GL96, Gui83, HH90, HW91, HT97a, HT97b, HP88, HMC^{+}93, Ize91, Jay96, JMV94, KR88, KS94, LT97, LK92b, Li95, LCWH97, MZ96, MEGN92, MN93, MS94, Nyc91, NGH^{+}95, Owe95, PO94, PF91, Por97, Qui98a, Raz90, Rup97, SG92, SC94, SK97, SL98b, Sto89, Suz85, TFP90, TA94a, TA94b, Tuk79a, Wan91, WS94, WR95, WC99, WG83, WZ97a, WK98, Yak85, ZDL96, Zha96, Zha99, BH88, Cro74, Die82, Dua83, HB88, JSG88, LQ85, Len88, Mül87, Noe87, Rom88, Sam93a, Sch88, ST60, SW87, Ste81, TS85, TJW88, Yan85]. Nonparametrics [SCCF93]. nonpositive [TKK89]. Nonrandom [Ros95a]. Nonrandomly [DS97a, GB97, LL97a, SBD97a, SBD97b]. Nonresponse [Con92, Kot94, PB94, Sta91, AG85, Lit82, RS86, Smo82]. nonstandard [SL87]. nonstationarity [Cre88]. Nonstationary [Ada98, AR90, CZ77, GM94, Koo97, Qui98b, Rob94, SG92, WPK99, YR95, BA83, Cha88a, Kit87, TT84]. nontrivial [Pic87]. norm [Ste87]. Normal [AL94, BH96, CCL90, CS90b, Cro97, Esc94, GHRPS90, Gle93, Hou95, Joe95, KM91a, Kun90, Lie94, LB93, Luc93a, OV95, OR92, Pes93, Rig97, Rol94, Sie93, SY95, WM98, dCP99a, Ber89, CH83, Che71, DB83, HG81, Hjo94, Hod87, Hoo89, KR76, Len88, LP86, Mal82, Mar64, Mar65, Ruk87, Sin86, SI85, SW86, Stu56, TC72]. normal-theory [Hoo89]. Normality [BG91a, BF93a, BH96, Cro74, LaR86, VJ88]. Normals [Roe94, RW97]. Norms [BM96]. Northeast [Sas97]. Northern [Sol91]. Note [Dor94, Hur85a, Kot94, BN74, CL85, CW85, Eub86, Fli84, HG82, HG83, Mar81, Won83]. Notion [Cha96, CS96a]. Notions [LS97]. Nuisance [BG91a, BB94, CW93, LK91, Sil96, TW89]. null [BS87, Bor87, Del89]. Number [BBC98, BF93b, CL92b, FM82, Gav78, GC95, HS98a, HW90a, HL97, Kie78, Mal78, PMW93, Roe94, SM78a, SM78b, Sim78, Vel98, Wil50, Che71, Col87, JR85, KOS85, PFV89, Wal86]. Numbers [Ben38, Gag43, JK93, KO95, LCG^{+}92, Cas85, Eth82, Pea61]. numerical [BHS81]. Numerically [GLT79]. Nursing [KR96]. Nutrient [KSJ94, MN93].

O [Bor87, DH91, Lav91b]. O-Ring [DH91, Lav91b]. O. [Bue82]. Objective [CC96a]. Objects [LJK93]. oblique [WR84]. observation [Zeh88]. Observational [RK90, Ros93a, Ros95a, SH92, Ros84b]. Observations [CLWZ95, Coo79, Fre94, KS97, Kus98, LW92, MBR91, OV95, Pet92, SB97, WSJ92, APS81, CM85, CM86, JG83, NWGP84, Tak81, WG83, Wei83, Wei84, WL84]. Observed [FSE98, GT98, RRZ94, Roc95, YKB97, Wal86]. Obtain [XLR91]. Obtained [RV90c, Tay85]. obtaining [PT83]. Occam [MR94]. occasion [NS93, Sti88]. Occupation [CRS^{+}91]. Occurred [GKB98]. Occurrence [BRR92, HW95b, Sol91]. Occurrence/Exposure [BRR92]. Occurs [Erk94, Hsi97]. Odds [Agr98, BB98, DD90, FS95b, HR97, MPS98, MRvdV97, RT96b, Shi91, SW98b, SW98c, VHE91, WKL95, YP99, DD88, MPG85, MPG86]. odds-rate [DD88]. Offered [CH91]. Official [AS90a]. Olkin [FM83, Ryu93]. Olkin-Sobel-Tong [FM83]. OLS [KD87]. Olympic [DAO94]. Omitted [MW98a, MW98b, Neu98, RW98, WPD98]. Omitted-Variables [MW98a, MW98b, RW98]. Omitting [BL93]. Oncology [GFT^{+}93]. One [BM99, CS92a, CF95, Fol96, HHS97, HY98, Hay90, HMO97a, HMO97b, KK84a, Leh93, MH99, PR90a, Sil92, SS95a, SRC92, Tab90, Tan94, WH90, CB87, DM88a, Die82, LL86, O’N84, PH86, Ros89, SW80, Sin86, Ste81, Won84]. one-parameter [SW80]. One-Sample [HMO97a, HMO97b, PR90a, WH90, Die82, O’N84, Ste81]. One-Sided [BM99, CS92a, CF95, Fol96, HHS97, HY98, Hay90, MH99, Sil92, SS95a, Tan94, KK84a, CB87, LL86]. One-Step [SRC92, Tab90]. one-way [PH86, Sin86]. Only [ASL96, HA94, Riz92, HA95]. Onset [BBH96]. Operating [Llo98]. Operations [McP42]. Opinion [GMD95]. opinions [LS86]. Optical [HYP97]. Optimal [BD89, BS85, Bly93, BCH95, CH88, CM93, CMST88, CW94c, DD81, DD82, Ede90, EK93, FE88, GLG92, GV96, Gri94, HTB94, HTP95, JMS97, KM91a, KP91, KPR91, Kus98, Min87, MP95, Par93b, Ros84a, SC96a, Spe85, Tam95, Tom94, Wan97, BF85, Bro81, GV86, HRR81, HPE81, LP86, Sin86, Stu82]. Optimality [BHS81, KW97, God82b, Jac89, Tar87]. Optimization [MLMT94, SS82b]. Optimum [AC95c, Fin90, GS97, HHM92, HYP97, MS88a, MS92, HHM88]. optional [Mar88, Sha83]. options [Ros84a]. Order [ACK87, AP99, DM95, Ede90, GT91, JS97a, MM93, Min98, MT95b, NSS97, Par96, QN98, RG93, RH96, Sch96, Shi91, Wan95, Zel93, Bha81, BHS81, HM88a, HG82, HG83, Lee87, Raf87, RW85, Ros42, Sol84, Vos86, WR84, Zac86, de 81, de 83a]. Order-Constrained [MM93]. Order-Restricted [MT95b, NSS97, ACK87, WR84]. Ordered [AMP90, AC95b, BM96, Cro95, DS97a, dE97, EGGS97, FT94, GB97, GM96, Hay90, JSKK98, LL97a, LGWC93, RK96, SBD97a, SBD97b, BS86a, Clo82, Dyk82, FD85, Gro81, KR84, Nai87c, Sim74, Tak81]. Ordering [AG96, DKR95, Muk96, Pan97, RS93, Wan96, Fra84, SW89]. Orderings [DF98, PW96]. orders [CJ88]. Ordeshook [Bow87]. Ordinal [AHB97, BZ99, CCC96, GKR88, Goo94, HZ96, Joh96a, KSW92, KC96, ML94, WKL95, Agr83, CD89, Sar89]. ordinary [Chr87, DP89, PP88]. organ [Cro74]. Organizing [Kit98]. oriented [Sch83a]. Orthogonal [AM92, CB90, Mar92, Rol94, Tan93, WW91a, Ye98b, BK85b, DH86, Vos86]. Orthogonalization [SG97]. Orthogonalized [CDP96]. Oscillation [KO96]. Other [BD91a, Bre92, Bre90, CS92a, SAK90, WM91, DH83]. Our [Wal93]. Out-Patients [TFAH94]. Outcome [FL95, GM96, VK92]. Outcomes [BRMZR97, CR92, DDV99, LLR95, LR97a, RRZ95, RRS98, SRR95]. Outer [Mey87]. Outlier [Bra90, BS98, Cha86, CLM93, GR92b, PY99, Pet92, HIT86, Sim87a]. Outliers [Atk94, BWC98, Fun93, HS93, Jen86b, LMR96, LRM96, MZ94, O’H90, RW96, Rv90a, Rv90b, BK85a, BK86, DG93, GH87a, Mar88]. Output [Chi95, DR91]. Ove [HL81]. Overdispersed [Bre90, DB97, KT94a]. Overdispersion [Dea92, Efr92, GS92, LR95b]. overparameterized [Sha86]. Oversmoothed [TS85]. Ozone [CCG^{+}97a, CCG^{+}97b, Cre97, GMS97, NT95, Niu96, SF97].

P [BC82a, Hoa73, Oak85]. P. [TV85]. Pain [PHL98]. Paired [AD90, FT94, LLE93, Alb88a, Wei83, WP85]. Paired-Comparison [AD90]. pairing [dJGM83]. Pairs [Da̧b90, WG88]. Pairwise [Hay89, HS89c, HVS84, SI85]. Panel [GH94, HT97c, LI98a, KL85]. Panels [PSS99]. Paper [Ano98d, MT49]. Paradigms [Fay96a, Ima95]. Paradox [He90, Sam93b, God82a, Gut82, KMT99, Mit91, Sha82]. Paradoxes [Hau92]. Parallel [Weg90, Bro81]. Parameter [BB94, CD95, Che90, EL97, Erk94, GSTM92, GMCM96, KW97, Lav91a, Lin91, NT92, O’H90, Sev91, SZJ87, SW98a, Tho91a, Ye93, BD81b, Cha88a, Cla87a, CC83, CW85, DM88c, FMD82, Lin85, LR87, Lou84, Moo81, Mor86, Oma81, Par93a, SM83, SW80, Sil84, Zac86]. parameter-effects [CW85]. Parameterizations [GPS95, Sla94, WJ93]. Parameters [BG91a, CLM93, CW93, FS95b, GT91, HHM92, KT95, KAT91, LK91, Pag90, Pol91, RTF92, Sil96, Sob93, Str93, WJ92, Yas95, ACK87, HHM88, Lee87, MS85a, Nyb89, Ole85, ST82, Sha85, SY84, Spi82, Spr83, TC72, TT84, TW89]. Parametric [CW94b, CS94, DN98, HMM98, HR94, Li96, Mei94, Mor83, NM98, YK96, DF98, FJV89]. parent [Ben83]. Pareto [AP89a, BVT96, CH97a]. part [Mat83, Akr91a]. Partial [AM92, CCD98, DS97b, Gar94, GMD95, GF96b, Gri92, Sas93, Hog74]. Partially [AR90, BM96, CFGW97, GT98, QN98, Qui98b, Rol94, YR95]. Participation [BD97]. Particle [PS99]. Particular [GK93, Mui85]. Partition [Cro97, GS97, Gut74]. Partitioning [Lan96, APS81]. Partitions [POH97]. Partly [PC97, SC97c]. Partner [SJ92]. Party [Gel94, Mer94a, Mer94b]. Parzen [Tuk79a]. Pascal [BCH95]. Past [Pol91]. Path [ME96]. Paths [ASL96, Joh96b]. Patient [SRR95]. Patients [KR96, LI98a, TFAH94, TDW95]. Pattern [GS98a, Lit93, KY86]. Pattern-Mixture [Lit93]. patterned [BK84, HN87]. Patterns [Bac90, BB90, CS91, DLB91, KO96, YK96, YP92]. Payment [PH90b]. PCP [RG94]. Peaks [Pan97]. Pearson [Akr88, ELR87, EL90, Goo96, Leh93]. Pearson-type [Akr88]. Pediatric [DXP92, DA96]. Penalized [PO94]. Penalty [GH95, SKF98]. People [EGG95]. percentage [JJS83]. Percentages [DF79, MYZ67]. percentiles [TT86]. Performance [GHMW94, HIT86, KAT91, PH90b, Sim87a, VY80, Las94]. performing [MP83]. Period [Bac90, DA96, SS90a]. Periodic [GHL96, LZ98, LR97b]. Periodograms [Ste95]. periods [NP83]. Permanent [Yam92]. Permutation [BPS96, ES99, FS98a, Ode91, RK90, Sch94a, SP91, Wel90a, PT83, Ros84b, Tri84, WG88]. permutational [MPS88]. Permutations [Ste90]. Permuted [LR96]. Person [GNK96]. Personal [NFL92]. Perspective [BCJ96, Hay94, Lju93, Wei95]. PES [EKWS91, MS91b, MS91c]. PET [OO97]. Peter [Bow87]. Pharmacokinetic [GBJ96, Wak96, WB96]. Pharmacokinetics [HDG97]. Phase [Hin71, Sit97, TS96]. Phenomena [LR97b, Sam93b, Fur46, Gei48]. Phi [EL90, ELR87]. Phi-Squared [EL90, ELR87]. philosophy [Jay79]. Photon [Wei97]. Physiological [GBJ96]. Piecewise [MPCB95, TZ81]. Pilot [CS96b]. Pine [RC94]. Pitman [GS89, MKSB90, RHS93]. Pivotal [FHJW96]. Placebo [GM96]. Placebo-Controlled [GM96]. Placement [AGM98]. placements [OW82]. Plackett [ML94]. Planning [AGM98]. Plans [FMR62, SB81]. plates [Cha88b]. players [Isr81]. Playing [Cre82]. Plebiscite [EV92, RSV95]. Plot [Hol95, MS92, MB88b]. Plots [BVT96, BH96, Coo94, CCD98, EM90, Gal94, Nai84, SS86]. Plurality [Est92]. Point [AF97, BH93, BR98c, CS95, CHP96, DR98, DLB91, Hös94a, Lon67, MH94a, MY91, RC94, Ros93a, SW95, SR94, SK93b, BS87, Bor87, MSW^{+}83, Wor74, YZ88]. Points [dE97, Fun93, Nau65, Rv90a, Rv90b, JJS83, Sim89]. Poisson [AY82, Bir54, Bol95, Bre90, CS95, CM97c, CZ77, CM85, Coo91, CHP96, Dea92, DB97, Efr92, HT81, KY96, Law87, NC99, VM96]. policies [Fox74]. Policy [Ano91, Kra90, Moy99, Nor90, Zel92, Sto89]. Poly [BS93]. Poly-Weibull [BS93]. Polychotomous [AC93, KBS97, BK85b]. polygons [Sco85]. Polynomial [AC97a, AM92, FHW95, Hal73, HW92a, Jay96, Rup97, PT83, Stu82]. Polynomials [SG96a]. Polytomous [For92, Hir92]. Pooled [PSS99]. Pooling [GS95, LTP94]. Poor [WT90b]. Population [ALM97, Bac90, BF92, Buc95a, CCG^{+}97a, CCG^{+}97b, Cre97, DKW97, EKWS91, FG91, GBJ96, GMS97, HS98a, LT94b, MV91, MR97a, MR97b, MS91b, MS91c, NSS97, RGZ95a, RGZ95b, RZ98, Sch95b, SS90a, SS95b, SF97, TL95, VL96, Wak96, WB96, WS96, Wol95, WC91, AK86, Cha86, CSC86, GM86a, Hil70, Isa74, KR82, Lin85, Lou84, MS85a, Mey87, NS93, PWM92, Raf87, RC81, SS82b, Smo82, Tam88a, Tuk85, Val85, Wil84, Wri83, EKT89]. Populations [Bir54, BBH94, CDW93, HS94, KBB93, LW91, Pol91, Rig97, Wan96, Ala84, Boo86, Joh78, MT81, MT82, SS85, Sin86]. Positive [DM95, BS84b, GA82, MZ89, Sha85]. Positively [SC97a, SC97b]. positivity [BJM81]. Positron [MO98, WEST91, VSK85]. Possible [BGH97]. Possibly [GS91b, SLS97]. Posterior [Che94a, Erk94, Fah92, Joh92, LR97c, PSS93, Sev91, WT90a, OB88, TW87, TK86, VW87]. Postrandomization [Roc95]. Poststratification [Val93]. Potency [KCHH93]. Potential [GJ94, KY86]. poverty [Sen86]. Power [BHS98, Bly93, CR84, DKMO89, EMM95, KW97, Kun90, LLR95, MS98b, MNS99, MLRR92, Ode91, RV90c, Car82, FW88, Gle85, Gro81, MS98a, Rea84, Tad80, TR87, WW87b]. Powerful [Tan94, Ber89, BM83]. powers [CG82, WA67]. pp [CMT55]. Practical [RW97, SCW96, SCW97, YB97]. pragmatic [Sin86]. Preaveraged [AM92]. Predict [Coo91, Man90]. Predictants [BS94]. Predicting [CKLH92, CWY97, GHCF94, HZS^{+}90, The93]. Prediction [Ber90a, BP96, Ber91b, BBC98, BH98, Bre92, BK99, Car82, Chr99, CDP96, CMMY91, DKS97, DH99, DLW95, GM94, Haa95, Han99, HJ92, HA94, HD97, KL96, LKCH99a, LKCH99b, Lee88, MP96, OKS97, PJ94, SC99, Ste99, TS90, WKY93, WW91b, AP89a, Bha81, BHT88, Efr83, Efr86b, Gui83, Har85, HA95, KA86, LP86, Raf87, Rei84, Roy86, Sch87a, Sti85, Val85, Yak85, Yan84, Zel86]. Predictions [CS95, DDV99, BD89]. Predictive [CO90, CS91, IL94, KO96, JG83, Las94, LF82, Len88]. Predictor [Ver93, BK85b, Nai87b]. Predictors [CS90a, SRR95, FH81a, FH81b, KY85, PH86]. Preliminary [GLG92]. Prepivoting [Ber88b]. Presence [AC95b, Car94, Fon92, JMS97, LTP94, RRZ95, Sil96, Wan92, Bow87, NM83b, Smo82, Sto86, TW89]. Present [Pol91]. Preserved [Hei95, Kor81]. Preserving [CGGM98a, CGGM98b, Joh98a, MT98, SHL98, BLE88, Hoe84]. Presidential [Moo98, Ket97]. Pressure [HL94]. Prevalent [HW95b]. Preventive [BHB96]. prevision [Gol83b]. Price [BG91b, CG92, DN98]. Prices [CG97]. Primary [BHB96]. Principal [BR92, HS89b, LT94a, Li92, Sch91b, Flu84, Krz79, Krz81, SHS86]. Principle [Bjø96, Coo98a, Coo98b, Li98b, Tar99, Ter90, God82a, Sch81, Sch82]. Principles [CMT54]. Prior [Cla96, DJ98, Esc94, GBJ96, Gus96, KW96a, LWW91, Lav91c, RH96, SHBF94, SY95, Hil70, IL91, Tei65]. priori [HT81]. priority [MS95]. Priors [BCJ96, CW93, Cla96, CV92, DG95, FS94, GS99, GB98, HC96, KM83, MR97a, MR97b, Ye93, BB89, Eub86, OB88, Oak85, Shi84]. Privacy [Lju93]. Probabilistic [GMD95, Gel94, Mer94a, Mer94b, SW90, Dic83]. Probabilities [AG96, CWY97, Ede90, GR96a, GFT^{+}93, GW91, He90, Joe95, Lau92, Li95, Mee90, Pes93, Riz92, Ros93b, Sta91, WK94, BE85, CGM83, MPS88, OB88, Sch83b, VW87, WF82]. Probability [BF90, BS98, GKB98, GJ84, HPS98, Hoe63, KO95, LPT91, Roj96, Sär96, SK93a, SMD93, SO94, Tuk49a, WK94, Xio95, Agn85, CF87, DZ82, FM83, HL81, JTB80, JTB81, KW88, KP83, Kim86, Kin85, MT49, Nai84, Sch84b, Sti86, VM83, Whi81c]. Probability-of-Exposure [SK93a]. Probing [VY80]. Probit [GHCF94, GWG98, LK92a, RK96, SA92, Ame74, GHP87, MS88b]. Problem [Akr91b, Bro97, DKR95, EL90, JK96, LLT54, Liu94, LWT94, PR90a, PSVS92, QTC96, SC96a, SM90b, Tan94, Vel98, Wel90b, WHGA96, AA87, AG85, BN74, CB87, Die82, Fab91, Fin93, GC82, HPE81, JW88, Sam85b, Tam88b, TJW88]. Problems [BH93, BJB95, Buo90, CM91a, CC96a, CS92a, DH91, EL97, GJ99, GSTM92, GK93, JMV94, KM91b, KLW94, Lav91b, LA93, LTW96, PY99, SCW96, SCW97, WPD98, Zha99, CCE85, CMT53, HN87, Tuk54]. Procedure [Akr91a, DT92, HOS94b, LW90, MBR98, PY99, Sit92, CS86, Cre88, EH87, HPE81, RC74, ST60]. Procedures [BW92, BW94, BD92, GS97, HS93, HH94, HH91, MH94a, MH99, SG96b, Sob91, Sob93, WH90, WC91, CS85, Fin93, GR93, HM88b, HM89, Hog74, IHC84, Isa74, MS85b, Ole85, Ruk84, Tam85, Wat87]. Process [AP96, CZ77, Coo91, CHP96, Esc94, GB98, Haa95, HOLDG98, Luc93b, MLMT94, MG97b, RC94, The93, Bha81, BHS81, Hol87, KK84b, Krä86, Law87]. Processes [AF97, Ber91a, Bir54, BR98c, CS95, DR98, DG92, Hof92, KR93, KY96, LRM96, Liu95, NC99, SK93b, Tel67, Krz87, Mac86, Mac87, Sto88, Tsa89]. Processing [CGGM98a, CGGM98b, Joh98a, MT98, SHL98]. Product [CTC96, Cro97, GS93, Pag90, BB89, Sha85]. Product-Moment [CTC96]. production [GA82]. Products [Bal95, SY95]. Profession [Hun94b, Ima95]. Profile [CJ96]. Profiles [Bal95, BS97c]. Profiling [NGG97]. Prognosis [Gra92]. Program [AGM98, Bak98, HZS^{+}90]. programming [O’G82, Vin69, Gol63]. Programs [Ano90, FIB90, GMWZ99, HT90, Lon67, Wam70]. Progress [Ste94, ZT97]. Progression [Bri92, DLD91, LCG^{+}92, YK96]. Projected [PML98]. Projectile [SM92]. Projecting [DKW97, DXP92]. Projection [Amm93, Dua90, FS81, FSS84, SS95b, YF95, Fri87]. Projection-method [YF95]. Projections [EV92, TA94a, DL88]. Promotional [BG91b]. Propagation [Lau92]. propensity [Ros84b]. Properties [AP99, BAP95, BW94, CW91, CD96, FH81a, FH81b, HH91, KY85, RS94a, SK90, Zha92, AK86, Bue82, FM83, GHP87, Law81, Law82, SL87, Sti87, Val87a]. Property [CG92, Sch87b]. Prophylaxis [RG94]. proportion [WPSL84, vd87]. Proportional [FG99, GR92a, HN92, HR97, MRvdV97, Pag90, Peñ98, RT96b, Ryu94, SDW98, YP99, AK86, LW89, Wei84]. Proportional-Odds [HR97]. Proportions [Amr95, HOS94b, Llo90, SG95b, SK90, WS96, FS87, HY86, SG95a, Val85]. Prospective [CWW95, DA96]. Prostate [MPCB95]. Protocol [Hei95]. Provided [PSVS92, Wil84]. Providers [NGG97]. Provisional [SK91]. Proxy [GBN96]. Pseudo [AS90b, KBB93, Pea61]. Pseudo-Maximum [AS90b, KBB93]. pseudo-random [Pea61]. Pseudolikelihood [SI90]. Pseudorandom [Gav78, Kie78, Mal78, SM78a, SM78b, Sim78]. Public [Ano91, CRS^{+}91, Moy99, Nor90, Zel92]. Public-use [CRS^{+}91]. publication [Bow87]. Punched [McP42]. Punishment [Gro90]. Purchase [AL94, ALJ99, GS93]. Purposive [AGM98]. Pursuit [Amm93, Dua90, FS81, FSS84, Fri87].

Quadratic [CV92, GK93, JS94, Lie94, Rol94, Sla94, KR86, Ste87]. Qualitative [Lam82, MT95a]. Quality [Ano90, BF92, FIB90, GJ94, GHMW94, HT90, JS97b, Kra90, LS93, NGH^{+}95, Riv98, Spe85]. Qualms [Sch85]. quantal [CC83, MS88b, Sch84a]. Quantile [BVT96, DG92, EM90, JSKK98, LaR91, LTW96, OR96, SM90a, YJ98, BK82, Mey87, Pad86, RS82, Yan85]. Quantile-Quantile [EM90]. Quantiles [Cha96, HK92, MBR91, Ros95a, RV90c, Sat95, SM92, OB88]. Quantitative [HTP95, LS87]. Quasi [Bre90, CS90a, CM98, Efr99, FHW95, GW88, Goo94, Jun96, Kul95, Rag93, Sar89, SS94, Sta94]. Quasi-Dependence [Goo94]. Quasi-Empirical [Rag93]. Quasi-Independence [Goo94, Sar89]. Quasi-Likelihood [Bre90, CM98, FHW95, GW88, CS90a, Jun96, SS94, Sta94]. Quasi-Linear [Efr99]. Quasi-Residuals [Kul95]. quasiunimodal [OB88]. Question [Fel72]. Questionnaire [RG95]. questions [APS81]. Quick [FS87]. quota [Kin85, Kre86].

R [Bue82, Hoa73]. Racial [YK96]. Racial/Ethnic [YK96]. Radial [Gal94, MG97a]. Radiation [PSVS92]. Radiotherapy [CNF95]. Rail [Sho93]. Railway [Ada93]. Rainfall [KP91]. Raking [Ski91]. Ramsey [Thu85]. Random [AP99, Amr95, AL95, BD91b, Che94b, Cro95, DKS97, dE97, FM82, Gag43, GHCF94, JTC99, KO96, LB94b, MS92, Pai97, RS96, SC96a, SLS97, SW97, Tad80, VL96, WKY93, Wan98b, Yas95, ZWWB96, CMS76, CMS87, Che71, CR89, Col87, DH93, Dyk82, Fat83, FD85, Hoe63, KH84, Khu81, KR76, Kot89, KP81b, KP82, Mar84, MZ89, NM83b, Pea61, PH86, Poo74, Pot87, Rei82, Rei84, Rel72, Rob87, Ros42, RS86, SCM91, Wor34, YD84, ZK91, dJGM83]. random-difference [Wor34]. Random-Effects [Cro95, GHCF94, VL96, ZWWB96, Fat83, Rei82]. Randomization [Raz90, Rom90, Hoo89]. Randomized [BC90, Bak98, Lju93, RG94, SO94, Tam81, CGG85, KR84, O’H87, Poo74, Sen74]. Randomly [CLWZ95, HP92, KS97, AA87, Alb88a, SM84a, Wan89]. Randomness [Bar82, SB90, Ali89, HM88a, Stu56]. Range [Fin90, Hay90, Hoe84, LR97b, Nyc91, SM81]. Ranges [OB88]. Rank [Akr90, Akr91a, AAB97, Bar82, BH94, BD92, CMNS99, CM97b, CM91a, CD96, Da̧b90, DM93, DM96, GL92, Gri92, HM88a, HMO97a, HMO97b, Hös94a, Kai94, KL99, KW96b, KY99, LD92, LS93, Mar92, MM95, MSH90, PR90a, Tab90, TA90, Tho91b, Won84, Yan97, AA87, Alb88a, BM83, Boi86, Boo86, BS86a, FHO87, GH87a, Hil74, HC84, KW85, KR84, MW81, MW82, MH82, Rin83a, Rin85, TA89]. Rank-Based [DM93, DM96, Hös94a, MSH90, HM88a]. rank-censored [KW85]. rank-frequency [Hil74]. Rank-one [Won84]. rank-score [HC84]. rank-transform [TA89]. Ranked [AD90, BW92, BW94, KS94, SS88]. Ranked-Set [BW92, BW94, SS88]. Ranking [Fon92, BD88, FV88]. Rankings [BW94, Tar87]. ranks [CSS89, HM83, Ros84a, ST60]. Raphson [LB94c, LB94d]. Rare [SH92, CSC86, KP83]. Rasch [LCG91]. Rate [AMCS98, BRR92, BPR99, DA96, GT98, GMCM96, HL96, HW95b, MN93, MZTT98, RZ98, DD88, Efr83, Efr86b, Gla80, Hay86, SM83, SW83, O’N94]. Rates [AKM94, BR95, BBH96, CNWW98, GS95, KR96, Min98, PH90b, RG94, Ros95c, WCXG97, YL97, BM85, BPN87, CH88, Ros95b]. Rather [DB97, SRR95]. Rating [BP94a]. Ratings [Ueb93]. Ratio [AME96, Agr98, Bar82, BB98, DD90, DKR95, FS95b, GR92b, HMY97, Jen91, Li95, LQT97, Lie94, MPS98, Mur95, Pen97, Ski91, SW98b, SW98c, TL97, Tin65, VW95, VHE91, VM96, Wan96, WM98, WKL95, Xio95, Yas95, Fra84, KC93, Kot89, MPG85, MPG86, Rob87, SL87, SW86, WR84]. Ratio-Based [Mur95]. Rationality [SSC94]. Ratios [Mar65, Shi91, WC92, Flu85, KM83, Loh86, Mal82, Wij85]. Raw [HYP97]. Ray [EP94]. Reach [LS91b]. Real [DR91]. Real-Time [DR91]. Really [CB94, Ben83]. reasonable [RHS93]. Reasoning [KSS96]. Recall [RMS90]. Recapture [Pol91, CM86, KMT99, MSW^{+}83]. Receiver [Llo98]. Recensoring [LLE93]. Recognition [GS98a, PJT96]. reconciled [BHT88]. Reconciliation [Goo96, SR94]. reconciliations [Sen86]. Reconciling [CB87]. Reconnaissance [Hod94]. Reconstruction [Kol96]. Reconstructions [Wei97]. Record [BR95, Sol95, Suz85]. Records [NFL92, TB85]. recover [CS89]. Recovering [GJP96]. Rectangle [Joe95]. Recurrent [HL96, WC99]. recursion [CKC83]. Recursive [Gri94, KS96, Abr88, SH81]. Recycling [NG94]. Red [RCH94]. Redescending [BH94, HRR81]. Redistricting [GK90]. Reduced [SS97]. Reduction [CL92a, Li92, Kan88, Li91]. Reference [Cla96, Goo94, GK93, KW95, LRR91, QH98, SY95, Ye93, BB89, CGG85]. Referendum [CH91]. Refinement [Bjö78]. Refinements [PWM92, Ber88b]. Refining [Ber90b]. Reflections [Nor90]. Reform [DN98]. regarding [WR84]. Regeneration [GRS98, Joh98b, MTY95]. Region [BO97, HW92a, PC96, SS96c]. Region-Based [BO97]. Regional [Oeh93]. Regions [Ber90a, CHP96, FHJW96, MZ97, Mee90, Tho90a, WCH98, FH89, FH90, IS87, LL81]. Regression [AL94, BRMZR97, BL92, BR94b, BRB76, BL93, BVT96, Ber77, BZ92, Bre92, Bre90, BGH93, BS94, CL92a, CRS99, Cha71, CMNS99, CJ95a, CJ95b, CTC96, CW94a, CM98, Cho92, CM97c, CRS^{+}91, Coo79, CW90, Coo94, CN94, CW97, DwSH94, Dea92, DB97, DC77, DBB^{+}94, Dor94, Dua90, DS97b, Dwy45, ES90a, EM92, ES90b, ES93, Fan92, FG94, FG95, FHW95, FS95a, FOS97, Fer98, FKYZ99, FL95, FS95b, FS81, FH78, FH81c, GW88, GMP97, Gey91, GS91a, GLG92, Gle92, GF96b, GK93, GL96, Gu90, GP92, HH90, HW91, HT97a, HT97b, Hal73, HM88b, HM89, HT90, HHM92, HW92a, Haw93, He99, HZ96, HZ98, Hin71, HR97, Hun94a, HZ90, Jay96, JTC99, JDP94, Jon96, KR96, KM91a, Koe99, KST95, KBS97, Kul95, KW97, Law87, LC95a, LT96a, LWY93, LS91a]. Regression [Lin91, LW92, LY93, LCWH97, Lit92, Liu96, LS99, LJK93, MW98a, MW98b, MY91, MEGN92, MSH90, MS99, MNS99, MT95b, NCC96, Nyc91, NGH^{+}95, OH99, PY99, Pol84, Por97, PM99, RMH97, RSY93, Raz90, RRZ94, RRZ95, RR95, RW98, Rol94, RT96b, RRS98, RH99, RHV99, RSW95, Rup97, SW95, SS97, Sch94a, Sch94b, Sch95a, Sev96, She93, She97, SAK90, SRC92, Sit97, SH90, SC80, SK97, SL98a, Spu99, SC92b, SS91b, SL98b, SB95, SCW96, SCW97, SR92, Str93, SGP98, Tab90, Tho90a, Tsa84, Vel98, VY80, WS94, WR95, WWZtO97, WJ92, WLW89, WT90a, WCR94, Wie98, WK98, Yam92, Yan97, Yan99, YP99, YL97, YJW95, YJ98, ZB86, Zha99, dHY93, Abr88, AP87, Bie83, Bin82, BF83, BF85, BP86, CD89, CGG85, Cas85, Cha84, Cla87a]. regression [Cla87b, Cla89, CH93, DW87, Dyk83, Efr86a, Efr88, Eri81, Esc86, Eub86, FP84, Gil84, GL83, GY93, Gut74, HB88, HHM88, HS89a, HW86, HMP87, Hod88, Hub83, IF82, JSG88, Jen86a, Jen86b, JG83, JV85, Jun96, KK84a, Kan88, KW85, KC93, KSY85, Krä86, KD87, KW82, KSC89, LQ85, Lea81, Li91, LZ89, LT88, MB88a, Mül87, Nai83, Nai87a, Nai87b, OYR86, Oma81, PH74, PC84, Rou84, RC81, Rup85, Sam85a, Sam93a, Sch87b, Shi84, Sil84, SS86, SS87b, Spo82, Sta89a, Sti85, SC85, Stu82, Swe84, TT86, TZ81, TJW88, TV85, Uus83, WA83, WM80, Whi81a, WM81, Whi81b, Whi81c, WK88, Wol87, Yak85, YZ88, dJdWW88]. regression-simple [Esc86]. Regressions [Ber50, Coo96, DM90, GLT79, KP90, VY80, Kob85, Kob86, PP88]. Regressor [CO90, But84]. Regressors [RRZ94, PP88]. regret [Mar88, Nai87a]. Regular [BH94]. regularity [Vos93]. Regularized [BC96, Fri89]. Regulation [LR96, Liu94]. Reinfarction [OMF^{+}93]. Reinterview [BF92, SG96b]. Rejection [LCW98, KP81b, KP82, SS80]. Rejoinder [AIR96, BR98a, CCG^{+}97b, CGM98b, CGGM98b, Coo98b, Efr96, Fay96b, GKL98b, HT97b, LKCH99b, MW98b, Mer94b, MS91c, RGZ95b, Rao96, Rv90a, RHV99, Rub96b, SM78b, SBD97b, SW98c, TW98b, BS87, Bue82, CB87, Daw82, Efr87, Efr94, Fab91, Goo91, HW94, Las94, LF82, Li91, PT82, SCM91, WMR91]. Related [Ber91a, Fer98, FMSP90, Fon92, LCG91, MO98, OK96, Sam93b, BD88, Bha85, Gla80, JV85, TV85]. relation [HS89c]. Relations [KSJ94, CKC83]. Relationship [DMLC97, HH94, KW95, SS90a, TDW95, Kun80, Thu85]. Relationships [GMP97, GR96b, HH91, PSS93]. Relative [GC95, Joe89, KCHH93, Wel90b, BR83, CD89, Cha88b, GG87, KD87, Ros84a, ST60]. Relatives [BGH97]. released [Hwa86]. relevance [Tei65, Tuk79b]. Relevant [HA94, HA95]. Reliability [AMCS98, CS96a, DC94, Gle92, KY96, LS86, Mei94, dCP99b, BJ74, WS89]. reliable [Ger82]. Remarks [DG95]. Remedian [RB90]. Remedication [PHL98]. REML [CD95]. Removal [BR98c, Pol91]. Removing [HW91]. renewable [Ros84a]. Renewal [DwSH94, GM86b]. Repair [AP99, PHS94, WS89]. Repeat [CG92]. Repeated [Agr97, Akr91a, AA94, AHB97, Bow87, Car94, Con92, JD91, JR96, JTC99, KW96b, Kus98, LD92, Lit95, Luc93b, MLRR92, OR96, RRZ95, Roc92, RRS98, SCC98, TA90, Tho91b, Tsi82, ZDL96, Abr88, Con89, HW86, KR88, LLS87, LD91, LB94c, LB94d, Rei82, SW82]. repeated-measurement [Rei82]. Repeated-Measures [JR96, KW96b, Lit95, KR88]. Replacement [BBC98, AK86]. Replicate [Chr91, SC96b, Vos93]. Replicated [DLB91].

Replicates [MNS99]. Replication [SCC98]. reply [BH88, Bor87, BF85, DG93, DH83, Efr96, Gew82, God82b, HHM88, HL81, Hol86, Kit87, LL87, LV88, MB88a, RHS93, Sha82, Sim87a, TW87]. Report [Ano99c, CCG90, LHA96, LCK97, LWG93, LHA94, LHAL95, Wam70, CMT53, CMT55]. Reported [GJDT98]. Reporting [CSC86, DXP92, Har90b, Wan92, BD89]. Reports [RBD93]. Representation [MFS84]. Representations [BM92]. Reproductive [SGC95]. Reprojecting [GT98]. Resampling [DC94, RW88, Sit92, Tro95]. Rescue [SMP95]. Residential [CG92]. Residual [CCD98, LR92, SG97, Chi74, Jen86b]. Residuals [BAP95, CHW92, KS96, Kul95, MSH93, PS86, Ran84]. Resistant [Bra90, GR92b, Jac93, GH87a, HIT86, JV85, Sim87a, TV85]. Resolution [BR94a]. Respect [GS94, Hsu82]. Response [Agr97, AC93, BD97, Buo96, CO90, CM97a, Coo96, Fyg97, GV96, GJP96, HYP97, Hir92, HB99, KPR91, KC96, Lan96, Lee96, LCG91, Lju93, PSVS92, SG97, CC83, Man88, O’H87, Poo74, RRS98, Sen74, Tam81, Whi81c]. Response-Dependent [GV96]. Responses [AHB97, BZ99, FLZ96, HZ98, LA94, RSV95, Zha98a, Hab82]. responsibility [Tuk79b]. Restoration [AGP91, O’S90]. Restricted [CD95, Kar87, KT95, MT95b, NSS97, PF97, Zuc98, ACK87, Car82, Dyk83, GH88, HP88, WR84]. restriction [Lee87]. Restrictions [GLG92, KT95, RH96, Shi91, HHO88]. Results [Ada93, CV95, LB94b, NM98, Whi89, BHS81, CHL82, DH83, Fab94, FP84, Lea84, Pri74, Tam88a, Tar87, TV85, Whi92]. retransformation [Dua83]. Retrospective [GV96, IT94]. Retrospectively [Wan92]. Returns [FP95]. Reverse [GS95]. Review [Ano98d, BF93b, CC96b, Dia78, Gei48, Gol63, Goo92, Hoa73, KS96, LKM98, Lit92, Met50, TW46b, Tuk47b, Tuk47c, Tuk47a, Tuk49a, Tuk49b, Tuk52, TW52, Hog74]. Reviews [Ano95b, Ano95c, Ano96a, Ano96b, Ano96c, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano99a, Ano99b]. Revising [YP92]. Revisions [Mar86]. Revisited [Coo98a, Coo98b, Li98b, BD81a, BC82a]. revival [Rom88]. Reweighted [DS91]. Reweighting [CN94]. Rheumatology [BS97b]. rho [GR81]. Rhythm [PF97]. Richardson [BY88]. Richness [HAS97]. Ridge [SC80, Cas85, Eri81, Law81, Law82, Lea81, Oma81]. Right [GL96, Gür96, LWT94, Qui98a, Gui88, Pad86]. Right-Censored [GL96, LWT94, Gui88, Pad86]. Right-Truncated [Gür96]. Rigid [PC96, Riv98]. Ring [DH91, Lav91b]. Risk [BRR92, BSB98, BF90, FG99, HM91, HTP95, BR83, GG87]. Risks [FW96, HW95b, Mei94, Pep91, GA82]. Robert [DH83]. robin [Won82a, Won84]. Robinson [Kot89]. Robust [Akr91b, Amm93, Atk94, BY86, CDW93, FKYZ99, GL87, HR94, Hoa73, LR95a, LLT89, Lav91a, Liu96, LM93, MH90, MSH93, Mor86, Rin83b, RS94b, RFB97, Rot83, RB90, Sil92, SRSD94, Tam95, TR95, Vel95, VY80, VFR97, WS94, WG88, WM90, WZ97b, Wie98, WK98, WR94, Cha86, Gas85, Hog74, LW89, MY95, MZ89, PMF88, Roy86, Stu82, Tam88a]. robust-type [Stu82].

Robustification [EMM95]. Robustness [BHS81, FOS97, HSP90, KC93, MS85b, Tsu84, Ber82, God82b, Lam82, Ols74, Pie87, Tsu86]. Rodent [HZS^{+}90]. Rogers [Hoa73]. Rohust [LRM96]. Roles [SB97]. Ronald [HL81]. Root [CJ96, Dwy45, MBL98, NP95, SL93a, TR97, jW97, Dav82]. Roots [GHL96, YR95, DHF84]. Rotation [CLY92, Cha88b]. Roughness [Hua97]. roulette [Eth82]. Round [Won82a, Won84]. Rounded [MYZ67]. Rounding [DF79]. Rows [HW90a]. Roy [Dav82]. Rubin [Efr94]. Rule [Pen97, Ye93, Efr83, Efr86b]. Rules [Bra90, KW96a, MG97a, Vid98, WSLP90, CSC86, HIT86, HP88, Sam81, Sim87a]. Run [Lou96, Sch83b, CJ88]. Runs [FK94, Lou96, McW90, RMS90].

S [Bow87, STM^{+}65]. Saddlepoint [BS98, GR96a, GJ99, Lie94, WBB93, ER86]. safe [Cre82]. Salamander [Shu97]. Sales [CG92]. Sample [Akr91b, Amr95, BW92, BPS95, CL92b, Chr91, CM91a, Cla96, DKR95, EL90, HMO97a, HMO97b, Joh96b, Kai94, KM91a, KSW92, Kot94, LRR91, LTW96, LWT94, MSL93, Noe87, PR90a, PWM92, RS96, RN97, Riz92, RK90, SW95, SG96a, SG95b, Sob91, SN90, SCW96, SCW97, TF90, WH90, Whi81b, Whi81c, Yan84, Zie90, AA87, APS81, BE85, BC82b, CT98, CHL82, CS86, CC83, DD88, Die82, FM83, GR81, GC82, GHP87, Gui83, HN87, Lin74, Lit82, MW81, MW82, MT81, MT82, O’N84, OW82, Rea84, Sch86b, SS87a, SG95a, SHS86, SW82, Ste81, SS88, TR87, TT84, VJ88, WL84, WM80, WM81, Zac86]. Sample-Admissible [Sob91]. sample-size [GC82]. Sampled [LW91]. Sampler [Cha93, KT94b, Liu94, RT92, ZM95, KT94c]. Samplers [MTY95]. Samples [AK92, BBH96, BW92, BW94, BR98a, BR98b, CH95b, CRS^{+}91, CS96b, FZ98, Fel74, GK95, Hau92, Kne98, KS93, KS94, LT97, Lan98, LTP94, Owe94, RS98, RSY93, Ros95a, Tel67, WW98, VS93, BK86, CM84a, Fay85, GL87, LaR86, LL87, MB82, RW85, RS86, SM84a, ST82, ST60, Tam85, Wei81]. Sampling [BBC98, Buo90, BS98, CLY92, DS92, DD90, Dor94, FMR62, Fis79, Gag43, Gav78, GHRPS90, GS90a, GSTM92, GS99, GR96b, GV96, Hei95, HC96, KM91a, Kie78, LKM98, LS91b, LNL95, LCW98, Mal78, MV91, MS96, OB93, Per97, RS96, Ros93b, Sär96, SM78a, SM78b, SS99, Sim78, Sit97, SO94, Tam95, Thé99, Tho90b, TS96, Tuk50a, Ver93, ZT97, Zha96, AK86, Bet89, Bow87, BS86b, Chr87, CMT54, CSC86, DD81, DD82, GM86a, God82b, Kin85, Kot89, Kre86, MS85a, MPS88, Rob87, Roy86, SA86, Tad80, Tam88a, Tar86, Tei65, TR87, Tuk50b, Val87a, Val87b, Wol84, Wri83, WHH88, ZK91]. Sampling-Based [GS90a]. Samuel [STM^{+}65]. Satellite [BR92, FS94, NT95]. Satisfaction [BZ99]. Saturated [BO95]. Savage [VW95]. Savage-Dickey [VW95]. Scalar [Ahn97]. Scale [AL95, CF95, Gle93, Har92, MY91, WM90, CC83, Esc86, MZ89, Swe84, WG83, YZ88, Zac86]. scale-change [WG83]. Scaled [CW94a]. scales [KR82]. Scan [HL97, NW97, POH97, BE85, GW84, Gla89, Nau82]. Scatterplot [CLNL87, CM84b]. Scheduling [LZ98]. Scheffé [Nai83].

Scheffé-type [Nai83]. Scheme [DC94, Joh98b]. Schemes [APR92, Chi74, Sha85]. Schervish [Oak85]. School [AK92, MSMP92]. Schwarz [KW95]. Science [Zel92, Jay79]. Scientific [Hur85a]. Score [NC99, SS95a, ACK87, HC84, MH82, Ros84b]. Scores [BH96, GS98b, Ste94, Tab90]. Scoring [HTB94]. Scorings [KSW92]. Scott [HHM88]. Screening [BC90, Bak98, GJ94, HTP95, LTP94, Par93b]. Sea [LR97b, PM97]. Search [CGM98a, CGM98b, DSM98, KKT98, MBL98, Zha98b]. Searching [GK95]. Seasonal [FMSP90, GHL96, PH90a, TR97, BPN87, DHF84, HT82, O’G82, Sch81, Sch82, Wat87]. Second [GT91, Tuk47c, RW85]. Second-Order [GT91, RW85]. Secondary [BHB96, WSJ92]. section [GA82, Sto82]. Sectional [Wan91]. sections [NWGP84]. Secular [HL94]. Seemingly [BG91b, Bin82]. Segmentation [BO97, Joh94]. Seismic [PB92]. Select [MNS99]. Selected [CB94, WK94]. Selecting [GS94, Wil84, Tam85, TA94b]. Selection [BP96, BZ92, Bre92, CKLH92, EL97, FMR62, GLB98, GMCM96, HZ90, JMS96, KW96a, KW97, LRM96, LA93, MR94, MBR98, NP95, OP96, OR98, Riz92, RFB97, SH92, Sha93, Sha96, Sil97, Sob91, Wan95, Wei95, Ye98a, YD84, Zha92, ZL95, BD88, CS87, DD81, DD82, DH86, EH87, FM83, FH86, FE88, GC82, HG81, Kim86, MB88a, Ros84a, SO83, Sta89b]. selectivity [Mil92]. Selector [RSW95]. Selectors [PM90, jW97, DM88c]. Self [BHB96, GJDT98, Kit98, Oak85, Tar99, WK94]. Self- [BHB96]. Self-calibrating [Oak85]. Self-Consistency [Tar99]. Self-Organizing [Kit98]. Self-Reported [GJDT98]. Self-Validating [WK94]. Sellke [Bor87, Bor87]. SEM [XLR91]. Semiconductor [HOLDG98]. Semilinear [ES97]. Semiparametric [CWY97, For94, GMP97, HMM98, HH96, Hun94a, JTC99, KS97, LC95a, LQT97, LCG91, NCC96, NCDF96, PS94a, RRZ95, RR95, RCL96, RT96b, RRS98, STR97, SKC94, SS94, SD97, WWZtO97, YP99, Zel93, ZLRS98, Hec88, OS87, Sta94, Wan89]. Sen [AML95]. sensitive [Tam81]. Sensitivity [BPS96, CH91, Gus96, Lav91c, QH98, RK90, Sev96, Wor74, Lea84]. Separate [BHB96, Ruk93, Loh85]. Separating [GS97, KW88]. Sequence [KM91b, LNL95, LNL99, YK96, DD81, DD82]. Sequences [CLWZ95, Sch83b, Yak85]. Sequential [EGMP93, GY93, HG81, HE83, HTP95, JD91, JT97, KLW94, LD91, LD92, LB94a, LC95b, LCW98, LC98a, LC98b, STR97, SMD93, Sha83, TS96, Xio95, FE88, Fin83, KP83, MS85b, Ros84a, SW80, Tak81, Tam85, Wu85]. Sequentially [FR90, HPS98]. Sera [LTP94]. Serial [AP96, Ber91a, Hei91, ZDL96, HM88a, PH74]. Series [Ada98, CMW97, Cha90, CL95, CW91, CLM93, CB90, DR96, EH91, GNK96, GM94, GS90b, KST98, KL90, Koo97, LK91, LMR96, LS91a, Li96, LT98, LL97b, Luc93b, MFS84, MP96, MSL93, Niu96, SAK90, SK91, The93, TA94a, Tuk52, VHE91, WPK99, dP98, Ali89, BPN87, Cha88a, Cre88, DHF84, DH86, FH81a, FH81b, Gew82, Gew84, GA82, HT87, Hol87, IS87, Kee82, Kit87, KY86, KY85, LZ89, Nef82, NP83, Nyb86, PB87, SU74, TB81, TA94b, Tsa84, Tsa87, Wor34, ZH87]. Service [GWG98]. Set [BB94, BW92, BW94, Daw95, FHTW97, Haw93, Jac93, KS94, Car82, CG84, GL83, MV85, SS88]. Sets [Ber90b, DD90, HH94, Hol99, Roe90, RB90, Ber88a, BCH95, CH83, Kob86, MSW^{+}83]. Several [Agr98, BB98, CV95, MPS98, MH99, SW98b, SW98c, TL97, Wan96, Kob85, Lin74, MPG85, MPG86, MRW87]. Severini [Sta94]. Severity [Wei92, YKB97]. Seymour [BH88]. Shape [WR94, SS85]. Shaping [Ket97]. Shelby [HL81]. Shift [FNP95, LWT94, KP83]. Shifts [The93]. shock [BS83b]. Shocks [dP98]. Short [KL90]. should [BF83, DM88a]. Shrinkage [BG91b, CKM97, DJ95, KCHH93, Kol96, Vid98, DB83, Geo86]. Shuttle [DH91, Lav91b]. Sichel [AY82, SZJ87]. Sided [BM99, CS92a, CF95, Fol96, HHS97, HY98, Hay90, MH99, Sil92, SS95a, Tan94, CB87, Fli84, KK84a, LL86]. Sides [WR95, KT94a]. Sieve [HR97]. Sigmoidally [Sch84a]. Sign [BLE88, BHS98, BHNO92, CS93, LQ85, ON89, Ran89, Spu88]. Sign-preserving [BLE88]. Signal [BSPK98, CCD97, KA91, Mar86, WA83]. Signed [CJ96, Da̧b90, HMO97a, HMO97b, PR90a]. Signed-Rank [Da̧b90, HMO97a, HMO97b, PR90a]. Significance [Fan96, FL98, Fin90, Fra91, LRR91, But84, PH81, SW82, Tsi82, TW89]. significant [Hay86, TA94b]. Simes [SC97a, SC97b]. SIMEX [CKLS96, WLGC98]. Similarities [HHH92]. Simple [Eas94, Fol96, Hay90, JR96, LCG91, MM93, Sär96, Wel90b, Dal82, Esc86, KC93, Kot89, LF82, MH82, NM83b, PMF88, PH74, Rel72, Rob87, RS86, Sin86, TR87]. simplest [Pic87]. Simplex [RS94a]. simplified [Tuk50b]. simplifying [PB87]. Simpson [Mit91, Sam93b]. simulated [JJS83, Rot83]. Simulating [Zel93, CMS76, CMS87]. Simulation [CS94, DKRW97, DD90, Gav78, Kie78, LR97c, Mal78, Ode91, PS99, SM78a, SM78b, Sim78, SC95, Tei65]. Simulation-Extrapolation [CS94, SC95]. Simultaneous [BS96a, Ber90b, BM96, FS95a, FNP95, Hsu82, HT81, IS87, Khu81, LT88, Mal82, MY91, Nai87b, Sat95, SG95a, SG95b, SA92, SN90, Sri72, Ber88a, BT81, DB83, DL74, FS87, HVS84, KK84a, Kun80, Nai87a, Pie87, Ric82, Sha83, SI85, Uus85]. Simultaneously [LA94]. Single [CFGW97, Goo96, HH96, Sev91, Wei97, Whi89, Fab94, Har81, Vos93, Whi92]. Single-Index [CFGW97, HH96]. Single-Parameter [Sev91]. Single-Photon [Wei97]. Singly [dCP99a, Sim74]. singly-ordered [Sim74]. Singular [Amm93, Mar64]. Situ [ST92]. six [Ols74]. Size [Amr95, Cla96, GLG92, Mil92, Nau65, PWM92, RZ98, SG95b, SN90, BE85, GMP97, GC82, Noe87, NWGP84, Raf87, SA86, SS85, SG95a, Whi81b, Whi81c]. Size-selectivity [Mil92]. sizes [GR81]. Skating [BP94a]. skew [Tay85]. Skewed [Che95]. skewness [FS98b]. Skilled [KR96]. Skin [BHB96]. Sliced [Fer98, Li91, Sch94b]. Sliding [FMSP90]. Sliding-Spans [FMSP90]. slopes [TV85]. Slovenian [RSV95]. Slowly [Ber91a, BG91a]. Small [CHL82, Chr91, CC83, GNSC98, GHP87, HOS94b, HH91, LR95a, MSML97, PR90b, Rag93, Rea84, Sri72, SCW96, SCW97, TR87, TF90, WC91, BOP88, FM83, GR81, Hab88, Vos86, Whi81c]. Small-Area [HH91, PR90b]. Small-Sample [Chr91, SCW96, SCW97, TF90, CHL82, CC83, Rea84, TR87, FM83]. Smaller [Fel74]. Smearing [Dua83]. Smirnov [Mas51]. Smith [DH83, Mar80]. Smooth [DLW95, GF96b, KL97, LaR91, Led94, MZ96, Peñ98, Raz90, Ter94, HS89a, PS82, Yan85]. Smoothed [EL97, Llo98, MBR91]. Smoothers [CGGM98a, CGGM98b, Joh98a, MT98, SC92a, SHL98]. Smoothing [Alt90, AM92, BF93a, BR98a, BR98b, Cho93, DJ98, EL97, FW96, FHMP96, FZ98, GKK91, GMCM96, Gu90, HHM92, HS98c, Kne98, KAT91, Koo97, KW97, Lan98, MT95a, MST97, NNR95, Nyc91, O’S90, O’S91, RS98, Ter90, Tho91a, Tuk52, WJ93, WW98, Wan98b, WCH98, dJ89, DM88c, HB88, HHM88, Nyc88, OYR86, Par93a, Sil84, VW87, WA83]. Smoothness [DJ95, Shi84, Eub86]. Sobel [FM83]. Soccer [RCH94]. Social [Ano91, GMWZ99, Pea91, SI90, Yam90, Fur46, Gei48]. Society [Wal93]. Socioeconomic [GBN96]. Software [AMCS98, KY96, DM88a]. Soldering [MLMT94]. solution [JW88]. Solutions [BRB76, Bjö78, SG96a]. Some [Ada93, Akr90, AKM94, Bue82, Buo90, Clo82, CS92b, DG95, DR96, Dwy67, Fab94, Fin90, Gal94, Gei92, Goo96, Hou95, HH91, KM91b, LH91, MS98b, Pot87, Pri74, Riv98, RRZ94, SC80, SK90, Stu82, Tam88a, Tam85, TC72, Tin65, Tuk50b, WM80, WM81, Whi89, Whi92, Wol86, WBB93, FM81, FP84, GKR88, HIT86, Hog74, Las94, MS98a, Nai84, Noe87, Rog88, Rom88, Sen86, Sim87a, Val87a, Wam70, Wol84]. Sort [Ver93]. Sounds [BSPK98]. Source [BR94a, GMWZ99, Var96]. Source-Destination [Var96]. sources [CM86]. Southern [RC94]. sozialen [Fur46, Gei48]. Space [ASL91, CPS92, Erk94, GL92, GS98b, Kit98, RC94, SW91a, Kit87, SS87b, Sto86, dJ89, DH91, Lav91b]. space-efficient [SS87b]. Space-Time [RC94, Sto86]. Spaces [CD95, Lav91a, Moo81]. Spacings [Kai94]. Spans [FMSP90]. Sparse [HT97a, HT97b, Por97, Koe86, Sim86, Zel87]. Spatial [AF97, BK99, BR98c, Chr99, CH92, DR98, DLB91, Gel94, Han99, HL98, HAS97, HOS95, HOLDG98, KP92, KO96, KD87, LKCH99a, LKCH99b, Mer94a, Mer94b, MST97, POH97, SG92, SC94, SC99, Ste95, Ste99, WEST91, YL97, HW94, Ste87]. Spatial-Temporal [KO96, HW94]. Spatially [BR94b]. Spatio [Haa95, WCXG97]. Spatio-Temporal [Haa95, WCXG97]. Spearman [GR81]. Special [Goo94, CGG85, Lee88]. Species [BBC98, BF93b, HAS97, CF87, DH83]. Specific [AKM94, BRR92, GFT^{+}93, PH90b, Sch91b, WM91]. Specification [DM90, GMD95, HN92, Nef82, Ryu94, Ter94, Pud82, Thu85]. Specificity [QH98]. Specified [Cha91, LWW91, RS90, Wie98, OB88]. Specimens [Hei95]. Spectra [SRSD94]. Spectral [Ada98, Efr98, PO94, SS90b, GW86, MG84]. spectrum [Hur85b]. Speech [PJT96, SB90]. Spherical [MG97a, FOS95]. Spherical-Radial [MG97a]. Spike [Bri92]. Spline [ASL96, BR98a, BR98b, Cho93, FZ98, Gu90, HS98c, HK92, Kne98, KAT91, Lan98, RS98, Tho91a, WW98, Wan98b, NWGP84, Sil84, WA83]. Splines [DS97b, Gra92, HK93, LS91a, LW97, SH90, WW83, Las94, Nyc88, VW87]. Split [LC93, MS92, RG95, MB88b]. Split-Plot [MS92, MB88b]. Sports [Ste94]. spread [ST60]. SPRT [Mar81]. Square [Dwy45, Ame74, Wil50]. Squared [EL90, HJ92, HP92, LR95a, PR90b, Bha81, ELR87, Fay85, Hab88, HV86, KOS85, LaR86, Law81, Law82, Lee87, Nai87c, Ole85, Ryu94, Sam85a, Tar86]. Squares [AI95, Bjö78, CD91, CW94a, CHW92, DS91, DS97b, ES90a, Gar94, GF96b, Hoo93, HM91, IT94, Lon67, RSY93, RSW95, Wam70, CR82, CGG85, CWR88, Chr87, DWL87, DP89, Dyk83, Krä86, Lea84, PP88, Rou84, RC80, RC82, SS86, SS87b]. Squeeze [SL80]. Stability [SRC92]. Stable [Buc95b, GLT79, CMS76, CMS87]. Stage [AI95, HOS94b, RSY93, Chr87, Roy86, SA86, Val87b, WHH88]. Stagewise [Hol99]. Stahel [MY95]. Stahel-Donoho [MY95]. Stamp [BD97, PH90b]. Standard [BH98, KH84, Wil84]. Standardized [LKC97]. Standards [PH90b, SCCF93]. Staniswalis [Sta94]. Stanley [HL81]. State [ASL91, CPS92, GL92, GS98b, Kit98, PH90b, SSK90, SW91a, Kit87, Zeh88, dJ89]. State-Dependent [SSK90, Zeh88]. State-Space [ASL91, CPS92, GL92, GS98b, SW91a, Kit87, dJ89]. State-Specific [PH90b]. States [LT94b, Ada93]. Stationarity [MS98b, MS98a]. Stationary [BM95, Mau95, PR94, Tel67, Tuk52, Mac86, Mac87, SU74, TT84]. Statisitics [CM91a]. Statist [BK86, Bor87, BC82a, CMS87, Cla89, DD82, Fab94, FH81a, FH81c, HG83, JTB81, Kot89, KP82, Krz81, Law82, LB94c, MW82, Mac87, MS98a, MPG86, MT82, MS95, NS86, Pra82, Rin85, RC82, Sch82, Sim87a, SG95a, Sta94, Wee94, Wes98, WM81, Whi92, YM86, de 83a]. Statistic [AP99, BW94, Ede90, HL97, McW90, Mee90, NW97, POH97, SC94, BE85, Cro74, Gla89, HC84, Raf87, SW82, WHH88]. Statistical [AASM93, Bir54, CMT53, CH92, DDD94, EFS92, FM82, Fra91, Gei92, GR93, GHMW94, Hau92, Kad90, KSJ94, KBB93, KM91b, NGG97, OKS97, RV90c, SJ92, Sto91, Tuk49a, Ueb93, Var88, Wal93, WSLP90, BJM81, CCE85, Dic83, DH93, Fel72, God82a, HW94, LLT89, Tuk79a, VSK85]. Statistically [Gav78, Kie78, Mal78, SM78a, SM78b, Sim78]. statistician [Tuk79b]. Statistics [Ada93, AAB97, Alb94, Ano91, Bre96, Gai96, GK93, Hol86, Hun94b, Ima95, Ket97, KY99, Kra90, Lan96, Lav91c, LRR91, Moo98, Nor90, Par96, Pea91, Pol91, Ruk93, RV98, SC97a, SC97b, SAK90, SK90, TA90, Wan95, Wel90b, WBB93, Zel92, Zel93, Ber88b, Boo86, Bue82, ER86, ELR87, FHO87, Fur46, GW84, Gei48, Hab88, HG82, HG83, Kru88, LL81, MH82, NS83, NS86, Nau82, PS82, RW85, Rea84, Sim86, Ste81, TR87, TA89, Tsi82, TW46a, Tuk54, VJ88, WW83, Wit88, Zac86, de 81, de 83a]. Statistik [Fur46, Gei48]. Statistische [TW52]. Status [RT96b, SK93b, vR98]. stayer [Fry84]. Steel [Spu88]. Stein [Ber82, GS91b, Gut82, Li92, LW90]. Step [DT92, HDG97, SRC92, Tab90, CMST88, Fin93, MH82]. step-down [Fin93]. step-type [CMST88]. Step-Up [DT92]. Stephen [HL81]. Stepwise [HH94, HB99, Tro95, CS85]. Stochastic [AG96, BH97, DF98, FT94, KR93, LT94b, LYY99, Muk96, NGC^{+}95, PS94a, POH97, RS93, She93, TCS94, Tel67, WW87a, Wan96, ZLRS98, Fra84, SW89]. Stochastically [EGGS97, Dyk82, FD85]. Stock [CG97]. stop [DM88a]. Stopper [RT92]. Stopping [SC96a, Ye93, Mar81, Mar88, Sam81, Sha83]. Straight [Mad59]. Strategies [EGG95, LNL95, Tam95, Agr83]. Strategy [BSB98]. Stratification [Jon96]. Stratified [BPS95, KM91a, LWY93, RS96, SCC98, WS96, Bet89, GL87, RW85, Val87a, Val87b]. stratum [KMT99]. Streaks [AASM93]. Strength [Pri96, BJ74]. strengths [CG82]. stress [BJ74]. stress- [BJ74]. Stretches [LMR96]. Stroke [KR96]. Strong [GR96b, HG82, HG83]. Stronger [HW75, Isr81]. Structural [AGP91, Sha86]. Structure [GK95, MSMP92, SG92, SV93, SW90, WJ92, WPK99, WHG98, YB97, CG84, GSS89, PB87, Rei82, Sam93a]. Structured [Seg92, SK91, LV88]. Structures [AS90b, FK99, GHL96, LNL99, Roc92, Vil90, GJ84, Gol86, Lee88, SB87, SB81]. Student [MH94b, PMF88]. Student-Athletes [MH94b]. Studentized [Hay90, SM81]. Studied [HZS^{+}90]. Studies [AGM98, Bak98, BP97, Bin98, BO95, BWC98, Bre98, CS91, CGL93, CWW95, CNWW98, DN98, FS95b, GKL98a, GKL98b, GJDT98, HB99, IRC98, JK96, Jud98, Kad98, LZ98, LB94b, Lit95, OO97, Pep91, QH98, RCL96, RK90, Ros93a, Ros95a, SH92, San98, STR97, WHGA96, ZWWB96, Cro74, DH83, JTB80, JTB81, KMT99, Ros84b]. Study [Akr91a, BHB96, Bre96, Coo91, DM97, DD90, DC94, DMLC97, EGMP93, GWG90, Hei91, Hod94, Sas97, SJ92, SHBF94, TFAH94, BT81, RC81, CMT55]. Studying [AHB97, Joh96b, DM89]. Sub [Bro83]. Sub-balanced [Bro83]. Subdistribution [FG99]. Subgroups [For94]. Subject [Con92, Mad59, PHL98, RG93, RS93, Gui93, HB86, Sto88, WS89]. Subjective [CKLH92, DM97, DZ82]. subpopulations [Mit91]. subsample [SA86]. subsamples [Ste81]. Subsampling [BPR99, BK99, Chr99, Han99, LKCH99a, LKCH99b, SC99, Ste99]. Subseries [She97]. Subset [Pan97, SH92]. Subspace [FBF97]. subspaces [BS84a]. Substitutions [KL90]. Subsurvival [TFP90]. Subtidal [PM97]. Success [HPS98, Ket97]. sufficiency [VM83]. Sufficient [MZ94]. suggestions [Hog74]. Sugihara [PT82]. Sulfate [Haa95, Oeh93]. sum [Kre86, ST60]. sum-quota [Kre86]. Summarize [Llo98]. Summarizing [BGH97]. Sums [IT94, MYZ67, Tel67, Hoe63, Mar65]. Super [PWM92]. Super- [PWM92]. Superclusters [Roe90]. Superior [BS97a]. Superleverage [SC92b]. superpopulation [IF82]. Supersmooth [Efr97]. Supplementing [Roc95]. Support [FHTW97, Gel94, Mer94a, Mer94b]. Suppression [Cox95]. Supremum [FHO87]. Surface [AM92, LR97b, Sho93]. Surfaces [LT94a]. Surprising [Sie93]. Surrogate [CS90a]. Surveillance [APR92, XXP93]. Survey [BF92, Bin96, BZ99, CCD97, DS92, Elt96, Fay96a, Fay96b, GK93, Hoa73, IF82, JMS96, Jud96, MSL93, PWM92, RG95, Rao96, RBD93, RSV95, Rub96b, SS96b, SCC98, SS99, SG96b, SR95, Sit92, Sta91, SJ90, Thé99, Zie90, Agr83, CSC86, God82b, LS87, RW88, MSML97, Sta91]. Surveys [ALM97, BP94b, Bin98, Bre98, CLY92, CM91b, GKL98a, GKL98b, GJDT98, Jud98, Kad98, Kin85, Kot94, PMW93, RN97, Riz92, San98, Ski91, SR96, BS85, CS86, Lit82]. Survival [ALHH92, AL95, AG96, BRR92, Bak98, BD93, BH97, Cha91, CWY97, DG92, EP94, FSE98, GWG90, Gra92, Gra95, GR92a, HMY97, HMC^{+}93, HW95b, LI98a, LC93, LYY99, LGWC93, MZ94, MSK96, Muk96, Mur95, Pru93, RC94, RS90, RS93, Sas97, SD97, TDW95, WH90, Wan91, WC99, XXP93, Yan99, YJW95, Zuc98, dCP99b, BH88, Dyk82, Efr88, FD85, Gas85, Oak89, Tsi82]. Surviving [Yam92]. Survivor [PSVS92, SW97]. Suspects [BGH97]. Sweep [ES90a]. Switching [SS91a]. Symmetric [McC82]. Symmetry [CNS90, Ede90, FG95, McW90, AKZ82, RA89]. symposium [CMT55, Tuk50a]. Synthesis [Goo96, KO96]. Syringe [NGH^{+}95]. System [MEGN92, RV90c, CT98, DL74, JR85]. Systematic [SO94, Wol84]. Systems [Ber91b, BD91b, FT94, GT98, Gri90, Lel94, LC98a, Llo98, LC98b, Mei94, Bro81, Ger82, Pud82, WS89].

T [Bor87, Sta94, ME96, Ede90, LCG^{+}92, Ver93]. T-Cell [LCG^{+}92]. Table [Alb97, CGM83, GSS89, KT88, O’N84, Sim74]. Tables [AMP90, Agr98, BS97a, BS97c, BB98, CS92b, DD90, DDD94, Fri94, Goo94, KP97, KA95, LW91, MPS98, Qui98a, RG93, SW98b, SW98c, VB97, VJ88, VHE91, YF90, ACK87, AG83, BOP88, CG84, Coh81, CG82, Dal82, GR81, Gas85, GH86, Gil86, GH88, GKR88, Gro81, Hab88, Koe86, MP83, MPG85, MPG86, Nai87c, PH81, Sar89, YF95]. Tactical [Hod94]. Tail [BVT96, Ede90, GR96a, GW91, Roj96, BE85, Sch84b]. tailed [Ben83, Riv86]. tails [FS98b]. Tanzania [AGM98]. Target [LW91]. tau [LQ85]. Tax [DN98]. Team [CH94]. Technique [Bri92, Joh92, Roe94, Zel93, Pot87]. Techniques [ES90b, FMR62, Jay96, KW96b, CLNL87, O’G82, Tam81]. Technological [KL90]. tectonic [Cha88b]. Telegraphic [Ano95b, Ano95c, Ano96a, Ano96b, Ano96c]. Telemetry [ASL91]. Temperatures [LR97b]. Template [PS94b]. Templates [AGP91]. Temporal [Haa95, KO96, WCXG97, HW94]. Temporally [DLW95]. Ten [RCH94]. tendencies [DR83]. Tensor [Tak83]. Teoria [Tuk47a]. Teratogenesis [LR97a]. Terms [Cho92, DH86]. Tessellation [AF97]. Test [Bar82, BPL97, DT92, ELR87, FG95, Fan96, FL98, Fol96, GR97a, GHMW94, GR97b, GLG92, GK93, HH90, HW92b, Hay90, HP92, HN92, Jac93, JSKK98, KW95, Led94, LH91, McW90, MH99, PR90a, PB92, PF97, QH98, RSY93, Ruk93, SC97a, SC97b, Sch91b, Shi91, SAK90, SS95a, Sil96, SK90, SW91b, TL97, VM96, Wan96, WM98, Wil50, YF90, dCP99a, Ben83, Ber88b, Boo82, BC82b, CMST88, Dav82, FMD82, Fay85, Fin93, Fra84, Hay86, HVY83, HW88, Jac89, JH82, KR84, KC93, Kob86, LQ85, LL81, MJ82, MH90, MP83, O’N84, PH74, Ran89, Sin86, Spu88, Ste81, SM81, WR84, Wei81, WW87b, Wol87]. tested [dAJ86]. Testing [AME96, AP99, Alb97, AC95b, AP96, AKZ82, BG91a, BS84a, BS87, BM99, BPR95, Boi86, Bor87, BDR92, BS98, Che95, CG97, CW94b, CB97, CF95, CM85, Dea92, DHF84, Eth82, ES90b, Eub97, GJ94, GJ99, GHL96, GS94, GL92, GK93, HMM98, HW92a, Hay90, HRW90, Hua97, HOS94b, Jay96, JB92, JK93, KW97, LTM92, Leh93, LGWC93, MZ94, MH94a, MM93, MBR98, Mun96, Nyb86, Nyb89, PHS94, Raz90, SSC94, SL93a, SC97a, SC97b, SMD93, SS82a, SW88, Tom96, Tro95, Tsa89, Tsa98, Wee91, Wei84, Wel90b, Wes97b, WHG98, YR95, dJGM83, BM81, CB87, CGM83, DD88, DM88a, DR83, HY86, HPE81, KY86, Loh85, MS85b, SD85, Tsi82, TW89, VJ88, Wes98]. Tests [AC97b, Ali89, AKM94, BOP92, BS97a, BY95, BS96b, BF93a, BWC98, Bre90, BHNO92, BH96, CS93, CM97b, Chr91, CS92b, CS92a, CD96, Da̧b90, DM90, EMM95, ES99, FS98a, Fin90, Gra95, HHS97, Hau92, HSP90, HR94, HMO97a, HN87, HMO97b, Hol99, IRC98, JGC95, JR96, Kai94, KL97, KL99, Kun90, LaR91, LK91, LD92, LLR95, LS93, Lou96, MM95, MS92, MS98b, MSW^{+}83, MC84, MNS99, MS91a, NP95, NC99, Ode91, OK96, OR92, Peñ98, QH98, Rob94, Rom90, RMS90, Ruk93, Ryu94, Sch94a, Sil92, ST92, TR97, Tan94, Tho91b, Wan97, Wei83, Wel90a, Xio95, ZM93, Akr88, AA87, Alb88a, BK85a, BK86, Ber89, BM83, BS86a, CS89, CG82, Die82, DKMO89, DvdM81, FW88, Gas85, Gle85, Gro81, Gui88, GV86, HM88a, HP88, Joh78, KOS85]. tests [KK84b, Koe86, LaR86, Lam82, Lee87, LL86, MW81, MW82, MS88a, MS98a, MS95, Nai87c, Noe87, NM83b, ON89, Ols74, OW82, Ran84, Rin83a, Rin85, Riv86, Rog88, Rom88, Ros84b, Rot83, Sch86b, SS87a, SL87, SS82a, Sim89, SW82, SW86, Stu56, Tar87, Thu85, Tri84, WL84, WG88, WM80, WM81, Zel87, ZC88]. Texas [CCG^{+}97a, CCG^{+}97b, Cre97, GMS97, SF97]. th [Zac86]. th-order [Zac86]. Theil [AML95]. Theil-Sen [AML95]. Their [BS97b, FG94, GC95, HHM92, Ye98b, Bue82, Cla87b, Cla89, Efr86a, HHM88, Las94]. theorem [Kar85, Leh81]. Theoretic [LB94a]. Theoretical [YB97, BHS81, CR84]. Theories [Leh93]. Theory [BS98, DDV99, dE97, FK94, Leh93, LYY99, Tuk49a, Tuk49b, dCP99b, CCE85, GM86b, Hog74, Hoo89, Mor83, Sco85, Sha89b, Sha86, Val85, VJ88, Vin69, Dia78]. Therapy [RG94]. There [Ber50, LTM92, WSJ92]. Think [Met50]. Thompson [SO94]. Thomson [Tuk47b, Tuk47c]. Those [Alb94]. Three [KA95, Kun90, Spu99, Tom94, NWGP84, PMF88, RW85, SI85, Uus85]. three-dimensional [NWGP84]. Three-Level [Tom94]. Three-Way [KA95]. Threshold [KP91, PB92, SGC95, Tsa98, Tsa89]. Thresholding [Fan96]. Tight [Ver93]. Time [AME96, Ada98, ALHH92, BD93, CM97a, CMW97, Cha90, CL95, CW91, CLM93, DR91, EH91, FT94, FW96, FSE98, GNK96, GS90b, Gri90, HK93, HDG97, Hol87, KST98, KL90, Koo97, KC96, LMR96, LWY93, LS91a, Li96, LT98, LW92, LL97b, MFS84, MP96, MSL93, Niu96, PS96, RC94, Roc92, SJ92, SAK90, SS96c, SS87b, SL98a, SW98a, The93, TA94a, Tuk52, WPK99, Yam92, dP98, vR98, Ali89, BHS81, BPN87, Cha88a, Cre88, DHF84, FH81a, FH81b, Gew82, Gew84, GA82, HT87, IS87, Kee82, Kit87, KY86, KY85, LZ89, Mar81, Nef82, NP83, Nyb86, Nyb89, PT83, PB87, Sch88, SU74, Sto88, Sto86, TB81, TA94b, Tsa84, Tsa87, WLW89, Wor34, Zac86, ZH87, HR97]. Time- [SS87b]. time-censored [Zac86]. Time-Dependent [AME96, Ada98, ALHH92, FT94, FSE98, HDG97, vR98]. time-series [GA82]. Time-Varying [FW96, Gri90, HK93]. Times [DB97, HW90b, MPCB95, SDW98, WH90]. Timing [ALJ99]. Tippett [Gag43]. Tobit [CCC96]. Tolerance [Ber77, Mee90, Atw84, LT88]. tolerate [Hjo94]. Tom [LF82]. Tomographic [Kol96]. Tomography [MO98, Var96, Wei97, WEST91, VSK85]. Tong [FM83]. Topological [ES90a]. Total [EKWS91, FG91, Mil92, MS91b, MS91c, BJM81, Smo82]. Totals [Qui98a, Zie90]. tournaments [Isr81]. Toxicity [BO95, HZS^{+}90, HB99]. Toxicological [HLC92]. Toxicology [HZS^{+}90]. Traces [WPK99]. Tracking [ASL96]. Tracts [SS95b]. Traffic [McC98, TW98a, TW98b, Var96, Var98]. Train [Bri92]. Training [Hod94, CM84a]. Traits [ZT97]. Transfer [Pap96]. Transform [Akr90, Akr91a, KT94a, WR95, TA89]. Transform-Both-Sides [WR95, KT94a]. Transformation [CWY97, KW96b, NCDF96, ST99, WR95, YM99, CS87, Spi82]. transformation-based [CS87]. Transformation-kernel [YM99]. Transformations [CJ95b, LCWH97, RA89, Vel95, WMR91, BD81a, BC82a, BF85, BJM81, Car82, CR84, Tay85]. Transformed [DKS97, HR84, HC84]. transients [Tan83]. Transition [LMR96, MPCB95, Sch99, Ter94, Wor74]. transitions [BP86]. transitivity [FH82]. Transmission [DA96, Wes97a]. Transplant [DwSH94, Cro74]. transportation [CCE85]. Trapping [WW91b]. travel [Sam85b]. Treat [Roc95]. Treatment [AI95, BC90, BP97, BS97a, Car94, Hol95, KA91, LTM92, MH99, OK96, RG94, Ros93a, ST99, TFAH94, Ueb93, WSLP90, BM83, EH83, EH84, Fli84, HP88, Hsu82, KR88, LR87, Tam85]. treatment-assignment [HP88]. Treatments [CS92a, DPWH99, GS97, Jac93, TL97, CMST88, Jac89, MV88, MRW87, Spu88]. treatments-versus-control [Spu88]. Treaty [PB92]. Tree [LV88, Ros93b, Seg92, LS86]. Tree-Structured [Seg92, LV88]. Trees [LC93, Zha98a]. Trend [CS92b, IRC98, Jac93, OK96, She93, SS96c, BPN87, CJ88, HY86, Nyb86, Rot83, YBN85]. trend-free [CJ88, YBN85]. Trend-Resistant [Jac93]. Trends [HL94, JMS97, Oeh93, SW88]. Trial [Bak98, CCC96, GM96, RG94, TFAH94]. Trials [BPS95, DS97a, EF91, EFS92, GB97, LD92, LB94a, LL97a, Roc95, SBD97a, SBD97b, Gui93, Sch83b]. Triangular [Goo94, Sar89]. Trimmed [RC80, RC82, SP91, dJdWW88]. Triogram [HKS98]. Trispectrum [WW93]. Troubles [SJ90]. True [EH91, GC95]. Truncated [CTC96, GSTM92, GMP97, GL96, Gür96, HL96, Li95, LQT97, dCP99a, Cha84, Gui88, Moo81, TJW88, Wan89]. Truncation [Fan96, NP95, DB83, Hod88]. Trussell [YP92]. Trying [CH94]. TSMARS [LR97b]. tsst [Ros42]. Tukey [Sim87a, CMT55, Hoa73, JV85, SMF94, Sim87a]. Tumor [CH92, Hei95, MC84, NWGP84]. Twins [HHH92]. Two [Akr90, Akr91b, Alb97, AK92, AI95, BHS98, Ber50, BW92, BDR92, BD92, BPS95, CM91a, CF95, DL92, DDD94, DKR95, GR97a, GR97b, GH90, HPS98, HDG97, Hin71, HOS94b, JMV94, Kai94, KSW92, Leh93, LTW96, Llo90, LWT94, MH99, NW97, O’S90, PB98, Pen97, Pes93, RSY93, RN97, RK90, Sit97, SW82, SK90, Tom94, TS96, WH90, WL84, AA87, BM81, Bin82, Cha88b, CJ88, Chr87, DD88, Dal82, Fli84, Fox74, FB83, Har87, HG81, HC84, Isa74, Kob86, KY86, MV88, MB82, NWGP84, OW82, Rin83a, Rin85, Roy86, SA86, SM83, Sch86b, TC72, TA89, Val87b, WR84, WHH88, dAJ86, vd87, NM98, Rig97]. Two- [CF95, Tom94]. two-arm [Fox74]. Two-Dimensional [DDD94, O’S90, Dal82, NWGP84]. two-equation [Bin82]. Two-Factor [Akr90, BD92, DL92, FB83]. two-level [CJ88]. two-parameter [SM83]. Two-Phase [Hin71, Sit97, TS96]. Two-Sample [Akr91b, BW92, BPS95, CM91a, DKR95, Kai94, KSW92, LTW96, LWT94, RK90, WH90, SW82, WL84, AA87, DD88, OW82, Sch86b]. Two-Sided [MH99, Fli84]. Two-Stage [AI95, HOS94b, RSY93, Chr87, Roy86, SA86, Val87b, WHH88]. Two-Step [HDG97]. Two-Way [Alb97, GH90, PB98, Pen97, HC84, Rin83a, Rin85, TA89]. Type [BDM97, GS91b, HW91, Jen91, JDP94, Sil92, Akr88, APS81, BM81, BM83, Bha85, Bof85, CMST88, Esc86, Gle85, LaR86, MB82, Nai87c, Pad86, Poo74, SY84, Stu82, Wit88, dJdWW88, BM92, Gui88, Nai83, Tsu84, Tsu86, Wes87b]. type- [MB82]. Types [NW97].

U.S. [GMWZ99, Hwa86, Joh90, LCMA92, MZTT98]. Umbrella [PW96, Pan97, MW81, MW82]. Unbalanced [AAB97, AD90, Hem79, MS92, SHH78, JK89, Wit88]. Unbiased [GS91a, GS91b, SW98a, Wan97, AP87, BLE88, Esc86, HV86, Hoe84, IS83, KR86, Kre86, KSC89, ST87, Sen74, Spo82]. Unbiasedness [NM83a, SSC94, GS89]. uncertain [Pud82]. Uncertainty [MR94, Wat87]. Unconditional [For94]. Uncontrolled [LB94b]. Uncorrelated [Kus98]. Uncovering [EGG95]. Underclass [Pea91]. Undercount [Sch91a, CT98]. Underground [KM96a]. underlying [Riv86, Sam81, Wes87a]. Understanding [Gil84, Gle93, Kra90]. Underwater [BSPK98]. Unemployment [MZTT98]. Unequal [Sär96, KR82, Kob86, Wil84]. Unified [DF98, LT98, Lju93, Tho91b, BM83, LR87]. Uniform [Bie83, CF95, KO95, Muk96, RS93, Mar65, Sam81]. Uniformity [HOLDG98, JK93, DvdM81]. Uniformly [Ber89, IS83, Tan94]. unifying [JSG88]. Unimodality [De 83b, HG82, HG83]. Union [BS86a]. Union-intersection [BS86a]. unions [BS84a]. Uniqueness [LK92a]. Unit [GHL96, NP95, SL93a, TR97, YR95, DHF84]. United [LT94b, Ada93]. Univariate [dE97, WK94, Akr88]. Unknown [BPR99, CL92a, CM98, DJ95, MPCB95, Pan97, Raf87]. Unmasking [Fun93, Rv90a, Rv90b]. Unobserved [KR96, KT94a]. unpaired [ST60]. Unrelated [BG91b, Bin82]. Unseen [BBC98]. Unskilled [Sam90]. Unsolved [Tuk54]. Unsupervised [FR99]. Untersuchung [Fur46, Gei48]. Updating [Cha71, DZ82, GLT79, NT92]. Upper [Sat95, BE85, WF82]. ups [Suz85]. Uptake [MN93]. Urban [Pea91]. Use [BP94b, CM91a, Dwy45, GFT^{+}93, GK93, Hod94, IT94, LW90, Man90, Mar92, MZ97, MSH93, NFL92, OMF^{+}93, Wei92, YK96, CRS^{+}91, Cro74, Efr86a, Gas85, Geb71, Hab88, Ros42, Sto82, Wor34]. Used [Cha91, RS90, Wam70, Tsi82]. Useful [Zuc98, BE85, NS83, NS86, Wij85]. usefulness [Cla87b, Cla89, MT49]. User [Lon67, Sch83a]. User-oriented [Sch83a]. uses [Dic83, MT49]. Using [AF97, ASL96, AL95, AIR96, BC90, Bal95, BR92, BBH96, BK99, CDW93, CLY92, Chr99, CRS^{+}91, CB90, CS96b, Coo91, CHW92, DM97, Eri81, EW95, FG91, For94, GHRPS90, GSTM92, GBJ96, GBN96, GB98, Gra92, GA82, GR92a, Han99, HK93, Hec96, HMC^{+}93, HL97, HOLDG98, IRC98, Jay96, Joh92, Joh96b, KO95, Kul95, KS93, LKCH99a, LKCH99b, LCG^{+}92, LK91, LMR96, LS95, LS91a, LR97b, LRR91, Llo98, MR94, MZ96, MKSB90, McC98, Mof96, ML94, MPCB95, PO94, PM97, Pet92, PWM92, Pri96, Riz92, RG96, RW97, Ros96, SK93a, SC99, Ste99, TW98a, TW98b, Var98, VW95, WPD98, Weg90, Wes97b, WKL95, WR94, XLR91, Yan99, AG83, Alb88b, BSPK98, BK89, BK85b, CSC86, FS98a, FH82, GN82, GM86b, HW86, HW88, IL91, Isa83]. using [LLT89, LR97a, LS86, Lou84, MRW87, Nai87b, Nef82, O’G82, Sam93a, SM83, Smo82, Tam85, Wes98, Zel86, de 81, de 83a]. Usual [NCDF96]. Utilitarian [Lju93]. Utilities [SSK90, KW88]. Utility [CP91]. Utilization [DG99, GWG98].

V [Daw82, Gei48]. Validated [Tho91a, NWGP84]. Validating [WK94, WC92]. Validation [ET97, HY98, KAT91, LS95, RFB97, SBS94, Sha93, Efr83, Hur85b, PC84, ST87, Sil84]. validatory [Par93a]. Validity [GBN96, Har90a, SMD93, Hoo89]. Valuation [CH91]. Value [Amm93, BF90, CG92, HW92b, Sch87a]. Values [BB94, Buo90, CH91, GC95, JB92, LS97, SMP95, Wee91, BS87, Bor87, CS86, Lou84, SO83, TW89, Uus85]. Variability [MO98, Per97, Wan97, Gui93]. Variable [AI95, BRMZR97, BZ92, BJB95, BZ99, BS96c, CR92, EF91, FH86, Hig98, LA93, LR97a, SB95, SO94, Wei95, ZL95, Bin82, But84, DM88b, Krä86, MB88a, Nai87b, Poo74, Pud82]. Variables [AK92, AIR96, BJB95, CH97b, EGGS97, Fel74, GH95, Hec96, Hou95, Jon96, LS95, Lie94, Mad59, MSML97, MW98a, MW98b, Mof96, Now92, PSVS92, RN97, RG96, RW98, Ros96, Sie93, WPD98, Agr83, BT81, BF83, BK85b, CGG85, CMS76, CMS87, DH93, DL74, Dyk82, FD85, Hoe63, Hwa86, KR76, KP81b, KP82, Mar64, Mar65, Mar84, MZ89, Rel72, Sch86a, Spl83, TH82, dAJ86]. Variance [Bin96, CD91, Cha90, CG97, CM98, CM97b, CM91a, CV92, DC87, DLB91, DBB^{+}94, Dor94, Elt96, Fay96b, GN82, GH90, GP92, HS94, IT94, Isa83, Jud96, KPR91, MM95, McC94, MP96, RS96, Rao96, Rub96b, Sär96, SG96a, SS99, Sit97, Sla94, Sri72, SO94, SW98a, SHGL96, Val93, Wee91, XLR91, BD88, Bet89, BS85, Bro83, DW87, Fab91, Flu85, Gol83a, HRR81, HV86, HS89c, IS83, JK89, Khu81, KR86, Loh86, Mar81, MS88a, MT81, MT82, Ols74, Pri74, RC81, Roy86, Ruk87, Sha89a, SM84b, TV85, Val87b, Wes87b, Wes87a, Wil84, Wol84, Won82a, Zeh88]. Variance-Covariance [XLR91]. Variances [FNP95, Lau92, Kob86, Lin74, Riv86, TKK89]. variant [KP81b, KP82]. Variate [DJ98, SS80]. Variates [GF96a, SMP95, SL80]. Variation [BJM81, Gra95, NM98, Pag90]. Variations [DG99]. variograms [Ste87]. Various [GMWZ99, SKF98]. Vary [SRR95]. Varying [FW96, Gri90, HK93, SW98a, WCH98]. Varying-Coefficient [WCH98]. Vector [Ahn97, Mau95, Qui98b, SG97, WM98, YR95, Spl83, Wor74]. Vectors [BDR92, GR97a, GR97b, SMF94]. Velleman [TV85]. Verification [LT97, PB92]. Versatility [KY99]. Version [Bar82, Led94, RS94b, Kar85]. Versions [JDP94, FHO87]. Versus [HMM98, RG94, VM96, Fli84, Spu88]. Vertical [DA96]. Very [Atk94]. Via [CDP96, ES90b, HR90, KT94b, NNR95, BPS96, CL92b, CCD97, DR98, DJ95, Gou97, KT94c, LR97c, LC95b, LV88, Moj99, MP95, PS94b, PS99, Sam85a, Wan98a, Won82a]. Viability [ZWWB96]. Vietnam [Ang91]. View [Lon67, Ber88b, JG83]. Views [Kra90]. Viral [ZDL96]. Visual [MT95a]. Visualization [CL92a, Li92]. Vitro [HZS^{+}90]. Voids [Roe90]. Volcanicity [GT91]. Volcanism [Sol91]. Volume [Ano97e]. Voluntary [Ang91]. Voronoi [AF97]. Vote [GKB98]. Votes [Est92]. voting [BP86]. Voyage [Bil97]. vs [Stu56].

W [HHM88, Hoa73, JV85, Sim87a, Tuk49b]. W. [God82b, Man94]. Wages [RBD93]. Waiting [HW90b]. Waksberg [Pot87]. Wald [LH91, Sil92]. Wald-Type [Sil92]. walk [NM83b]. Wallis [Wei81]. Walsh [Sto91]. Walsh-Fourier [Sto91]. Warning [APR92, Hab88]. Wasserman [HL81]. Water [KM96a]. Waters [KSJ94]. Watson [SAK90]. Wave [MLMT94]. Wave-Soldering [MLMT94]. Wavelet [AGM94, CKM97, DJ95, Efr99, Fan96, Kol96, Vid98]. Wavelets [BSPK98, PM97, Wan98a]. Way [Alb97, Fri94, GH90, KA95, PB98, Pen97, HC84, PH86, Rin83a, Rin85, Sin86, TA89]. Weak [BJB95, Cha91, DG92, RT96a]. Weibull [BS93, MSK96, SY84]. weighing [GM86b]. weight [PS82]. Weighted [Che94a, CC96a, Hoo93, HM91, KY99, MBL98, Mül87, RV98, RP96, Sas93, WWZtO97, Yan97, Yan99, CWR88, GL83, Mac86, Mac87, Tar87]. Weighting [PMW93, Zie90]. Weights [Bly93, Per97, PWM92, Wie98, CWR88]. Welch [Rot83]. Welfare [Ano90, FIB90, HT90, Kra90]. well [Daw82]. well-calibrated [Daw82]. Welsch [Rup85]. Wet [Haa95, Oeh93]. Whale [RZ98]. Whales [Buc95a, RGZ95a, RGZ95b, Sch95b, TL95, Wol95]. wheel [Eth82]. Where [Tuk60]. Which [Est92, SRR95]. White [Fab94, SB81]. Whitney [BW94, Mee90]. Whittle [PO94]. Widely [Wam70]. Width [Wil50, Kan88, Nai83]. Wiener [Tuk52]. Wilcoxon [BW94, FHO87, SW82, Tab90]. Wildlife [ASL91, Pol91]. Wilks [STM^{+}65]. win [Isr81]. Windfield [CB94]. Window [Haa90, WHGA96, MR94]. Winner [Est92]. Within [WJ92, Gui93]. within-subject [Gui93]. Without [BC90, EV92, Rom90, AK86, HB99]. word [Che71]. Worked [RBD93].

Workers [Sam90]. Working [CKLH92]. World [Nor90].

x [DD90, Bre92]. X-Fixed [Bre92]. X’s [Lit92].

year [Tuk85]. Years [Rub96a]. York [DXP92]. Yule [Goo96, KMT99].

Zellner [Kar81]. Zero [BS96a]. Zidek [Tsu84, Tsu86]. Zipf [Hil70, Hil74, HW75]. zonoid [KM96b].

References

Albers:1987:CRT

[AA87] Willem Albers and Michael G. Akritas. Combined rank tests for the two-sample problem with randomly censored data. Journal of the American Statistical Association, 82(398):648–655, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Akritas:1994:FNH

[AA94] Michael G. Akritas and Steven F. Arnold. Fully nonparametric hypotheses for factorial designs I: Multivariate repeated measures designs. Journal of the American Statistical Association, 89(425): 336–343, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Akritas:1997:NHR

[AAB97] Michael G. Akritas, Steven F. Arnold, and Edgar Brunner. Nonparametric hypotheses and rank statistics for unbalanced factorial designs. Journal of the American Statistical Association, 92 (437):258–265, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/march/akritas.htm.

Albright:1993:SAH

[AASM93] S. Christian Albright, Jim Albert, Hal S. Stern, and Carl N. Morris. A statistical analysis of hitting streaks in baseball. Journal of the American Statistical Association, 88(424):1175–??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Akritas:1997:NMF

[AB97] Michael G. Akritas and Edgar Brunner. Nonparametric methods for factorial designs with censored data. Journal of the American Statistical Association, 92(438):568–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Abramson:1988:RRH

[Abr88] Ian Abramson. A recursive regression for high-dimensional models, with application to growth curves and repeated measures. Journal of the American Statistical Association, 83(404):1073–1077, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Albert:1993:BAB

[AC93] James H. Albert and Siddhartha Chib. Bayesian analysis of binary and polychotomous response data. Journal of the American Statistical Association, 88(422):669–??, June 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Altman:1995:NEB

[AC95a] N. S. Altman and G. Casella. Nonparametric empirical Bayes growth curve analysis. Journal of the American Statistical Association, 90(430):508–515, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Alvo:1995:TOA

[AC95b] Mayer Alvo and Paul Cabilio. Testing ordered alternatives in the presence of incomplete data. Journal of the American Statistical Association, 90(431):1015–1024, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Atkinson:1995:ODH

[AC95c] A. C. Atkinson and R. D. Cook. D-optimum designs for heteroscedastic linear models. Journal of the American Statistical Association, 90(429):204–212, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Aerts:1997:LPE

[AC97a] Marc Aerts and Gerda Claeskens. Local polynomial estimation in multiparameter likelihood models. Journal of the American Statistical Association, 92(440):1536–1545, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/AERTS.HTM.

Albert:1997:BTM

[AC97b] Jim Albert and Siddhartha Chib. Bayesian tests and model diagnostics in conditionally independent hierarchical models. Journal of the American Statistical Association, 92(439):916– 925, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/sept/albert.htm.

Agresti:1987:ORS

[ACK87] Alan Agresti, Christy Chuang, and Abbas Kezouh. Orderrestricted score parameters in association models for contingency tables. Journal of the American Statistical Association, 82(398): 619–623, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Andrews:1990:NAU

[AD90] D. M. Andrews and H. A. David. Nonparametric analysis of unbalanced paired-comparison or ranked data. Journal of the American Statistical Association, 85(412):1140–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Adams:1893:SRR

[Ada93] H. C. Adams. Some recent results in railway statistics in the United States. Journal of the American Statistical Association, 3(??):501–512, ???? 1893. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Adak:1998:TDS

[Ada98] Sudeshna Adak. Time-dependent spectral analysis of nonstationary time series. Journal of the American Statistical Association, 93 (444):1488–1501, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/dec/adak.html.

Allard:1997:NML

[AF97] Denis Allard and Chris Fraley. Nonparametric maximum likelihood estimation of features in spatial point processes using Voronoi tessellation. Journal of the American Statistical Association, 92(440):1485–1493, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/ALLARD.HTM.

Albert:1983:BEM

[AG83] James H. Albert and Arjun K. Gupta. Bayesian estimation methods for 2 × 2 contingency tables using mixtures of Dirichlet distributions. Journal of the American Statistical Association, 78(383): 708–717, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Albert:1985:BMB

[AG85] James H. Albert and Arjun K. Gupta. Bayesian methods for binomial data with applications to a nonresponse problem. Journal of the American Statistical Association, 80(389):167–174, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Arjas:1996:BIS

[AG96] Elja Arjas and Dario Gasbarra. Bayesian inference of survival probabilities, under stochastic ordering constraints. Journal of the American Statistical Association, 91(435):1101–1109, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Antoniadis:1994:WMC

[AGM94] A. Antoniadis, G. Grégoire, and I. W. McKeague. Wavelet methods for curve estimation. Journal of the American Statistical Association, 89(428):1340–1353, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Angeles:1998:ACS

[AGM98] Gustavo Angeles, David K. Guilkey, and Thomas A. Mroz. Applications and case studies — purposive program placement and the estimation of family planning program effects in Tanzania. Journal of the American Statistical Association, 93(443): 884–899, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/angeles.html.

Agnew:1985:MPA

[Agn85] Carson E. Agnew. Multiple probability assessments by dependent experts. Journal of the American Statistical Association, 80(390):

343–347, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Amit:1991:SIR

[AGP91] Yali Amit, Ulf Grenander, and Mauro Piccioni. Structural image restoration through deformable templates. Journal of the American Statistical Association, 86(414):376–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Agresti:1983:SSM

[Agr83] Alan Agresti. A survey of strategies for modeling crossclassifications having ordinal variables. Journal of the American Statistical Association, 78(381):184–198, ???? 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Agresti:1997:MRM

[Agr97] Alan Agresti. A model for repeated measurements of a multivariate binary response. Journal of the American Statistical Association, 92(437):315–??, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/agresti.htm.

Agresti:1998:CAE

[Agr98] Alan Agresti. Comment: Approximately exact inference for the common odds ratio in several 2×2 tables. Journal of the American Statistical Association, 93(444):1307–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson:1981:EDM

[AH81] T. W. Anderson and Cheng Hsiao. Estimation of dynamic models with error components. Journal of the American Statistical Association, 76(375):598–606, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Albert:1997:MRM

[AHB97] Paul S. Albert, Sally A. Hunsberger, and Frank M. Biro. Modeling repeated measures with monotonic ordinal responses and misclassification, with applications to studying maturation. Journal of the American Statistical Association, 92(440):1304–1311, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/ALBERTPS.HTM.

Ahn:1997:IVA

[Ahn97] Sung K. Ahn. Inference of vector autoregressive models with cointegration and scalar components. Journal of the American Statistical Association, 92(437):350–356, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/ahn.htm.

Angrist:1995:TSL

[AI95] Joshua D. Angrist and Guido W. Imbens. Two-stage least squares estimation of average causal effects in models with variable treatment intensity. Journal of the American Statistical Association, 90(430):431–442, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Angrist:1996:RIC

[AIR96] Joshua D. Angrist, Guido W. Imbens, and Donald B. Rubin. Rejoinder: Identification of causal effects using instrumental variables. Journal of the American Statistical Association, 91(434): 444–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Andreatta:1986:EFP

[AK86] G. Andreatta and G. M. Kaufman. Estimation of finite population properties when sampling is without replacement and proportional to magnitude. Journal of the American Statistical Association, 81 (395):657–666, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Angrist:1992:EAS

[AK92] Joshua D. Angrist and Alan B. Krueger. The effect of age at school entry on educational attainment: An application of instrumental variables with moments from two samples. Journal of the American Statistical Association, 87(418):328–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Aly:1994:STC

[AKM94] Emad-Eldin A. A. Aly, Subhash Kochar, and Ian W. McKeague. Some tests for comparing cumulative incidence functions and cause-specific hazard rates. Journal of the American Statistical Association, 89(427):994–999, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Akritas:1986:BKM

[Akr86] Michael G. Akritas. Bootstrapping the Kaplan–Meier estimator. Journal of the American Statistical Association, 81(396):1032– 1038, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Akritas:1988:PTG

[Akr88] Michael G. Akritas. Pearson-type goodness-of-fit tests: the univariate case. Journal of the American Statistical Association, 83 (401):222–230, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Akritas:1990:RTM

[Akr90] Michael G. Akritas. The rank transform method in some twofactor designs. Journal of the American Statistical Association, 85(409):73–78, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Akritas:1991:LRT

[Akr91a] Michael G. Akritas. Limitations of the rank transform procedure: A study of repeated measures designs, Part I. Journal of the American Statistical Association, 86(414):457–460, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Akritas:1991:RET

[Akr91b] Michael G. Akritas. Robust M estimation in the two-sample problem. Journal of the American Statistical Association, 86(413): 201–204, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Antille:1982:TS

[AKZ82] A. Antille, G. Kersting, and W. Zucchini. Testing symmetry. Journal of the American Statistical Association, 77(379):639–646, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Allenby:1994:MHP

[AL94] Greg M. Allenby and Peter J. Lenk. Modeling household purchase behavior with logistic normal regression. Journal of the American Statistical Association, 89(428):1218–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson:1995:SAU

[AL95] Jon E. Anderson and Thomas A. Louis. Survival analysis using a scale change random effects model. Journal of the American Statistical Association, 90(430):669–679, ???? 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Alam:1984:AGD

[Ala84] Khursheed Alam. Analysis of grouped data from multinomial populations. Journal of the American Statistical Association, 79(385): 132–135, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Albers:1988:CRT

[Alb88a] Willem Albers. Combined rank tests for randomly censored paired data. Journal of the American Statistical Association, 83(404): 1159–1162, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Albert:1988:CMU

[Alb88b] James H. Albert. Computational methods using a Bayesian hierarchical generalized linear model. Journal of the American Statistical Association, 83(404):1037–1044, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Albert:1994:EBH

[Alb94] Jim Albert. Exploring baseball hitting data: What about those breakdown statistics? Journal of the American Statistical Association, 89(427):1066–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Albert:1997:BTE

[Alb97] James H. Albert. Bayesian testing and estimation of association in a two-way contingency table. Journal of the American Statistical Association, 92(438):685–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson:1992:TDA

[ALHH92] Jon E. Anderson, Thomas A. Louis, Niels V. Holm, and Bent Harvald. Time-dependent association measures for bivariate survival distributions. Journal of the American Statistical Association, 87 (419):641–650, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ali:1989:TAR

[Ali89] Mukhtar M. Ali. Tests for autocorrelation and randomness in multiple time series. Journal of the American Statistical Association, 84(406):533–540, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Allenby:1999:DMP

[ALJ99] Greg M. Allenby, Robert P. Leone, and Lichung Jen. A dynamic model of purchase timing with application to direct marketing. Journal of the American Statistical Association, 94 (446):365–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/allenby.pdf.

Arora:1997:EBE

[ALM97] Vipin Arora, P. Lahiri, and Kanchan Mukherjee. Empirical Bayes estimation of finite population means from complex surveys. Journal of the American Statistical Association, 92(440): 1555–1562, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/ARORA.HTM.

Altman:1990:KSD

[Alt90] N. S. Altman. Kernel smoothing of data with correlated errors. Journal of the American Statistical Association, 85(411):749–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Azari:1992:PLO

[AM92] A. S. Azari and Hans-Georg Müller. Preaveraged localized orthogonal polynomial estimators for surface smoothing and partial differentiation. Journal of the American Statistical Association, 87 (420):1005–1017, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Al-Mutairi:1998:ACF

[AMCS98] Dhaifalla Al-Mutairi, Yiping Chen, and Nozer D. Singpurwalla. An adaptive concatenated failure rate model for software reliability. Journal of the American Statistical Association, 93(443): 1150–1163, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/al-mutairi.html.

Amemiya:1974:BPA

[Ame74] T. Amemiya. Bivariate probit analysis: minimum chi-square methods. Journal of the American Statistical Association, 69(348):940– 944, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Abrahamowicz:1996:TDH

[AME96] Michal Abrahamowicz, Todd MacKenzie, and John M. Esdaile. Time-dependent hazard ratio: Modeling and hypothesis testing with application in lupus nephritis. Journal of the American Statistical Association, 91(436):1432–??, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Akritas:1995:TSE

[AML95] Michael G. Akritas, Susan A. Murphy, and Michael P. LaValley. The Theil-Sen estimator with doubly censored data and applications to astronomy. Journal of the American Statistical Association, 90(429):170–177, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ammann:1993:RSV

[Amm93] Larry P. Ammann. Robust singular value decompositions: A new approach to projection pursuit. Journal of the American Statistical Association, 88(422):505–514, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Agresti:1990:EIC

[AMP90] Alan Agresti, Cyrus R. Mehta, and Nitin R. Patel. Exact inference for contingency tables with ordered categories. Journal of the American Statistical Association, 85(410):453–458, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Amrhein:1995:MEP

[Amr95] Peter Amrhein. Minimax estimation of proportions under random sample size. Journal of the American Statistical Association, 90 (431):1107–1111, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Angrist:1991:DLV

[Ang91] Joshua D. Angrist. The draft lottery and voluntary enlistment in the Vietnam era. Journal of the American Statistical Association, 86(415):584–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1990:QCW

[Ano90] Anonymous. Quality control of welfare programs. Journal of the American Statistical Association, 85(411):848–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1991:SSP

[Ano91] Anonymous. Social statistics and public policy for the 1990s. Journal of the American Statistical Association, 86(414):486–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1995:EC

[Ano95a] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 90(432):1495–??, ???? 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1995:TRa

[Ano95b] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 90(430):809–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1995:TRb

[Ano95c] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 90(432):1492–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1996:TRa

[Ano96a] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 91(433):440–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1996:TRb

[Ano96b] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 91(434):915–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1996:TRc

[Ano96c] Anonymous. Telegraphic reviews. Journal of the American Statistical Association, 91(435):1380–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1997:EC

[Ano97a] Anonymous. 1997 editorial collaborators. Journal of the American Statistical Association, 92(440):1658–??, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/97editors.
html.

Anonymous:1997:BRa

[Ano97b] Anonymous. Book reviews. Journal of the American Statistical Association, 92(437):383–??, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1997:BRb

[Ano97c] Anonymous. Book reviews. Journal of the American Statistical Association, 92(438):787–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1997:BRc

[Ano97d] Anonymous. Book reviews. Journal of the American Statistical Association, 92(439):1213–??, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1997:IV

[Ano97e] Anonymous. Index to volume 92. Journal of the American Statistical Association, 92(440):1662–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1998:BRa

[Ano98a] Anonymous. Book reviews. Journal of the American Statistical Association, 93(441):397–??, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1998:BRb

[Ano98b] Anonymous. Book reviews. Journal of the American Statistical Association, 93(442):832–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1998:GEC

[Ano98c] Anonymous. General — editorial collaborators. Journal of the American Statistical Association, 93(444):1531–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1998:GRP

[Ano98d] Anonymous. General — review paper. Journal of the American Statistical Association, 93(443):1216–??, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1999:BRa

[Ano99a] Anonymous. Book reviews. Journal of the American Statistical Association, 94(445):333–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1999:BRb

[Ano99b] Anonymous. Book reviews. Journal of the American Statistical Association, 94(446):645–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1999:ER

[Ano99c] Anonymous. Editors’ report for 1998. Journal of the American Statistical Association, 94(446):661–??, June 1999. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/index.html#98tm.

Anonymous:1999:LEa

[Ano99d] Anonymous. Letters to the Editor. Journal of the American Statistical Association, 94(445):351–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:1999:LEb

[Ano99e] Anonymous. Letters to the Editor. Journal of the American Statistical Association, 94(446):659–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Abdelbasit:1983:EDB

[AP83] K. M. Abdelbasit and R. L. Plackett. Experimental design for binary data. Journal of the American Statistical Association, 78 (381):90–98, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Andrews:1987:BMU

[AP87] Donald W. K. Andrews and Peter C. B. Phillips. Best medianunbiased estimation in linear regression with bounded asymmetric loss functions. Journal of the American Statistical Association, 82 (399):886–893, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Arnold:1989:BEP

[AP89a] Barry C. Arnold and S. James Press. Bayesian estimation and prediction for Pareto data. Journal of the American Statistical Association, 84(408):1079–1084, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Arnold:1989:CCD

[AP89b] Barry C. Arnold and S. James Press. Compatible conditional distributions. Journal of the American Statistical Association, 84 (405):152–156, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Andrews:1996:TSC

[AP96] Donald W. K. Andrews and Werner Ploberger. Testing for serial correlation against an ARMA(1,1) process. Journal of the American Statistical Association, 91(435):1331–1342, ???? 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Agustin:1999:OSP

[AP99] Ma. Zenia N. Agustin and Edsel A. Peña. Order statistic properties, random generation, and goodness-of-fit testing for a minimal repair model. Journal of the American Statistical Association, 94 (445):266–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Assaf:1992:NLW

[APR92] David Assaf, Moshe Pollak, and Ya’acov Ritov. A new look at warning and action lines of surveillance schemes. Journal of the American Statistical Association, 87(419):889–??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Arrow:1981:PSB

[APS81] Kenneth J. Arrow, Leon Pesotchinsky, and Milton Sobel. On partitioning a sample with binary-type questions in lieu of collecting observations. Journal of the American Statistical Association, 76 (374):402–409, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ahn:1990:EPN

[AR90] Sung K. Ahn and Gregory C. Reinsel. Estimation for partially nonstationary multivariate autoregressive models. Journal of the American Statistical Association, 85(411):813–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Alho:1990:EMO

[AS90a] Juha M. Alho and Bruce D. Spencer. Error models for official mortality forecasts. Journal of the American Statistical Association, 85(411):609–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Arminger:1990:PML

[AS90b] Gerhard Sobel Arminger and Michael E. Sobel. Pseudo-maximum likelihood estimation of mean and covariance structures with missing data. Journal of the American Statistical Association, 85(409): 195–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson-Sprecher:1991:SSA

[ASL91] Richard Anderson-Sprecher and Johannes Ledolter. State-space analysis of wildlife telemetry data. Journal of the American Statistical Association, 86(415):595–??, September 1991. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson-Sprecher:1996:SEP

[ASL96] Richard Anderson-Sprecher and Russell V. Lenth. Spline estimation of paths using bearings-only tracking data. Journal of the American Statistical Association, 91(433):276–283, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Atkinson:1994:FVR

[Atk94] A. C. Atkinson. Fast very robust methods for the detection of multiple outliers. Journal of the American Statistical Association, 89(428):1329–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Atwood:1984:ATI

[Atw84] Corwin L. Atwood. Approximate tolerance intervals, based on maximum likelihood estimates. Journal of the American Statistical Association, 79(386):459–465, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Atkinson:1982:ISC

[AY82] A. C. Atkinson and Lam Yeh. Inference for Sichel’s compound Poisson distribution. Journal of the American Statistical Association, 77(377):153–158, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Byrne:1983:IAM

[BA83] Philip J. Byrne and Steven F. Arnold. Inference about multivariate means for a nonstationary autoregressive model. Journal of the American Statistical Association, 78(384):850–855, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bacchetti:1989:AIM

[Bac89] Peter Bacchetti. Additive isotonic models. Journal of the American Statistical Association, 84(405):289–294, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bacchetti:1990:EIP

[Bac90] Peter Bacchetti. Estimating the incubation period of AIDS by comparing population infection and diagnosis patterns. Journal of the American Statistical Association, 85(412):1002–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Baker:1998:ACS

[Bak98] Stuart G. Baker. Applications and case studies — analysis of survival data from A randomized trial with all-or-none compliance: Estimating the cost-effectiveness of a cancer screening program. Journal of the American Statistical Association, 93(443): 929–934, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/baker.html.

Balding:1995:EPF

[Bal95] David J. Balding. Estimating products in forensic identification using DNA profiles. Journal of the American Statistical Association, 90(431):839–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Baltazar-Aban:1995:PHB

[BAP95] Inmaculada Baltazar-Aban and Edsel A. Peña. Properties of hazard-based residuals and implications in model diagnostics. Journal of the American Statistical Association, 90(429):185–197, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bartels:1982:RVN

[Bar82] Robert Bartels. The rank version of von Neumann’s ratio test for randomness. Journal of the American Statistical Association, 77(377):40–46, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287767.

Barabba:1991:TGL

[Bar91] Vincent P. Barabba. Through a glass less darkly. Journal of the American Statistical Association, 86(413):1–??, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Block:1974:CBE

[BB74] H. W. Block and A. P. Basu. A continuous bivariate exponential extension. Journal of the American Statistical Association, 69 (348):1031–1037, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bjornstad:1988:EBE

[BB88] Jan F. Bjørnstad and Ronald W. Butler. The equivalence of backward elimination and multiple comparisons. Journal of the American Statistical Association, 83(401):136–144, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1989:EPM

[BB89] James O. Berger and Jose-M. Bernardo. Estimating a product of means: Bayesian analysis with reference priors. Journal of the American Statistical Association, 84(405):200–207, ???? 1989.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bloom:1990:MAM

[BB90] David E. Bloom and Neil G. Bennett. Modeling American marriage patterns. Journal of the American Statistical Association, 85(412):1009–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1994:VMC

[BB94] Roger L. Berger and Dennis D. Boos. P values maximized over a confidence set for the nuisance parameter. Journal of the American Statistical Association, 89(427):1012–1016, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Booth:1998:CAE

[BB98] James G. Booth and Ronald W. Butler. Comment: Approximately exact inference for the common odds ratio in several 2 × 2 tables. Journal of the American Statistical Association, 93(444):1310–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boneh:1998:EPF

[BBC98] Shahar Boneh, Arnon Boneh, and R. J. Caron. Estimating the prediction function and the number of unseen species in sampling with replacement. Journal of the American Statistical Association, 93(441):372–379, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/BONEH.HTM.

Berzuini:1997:DCI

[BBGL97] Carlo Berzuini, Nicola G. Best, Walter R. Gilks, and Cristiana Larizza. Dynamic conditional independence models and Markov chain Monte Carlo methods. Journal of the American Statistical Association, 92(440):1403–1412, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/BERZUINI.HTM.

Booth:1994:BMF

[BBH94] James G. Booth, Ronald W. Butler, and Peter Hall. Bootstrap methods for finite populations. Journal of the American Statistical Association, 89(428):1282–1289, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1996:EIR

[BBH96] Agnes Berger, Carol A. Bodian, and Warren M. Hirsch. On estimating incidence rates of diseases with delayed onset using biased samples. Journal of the American Statistical Association, 91(434): 831–840, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Box:1982:CAT

[BC82a] G. E. P. Box and D. R. Cox. Comment on: “An analysis of transformations revisited” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 296–311, MR 83b:62048] by P. J. Bickel and K. A. Doksum. Journal of the American Statistical Association, 77(377):209–210, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [BD81a].

Brookmeyer:1982:SMT

[BC82b] Ron Brookmeyer and John Crowley. A k-sample median test for censored data. Journal of the American Statistical Association, 77(378):433–440, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Baker:1990:ESE

[BC90] Stuart G. Baker and Kenneth C. Chu. Evaluating screening for the early detection and treatment of cancer without using a randomized control group. Journal of the American Statistical Association, 85(410):321–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Breslow:1993:AIG

[BC93] N. E. Breslow and D. G. Clayton. Approximate inference in generalized linear mixed models. Journal of the American Statistical Association, 88(421):9–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bensmail:1996:RGD

[BC96] Halima Bensmail and Gilles Celeux. Regularized Gaussian discriminant analysis through eigenvalue decomposition. Journal of the American Statistical Association, 91(436):1743–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:1995:OCS

[BCH95] Lawrence D. Brown, George Casella, and J. T. Gene Hwang. Optimal confidence sets, bioequivalence, and the limaçon of Pascal. Journal of the American Statistical Association, 90(431):880–889, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bedrick:1996:NPP

[BCJ96] Edward J. Bedrick, Ronald Christensen, and Wesley Johnson. A new perspective on priors for generalized linear models. Journal of the American Statistical Association, 91(436):1450–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bickel:1981:ATR

[BD81a] Peter J. Bickel and Kjell A. Doksum. An analysis of transformations revisited. Journal of the American Statistical Association, 76(374):296–311, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comment [BC82a].

Blumenthal:1981:EBP

[BD81b] Saul Blumenthal and Ram C. Dahiya. Estimating the binomial parameter n. Journal of the American Statistical Association, 76 (376):903–909, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1988:BAR

[BD88] James O. Berger and John Deely. A Bayesian approach to ranking and selection of related means with alternatives to analysisof-variance methodology. Journal of the American Statistical Association, 83(402):364–373, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bayarri:1989:ORP

[BD89] M. J. Bayarri and M. H. DeGroot. Optimal reporting of predictions. Journal of the American Statistical Association, 84(405): 214–222, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bayarri:1991:WBE

[BD91a] M. J. Bayarri and M. H. DeGroot. What Bayesians expect of each other. Journal of the American Statistical Association, 86(416):

924–932, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bryant:1991:EBA

[BD91b] John Bryant and Roger Day. Empirical Bayes analysis for systems of mixed models with linked autocorrelated random effects. Journal of the American Statistical Association, 86(416):1007–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brunner:1992:RPT

[BD92] Edgar Brunner and Holger Dette. Rank procedures for the twofactor mixed model. Journal of the American Statistical Association, 87(419):884–888, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Burr:1993:CBM

[BD93] Deborah Burr and Hani Doss. Confidence bands for the median survival time as a function of the covariates in the Cox model. Journal of the American Statistical Association, 88(424):1330– 1340, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bollinger:1997:MDC

[BD97] Christopher R. Bollinger and Martin H. David. Modeling discrete choice with response error: Food stamp participation. Journal of the American Statistical Association, 92(439):827–835, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/bollinger.htm.

Brunner:1997:BTA

[BDM97] Edgar Brunner, Holger Dette, and Axel Munk. Box-type approximations in nonparametric factorial designs. Journal of the American Statistical Association, 92(440):1494–??, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boudjellaba:1992:TCB

[BDR92] Hafida Boudjellaba, Jean-Marie Dufour, and Roch Roy. Testing causality between two vectors in multivariate autoregressive moving average models. Journal of the American Statistical Association, 87(420):1082–1090, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berman:1985:UUB

[BE85] M. Berman and G. K. Eagleson. A useful upper bound for the tail probabilities of the scan statistic when the sample size is large. Journal of the American Statistical Association, 80(392):886–889, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Benford:1938:LAN

[Ben38] Frank Benford. The law of anomalous numbers. Proceedings of the American Philosophical Society, 78(4):551–572, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL http://links.jstor.org/sici?sici=0003-049X%
2819380331%2978%3A4%3C551%3ATLOAN%3E2.0.CO%3B2-G. See

comments about Benford’s biased rounding practices [DF79].

Benjamini:1983:TRC

[Ben83] Yoav Benjamini. Is the t test really conservative when the parent distribution is long-tailed. Journal of the American Statistical Association, 78(383):645–654, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berkson:1950:TTR

[Ber50] J. Berkson. Are there two regressions. Journal of the American Statistical Association, 45(??):164–180, ???? 1950. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berk:1977:TCR

[Ber77] K. N. Berk. Tolerance and condition in regression computations. Journal of the American Statistical Association, 72(??):863–866, ???? 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1982:BRS

[Ber82] James Berger. Bayesian robustness and the Stein effect. Journal of the American Statistical Association, 77(378):358–368, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Beran:1988:BSC

[Ber88a] Rudolf Beran. Balanced simultaneous confidence sets. Journal of the American Statistical Association, 83(403):679–686, ????

1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Beran:1988:PTS

[Ber88b] Rudolf Beran. Prepivoting test statistics: a bootstrap view of asymptotic refinements. Journal of the American Statistical Association, 83(403):687–697, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1989:UMP

[Ber89] Roger L. Berger. Uniformly more powerful tests for hypotheses concerning linear inequalities and normal means. Journal of the American Statistical Association, 84(405):192–199, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Beran:1990:CPR

[Ber90a] Rudolf Beran. Calibrating prediction regions. Journal of the American Statistical Association, 85(411):715–723, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Beran:1990:RBS

[Ber90b] Rudolf Beran. Refining bootstrap simultaneous confidence sets. Journal of the American Statistical Association, 85(410):417–426, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Beran:1991:ELG

[Ber91a] Jan Beran. M estimators of location for Gaussian and related processes with slowly decaying serial correlations. Journal of the American Statistical Association, 86(415):704–708, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berliner:1991:LBP

[Ber91b] L. Mark Berliner. Likelihood and Bayesian prediction of chaotic systems. Journal of the American Statistical Association, 86(416): 938–952, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bethel:1989:MVE

[Bet89] James Bethel. Minimum variance estimation in stratified sampling. Journal of the American Statistical Association, 84(405): 260–265, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Breiman:1983:HMV

[BF83] L. Breiman and D. Freedman. How many variables should be entered in a regression equation. Journal of the American Statistical Association, 78(381):131–136, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Breiman:1985:EOT

[BF85] Leo Breiman and Jerome H. Friedman. Estimating optimal transformations for multiple regression and correlation. with discussion and with a reply by the authors. Journal of the American Statistical Association, 80(391):580–619, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:1990:LBR

[BF90] Lawrence D. Brown and Roger H. Farrell. A lower bound for the risk in estimating the value of a probability density. Journal of the American Statistical Association, 85(412):1147–1153, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Biemer:1992:QRD

[BF92] Paul P. Biemer and Gösta Forsman. On the quality of reinterview data with application to the current population survey. Journal of the American Statistical Association, 87(420):915–923, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bowman:1993:ASD

[BF93a] A. W. Bowman and P. J. Foster. Adaptive smoothing and densitybased tests of multivariate normality. Journal of the American Statistical Association, 88(422):529–537, June 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bunge:1993:ENS

[BF93b] J. Bunge and M. Fitzpatrick. Estimating the number of species: A review. Journal of the American Statistical Association, 88(421):

364–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Banerjee:1997:IDL

[BF97] Mousumi Banerjee and Edward W. Frees. Influence diagnostics for linear longitudinal models. Journal of the American Statistical Association, 92(439):999–1005, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/banerjee.htm.

Beran:1991:SDC

[BG91a] Jan Beran and Sucharita Ghosh. Slowly decaying correlations, testing normality, nuisance parameters. Journal of the American Statistical Association, 86(415):785–791, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Blattberg:1991:SEP

[BG91b] Robert C. Blattberg and Edward I. George. Shrinkage estimation of price and promotional elasticities: Seemingly unrelated equations. Journal of the American Statistical Association, 86(414): 304–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brockmann:1993:LAB

[BGH93] Michael Brockmann, Theo Gasser, and Eva Herrmann. Locally adaptive bandwidth choice for kernel regression estimators. Journal of the American Statistical Association, 88(424):1302–1309, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Belin:1997:SDE

[BGH97] Thomas R. Belin, David W. Gjertson, and Ming Yi Hu. Summarizing DNA evidence when relatives are possible suspects. Journal of the American Statistical Association, 92(438):706–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berliner:1988:BNS

[BH88] L. Mark Berliner and Bruce M. Hill. Bayesian nonparametric survival analysis. with comments by Christian Genest, Jack Kalbfleisch and Seymour Geisser and with a reply by the authors.

Journal of the American Statistical Association, 83(403):772–784, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barry:1993:BAC

[BH93] Daniel Barry and J. A. Hartigan. A Bayesian analysis for change point problems. Journal of the American Statistical Association, 88(421):309–319, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:1994:RRR

[BH94] B. M. Brown and T. P. Hettmansperger. Regular redescending rank estimates. Journal of the American Statistical Association, 89(426):538–542, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:1996:NSN

[BH96] B. M. Brown and T. P. Hettmansperger. Normal scores, normal plots, and tests for normality. Journal of the American Statistical Association, 91(436):1668–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Butler:1997:SNM

[BH97] Ronald W. Butler and Aparna V. Huzurbazar. Stochastic network models for survival analysis. Journal of the American Statistical Association, 92(437):246–257, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/butler.htm.

Booth:1998:SEP

[BH98] James G. Booth and James P. Hobert. Standard errors of prediction in generalized linear mixed models. Journal of the American Statistical Association, 93(441):262–272, March 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/BOOTH.HTM.

Bhansali:1981:EKO

[Bha81] R. J. Bhansali. Effects of not knowing the order of an autoregressive process on the mean squared error of prediction. I. Journal of the American Statistical Association, 76(375):588–597, ????

1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhattacharyya:1985:AML

[Bha85] Gouri K. Bhattacharyya. The asymptotics of maximum likelihood and related estimators based on type ii censored data. Journal of the American Statistical Association, 80(390):398–404, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Begg:1996:SEP

[BHB96] Colin B. Begg, Ying Huang, and Marianne Berwick. Separate estimation of primary and secondary cancer preventive impact: Analysis of a case-control study of skin self- examination and melanoma. Journal of the American Statistical Association, 91 (436):1381–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:1992:CBS

[BHNO92] B. M. Brown, T. P. Hettmansperger, J. Nyblom, and H. Oja. On certain bivariate sign tests and medians. Journal of the American Statistical Association, 87(417):127–135, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bickel:1981:RDA

[BHS81] P. J. Bickel, Agnes M. Herzberg, and M. F. Schilling. Robustness of design against autocorrelation in time. II. optimality, theoretical and numerical results for the first-order autoregressive process. Journal of the American Statistical Association, 76(376):870–877, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Benjamini:1998:CIM

[BHS98] Y. Benjamini, Y. Hochberg, and P. B. Stark. Confidence intervals with more power to determine the sign: Two ends constrain the means. Journal of the American Statistical Association, 93 (441):309–317, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/HOCHBERG.HTM.

Brewer:1988:HNC

[BHT88] K. R. W. Brewer, Muhammad Hanif, and S. M. Tam. How nearly can model-based prediction and design-based estimation be reconciled. Journal of the American Statistical Association, 83(401): 128–132, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bierens:1983:UCK

[Bie83] Herman J. Bierens. Uniform consistency of kernel estimators of a regression function under generalized conditions. Journal of the American Statistical Association, 78(383):699–707, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Billard:1997:VD

[Bil97] Lynne Billard. A voyage of discovery. Journal of the American Statistical Association, 92(437):1–??, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/Billard/billard.html.

Binkley:1982:EVC

[Bin82] James K. Binkley. The effect of variable correlation on the efficiency of seemingly unrelated regression in a two-equation model. Journal of the American Statistical Association, 77(380):890–895, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Binder:1996:CVE

[Bin96] David A. Binder. Comment: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91 (434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Binder:1998:ACS

[Bin98] David A. Binder. Applications and case studies — comment: Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443):858–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Birnbaum:1954:SMP

[Bir54] A. Birnbaum. Statistical methods for Poisson processes and exponential populations. Journal of the American Statistical Association, ??(??):254–266, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhattacharyya:1974:ERM

[BJ74] G. K. Bhattacharyya and R. A. Johnson. Estimation of reliability in a multicomponent stress- strength model. Journal of the American Statistical Association, 69(348):966–970, December 1974. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bound:1995:PIV

[BJB95] John Bound, David A. Jaeger, and Regina M. Baker. Problems with instrumental variables estimation when the correlation between the instruments and the endogenous explanatory variable is weak. Journal of the American Statistical Association, 90(430): 443–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:1981:VDT

[BJM81] Lawrence D. Brown, Iain M. Johnstone, and K. Brenda MacGibbon. Variation diminishing transformations: a direct approach to total positivity and its statistical applications. Journal of the American Statistical Association, 76(376):824–832, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bjorck:1978:CIR

[Bjö78] Å. Björck. Comment on the iterative refinement of least squares solutions. Journal of the American Statistical Association, 73(??): 161–166, ???? 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bjornstad:1996:GLF

[Bjø96] Jan F. Bjørnstad. On the generalization of the likelihood function and the likelihood principle. Journal of the American Statistical Association, 91(434):791–806, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bassett:1982:EQF

[BK82] Gilbert Bassett, Jr. and Roger Koenker. An empirical quantile function for linear models with iid errors. Journal of the American Statistical Association, 77(378):407–415, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bonney:1984:IPC

[BK84] G. E. Bonney and G. E. Kissling. On the inverse of a patterned covariance matrix. Journal of the American Statistical Association, 79(387):722–723, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bendre:1985:MET

[BK85a] S. M. Bendre and B. K. Kale. Masking effect on tests for outliers in exponential models. Journal of the American Statistical Association, 80(392):1020–1025, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [BK86].

Butler:1985:DPP

[BK85b] William J. Butler and Richard A. Kronmal. Discrimination with polychotomous predictor variables using orthogonal functions. Journal of the American Statistical Association, 80(390): 443–448, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bendre:1986:CME

[BK86] S. M. Bendre and B. K. Kale. Correction: “Masking effect on tests for outliers in exponential samples” [J. Amer. Statist. Assoc. 80 (1985), no. 392, 1020–1025, MR 87c:62036]. Journal of the American Statistical Association, 81(396):1132, ???? 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [BK85a].

Bickel:1989:CBD

[BK89] P. J. Bickel and A. M. Krieger. Confidence bands for a distribution function using the bootstrap. Journal of the American Statistical Association, 84(405):95–100, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buhlmann:1999:CPS

[BK99] Peter Bühlmann and Hans R. Künsch. Comment: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):97–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bethlehem:1990:DCM

[BKP90] Jelke G. Bethlehem, Wouter J. Keller, and Jeroen Pannekoek. Disclosure control of microdata. Journal of the American Statistical Association, 85(409):38–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barrett:1992:GCI

[BL92] Bruce E. Barrett and Robert F. Ling. General classes of influence measures for multivariate regression. Journal of the American Statistical Association, 87(417):184–191, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Begg:1993:LEC

[BL93] Melissa Dowd Begg and Stephen Lagakos. Loss in efficiency caused by omitting covariates and misspecifying exposure in logistic regression models. Journal of the American Statistical Association, 88(421):166–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bar-Lev:1988:SPU

[BLE88] Shaul K. Bar-Lev and Peter Enis. Sign-preserving unbiased estimators in linear exponential families. Journal of the American Statistical Association, 83(404):1187–1189, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Blyth:1986:ABC

[Bly86] Colin R. Blyth. Approximate binomial confidence limits. Journal of the American Statistical Association, 81(395):843–855, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Blyth:1993:OKW

[Bly93] Stephen Blyth. Optimal kernel weights under a power criterion. Journal of the American Statistical Association, 88(424):1284–??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bradu:1970:ELL

[BM70] D. Bradu and Y. Mundlak. Estimation in lognormal linear models. Journal of the American Statistical Association, 65(329):198–211, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhattacharyya:1981:TET

[BM81] Gouri K. Bhattacharyya and Kishan G. Mehrotra. On testing equality of two exponential distributions under combined type ii censoring. Journal of the American Statistical Association, 76 (376):886–894, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhattacharyya:1983:UTL

[BM83] Gouri K. Bhattacharyya and Kishan G. Mehrotra. A unified treatment of locally most powerful rank tests under type ii censoring. Journal of the American Statistical Association, 78(382):375–381, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Basford:1985:EAR

[BM85] K. E. Basford and G. J. McLachlan. Estimation of allocation rates in a cluster analysis context. Journal of the American Statistical Association, 80(390):286–293, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barlow:1992:FTR

[BM92] Richard E. Barlow and Max B. Mendel. De Finetti-type representations for life distributions. Journal of the American Statistical Association, 87(420):1116–1122, December 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Billard:1995:ESM

[BM95] L. Billard and M. R. Meshkani. Estimation of a stationary Markov chain. Journal of the American Statistical Association, 90(429): 307–315, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berk:1996:DCD

[BM96] Robert Berk and Ruth Marcus. Dual cones, dual norms, and simultaneous inference for partially ordered means. Journal of the American Statistical Association, 91(433):318–328, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1999:DBF

[BM99] J. O. Berger and J. Mortera. Default Bayes factors for one-sided hypothesis testing. Journal of the American Statistical Association, 94(446):542–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berkson:1974:NME

[BN74] J. Berkson and B. N. Nagnur. A note on the minimum chi_{1}^{2} estimate and a L.A.M.S.T. chi^{2} in the no interaction problem. Journal of the American Statistical Association, 69(348):1038–1040, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bowman:1995:SMA

[BO95] Dale Bowman and E. Olusegun George. A saturated model for analyzing exchangeable binary data: Applications to clinical and developmental toxicity studies. Journal of the American Statistical Association, 90(431):871–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bose:1997:RBS

[BO97] Smarajit Bose and Finbarr O’Sullivan. A region-based segmentation method for multichannel image data. Journal of the American Statistical Association, 92(437):92–??, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/bose.htm.

Bofinger:1985:MCT

[Bof85] Eve Bofinger. Multiple comparisons and type iii errors. Journal of the American Statistical Association, 80(390):433–437, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boik:1986:TRM

[Boi86] Robert J. Boik. Testing the rank of a matrix with applications to the analysis of interaction in ANOVA. Journal of the American Statistical Association, 81(393):243–248, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bolstad:1995:MDP

[Bol95] William M. Bolstad. The multiprocess dynamic Poisson model. Journal of the American Statistical Association, 90(429):227–232, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boos:1981:MDE

[Boo81] Dennis D. Boos. Minimum distance estimators for location and goodness of fit. Journal of the American Statistical Association, 76(375):663–670, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boos:1982:TAA

[Boo82] Dennis D. Boos. A test for asymmetry associated with the Hodges-Lehmann estimator. Journal of the American Statistical Association, 77(379):647–651, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boos:1986:CPL

[Boo86] Dennis D. Boos. Comparing k populations with linear rank statistics. Journal of the American Statistical Association, 81(396): 1018–1025, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Baglivo:1988:MAC

[BOP88] Jenny Baglivo, Donald Olivier, and Marcello Pagano. Methods for the analysis of contingency tables with large and small cell counts. Journal of the American Statistical Association, 83(404): 1006–1013, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Baglivo:1992:MEG

[BOP92] Jenny Baglivo, Donald Olivier, and Marcello Pagano. Methods for exact goodness-of-fit tests. Journal of the American Statistical Association, 87(418):464–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bornholt:1987:CJB

[Bor87] Graham Bornholt. Comment on J. O. Berger and T. Sellke: “Testing a point null hypothesis: irreconcilability of P values and evidence” [J. Amer. Statist. Assoc. 82 (1987), no. 397, 112–139, MR 89a:62009a]. with a reply by Sellke. Journal of the American Statistical Association, 82(400):1201, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [BS87].

Bowden:1987:RSP

[Bow87] Roger J. Bowden. Repeated sampling in the presence of publication effects. with comments by Lawrence S. Cahoon and Peter C. Ordeshook. Journal of the American Statistical Association, 82(398):476–491, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:1986:ADE

[BP86] Philip J. Brown and Clive D. Payne. Aggregate data, ecological regression, and voting transitions. Journal of the American Statistical Association, 81(394):452–460, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bassett:1994:RS

[BP94a] Gilbert W. Bassett, Jr. and Joseph Persky. Rating skating. Journal of the American Statistical Association, 89(427):1075–1079, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Binder:1994:UEF

[BP94b] David A. Binder and Zdenek Patak. Use of estimating functions for estimation from complex surveys. Journal of the American Statistical Association, 89(427):1035–1043, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1996:IBF

[BP96] James O. Berger and Luis R. Pericchi. The intrinsic Bayes factor for model selection and prediction. Journal of the American Statistical Association, 91(433):109–122, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Balke:1997:BTE

[BP97] Alexander Balke and Judea Pearl. Bounds on treatment effects from studies with imperfect compliance. Journal of the American Statistical Association, 92(439):1171–1176, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/balke.htm.

Bartoszynski:1997:MGF

[BPL97] Robert Bartoszynski, Dennis K. Pearl, and John Lawrence. A multidimensional goodness-of-fit test based on interpoint distances. Journal of the American Statistical Association, 92(438):577–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Box:1987:ETG

[BPN87] George E. P. Box, David A. Pierce, and Paul Newbold. Estimating trend and growth rates in seasonal time series. Journal of the American Statistical Association, 82(397):276–282, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bertolino:1995:MBF

[BPR95] Francesco Bertolino, Ludovico Piccinato, and Walter Racugno. Multiple Bayes factors for testing hypotheses. Journal of the American Statistical Association, 90(429):213–219, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bertail:1999:SEU

[BPR99] Patrice Bertail, Dimitris N. Politis, and Joseph P. Romano. On subsampling estimators with unknown rate of convergence. Journal of the American Statistical Association, 94(446):569–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brunner:1995:NMS

[BPS95] Edgar Brunner, Madan L. Puri, and Shan Sun. Nonparametric methods for stratified two-sample designs with application to multiclinic trials. Journal of the American Statistical Association, 90 (431):1004–1014, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Baglivo:1996:PDG

[BPS96] Jenny Baglivo, Marcello Pagano, and Cathie Spino. Permutation distributions via generating functions with applications to sensitivity analysis of discrete data. Journal of the American Statistical Association, 91(435):1037–1046, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Begun:1983:ERR

[BR83] J. M. Begun and N. Reid. Estimating the relative risk with censored data. Journal of the American Statistical Association, 78 (382):337–341, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Banfield:1992:IFI

[BR92] Jeffrey D. Banfield and Adrian E. Raftery. Ice floe identification in satellite images using mathematical morphology and clustering about principal curves. Journal of the American Statistical Association, 87(417):7–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bandeen-Roche:1994:RAM

[BR94a] Karen Bandeen-Roche. Resolution of additive mixtures into source components and contributions: A compositional approach. Journal of the American Statistical Association, 89(428):1450–1458, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Basu:1994:RMS

[BR94b] Sabyasachi Basu and Gregory C. Reinsel. Regression models with spatially correlated errors. Journal of the American Statistical Association, 89(425):88–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Belin:1995:MCF

[BR95] Thomas R. Belin and Donald B. Rubin. A method for calibrating false-match rates in record linkage. Journal of the American Statistical Association, 90(430):694–??, ???? 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brumback:1998:RSS

[BR98a] Babette A. Brumback and John A. Rice. Rejoinder: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443): 991–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brumback:1998:SSM

[BR98b] Babette A. Brumback and John A. Rice. Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443):961– 994, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/brumback.html. With

comments and a rejoinder by the authors.

Byers:1998:NNC

[BR98c] Simon Byers and Adrian E. Raftery. Nearest neighbor clutter removal for estimating features in spatial point processes. Journal of the American Statistical Association, 93(442):577–584, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/june/BYERS.HTM.

Brant:1990:CCR

[Bra90] Rollin Brant. Comparing classical and resistant outlier rules. Journal of the American Statistical Association, 85(412):1083–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Beaton:1976:ARS

[BRB76] Albert E. Beaton, Donald B. Rubin, and John L. Barone. The acceptability of regression solutions: Another look at computational accuracy. Journal of the American Statistical Association, 71(353):158–168, ???? 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Breslow:1990:THO

[Bre90] Norman Breslow. Tests of hypotheses in overdispersed Poisson regression and other quasi-likelihood models. Journal of the American Statistical Association, 85(410):565–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Breiman:1992:LBO

[Bre92] Leo Breiman. The little bootstrap and other methods for dimensionality selection in regression: X-fixed prediction error. Journal of the American Statistical Association, 87(419):738–754, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Breslow:1996:SEC

[Bre96] N. E. Breslow. Statistics in epidemiology: The case-control study. Journal of the American Statistical Association, 91(433):14–28, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brehm:1998:ACS

[Bre98] John Brehm. Applications and case studies — comment: Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443):859–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brillinger:1992:NCS

[Bri92] David R. Brillinger. Nerve cell spike train data analysis: A progression of technique. Journal of the American Statistical Association, 87(418):260–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bandeen-Roche:1997:LVR

[BRMZR97] Karen Bandeen-Roche, Diana L. Miglioretti, Scott L. Zeger, and Paul J. Rathouz. Latent variable regression for multiple discrete outcomes. Journal of the American Statistical Association, 92(440):1375–1386, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/BANDEEN.HTM.

Brostrom:1981:OCC

[Bro81] Göran Broström. On optimal choice of components for parallel systems. Journal of the American Statistical Association, 76(374): 486–490, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:1983:SBD

[Bro83] Kenneth G. Brown. Sub-balanced data and the mixed analysis of variance. Journal of the American Statistical Association, 78(381): 162–167, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brody:1992:CMD

[Bro92] Janet Brody. Comment on Maxwell and Delaney. Journal of the American Statistical Association, 87(417):57–??, March 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brostrom:1997:MAC

[Bro97] Goran Brostrom. A martingale approach to the changepoint problem. Journal of the American Statistical Association, 92(439):

1177–1183, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/sept/brostrom.htm.

Babu:1992:NES

[BRR92] Gutti Jogesh Babu, C. Radhakrishna Rao, and M. Bhaskara Rao. Nonparametric estimation of specific occurrence/exposure rate in risk and survival analysis. Journal of the American Statistical Association, 87(417):84–89, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Blyth:1983:BCI

[BS83a] Colin R. Blyth and Harold A. Still. Binomial confidence intervals. Journal of the American Statistical Association, 78(381):108–116, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boyles:1983:MLE

[BS83b] Russell A. Boyles and Francisco J. Samaniego. Maximum likelihood estimation for a discrete multivariate shock model. Journal of the American Statistical Association, 78(382):445–448, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1984:THC

[BS84a] Roger L. Berger and Dennis F. Sinclair. Testing hypotheses concerning unions of linear subspaces. Journal of the American Statistical Association, 79(385):158–163, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boyles:1984:MIM

[BS84b] Russell A. Boyles and Francisco J. Samaniego. Modeling and inference for multivariate binary data with positive dependence. Journal of the American Statistical Association, 79(385):188–193, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Biemer:1985:ODI

[BS85] Paul P. Biemer and S. Lynne Stokes. Optimal design of interviewer variance experiments in complex surveys. Journal of the American Statistical Association, 80(389):158–166, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boyd:1986:UIR

[BS86a] Michael N. Boyd and Pranab K. Sen. Union-intersection rank tests for ordered alternatives in ANOCOVA. Journal of the American Statistical Association, 81(394):526–532, ???? 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boyles:1986:EDF

[BS86b] Russell A. Boyles and Francisco J. Samaniego. Estimating a distribution function based on nomination sampling. Journal of the American Statistical Association, 81(396):1039–1045, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1987:TPN

[BS87] James O. Berger and Thomas Sellke. Testing a point null hypothesis: irreconcilability of P values and evidence. with comments and a rejoinder by the authors. Journal of the American Statistical Association, 82(397):112–139, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comment [Bor87].

Bradley:1991:IAD

[BS91] Ralph A. Bradley and Frances P. Stewart. Intrablock analysis of designs with multiple blocking criteria. Journal of the American Statistical Association, 86(415):792–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1993:BAP

[BS93] James O. Berger and Dongchu Sun. Bayesian analysis for the Poly-Weibull distribution. Journal of the American Statistical Association, 88(424):1412–1418, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brooks:1994:JCR

[BS94] Rodney Brooks and Mervyn Stone. Joint continuum regression for multiple predictants. Journal of the American Statistical Association, 89(428):1374–1377, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Benjamini:1996:NSC

[BS96a] Yoav Benjamini and Philip B. Stark. Nonequivariant simultaneous confidence intervals less likely to contain zero. Journal of the American Statistical Association, 91(433):329–337, ???? 1996.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bolviken:1996:CIM

[BS96b] Erik Bølviken and Eva Skovlund. Confidence intervals from Monte Carlo tests. Journal of the American Statistical Association, 91 (435):1071–1078, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buzas:1996:IVE

[BS96c] Jeffrey S. Buzas and Leonard A. Stefanski. Instrumental variable estimation in generalized linear measurement error models. Journal of the American Statistical Association, 91(435):999–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1997:ITS

[BS97a] Vance Berger and Harold Sackrowitz. Improving tests for superior treatment in contingency tables. Journal of the American Statistical Association, 92(438):700–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bloch:1997:MDF

[BS97b] Daniel A. Bloch and Bernard W. Silverman. Monotone discriminant functions and their applications in rheumatology. Journal of the American Statistical Association, 92(437):144–153, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/bloch.htm.

Blyth:1997:HEA

[BS97c] Colin R. Blyth and Robert G. Staudte. Hypothesis estimates and acceptability profiles for 2 × 2 contingency tables. Journal of the American Statistical Association, 92(438):694–??, June 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Butler:1998:SAM

[BS98] Ronald W. Butler and Richard K. Sutton. Saddlepoint approximation for multivariate cumulative distribution functions and probability computations in sampling theory and outlier testing. Journal of the American Statistical Association, 93 (442):596–604, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/BUTLER.HTM.

Begg:1998:NSE

[BSB98] Colin B. Begg, Jaya M. Satagopan, and Marianne Berwick. A new strategy for evaluating the impact of epidemiologic risk factors for cancer with application to melanoma. Journal of the American Statistical Association, 93(442):415–426, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/june/BEGG.HTM.

Bailey:1998:SDU

[BSPK98] Trevor C. Bailey, Theofanis Sapatinas, Kenneth J. Powell, and Wojtek J. Krzanowski. Signal detection in underwater sounds using wavelets. Journal of the American Statistical Association, 93(441):73–83, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/BAILEY.HTM.

Bowden:1981:CSI

[BT81] R. J. Bowden and D. A. Turkington. A comparative study of instrumental variables estimators for nonlinear simultaneous models. Journal of the American Statistical Association, 76(376):988–995, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buckland:1995:CID

[Buc95a] Stephen T. Buckland. Comment: Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90(430):417–??, ???? 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buckle:1995:BIS

[Buc95b] D. J. Buckle. Bayesian inference for stable distributions. Journal of the American Statistical Association, 90(430):605–613, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buehler:1982:SAS

[Bue82] Robert J. Buehler. Some ancillary statistics and their properties. with comments by O. Barndorff-Nielsen, D. R. Cox and David Hinkley and a rejoinder by the author. Journal of the American Statistical Association, 77(379):581–594, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buonaccorsi:1990:DSE

[Buo90] John P. Buonaccorsi. Double sampling for exact values in some multivariate measurement error problems. Journal of the American Statistical Association, 85(412):1075–??, December 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buonaccorsi:1996:MER

[Buo96] John P. Buonaccorsi. Measurement error in the response in the general linear model. Journal of the American Statistical Association, 91(434):633–642, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Burr:1994:CCB

[Bur94] Deborah Burr. A comparison of certain bootstrap confidence intervals in the Cox model. Journal of the American Statistical Association, 89(428):1290–1302, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Butler:1984:SAB

[But84] Ronald W. Butler. The significance attained by the best-fitting regressor variable. Journal of the American Statistical Association, 79(386):341–348, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Beirlant:1996:TIE

[BVT96] Jan Beirlant, Petra Vynckier, and Jozef L. Teugels. Tail index estimation, Pareto quantile plots, and regression diagnostics. Journal of the American Statistical Association, 91(436):1659–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bohn:1992:NTS

[BW92] Lora L. Bohn and Douglas A. Wolfe. Nonparametric two-sample procedures for ranked-set samples data. Journal of the American Statistical Association, 87(418):552–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bohn:1994:EIJ

[BW94] Lora L. Bohn and Douglas A. Wolfe. The effect of imperfect judgment rankings on properties of procedures based on the ranked-set samples analog of the Mann–Whitney–Wilcoxon statistic. Journal of the American Statistical Association, 89(425):168–176, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bradlow:1998:ACS

[BWC98] Eric T. Bradlow, Robert E. Weiss, and Meehyung Cho. Applications and case studies — Bayesian identification of outliers in computerized adaptive tests. Journal of the American Statistical Association, 93(443):910–919, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/bradlow.html.

Bustos:1986:REA

[BY86] Oscar H. Bustos and Víctor J. Yohai. Robust estimates for ARMA models. Journal of the American Statistical Association, 81(393): 155–168, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bickel:1988:REB

[BY88] Peter J. Bickel and Joseph A. Yahav. Richardson extrapolation and the bootstrap. Journal of the American Statistical Association, 83(402):387–393, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bewley:1995:TCB

[BY95] Ronald Bewley and Minxian Yang. Tests for cointegration based on canonical correlation analysis. Journal of the American Statistical Association, 90(431):990–996, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bickel:1992:VSN

[BZ92] Peter Bickel and Ping Zhang. Variable selection in nonparametric regression with categorical covariates. Journal of the American Statistical Association, 87(417):90–97, March 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bradlow:1999:HLV

[BZ99] Eric T. Bradlow and Alan M. Zaslavsky. A hierarchical latent variable model for ordinal data from a customer satisfaction survey with “no answer” responses. Journal of the American Statistical Association, 94(445):43–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1982:PPT

[Car82] Raymond J. Carroll. Prediction and power transformations when the choice of power is restricted to a finite set. Journal of the American Statistical Association, 77(380):908–915, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carriere:1994:IRM

[Car94] Keumhee Chough Carriere. Incomplete repeated measures data analysis in the presence of treatment effects. Journal of the American Statistical Association, 89(426):680–??, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Casella:1985:CNM

[Cas85] George Casella. Condition numbers and minimax ridge regression estimators. Journal of the American Statistical Association, 80 (391):753–758, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Casella:1987:RBF

[CB87] George Casella and Roger L. Berger. Reconciling Bayesian and frequentist evidence in the one-sided testing problem. with comments and a rejoinder by the authors. Journal of the American Statistical Association, 82(397):106–111, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clutton-Brock:1990:DEU

[CB90] M. Clutton-Brock. Density estimation using exponentials of orthogonal series. Journal of the American Statistical Association, 85(411):760–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Casella:1994:ESB

[CB94] George Casella and Roger L. Berger. Estimation with selected binomial information or do you really believe that Dave Windfield is batting .471? Journal of the American Statistical Association, 89(427):1080–1090, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Christensen:1997:TIA

[CB97] Ronald Christensen and Edward J. Bedrick. Testing the independence assumption in linear models. Journal of the American Statistical Association, 92(439):1006–1016, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/christensen.htm.

Chan:1974:ADA

[CC74] L. K. Chan and S. W. Cheng. An algorithm for determining the asymptotically best linear estimate of the mean from multiply censored logistic data. Journal of the American Statistical Association, 69(348):1027–1030, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cobb:1983:SSQ

[CC83] E. Benton Cobb and J. D. Church. Small-sample quantal response methods for estimating the location parameter for a location-scale family of dose-response curves. Journal of the American Statistical Association, 78(381):99–107, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clyde:1996:ECW

[CC96a] Merlise Clyde and Kathryn Chaloner. The equivalence of constrained and weighted designs in multiple objective design problems. Journal of the American Statistical Association, 91(435): 1236–1244, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cowles:1996:MCM

[CC96b] Mary Kathryn Cowles and Bradley P. Carlin. Markov chain Monte Carlo convergence diagnostics: a comparative review. Journal of the American Statistical Association, 91(434):883–904, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cutler:1996:MHD

[CCB96] Adele Cutler and Olga I. Cordero-Brana. Minimum Hellinger distance estimation for finite mixture models. Journal of the American Statistical Association, 91(436):1716–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cowles:1996:BTM

[CCC96] Mary Kathryn Cowles, Bradley P. Carlin, and John E. Connett. Bayesian tobit modeling of longitudinal ordinal clinical trial compliance data with nonignorable missingness. Journal of the American Statistical Association, 91(433):86–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1997:NMB

[CCD97] Zhao-Guo Chen, Pierre A. Cholette, and Estela Bee Dagum. A nonparametric method for benchmarking survey data via signal extraction. Journal of the American Statistical Association, 92(440):1563–1571, December 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.amstat.org/publications/jasa/abstracts_97/
december/CHEN_ZG.HTM.

Cook:1998:PRP

[CCD98] R. Dennis Cook and Rodney Croos-Dabrera. Partial residual plots in generalized linear models. Journal of the American Statistical Association, 93(442):730–739, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/june/COOK.HTM.

Causey:1985:ATT

[CCE85] Beverley D. Causey, Lawrence H. Cox, and Lawrence R. Ernst. Applications of transportation theory to statistical problems. Journal of the American Statistical Association, 80(392):903–909, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clogg:1990:ER

[CCG90] Clifford C. Clogg, Raymond J. Carroll, and Donald Guthrie. Editors’ report for 1989. Journal of the American Statistical Association, 85(410):273–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1997:OEP

[CCG^{+}97a] R. J. Carroll, R. Chen, E. I. George, T. H. Li, H. J. Newton, H. Schmiediche, and N. Wang. Ozone exposure and population density in Harris County, Texas. Journal of the American Statistical Association, 92(438):392–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1997:ROE

[CCG^{+}97b] R. J. Carroll, R. Chen, E. I. George, T. H. Li, H. J. Newton, H. Schmiediche, and N. Wang. Rejoinder: Ozone exposure and population density in Harris County, Texas. Journal of the American Statistical Association, 92(438):414–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1990:DAA

[CCL90] R. Dennis Cook and Tsai Chih-Ling. Diagnostics for assessing the accuracy of normal approximations in exponential family nonlinear models. Journal of the American Statistical Association, 85 (411):770–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Campbell:1989:CEM

[CD89] M. Karen Campbell and Allan Donner. Classification efficiency of multinomial logistic regression relative to ordinal logistic regression. Journal of the American Statistical Association, 84(406): 587–591, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Calvin:1991:LSE

[CD91] James A. Calvin and Richard L. Dykstra. Least squares estimation of covariance matrices in balanced multivariate variance components models. Journal of the American Statistical Association, 86 (414):388–395, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Calvin:1995:REC

[CD95] James A. Calvin and Richard L. Dykstra. REML estimation of covariance matrices with restricted parameter spaces. Journal of the American Statistical Association, 90(429):321–329, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cragg:1996:APL

[CD96] John G. Cragg and Stephen G. Donald. On the asymptotic properties of LDU-Based tests of the rank of a matrix. Journal of the American Statistical Association, 91(435):1301–1309, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clyde:1996:POM

[CDP96] Merlise Clyde, Heather Desimone, and Giovanni Parmigiani. Prediction via orthogonalized model mixing. Journal of the American Statistical Association, 91(435):1197–??, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chambers:1993:BRE

[CDW93] Raymond L. Chambers, Alan H. Dorfman, and Thomas E. Wehrly. Bias robust estimation in finite populations using nonparametric calibration. Journal of the American Statistical Association, 88 (421):268–277, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clayton:1987:NEP

[CF87] Murray K. Clayton and Edward W. Frees. Nonparametric estimation of the probability of discovering a new species. Journal of the American Statistical Association, 82(397):305–311, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cohen:1995:THU

[CF95] Arthur Cohen and Mendel Fygenson. Testing homogeneity of uniform scale distributions against two- sided and one-sided alternatives. Journal of the American Statistical Association, 90(431): 1062–1067, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1997:GPL

[CFGW97] R. J. Carroll, Jianqing Fan, Irene Gijbels, and M. P. Wand. Generalized partially linear single-index models. Journal of the American Statistical Association, 92(438):477–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crook:1982:PST

[CG82] James Flinn Crook and Irving John Good. The powers and strengths of tests for multinomials and contingency tables. Journal of the American Statistical Association, 77(380):793–802, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clogg:1984:LSA

[CG84] Clifford C. Clogg and Leo A. Goodman. Latent structure analysis of a set of multidimensional contingency tables. Journal of the American Statistical Association, 79(388):762–780, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carlin:1990:AEB

[CG90] Bradley P. Carlin and Alan E. Gelfand. Approaches for empirical Bayes confidence intervals. Journal of the American Statistical Association, 85(409):105–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clapp:1992:EPI

[CG92] John M. Clapp and Carmelo Giaccotto. Estimating price indices for residential property: A comparison of repeat sales and assessed value methods. Journal of the American Statistical Association, 87(418):300–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1997:TLV

[CG97] Jie Chen and A. K. Gupta. Testing and locating variance changepoints with application to stock prices. Journal of the American Statistical Association, 92(438):738–??, June 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1985:CLS

[CGG85] Raymond J. Carroll, Paul Gallo, and Leon Jay Gleser. Comparison of least squares and errors-in-variables regression, with special reference to randomized analysis of covariance. Journal of the American Statistical Association, 80(392):929–932, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chu:1998:EPS

[CGGM98a] C. K. Chu, I. K. Glad, F. Godtliebsen, and J. S. Marron. Edge preserving smoothers for image processing. Journal of the American Statistical Association, 93(442):526–556, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_

98/june/CHU.HTM. With comments and a rejoinder by the authors.

Chu:1998:REP

[CGGM98b] C. K. Chu, I. K. Glad, F. Godtliebsen, and J. S. Marron. Rejoinder: Edge preserving smoothers for image processing. Journal of the American Statistical Association, 93(442):553–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1993:CCS

[CGL93] R. J. Carroll, M. H. Gail, and J. H. Lubin. Case-control studies with errors in covariates. Journal of the American Statistical Association, 88(421):185–199, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cohen:1983:HTM

[CGM83] Arthur Cohen, Constantine Gatsonis, and John I. Marden. Hypothesis testing for marginal probabilities in a 2 × 2 × 2 contingency table with conditional independence. Journal of the American Statistical Association, 78(384):920–929, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chipman:1998:BCM

[CGM98a] Hugh A. Chipman, Edward I. George, and Robert E. McCulloch. Bayesian CART model search. Journal of the American Statistical Association, 93(443):935–947, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/chipman.html.

Chipman:1998:RBC

[CGM98b] Hugh A. Chipman, Edward I. George, and Robert E. McCulloch. Rejoinder: Bayesian CART model search. Journal of the American Statistical Association, 93(443):957–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Casella:1983:EBC

[CH83] George Casella and Jiunn Tzon Hwang. Empirical Bayes confidence sets for the mean of a multivariate normal distribution. Journal of the American Statistical Association, 78(383):688–698, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1988:ORC

[CH88] Raymond J. Carroll and Peter Hall. Optimal rates of convergence for deconvolving a density. Journal of the American Statistical Association, 83(404):1184–1186, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cameron:1991:RCV

[CH91] Trudy Ann Cameron and D. D. Huppert. Referendum contingent valuation estimates: Sensitivity to the assignment of offered values. Journal of the American Statistical Association, 86(416):910– ??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cressie:1992:SSA

[CH92] Noel Cressie and Frederick L. Hulting. A spatial statistical analysis of tumor growth. Journal of the American Statistical Association, 87(418):272–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Coakley:1993:BIH

[CH93] Clint W. Coakley and Thomas P. Hettmansperger. A bounded influence, high breakdown, efficient regression estimator. Journal of the American Statistical Association, 88(423):872–880, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Craig:1994:TTD

[CH94] Lee A. Craig and Alastair R. Hall. Trying out for the team: Do exhibitions matter? evidence from the National Football League. Journal of the American Statistical Association, 89(427):1091–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Castillo:1995:MLD

[CH95a] Enrique Castillo and Ali S. Hadi. Modeling lifetime data with application to fatigue models. Journal of the American Statistical Association, 90(431):1041–1054, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clayton:1995:CIA

[CH95b] Murray K. Clayton and Brian D. Hudelson. Confidence intervals for autocorrelations based on cyclic samples. Journal of the American Statistical Association, 90(430):753–757, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Castillo:1997:FGP

[CH97a] Enrique Castillo and Ali S. Hadi. Fitting the generalized Pareto distribution to data. Journal of the American Statistical Association, 92(440):1609–1620, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/CASTILLO.HTM.

Chang:1997:ILE

[CH97b] Yi-Ping Chang and Wen-Tao Huang. Inferences for the linear errors-in-variables with changepoint models. Journal of the American Statistical Association, 92(437):171–178, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/chang.htm.

Chambers:1971:RU

[Cha71] J. M. Chambers. Regression updating. Journal of the American Statistical Association, 66(??):744–748, ???? 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chang:1984:TML

[Cha84] Y. C. Chang. A truncated maximum likelihood estimator of a constrained bivariate linear regression coefficient. Journal of the American Statistical Association, 79(386):454–458, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chambers:1986:ORF

[Cha86] Raymond L. Chambers. Outlier robust finite population estimation. Journal of the American Statistical Association, 81(396): 1063–1069, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:1988:PIN

[Cha88a] Ngai Hang Chan. The parameter inference for nearly nonstationary time series. Journal of the American Statistical Association, 83(403):857–862, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chang:1988:ERR

[Cha88b] Ted Chang. Estimating the relative rotation of two tectonic plates from boundary crossings. Journal of the American Statistical Association, 83(404):1178–1183, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:1990:INI

[Cha90] Ngai Hang Chan. Inference for near-integrated time series with infinite variance. Journal of the American Statistical Association, 85(412):1069–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chang:1991:WCE

[Cha91] Myron N. Chang. On weak convergence of an estimator of the survival function when new is better than used of a specified age. Journal of the American Statistical Association, 86(413):173–178, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:1993:ABG

[Cha93] K. S. Chan. Asymptotic behavior of the Gibbs sampler. Journal of the American Statistical Association, 88(421):320–326, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chaudhuri:1996:GNQ

[Cha96] Probal Chaudhuri. On a geometric notion of quantiles for multivariate data. Journal of the American Statistical Association, 91 (434):862–872, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chew:1970:CME

[Che70] V. Chew. Covariance matrix estimation in linear models. Journal of the American Statistical Association, 65(329):173–181, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1971:RNN

[Che71] E. H. Chen. A random normal number generator for 32-bit-word computers. Journal of the American Statistical Association, 66 (334):400–403, ???? 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1990:AAI

[Che90] Han Feng Chen. The accuracy of approximate intervals for a binomial parameter. Journal of the American Statistical Association, 85(410):514–518, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1994:IWM

[Che94a] Ming-Hui Chen. Importance-weighted marginal Bayesian posterior density estimation. Journal of the American Statistical Association, 89(427):818–824, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:1994:NEM

[Che94b] Philip E. Cheng. Nonparametric estimation of mean functionals with data missing at random. Journal of the American Statistical Association, 89(425):81–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1995:TMS

[Che95] Ling Chen. Testing the mean of skewed distributions. Journal of the American Statistical Association, 90(430):767–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chitturi:1974:DRA

[Chi74] R. V. Chitturi. Distribution of residual autocorrelations in multiple autoregressive schemes. Journal of the American Statistical Association, 69(348):928–934, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chib:1995:MLG

[Chi95] Siddhartha Chib. Marginal likelihood from the Gibbs output. Journal of the American Statistical Association, 90(432):1313– 1321, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1982:SSR

[CHL82] Y. Y. Chen, M. Hollander, and N. A. Langberg. Small-sample results for the Kaplan–Meier estimator. Journal of the American Statistical Association, 77(377):141–144, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chong:1992:DNR

[Cho92] Gu Chong. Diagnostics for nonparametric regression models with additive terms. Journal of the American Statistical Association, 87(420):1051–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chong:1993:SSD

[Cho93] Gu Chong. Smoothing spline density estimation: A dimensionless automatic algorithm. Journal of the American Statistical Association, 88(422):495–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cowling:1996:BCR

[CHP96] Ann Cowling, Peter Hall, and Michael J. Phillips. Bootstrap confidence regions for the intensity of a Poisson point process. Journal of the American Statistical Association, 91(436):1516–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Christensen:1987:ATS

[Chr87] Ronald Christensen. The analysis of two-stage sampling data by ordinary least squares. Journal of the American Statistical Association, 82(398):492–498, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Christensen:1991:SSC

[Chr91] Ronald Christensen. Small-sample characterizations of near replicate lack-of-fit tests. Journal of the American Statistical Association, 86(415):752–756, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1994:ELM

[CHR94] R. J. Carroll, Peter Hall, and D. Ruppert. Estimation of lag in misregistered, averaged images. Journal of the American Statistical Association, 89(425):219–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Christensen:1999:CPS

[Chr99] Ronald Christensen. Comment: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):99–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1992:CMM

[CHW92] R. D. Cook, D. M. Hawkins, and S. Weisberg. Comparison of model misspecification diagnostics using residuals from least mean of squares and least median of squares fits. Journal of the American Statistical Association, 87(418):419–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:1988:CTF

[CJ88] Ching-Shui Cheng and Mike Jacroux. The construction of trendfree run orders of two-level factorial designs. Journal of the American Statistical Association, 83(404):1152–1158, ???? 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1995:DLC

[CJ95a] Jian Shen Chen and Robert I. Jennrich. Diagnostics for linearization confidence intervals in nonlinear regression. Journal of the American Statistical Association, 90(431):1068–1074, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1995:TIL

[CJ95b] Jian-Shen Chen and Robert I. Jennrich. Transformations for improving linearization confidence intervals in nonlinear regression. Journal of the American Statistical Association, 90(432):1271– 1276, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1996:SRD

[CJ96] Jian-Shen Chen and Robert I. Jennrich. The signed root deviance profile and confidence intervals in maximum likelihood analysis. Journal of the American Statistical Association, 91(435):993–998, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carlstein:1992:BE

[CK92] E. Carlstein and C. Krishnamoorthy. Boundary estimation. Journal of the American Statistical Association, 87(418):430–438, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cobb:1983:EMR

[CKC83] Loren Cobb, Peter Koppstein, and Neng Hsin Chen. Estimation and moment recursion relations for multimodal distributions of the exponential family. Journal of the American Statistical Association, 78(381):124–130, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carlin:1992:PWM

[CKLH92] Bradley P. Carlin, Robert E. Kass, F. Javier Lerch, and Brian R. Huguenard. Predicting working memory failure: A subjective Bayesian approach to model selection. Journal of the American Statistical Association, 87(418):319–??, June 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1996:ASE

[CKLS96] Raymond J. Carroll, Helmut Küchenhoff, F. Lombard, and Leonard A. Stefanski. Asymptotics for the SIMEX estimator in nonlinear measurement error models. Journal of the American Statistical Association, 91(433):242–250, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chipman:1997:ABW

[CKM97] Hugh A. Chipman, Eric D. Kolaczyk, and Robert E. McCulloch. Adaptive Bayesian wavelet shrinkage. Journal of the American Statistical Association, 92(440):1413–1421, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/CHIPMAN.HTM.

Carroll:1985:NEB

[CL85] Raymond J. Carroll and F. Lombard. A note on N estimators for the binomial distribution. Journal of the American Statistical Association, 80(390):423–426, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1992:MER

[CL92a] Raymond J. Carroll and Ker-Chau Li. Measurement error regression with unknown link: Dimension reduction and data visualization. Journal of the American Statistical Association, 87(420): 1040–1050, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chao:1992:ENC

[CL92b] Anne Chao and Shen-Ming Lee. Estimating the number of classes via sample coverage. Journal of the American Statistical Association, 87(417):210–217, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:1995:MCE

[CL95] K. S. Chan and Johannes Ledolter. Monte Carlo EM estimation for time series models involving counts. Journal of the American Statistical Association, 90(429):242–252, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clarke:1987:ACL

[Cla87a] G. P. Y. Clarke. Approximate confidence limits for a parameter function in nonlinear regression. Journal of the American Statistical Association, 82(397):221–230, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clarke:1987:MCT

[Cla87b] G. P. Y. Clarke. Marginal curvatures and their usefulness in the analysis of nonlinear regression models. Journal of the American Statistical Association, 82(399):844–850, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [Cla89].

Clarke:1989:CMC

[Cla89] G. P. Y. Clarke. Correction to: “Marginal curvatures and their usefulness in the analysis of nonlinear regression models” [J. Amer. Statist. Assoc. 82 (1987), no. 399, 844–850, MR 88k:62104]. Journal of the American Statistical Association, 84(405):352, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Cla87b].

Clarke:1996:IRP

[Cla96] Bertrand Clarke. Implications of reference priors for prior information and for sample size. Journal of the American Statistical Association, 91(433):173–184, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chung:1993:JEM

[CLM93] Chen Chung and Liu Lon-Mu. Joint estimation of model parameters and outlier effects in time series. Journal of the American Statistical Association, 88(421):284–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carr:1987:SMT

[CLNL87] D. B. Carr, R. J. Littlefield, W. L. Nicholson, and J. S. Littlefield. Scatterplot matrix techniques for large N. Journal of the American Statistical Association, 82(398):424–436, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clogg:1982:SMA

[Clo82] Clifford C. Clogg. Some models for the analysis of association in multiway cross-classifications having ordered categories. Journal of the American Statistical Association, 77(380):803–815, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chiu:1996:MLC

[CLT96] Tom Y. M. Chiu, Tom Leonard, and Kam-Wah Tsui. The matrixlogarithmic covariance model. Journal of the American Statistical Association, 91(433):198–210, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cole:1995:EBM

[CLWZ95] Bernard F. Cole, Mei-Ling T. Lee, G. Alex Whitmore, and Alan M. Zaslavsky. An empirical Bayes model for Markov-dependent binary sequences with randomly missing observations. Journal of the American Statistical Association, 90(432):1364–1372, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chhikara:1992:EUM

[CLY92] Raj S. Chhikara and Deng Lih-Yuan. Estimation using multiyear rotation design sampling in agricultural surveys. Journal of the American Statistical Association, 87(420):924–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chhikara:1984:LDA

[CM84a] Raj S. Chhikara and Jim McKeon. Linear discriminant analysis with misallocation in training samples. Journal of the American Statistical Association, 79(388):899–906, ???? 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cleveland:1984:MFS

[CM84b] William S. Cleveland and Robert McGill. The many faces of a scatterplot. Journal of the American Statistical Association, 79 (388):807–822, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Collings:1985:TGF

[CM85] Bruce J. Collings and Barry H. Margolin. Testing goodness of fit for the Poisson assumption when observations are not identically distributed. Journal of the American Statistical Association, 80 (390):411–418, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cowan:1986:CRM

[CM86] Charles D. Cowan and Donald Malec. Capture-recapture models when both sources have clustered observations. Journal of the American Statistical Association, 81(394):347–353, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chung:1991:UNM

[CM91a] Lyinn Chung and John I. Marden. Use of nonnull models for rank statisitics in bivariate, two-sample, and analysis of variance problems. Journal of the American Statistical Association, 86(413): 188–200, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Corder:1991:NSH

[CM91b] Larry S. Corder and Kenneth G. Manton. National surveys and the health and functioning of the elderly: The effects of design and content. Journal of the American Statistical Association, 86(414): 513–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chaudhuri:1993:NEO

[CM93] Probal Chaudhuri and Per A. Mykland. Nonlinear experiments: Optimal design and inference based on likelihood. Journal of the American Statistical Association, 88(422):538–546, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Capra:1997:ATM

[CM97a] William B. Capra and Hans-Georg Müller. An acceleratedtime model for response curves. Journal of the American Statistical Association, 92(437):72–83, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/capra.htm.

Choi:1997:AMR

[CM97b] K. Choi and J. Marden. An approach to multivariate rank tests in multivariate analysis of variance. Journal of the American Statistical Association, 92(440):1581–1590, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/CHOI_K.HTM.

Christiansen:1997:HPR

[CM97c] Cindy L. Christiansen and Carl N. Morris. Hierarchical Poisson regression modeling. Journal of the American Statistical Association, 92(438):618–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chiou:1998:QLR

[CM98] Jeng-Min Chiou and Hans-Georg Müller. Quasi-likelihood regression with unknown link and variance functions. Journal of the American Statistical Association, 93(444):1376–1387, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/dec/chiou.html.

Currin:1991:BPD

[CMMY91] Carla Currin, Toby Mitchell, Max Morris, and Don Ylvisaker. Bayesian prediction of deterministic functions, with applications to the design and analysis of computer experiments. Journal of the American Statistical Association, 86(416):953–963, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chang:1999:HBR

[CMNS99] William H. Chang, Joseph W. McKean, Joshua D. Naranjo, and Simon J. Sheather. High-breakdown rank regression. Journal of the American Statistical Association, 94(445):205–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chambers:1976:MSS

[CMS76] J. M. Chambers, C. L. Mallows, and B. W. Stuck. A method for simulating stable random variables. Journal of the American Statistical Association, 71(354):340–344, ???? 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [CMS87].

Chambers:1987:CMS

[CMS87] J. M. Chambers, C. L. Mallows, and B. W. Stuck. Correction to: “A method for simulating stable random variables” [J. Amer. Statist. Assoc. 71 (1976), no. 354, 340–344, MR 54 #4059]. Journal of the American Statistical Association, 82(398):704, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CMS76].

Cheng:1988:OST

[CMST88] Ching-Shui Cheng, Dibyen Majumdar, John Stufken, and Tahsin Erkan Türe. Optimal step-type designs for comparing test treatments with a control. Journal of the American Statistical Association, 83(402):477–482, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cochran:1953:SPK

[CMT53] W. G. Cochran, F. Mosteller, and John W. Tukey. Statistical problems of the Kinsey report. Journal of the American Statistical Association, 48(??):673–716, ???? 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cochran:1954:PS

[CMT54] W. G. Cochran, F. Mosteller, and John W. Tukey. Principles of sampling. Journal of the American Statistical Association, 49(??): 13–35, ???? 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cochran:1955:CMT

[CMT55] W. G. Cochran, F. Mosteller, and John W. Tukey. The Cochran– Mosteller–Tukey report on the Kinsey Study: A symposium (pp. 811–829). Journal of the American Statistical Association, 50(??): 829, ???? 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cargnoni:1997:BFM

[CMW97] Claudia Cargnoni, Peter Muller, and Mike West. Bayesian forecasting of multinomial time series through conditionally Gaussian dynamic models. Journal of the American Statistical Association, 92(438):640–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1994:RAE

[CN94] R. Dennis Cook and Christopher J. Nachtsheim. Reweighting to achieve elliptically contoured covariates in regression. Journal of the American Statistical Association, 89(426):592–??, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chappell:1995:MDL

[CNF95] Rick Chappell, David M. Nondahl, and John F. Fowler. Modeling dose and local control in radiotherapy. Journal of the American Statistical Association, 90(431):829–??, ???? 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cryer:1990:FES

[CNS90] Jonathan D. Cryer, John C. Nankervis, and N. E. Savin. Forecast error symmetry in ARIMA models. Journal of the American Statistical Association, 85(411):724–728, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cohen:1998:ACS

[CNWW98] Jacqueline Cohen, Daniel Nagin, Garrick Wallstrom, and Larry Wasserman. Applications and case studies — hierarchical Bayesian analysis of arrest rates. Journal of the American Statistical Association, 93(444):1260–1270, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/dec/cohen.html.

Cain:1990:RLB

[CO90] Michael Cain and Roger J. Owen. Regressor levels for Bayesian predictive response. Journal of the American Statistical Association, 85(409):228–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cohen:1981:IMM

[Coh81] Arthur Cohen. Inference for marginal means in contingency tables with conditional independence. Journal of the American Statistical Association, 76(376):895–902, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Collings:1987:CRN

[Col87] Bruce Jay Collings. Compound random number generators. Journal of the American Statistical Association, 82(398):525–527, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289456.

Conaway:1989:ARC

[Con89] Mark R. Conaway. Analysis of repeated categorical measurements with conditional likelihood methods. Journal of the American Statistical Association, 84(405):53–62, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Conaway:1992:ARC

[Con92] Mark R. Conaway. The analysis of repeated categorical measurements subject to nonignorable nonresponse. Journal of the American Statistical Association, 87(419):817–??, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1979:IOL

[Coo79] R. D. Cook. Influential observations in linear regression. Journal of the American Statistical Association, 74(365):169–174, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cooil:1991:UMM

[Coo91] Bruce Cooil. Using medical malpractice data to predict the frequency of claims: A study of Poisson process models. Journal of the American Statistical Association, 86(414):285–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1994:IRP

[Coo94] R. Dennis Cook. On the interpretation of regression plots. Journal of the American Statistical Association, 89(425):177–189, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1996:GRB

[Coo96] R. Dennis Cook. Graphics for regressions with a binary response. Journal of the American Statistical Association, 91(435):983–992, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1998:PHD

[Coo98a] R. Dennis Cook. Principle Hessian directions revisited. Journal of the American Statistical Association, 93(441):84–100, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/

abstracts_98/COOK.HTM. With comments by Ker-Chau Li and a rejoinder by the author.

Cook:1998:RPH

[Coo98b] R. Dennis Cook. Rejoinder: Principle Hessian directions revisited. Journal of the American Statistical Association, 93(441): 98–??, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cox:1995:NMC

[Cox95] Lawrence H. Cox. Network models for complementary cell suppression. Journal of the American Statistical Association, 90(432): 1453–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carlin:1991:EUA

[CP91] Bradley P. Carlin and Nicholas G. Polson. An expected utility approach to influence diagnostics. Journal of the American Statistical Association, 86(416):1013–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carota:1996:DMM

[CPP96] Cinzia Carota, Giovanni Parmigiani, and Nicholas G. Polson. Diagnostic measures for model criticism. Journal of the American Statistical Association, 91(434):753–762, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carlin:1992:MCA

[CPS92] Bradley P. Carlin, Nicholas G. Polson, and David S. Stoffer. A Monte Carlo approach to nonnormal and nonlinear state-space modeling. Journal of the American Statistical Association, 87 (418):493–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1982:CBM

[CR82] R. J. Carroll and David Ruppert. A comparison between maximum likelihood and generalized least squares in a heteroscedastic linear model. Journal of the American Statistical Association, 77 (380):878–882, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1984:PTW

[CR84] Raymond J. Carroll and David Ruppert. Power transformations when fitting theoretical models to data. Journal of the American Statistical Association, 79(386):321–328, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chi:1989:MLD

[CR89] Eric M. Chi and Gregory C. Reinsel. Models for longitudinal data with random effects and ar(1) errors. Journal of the American Statistical Association, 84(406):452–459, ???? 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Catalano:1992:BLV

[CR92] Paul J. Catalano and Louise M. Ryan. Bivariate latent variable models for clustered discrete and continuous outcomes. Journal of the American Statistical Association, 87(419):651–??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crawford:1994:ALM

[Cra94] Sybil L. Crawford. An application of the Laplace method to finite mixture distributions. Journal of the American Statistical Association, 89(425):259–267, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cressie:1982:PSM

[Cre82] Noel Cressie. Playing safe with misweighted means. Journal of the American Statistical Association, 77(380):754–759, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cressie:1988:GPD

[Cre88] Noel Cressie. A graphical procedure for determining nonstationarity in time series. Journal of the American Statistical Association, 83(404):1108–1116, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cressie:1997:COE

[Cre97] Noel Cressie. Comment: Ozone exposure and population density in Harris County, Texas. Journal of the American Statistical Association, 92(438):411–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Croux:1994:GE

[CRH94] Christophe Croux, Peter J. Rousseeuw, and Ola Hössjer. Generalized S-estimators. Journal of the American Statistical Association, 89(428):1271–1281, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crowley:1974:ANN

[Cro74] J. Crowley. Asymptotic normality of a new nonparametric statistic for use in organ transplant studies. Journal of the American Statistical Association, 69(348):1006–1011, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crockett:1985:ADH

[Cro85] Patrick W. Crockett. Asymptotic distribution of the Hildreth-Houck estimator. Journal of the American Statistical Association, 80(389):202–204, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crouchley:1995:REM

[Cro95] Robert Crouchley. A random-effects model for ordered categorical data. Journal of the American Statistical Association, 90(430): 489–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crowley:1997:PPM

[Cro97] Evelyn M. Crowley. Product partition models for normal means. Journal of the American Statistical Association, 92 (437):192–??, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/march/crowley.htm.

Clogg:1991:MII

[CRS^{+}91] Clifford C. Clogg, Donald B. Rubin, Nathaniel Schenker, Bradley Schultz, and Lynn Weidman. Multiple imputation of industry and occupation codes in census public-use samples using Bayesian logistic regression. Journal of the American Statistical Association, 86(413):68–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1999:DRD

[CRS99] Raymond J. Carroll, David Ruppert, and Leonard A. Stefanski. Discussion: Regression depth. Journal of the American Statistical Association, 94(446):410–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1998:LEE

[CRW98] Raymond J. Carroll, David Ruppert, and Alan H. Welsh. Local estimating equations. Journal of the American Statistical Association, 93(441):214–227, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/CARROLL.HTM.

Campbell:1985:NSM

[CS85] Gregory Campbell and John H. Skillings. Nonparametric stepwise multiple comparison procedures. Journal of the American Statistical Association, 80(392):998–1003, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chiu:1986:BPI

[CS86] H. Y. Chiu and J. Sedransk. A Bayesian procedure for imputing missing values in sample surveys. Journal of the American Statistical Association, 81(395):667–676, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cohen:1987:AIF

[CS87] Arthur Cohen and H. B. Sackrowitz. An approach to inference following model selection with applications to transformation-based and adaptive inference. Journal of the American Statistical Association, 82(400):1123–1130, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cohen:1989:ETR

[CS89] Arthur Cohen and H. B. Sackrowitz. Exact tests that recover interblock information in balanced incomplete blocks designs. Journal of the American Statistical Association, 84(406):556–559, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1990:AQL

[CS90a] Raymond J. Carroll and Leonard A. Stefanski. Approximate quasilikelihood estimation in models with surrogate predictors. Journal of the American Statistical Association, 85(411):652–663, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crouch:1990:EIF

[CS90b] Edmund A. C. Crouch and Donna Spiegelman. The evaluation of integrals of the form ∫_{−∞}^{+∞} f(t)exp(−t^{2})dt: Application to logisticnormal models. Journal of the American Statistical Association, 85(410):464–469, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Calvin:1991:BFP

[CS91] James A. Calvin and J. Sedransk. Bayesian and frequentist predictive inference for the patterns of care studies. Journal of the American Statistical Association, 86(413):36–??, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cohen:1992:ITC

[CS92a] Arthur Cohen and H. B. Sackrowitz. Improved tests for comparing treatments against a control and other one-sided problems. Journal of the American Statistical Association, 87(420):1137–1144, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cohen:1992:EST

[CS92b] Arthur Cohen and Harold B. Sackrowitz. An evaluation of some tests of trend in contingency tables. Journal of the American Statistical Association, 87(418):470–??, June 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chaudhuri:1993:STM

[CS93] Probal Chaudhuri and Debapriya Sengupta. Sign tests in multidimension: Inference based on the geometry of the data cloud. Journal of the American Statistical Association, 88(424):1363–1370, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1994:SEE

[CS94] J. R. Cook and L. A. Stefanski. Simulation-extrapolation estimation in parametric measurement error models. Journal of the American Statistical Association, 89(428):1314–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Campodonico:1995:IPP

[CS95] Sylvia Campodonico and Nozer D. Singpurwalla. Inference and predictions from Poisson point processes incorporating expert knowledge. Journal of the American Statistical Association, 90 (429):220–??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1996:NCR

[CS96a] Jingxian Chen and Nozer D. Singpurwalla. The notion of “composite reliability” and its hierarchical Bayes estimation. Journal of the American Statistical Association, 91(436):1474–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cohen:1996:LCB

[CS96b] Arthur Cohen and H. B. Sackrowitz. Lower confidence bounds using pilot samples with an application to auditing. Journal of the American Statistical Association, 91(433):338–342, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Czaja:1986:RBS

[CSC86] Ronald F. Czaja, Cecelia B. Snowden, and Robert J. Casady. Reporting bias and sampling errors in a survey of a rare population using multiplicity counting rules. Journal of the American Statistical Association, 81(394):411–419, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chinchilli:1989:IBR

[CSS89] Vernon M. Chinchilli, Barry H. Schwab, and Pranab K. Sen. Inference based on ranks for the multiple-design multivariate linear model. Journal of the American Statistical Association, 84(406): 517–524, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:1993:FCA

[CT93] Rong Chen and Ruey S. Tsay. Functional-coefficient autoregressive models. Journal of the American Statistical Association, 88(421): 298–308, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cordy:1997:DDF

[CT97] Clifford B. Cordy and David R. Thomas. Deconvolution of a distribution function. Journal of the American Statistical Association, 92(440):1459–1465, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/CORDY.HTM.

Chao:1998:SCA

[CT98] Anne Chao and P. K. Tsay. A sample coverage approach to multiple-system estimation with application to census undercount. Journal of the American Statistical Association, 93(441): 283–293, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/CHAO.HTM.

Chen:1996:PMC

[CTC96] Chen-Hsin Chen, Wei-Yann Tsai, and Wei-Hsiung Chao. The product-moment correlation coefficient and linear regression for truncated data. Journal of the American Statistical Association, 91(435):1181–1186, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Consonni:1992:CPE

[CV92] Guido Consonni and Piero Veronese. Conjugate priors for exponential families having quadratic variance functions. Journal of the American Statistical Association, 87(420):1123–1127, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Consonni:1995:BMC

[CV95] Guido Consonni and Piero Veronese. A Bayesian method for combining results from several binomial experiments. Journal of the American Statistical Association, 90(431):935–944, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1985:NPE

[CW85] R. Dennis Cook and Jeffrey A. Witmer. A note on parametereffects curvature. Journal of the American Statistical Association, 80(392):872–878, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1990:CCN

[CW90] R. Dennis Cook and Sanford Weisberg. Confidence curves in nonlinear regression. Journal of the American Statistical Association, 85(410):544–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chang:1991:ETS

[CW91] Wei-Chien Chang and Donald E. Weiss. An examination of the time series properties of beta in the market model. Journal of the American Statistical Association, 86(416):883–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clarke:1993:NPN

[CW93] Bertrand Clarke and Larry Wasserman. Noninformative priors and nuisance parameters. Journal of the American Statistical Association, 88(424):1427–1432, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:1994:ALS

[CW94a] K. F. Cheng and J. W. Wu. Adjusted least squares estimates for the scaled regression coefficients with censored data. Journal of the American Statistical Association, 89(428):1483–1491, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:1994:TGF

[CW94b] K. F. Cheng and J. W. Wu. Testing goodness of fit for a parametric family of link functions. Journal of the American Statistical Association, 89(426):657–664, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1994:ECC

[CW94c] R. Dennis Cook and Weng Kee Wong. On the equivalence of constrained and compound optimal designs. Journal of the American Statistical Association, 89(426):687–692, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:1997:GAA

[CW97] R. Dennis Cook and Sanford Weisberg. Graphics for assessing the adequacy of regression models. Journal of the American Statistical Association, 92(438):490–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1988:EEW

[CWR88] Raymond J. Carroll, C.-F. Jeff Wu, and David Ruppert. The effect of estimating weights in weighted least squares. Journal of the American Statistical Association, 83(404):1045–1054, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:1995:PAL

[CWW95] R. J. Carroll, Suojin Wang, and C. Y. Wang. Prospective analysis of logistic case-control studies. Journal of the American Statistical Association, 90(429):157–169, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:1997:PSP

[CWY97] S. C. Cheng, L. J. Wei, and Z. Ying. Predicting survival probabilities with semiparametric transformation models. Journal of the American Statistical Association, 92(437):227–235, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/cheng.htm.

Clevenson:1977:BLE

[CZ77] M. L. Clevenson and J. V. Zidek. Bayes linear estimators of the intensity function of the nonstationary Poisson process. Journal of the American Statistical Association, 72(357):112–120, ???? 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dunn:1996:EHV

[DA96] D. T. Dunn and A. E. Ades. Estimating the HIV vertical transmission rate and the pediatric AIDS incubation period from prospective data. Journal of the American Statistical Association, 91(435): 935–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dabrowska:1990:SRT

[Da̧b90] Dorota M. Da̧browska. Signed-rank tests for censored matched pairs. Journal of the American Statistical Association, 85(410): 478–485, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

deAmorim:1986:EDE

[dAJ86] Sebastiao de Amorim and Richard A. Johnson. Experimental designs for estimating the correlation between two destructively tested variables. Journal of the American Statistical Association, 81(395):807–812, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dallal:1982:SIM

[Dal82] Gerard E. Dallal. A simple interaction model for two-dimensional contingency tables. Journal of the American Statistical Association, 77(378):425–432, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dawkins:1994:AOH

[DAO94] Brian P. Dawkins, Peter M. Andreae, and Paul M. O’Connor. Analysis of Olympic heptathlon data. Journal of the American Statistical Association, 89(427):1100–??, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davis:1982:EMM

[Dav82] A. W. Davis. On the effects of moderate multivariate nonnormality on Roy’s largest root test. Journal of the American Statistical Association, 77(380):896–900, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dawid:1982:WCB

[Daw82] A. P. Dawid. The well-calibrated Bayesian. with comments by Joseph B. Kadane and D. V. Lindley and a rejoinder by the author. Journal of the American Statistical Association, 77(379):605–613 (or 605–610??), September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dawkins:1995:IGD

[Daw95] Brian P. Dawkins. Investigating the geometry of a p-dimensional data set. Journal of the American Statistical Association, 90(429): 350–??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dey:1983:TSE

[DB83] Dipak K. Dey and James O. Berger. On truncation of shrinkage estimators in simultaneous estimation of normal means. Journal of the American Statistical Association, 78(384):865–869, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dean:1997:ELA

[DB97] C. B. Dean and R. Balshaw. Efficiency lost by analyzing counts rather than event times in Poisson and overdispersed Poisson regression models. Journal of the American Statistical Association, 92(440):1387–1398, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/DEAN.HTM.

Doksum:1994:CCL

[DBB^{+}94] Kjell Doksum, Stephen Blyth, Eric Bradlow, Xiao-Li Meng, and Hongyu Zhao. Correlation curves as local measures of variance explained by regression. Journal of the American Statistical Association, 89(426):571–582, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dent:1977:MCA

[DC77] Warren T. Dent and David C. Cavander. More on computational accuracy in regression. Journal of the American Statistical Association, 72(359):598–600, ???? 1977. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davidian:1987:VFE

[DC87] M. Davidian and R. J. Carroll. Variance function estimation. Journal of the American Statistical Association, 82(400):1079–1091, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Doss:1994:CRS

[DC94] Hani Doss and Yuang Chin Chiang. Choosing the resampling scheme when bootstrapping: A case study in reliability. Journal of the American Statistical Association, 89(425):298–308, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

delCastillo:1999:BTE

[dCP99a] Joan del Castillo and Pedro Puig. The best test of exponentiality against singly truncated normal alternatives. Journal of the American Statistical Association, 94(446):529–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

delCastillo:1999:IEM

[dCP99b] Joan del Castillo and Pedro Puig. Invariant exponential models applied to reliability theory and survival analysis. Journal of the American Statistical Association, 94(446):522–??, June 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dhariyal:1981:OSF

[DD81] Ishwari D. Dhariyal and Edward J. Dudewicz. Optimal selection from a finite sequence with sampling cost. Journal of the American Statistical Association, 76(376):954–959, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See corrigendum [DD82].

Dhariyal:1982:COS

[DD82] Ishwari D. Dhariyal and Edward J. Dudewicz. Corrigendum: “Optimal selection from a finite sequence with sampling cost” [J. Amer. Statist. Assoc. 76 (1981), no. 376, 954–959, MR 83e:62111]. Journal of the American Statistical Association, 77(380):955, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [DD81].

Dabrowska:1988:ETT

[DD88] Dorota M. Da̧browska and Kjell A. Doksum. Estimation and testing in a two-sample generalized odds-rate model. Journal of the American Statistical Association, 83(403):744–749, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Donald:1990:SSA

[DD90] Alan Donald and Allan Donner. A simulation study of the analysis of sets of 2 x 2 contingency tables under cluster sampling: Estimation of a common odds ratio. Journal of the American Statistical Association, 85(410):537–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DuarteDeCarvalho:1994:SDT

[DDD94] Filipa Duarte De Carvalho, Nico P. Dellaert, and Margarida De Sanches Osorio. Statistical disclosure in two-dimensional tables: General tables. Journal of the American Statistical Association, 89 (428):1547–1557, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Das:1999:CPO

[DDV99] Marcel Das, Jeff Dominitz, and Arthur Van Soest. Comparing predictions and outcomes: Theory and application to income changes. Journal of the American Statistical Association, 94(445):75–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

deHaan:1981:EMF

[de 81] Laurens de Haan. Estimation of the minimum of a function using order statistics. Journal of the American Statistical Association, 76(374):467–469, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See corrigenda [de 83a].

deHaan:1983:CEM

[de 83a] Laurens de Haan. Corrigenda: “Estimation of the minimum of a function using order statistics” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 467–469, MR 83a:62103]. Journal of the American Statistical Association, 78(384):1008, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [de 81].

DeRiggi:1983:ULF

[De 83b] Dennis F. De Riggi. Unimodality of likelihood functions for the binomial distribution. Journal of the American Statistical Association, 78(381):181–183, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Edwardes:1997:URC

[dE97] Michael D. deB. Edwardes. Univariate random cut-points theory for the analysis of ordered categorical data. Journal of the American Statistical Association, 92(439):1114–1123, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/edwardes.htm.

Dean:1992:TOP

[Dea92] C. B. Dean. Testing for overdispersion in Poisson and binomial regression models. Journal of the American Statistical Association, 87(418):451–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Delampady:1989:LBB

[Del89] Mohan Delampady. Lower bounds on Bayes factors for interval null hypotheses. Journal of the American Statistical Association, 84(405):120–124, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diaconis:1979:RP

[DF79] Persi Diaconis and David Freedman. On rounding percentages. Journal of the American Statistical Association, 74(366):359–364, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284288. The

authors extend prior work on correctness of sums of rounded percentages [MYZ67], and criticize biased rounding practices in [Ben38].

Dardanoni:1998:UAL

[DF98] Valentino Dardanoni and Antonio Forcina. A unified approach to likelihood inference on stochastic orderings in a non-parametric context. Journal of the American Statistical Association, 93(443): 1112–1123, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/dardanoni.html.

Doss:1992:EAW

[DG92] Hani Doss and Richard D. Gill. An elementary approach to weak convergence for quantile processes, with applications to censored survival data. Journal of the American Statistical Association, 87 (419):869–877, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davies:1993:IMO

[DG93] Laurie Davies and Ursula Gather. The identification of multiple outliers. with discussion and a reply by the authors. Journal of the American Statistical Association, 88(423):782–801, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Datta:1995:SRN

[DG95] Gauri Sankar Datta and Malay Ghosh. Some remarks on noninformative priors. Journal of the American Statistical Association, 90 (432):1357–1363, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Daniels:1999:HGL

[DG99] Michael J. Daniels and Constantine Gatsonis. Hierarchical generalized linear models in the analysis of variations in health care utilization. Journal of the American Statistical Association, 94(445): 29–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DuMouchel:1983:BMC

[DH83] William H. DuMouchel and Jeffrey E. Harris. Bayes methods for combining the results of cancer studies in humans and other species. with comments by Daniel Krewski, Adrian F. M. Smith and Robert E. Kass, and with a reply by the authors. Journal of the American Statistical Association, 78(382):293–315, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diggle:1986:STO

[DH86] Peter J. Diggle and Peter Hall. The selection of terms in an orthogonal series density estimator. Journal of the American Statistical Association, 81(393):230–233, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dalal:1991:PEI

[DH91] Siddhartha R. Dalal and Bruce Hoadley. Problems in extrapolation illustrated with Space Shuttle O-ring data comment. Journal of the American Statistical Association, 86(416):921–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dufour:1993:IEB

[DH93] Jean-Marie Dufour and Marc Hallin. Improved Eaton bounds for linear combinations of bounded random variables, with statistical applications. Journal of the American Statistical Association, 88 (423):1026–1033, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ding:1999:PIF

[DH99] A. Adam Ding and J. T. Gene Hwang. Prediction intervals, factor analysis models and HELP. Journal of the American Statistical Association, 94(446):446–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/ding.pdf.

Dhar:1993:CCM

[Dha93] Sunil K. Dhar. Computation of certain minimum distance estimators in ar[k] model. Journal of the American Statistical Association, 88(421):278–283, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dickey:1984:TUR

[DHF84] D. A. Dickey, D. P. Hasza, and W. A. Fuller. Testing for unit roots in seasonal time series. Journal of the American Statistical Association, 79(386):355–367, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

deAngelis:1993:ABA

[dHY93] D. de Angelis, Peter Hall, and G. A. Young. Analytical and bootstrap approximations to estimator distributions in L^{1} regression. Journal of the American Statistical Association, 88(424): 1310–1316, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diaconis:1978:RMT

[Dia78] P. Diaconis. Review of “A Mathematical Theory of Evidence”. Journal of the American Statistical Association, 73(363):677–678, ???? 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dickey:1983:MHF

[Dic83] James M. Dickey. Multiple hypergeometric functions: probabilistic interpretations and statistical uses. Journal of the American Statistical Association, 78(383):628–637, ???? 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dietz:1982:BNT

[Die82] E. Jacquelin Dietz. Bivariate nonparametric tests for the onesample location problem. Journal of the American Statistical Association, 77(377):163–169, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

deJong:1989:SIS

[dJ89] Piet de Jong. Smoothing and interpolation with the state-space model. Journal of the American Statistical Association, 84(408): 1085–1088, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Donoho:1995:AUS

[DJ95] David L. Donoho and Iain M. Johnstone. Adapting to unknown smoothness via wavelet shrinkage. Journal of the American Statistical Association, 90(432):1200–1224, ???? 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dickey:1998:FVP

[DJ98] James M. Dickey and Thomas J. Jiang. Filtered-variate prior distributions for histogram smoothing. Journal of the American Statistical Association, 93(442):651–662, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/june/DICKEY.HTM.

deJongh:1988:MTB

[dJdWW88] P. J. de Jongh, T. de Wet, and A. H. Welsh. Mallows-type bounded-influence-regression trimmed means. Journal of the American Statistical Association, 83(403):805–810, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

deJong:1983:TRP

[dJGM83] Piet de Jong, Malcolm Greig, and Dilip Madan. Testing for random pairing. Journal of the American Statistical Association, 78 (382):332–336, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dickey:1987:BMC

[DJK87] James M. Dickey, Jyh-Ming Jiang, and Joseph B. Kadane. Bayesian methods for censored categorical data. Journal of the American Statistical Association, 82(399):773–781, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Drost:1989:PAM

[DKMO89] F. C. Drost, W. C. M. Kallenberg, D. S. Moore, and J. Oosterhoff. Power approximations to multinomial tests of fit. Journal of the American Statistical Association, 84(405):130–141, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dykstra:1995:ILR

[DKR95] Richard Dykstra, Subhash Kochar, and Tim Robertson. Inference for likelihood ratio ordering in the two-sample problem. Journal of the American Statistical Association, 90(431):1034–1040, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DiCiccio:1997:CBF

[DKRW97] Thomas J. DiCiccio, Robert E. Kass, Adrian Raftery, and Larry Wasserman. Computing Bayes factors by combining simulation and asymptotic approximations. Journal of the American Statistical Association, 92(439):903–915, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/diciccio.htm.

DeOliveira:1997:BPT

[DKS97] Victor De Oliveira, Benjamin Kedem, and David A. Short. Bayesian prediction of transformed Gaussian random fields. Journal of the American Statistical Association, 92(440):1422– 1433, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/DEOLIVEI.HTM.

Daponte:1997:BDP

[DKW97] Beth Osborne Daponte, Joseph B. Kadane, and Lara J. Wolfson. Bayesian demography: Projecting the Iraqi Kurdish population, 1977–1990. Journal of the American Statistical Association, 92(440):1256–1267, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/DAPONTE.HTM.

Dutta:1974:IIV

[DL74] M. Dutta and E. Lyttkens. Iterative instrumental variables method and estimation of a large simultaneous system. Journal of the American Statistical Association, 69(348):977–986, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Deely:1981:BEB

[DL81] J. J. Deely and D. V. Lindley. Bayes empirical Bayes. Journal of the American Statistical Association, 76(376):833–841, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dykstra:1988:DPM

[DL88] Richard L. Dykstra and Jon H. Lemke. Duality of I projections and maximum likelihood estimation for log-linear models under cone constraints. Journal of the American Statistical Association, 83(402):546–554, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dean:1992:MDT

[DL92] A. M. Dean and S. M. Lewis. Multidimensional designs for twofactor experiments. Journal of the American Statistical Association, 87(420):1158–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diggle:1991:AVR

[DLB91] Peter J. Diggle, Nicholas Lange, and Francine M. Benes. Analysis of variance for replicated spatial point patterns in clinical neuroanatomy. Journal of the American Statistical Association, 86 (415):618–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DeGruttola:1991:MPH

[DLD91] Victor DeGruttola, Nicholas Lange, and Urania Dafni. Modeling the progression of HIV infection. Journal of the American Statistical Association, 86(415):569–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Donnelly:1995:PCS

[DLW95] Christl A. Donnelly, Nan M. Laird, and James H. Ware. Prediction and creation of smooth curves for temporally correlated longitudinal data. Journal of the American Statistical Association, 90(431): 984–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dalal:1988:WSO

[DM88a] S. R. Dalal and C. L. Mallows. When should one stop testing software. Journal of the American Statistical Association, 83(403): 872–879, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dayton:1988:CVL

[DM88b] C. Mitchell Dayton and George B. Macready. Concomitantvariable latent-class models. Journal of the American Statistical Association, 83(401):173–178, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diggle:1988:ESP

[DM88c] Peter Diggle and J. S. Marron. Equivalence of smoothing parameter selectors in density and intensity estimation. Journal of the American Statistical Association, 83(403):793–800, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diaconis:1989:MSC

[DM89] Persi Diaconis and Frederick Mosteller. Methods for studying coincidences. Journal of the American Statistical Association, 84 (408):853–861, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davidson:1990:STB

[DM90] Russell Davidson and James G. MacKinnon. Specification tests based on artificial regressions. Journal of the American Statistical Association, 85(409):220–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davis:1993:RBM

[DM93] James B. Davis and Joseph W. McKean. Rank-based methods for multivariate linear models. Journal of the American Statistical Association, 88(421):245–251, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Datta:1995:BIF

[DM95] Somnath Datta and William P. McCormick. Bootstrap inference for a first-order autoregression with positive innovations. Journal of the American Statistical Association, 90(432):1289–1300, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dixon:1996:RBA

[DM96] Sherry L. Dixon and Joseph W. McKean. Rank-based analysis of the heteroscedastic linear model. Journal of the American Statistical Association, 91(434):699–712, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dominitz:1997:UED

[DM97] Jeff Dominitz and Charles F. Manski. Using expectations data to study subjective income expectations. Journal of the American Statistical Association, 92(439):855–867, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/dominitz.htm.

Durazo-Arvizu:1997:ENB

[DMLC97] Ramon Durazo-Arvizu, Daniel McGee, Zhaohai Li, and Richard Cooper. Establishing the nadir of the body mass index-mortality relationship: A case study. Journal of the American Statistical Association, 92(440):1312–1319, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/DURAZO.HTM.

Deaton:1998:ACS

[DN98] Angus Deaton and Serena Ng. Applications and case studies — parametric and nonparametric approaches to price and tax reform. Journal of the American Statistical Association, 93(443):

900–909, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/deaton.html.

Dorfman:1994:NVE

[Dor94] Alan H. Dorfman. A note on variance estimation for the regression estimator in double sampling. Journal of the American Statistical Association, 89(425):137–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dielman:1989:EOL

[DP89] Terry E. Dielman and Roger C. Pfaffenberger. Efficiency of ordinary least squares for linear models with autocorrelation. Journal of the American Statistical Association, 84(405):248, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

deJong:1998:DST

[dP98] Piet de Jong and Jeremy Pinzer. Diagnosing shocks in time series. Journal of the American Statistical Association, 93 (442):796–806, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/DE_JONG.HTM.

Dominici:1999:MAM

[DPWH99] Francesca Dominici, Giovanni Parmigiani, Robert L. Wolpert, and Vic Hasselblad. Meta-analysis of migraine headache treatments: Combining information from heterogeneous designs. Journal of the American Statistical Association, 94(445):16–??, March 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dykstra:1983:TMT

[DR83] Richard L. Dykstra and Tim Robertson. On testing monotone tendencies. Journal of the American Statistical Association, 78 (382):342–350, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diebold:1991:FOC

[DR91] Francis X. Diebold and Glenn D. Rudebusch. Forecasting output with the composite leading index: a real-time analysis. Journal of the American Statistical Association, 86(415):603–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davison:1996:SMD

[DR96] A. C. Davison and N. I. Ramesh. Some models for discretized series of events. Journal of the American Statistical Association, 91(434):601–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dasgupta:1998:DFS

[DR98] Abhijit Dasgupta and Adrian E. Raftery. Detecting features in spatial point processes with clutter via model-based clustering. Journal of the American Statistical Association, 93(441): 294–302, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/DASGUPTA.HTM.

Devlin:1992:FID

[DRR92] B. Devlin, Neil Risch, and Kathryn Roeder. Forensic inference from DNA fingerprints. Journal of the American Statistical Association, 87(418):337–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dollinger:1991:IFI

[DS91] Michael B. Dollinger and Robert G. Staudte. Influence functions of iteratively reweighted least squares estimators. Journal of the American Statistical Association, 86(415):709–716, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Deville:1992:CES

[DS92] Jean-Claude Deville and Carl-Erik Särndal. Calibration estimators in survey sampling. Journal of the American Statistical Association, 87(418):376–382, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diggle:1997:CAN

[DS97a] Peter J. Diggle and David M. Smith. Comment: Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials. Journal of the American Statistical Association, 92(440):1251–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Durand:1997:ASP

[DS97b] Jean-François Durand and Robert Sabatier. Additive splines for partial least squares regression. Journal of the American Statistical Association, 92(440):1546–1554, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/DURAND.HTM.

Denison:1998:CBC

[DSM98] D. G. T. Denison, A. F. M. Smith, and B. K. Mallick. Comment: Bayesian CART model search. Journal of the American Statistical Association, 93(443):954–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DiCiccio:1987:BCI

[DT87] Thomas DiCiccio and Robert Tibshirani. Bootstrap confidence intervals and bootstrap approximations. Journal of the American Statistical Association, 82(397):163–170, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dunnett:1992:SMT

[DT92] Charles W. Dunnett and Ajit C. Tamhane. A step-up multiple test procedure. Journal of the American Statistical Association, 87(417):162–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Duan:1983:SEN

[Dua83] Naihua Duan. Smearing estimate: a nonparametric retransformation method. Journal of the American Statistical Association, 78 (383):605–610, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Duan:1990:APP

[Dua90] Naihua Duan. The adjoint projection pursuit regression. Journal of the American Statistical Association, 85(412):1029–1038, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dudewicz:1981:EBT

[DvdM81] Edward J. Dudewicz and Edward C. van der Meulen. Entropybased tests of uniformity. Journal of the American Statistical Association, 76(376):967–974, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Deng:1987:EVR

[DW87] Lih-Yuan Deng and C.-F. J. Wu. Estimation of variance of the regression estimator. Journal of the American Statistical Association, 82(398):568–576, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DeGruttola:1987:IAG

[DWL87] Victor De Gruttola, James H. Ware, and Thomas A. Louis. Influence analysis of generalized least squares estimators. Journal of the American Statistical Association, 82(399):911–917, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dabrowska:1994:CRM

[DwSH94] Dorota M. Dabrowska, Guo wen Sun, and Mary M. Horowitz. Cox regression in a Markov renewal model: An application to the analysis of bone marrow transplant data. Journal of the American Statistical Association, 89(427):867–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dwyer:1945:SRM

[Dwy45] P. Dwyer. The square root method and its use in correlation and regression. Journal of the American Statistical Association, 40(??): 493–503, ???? 1945. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dwyer:1967:SAM

[Dwy67] P. S. Dwyer. Some applications of matrix derivatives in multivariate analysis. Journal of the American Statistical Association, 62(??):607–625, ???? 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Degruttola:1992:PAN

[DXP92] Victor Degruttola, Ming Tu Xin, and Marcello Pagano. Pediatric AIDS in New York City: Estimating the distributions of infection, latency, and reporting delay and projecting future incidence. Journal of the American Statistical Association, 87(419):633–??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dyer:1982:CGD

[Dye82] Danny Dyer. The convolution of generalized F distributions. Journal of the American Statistical Association, 77(377):184–189, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dykstra:1982:MLE

[Dyk82] Richard L. Dykstra. Maximum likelihood estimation of the survival functions of stochastically ordered random variables. Journal of the American Statistical Association, 77(379):621–628, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dykstra:1983:ARL

[Dyk83] Richard L. Dykstra. An algorithm for restricted least squares regression. Journal of the American Statistical Association, 78 (384):837–842, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diaconis:1982:USP

[DZ82] Persi Diaconis and Sandy L. Zabell. Updating subjective probability. Journal of the American Statistical Association, 77(380): 822–830, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Easton:1991:CML

[Eas91] George S. Easton. Compromise maximum likelihood estimators for location. Journal of the American Statistical Association, 86 (416):1051–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Easton:1994:SDG

[Eas94] George S. Easton. A simple dynamic graphical diagnostic method for almost any model. Journal of the American Statistical Association, 89(425):201–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Edelman:1990:IOO

[Ede90] David Edelman. An inequality of optimal order for the tail probabilities of the T statistic under symmetry. Journal of the American Statistical Association, 85(409):120–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1991:CEV

[EF91] B. Efron and D. Feldman. Compliance as an explanatory variable in clinical trials. Journal of the American Statistical Association, 86(413):9–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1983:EER

[Efr83] Bradley Efron. Estimating the error rate of a prediction rule: improvement on cross-validation. Journal of the American Statistical Association, 78(382):316–331, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1984:CNL

[Efr84] Bradley Efron. Comparing nonnested linear models. Journal of the American Statistical Association, 79(388):791–803, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1986:DEF

[Efr86a] Bradley Efron. Double exponential families and their use in generalized linear regression. Journal of the American Statistical Association, 81(395):709–721, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1986:HBA

[Efr86b] Bradley Efron. How biased is the apparent error rate of a prediction rule. Journal of the American Statistical Association, 81 (394):461–470, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1987:BBC

[Efr87] Bradley Efron. Better bootstrap confidence intervals. with comments and a rejoinder by the author. Journal of the American Statistical Association, 82(397):171–200, ???? 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1988:LRS

[Efr88] Bradley Efron. Logistic regression, survival analysis, and the Kaplan–Meier curve. Journal of the American Statistical Association, 83(402):414–425, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1990:MEB

[Efr90] Bradley Efron. More efficient bootstrap computations. Journal of the American Statistical Association, 85(409):79–89, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1992:POE

[Efr92] Bradley Efron. Poisson overdispersion estimates based on the method of asymmetric maximum likelihood. Journal of the American Statistical Association, 87(417):98–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1994:MDI

[Efr94] Bradley Efron. Missing data, imputation, and the bootstrap. with a comment by Donald B. Rubin and a rejoinder by the author. Journal of the American Statistical Association, 89(426):463–479, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1996:REB

[Efr96] Bradley Efron. Rejoinder: Empirical Bayes methods for combining likelihoods. with discussion and a reply by the author. Journal of the American Statistical Association, 91(434):538–565, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efromovich:1997:DEC

[Efr97] Sam Efromovich. Density estimation for the case of supersmooth measurement error. Journal of the American Statistical Association, 92(438):526–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efromovich:1998:DDE

[Efr98] Sam Efromovich. Data-driven efficient estimation of the spectral density. Journal of the American Statistical Association, 93 (442):762–769, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/EFROMOVI.HTM.

Efromovich:1999:QLW

[Efr99] Sam Efromovich. Quasi-linear wavelet emission. Journal of the American Statistical Association, 94(445):189–??, March 1999.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ellenberg:1992:SIA

[EFS92] Susan S. Ellenberg, Dianne M. Finkelstein, and David A. Schoenfeld. Statistical issues arising in AIDS clinical trials. Journal of the American Statistical Association, 87(418):562–??, June 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

El-Gamal:1995:PBU

[EGG95] Mahmoud A. El-Gamal and David M. Grether. Are people Bayesian? uncovering behavioral strategies. Journal of the American Statistical Association, 90(432):1137–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Evans:1997:BAS

[EGGS97] Michael Evans, Zvi Gilula, Irwin Guttman, and Tim Swartz. Bayesian analysis of stochastically ordered distributions of categorical variables. Journal of the American Statistical Association, 92(437):208–214, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/march/evans.htm.

El-Gamal:1993:BSE

[EGMP93] Mahmoud A. El-Gamal, Richard D. McKelvey, and Thomas R. Palfrey. A Bayesian sequential experimental study of learning in games. Journal of the American Statistical Association, 88(422): 428–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Edwards:1983:MCB

[EH83] Donald G. Edwards and Jason C. Hsu. Multiple comparisons with the best treatment. Journal of the American Statistical Association, 78(384):965–971, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [EH84].

Edwards:1984:CMC

[EH84] Donald G. Edwards and Jason C. Hsu. Corrigenda: “Multiple comparisons with the best treatment”. Journal of the American Statistical Association, 79(388):965, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [EH83].

Edwards:1987:FMS

[EH87] David Edwards and Tomáš Havránek. A fast model selection procedure for large families of models. Journal of the American Statistical Association, 82(397):205–213, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Edwards:1991:TTS

[EH91] Cheryl L. Edwards and E. Philip Howrey. A “true” time series and its indicators: An alternative approach. Journal of the American Statistical Association, 86(416):878–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Etzioni:1993:OED

[EK93] Ruth Etzioni and Joseph B. Kadane. Optimal experimental design for another’s analysis. Journal of the American Statistical Association, 88(424):1404–1411, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ericksen:1989:ACH

[EKT89] E. Ericksen, J. B. Kadane, and John W. Tukey. Adjusting the 1980 Census of Housing and Population. Journal of the American Statistical Association, 84(??):927–944, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ericksen:1991:TEP

[EKWS91] Eugene P. Ericksen, Joseph B. Kadane, Kenneth W. Wachter, and Terence P. Speed. Total error in PES estimates of population comments. Journal of the American Statistical Association, 86 (416):855–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Eubank:1990:CPP

[EL90] R. L. Eubank and V. N. LaRiccia. Components of Pearson’s phisquared distance measure for the k-sample problem. Journal of the American Statistical Association, 85(410):441–??, June 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Eggermont:1997:NSE

[EL97] P. P. B. Eggermont and V. N. Lariccia. Nonlinearly smoothed EM density estimation with automated smoothing parameter selection for nonparametric deconvolution problems. Journal of the American Statistical Association, 92(440):1451–1458, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/EGERMONT.HTM.

Eubank:1987:TSD

[ELR87] R. L. Eubank, V. N. LaRiccia, and R. B. Rosenstein. Test statistics derived as components of Pearson’s phi-squared distance measure. Journal of the American Statistical Association, 82(399):816–825, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Eltinge:1996:CVE

[Elt96] John L. Eltinge. Comment: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91 (434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Easton:1990:MGQ

[EM90] George S. Easton and Robert E. McCulloch. A multivariate generalization of quantile-quantile plots. Journal of the American Statistical Association, 85(410):376–??, June 1990. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ellis:1992:LBR

[EM92] Steven P. Ellis and Stephan Morgenthaler. Leverage and breakdown in L_{1} regression. Journal of the American Statistical Association, 87(417):143–148, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ehm:1995:PRA

[EMM95] W. Ehm, E. Mammen, and D. W. Müller. Power robustification of approximately linear tests. Journal of the American Statistical Association, 90(431):1025–1033, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1994:SAG

[EP94] Bradley Efron and Vahe Petrosian. Survival analysis of the gamma-ray burst data. Journal of the American Statistical Association, 89(426):452–??, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Easton:1986:GSA

[ER86] George S. Easton and Elvezio Ronchetti. General saddlepoint approximations with applications to L statistics. Journal of the American Statistical Association, 81(394):420–430, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Erickson:1981:URR

[Eri81] Gary M. Erickson. Using ridge regression to estimate directly lagged effects in marketing. Journal of the American Statistical Association, 76(376):766–773, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Erkanli:1994:LAP

[Erk94] Alaattin Erkanli. Laplace approximations for posterior expectations when the mode occurs at the boundary of the parameter space. Journal of the American Statistical Association, 89(425): 250–258, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Edelsbrunner:1990:CLM

[ES90a] Herbert Edelsbrunner and Diane L. Souvaine. Computing least median of squares regression lines and guided topological sweep. Journal of the American Statistical Association, 85(409):115–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Eubank:1990:TGF

[ES90b] R. L. Eubank and C. H. Spiegelman. Testing the goodness of fit of a linear model via nonparametric regression techniques. Journal of the American Statistical Association, 85(410):387–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Eubank:1993:CBN

[ES93] R. L. Eubank and P. L. Speckman. Confidence bands in nonparametric regression. Journal of the American Statistical Association, 88(424):1287–1301, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Emond:1997:EEG

[ES97] Mary Jane Emond and Steven G. Self. An efficient estimator for the generalized semilinear model. Journal of the American Statistical Association, 92(439):1033–1040, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/emond.htm.

Ernst:1999:CPT

[ES99] Michael D. Ernst and William R. Schucany. A class of permutation tests of bivariate interchangeability. Journal of the American Statistical Association, 94(445):273–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Escobar:1986:ERS

[Esc86] Luis A. Escobar. The equivalence of regression-simple and bestlinear-unbiased estimators with type ii censored data from a location scale distribution. Journal of the American Statistical Association, 81(393):210–214, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Escobar:1994:ENM

[Esc94] Michael D. Escobar. Estimating normal means with a Dirichlet process prior. Journal of the American Statistical Association, 89 (425):268–277, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Esty:1992:VCW

[Est92] Warren W. Esty. Votes or competitions which determine a winner by estimating expected plurality. Journal of the American Statistical Association, 87(418):373–375, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:1997:ICV

[ET97] Bradley Efron and Robert Tibshirani. Improvements on crossvalidation: The .632+ bootstrap method. Journal of the American Statistical Association, 92(438):548–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ethier:1982:TFN

[Eth82] S. N. Ethier. Testing for favorable numbers on a roulette wheel. Journal of the American Statistical Association, 77(379):660–665, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Eubank:1986:NSP

[Eub86] R. L. Eubank. A note on smoothness priors and nonlinear regression. Journal of the American Statistical Association, 81(394): 514–517, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Eubank:1997:TGF

[Eub97] R. L. Eubank. Testing goodness of fit with multinomial data. Journal of the American Statistical Association, 92(439):1084– 1093, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/sept/eubank.htm.

Engel:1992:CPP

[EV92] Eduardo Engel and Achilles Venetoulias. The Chilean plebiscite: Projections without historic data. Journal of the American Statistical Association, 87(420):933–??, December 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Escobar:1995:BDE

[EW95] Michael D. Escobar and Mike West. Bayesian density estimation and inference using mixtures. Journal of the American Statistical Association, 90(430):577–588, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fabian:1991:PIA

[Fab91] Václav Fabian. On the problem of interactions in the analysis of variance. with discussion and a rejoinder by the author. Journal of the American Statistical Association, 86(414):362–375, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fabian:1994:CSA

[Fab94] Václav Fabian. Comment on: “Some asymptotic results for learning in single hidden-layer feedforward network models” [J. Amer. Statist. Assoc. 84 (1989), no. 408, 1003–1013, MR 92e:62119] by H. White. Journal of the American Statistical Association, 89(428): 1571, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Whi89, Whi92].

Fahrmeir:1992:PME

[Fah92] Ludwig Fahrmeir. Posterior mode estimation by extended Kalman filtering for multivariate dynamic generalized linear models. Journal of the American Statistical Association, 87(418):501–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:1992:DAN

[Fan92] Jianqing Fan. Design-adaptive nonparametric regression. Journal of the American Statistical Association, 87(420):998–1004, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:1996:TSB

[Fan96] Jianqing Fan. Test of significance based on wavelet thresholding and neyman’s truncation. Journal of the American Statistical Association, 91(434):674–688, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fatti:1983:REM

[Fat83] L. P. Fatti. The random-effects model in discriminant analysis. Journal of the American Statistical Association, 78(383):679–687, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fay:1985:JCS

[Fay85] Robert E. Fay. A jackknifed chi-squared test for complex samples. Journal of the American Statistical Association, 80(389):148–157, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fay:1996:APA

[Fay96a] Robert E. Fay. Alternative paradigms for the analysis of imputed survey data. Journal of the American Statistical Association, 91 (434):490–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fay:1996:RVE

[Fay96b] Robert E. Fay. Rejoinder: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91 (434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fries:1983:ATF

[FB83] Arthur Fries and Gouri K. Bhattacharyya. Analysis of two-factor experiments under an inverse Gaussian model. Journal of the American Statistical Association, 78(384):820–826, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Flury:1997:DSM

[FBF97] Leah Flury, Benzion Boukai, and Bernard D. Flury. The discrimination subspace model. Journal of the American Statistical Association, 92(438):757–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feltz:1985:MLE

[FD85] Carol J. Feltz and Richard L. Dykstra. Maximum likelihood estimation of the survival functions of N stochastically ordered random variables. Journal of the American Statistical Association, 80 (392):1012–1019, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ferenstein:1988:OSS

[FE88] E. Z. Ferenstein and E. G. Enns. Optimal sequential selection from a known distribution with holding costs. Journal of the American Statistical Association, 83(402):382–386, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fellagi:1972:QSC

[Fel72] I. Fellagi. On the question of statistical confidentiality. Journal of the American Statistical Association, 61(337):??, March 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feldstein:1974:EVC

[Fel74] M. Feldstein. Errors in variables: A consistent estimator with smaller MSE in finite samples. Journal of the American Statistical Association, 69(348):990–996, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ferguson:1982:IML

[Fer82] Thomas S. Ferguson. An inconsistent maximum likelihood estimate. Journal of the American Statistical Association, 77(380): 831–834, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ferre:1998:DDS

[Fer98] Louis Ferré. Determining the dimension in sliced inverse regression and related methods. Journal of the American Statistical Association, 93(441):132–140, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/FERRE.HTM.

Faulkenberry:1991:EPT

[FG91] G. David Faulkenberry and Abderrazak Garoui. Estimating a population total using an area frame. Journal of the American Statistical Association, 86(414):445–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:1994:CRL

[FG94] Jianqing Fan and Irène Gijbels. Censored regression: Local linear approximations and their applications. Journal of the American Statistical Association, 89(426):560–570, June 1994. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:1995:CNT

[FG95] Yanqin Fan and Ramazan Gencay. A consistent nonparametric test of symmetry in linear regression models. Journal of the American Statistical Association, 90(430):551–557, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fine:1999:PHM

[FG99] Jason P. Fine and Robert J. Gray. A proportional hazards model for the subdistribution of a competing risk. Journal of the American Statistical Association, 94(446):496–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fuller:1978:REA

[FH78] Wayne A. Fuller and Michael A. Hidiroglou. Regression estimation after correcting for attenuation. Journal of the American Statistical Association, 73(361):99–104, ???? 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [FH81c].

Fuller:1981:CPP

[FH81a] Wayne A. Fuller and David P. Hasza. Correction: “Properties of predictors for autoregressive time series” [J. Amer. Statist. Assoc. 76 (1981), no. 373, 155–161, MR 82c:62132]. Journal of the American Statistical Association, 76(376):1023, ???? 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [FH81b].

Fuller:1981:PPA

[FH81b] Wayne A. Fuller and David P. Hasza. Properties of predictors for autoregressive time series. Journal of the American Statistical Association, 76(373):155–161, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [FH81a].

Fuller:1981:CRE

[FH81c] Wayne A. Fuller and Michael A. Hidiroglou. Correction: “Regression estimation after correcting for attenuation” [J. Amer. Statist. Assoc. 73 (1978), no. 361, 99–104, MR 80k:62087]. Journal of the American Statistical Association, 76(376):1022, ???? 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [FH78, JTB81].

Frank:1982:CIU

[FH82] Ove Frank and Frank Harary. Cluster inference by using transitivity indices in empirical graphs. Journal of the American Statistical Association, 77(380):835–840, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fatti:1986:VSH

[FH86] L. Paul Fatti and Douglas M. Hawkins. Variable selection in heteroscedastic discriminant analysis. Journal of the American Statistical Association, 81(394):494–500, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fisher:1989:BCR

[FH89] Nicholas I. Fisher and Peter Hall. Bootstrap confidence regions for directional data. Journal of the American Statistical Association, 84(408):996–1002, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [FH90].

Fisher:1990:CBC

[FH90] Nicholas I. Fisher and Peter Hall. Correction: “Bootstrap confidence regions for directional data”. Journal of the American Statistical Association, 85(410):608, ???? 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [FH89].

Fisher:1996:IPM

[FHJW96] Nicholas I. Fisher, Peter Hall, Bing-Yi Jing, and Andrew T. A. Wood. Improved pivotal methods for constructing confidence regions with directional data. Journal of the American Statistical Association, 91(435):1062–1070, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:1996:LSN

[FHMP96] Jianqing Fan, Peter Hall, Michael A. Martin, and Prakash Patil. On local smoothing of nonparametric curve estimators. Journal of the American Statistical Association, 91(433):258–266, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fleming:1987:SVL

[FHO87] Thomas R. Fleming, David P. Harrington, and Margaret O’Sullivan. Supremum versions of the log-rank and generalized Wilcoxon statistics. Journal of the American Statistical Association, 82(397):312–320, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fisher:1997:ECS

[FHTW97] Nicholas I. Fisher, Peter Hall, Berwin A. Turlach, and G. S. Watson. On the estimation of a convex set from noisy data on its support function. Journal of the American Statistical Association, 92(437):84–91, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/march/fisher.htm.

Fan:1995:LPK

[FHW95] Jianqing Fan, Nancy E. Heckman, and M. P. Wand. Local polynomial kernel regression for generalized linear models and quasilikelihood functions. Journal of the American Statistical Association, 90(429):141–150, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fairley:1990:IWQ

[FIB90] William B. Fairley, Alan J. Izenman, and Partha Bagchi. Inference for welfare quality control programs. Journal of the American Statistical Association, 85(411):874–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Finster:1983:FAS

[Fin83] Mark Finster. A frequentistic approach to sequential estimation in the general linear model. Journal of the American Statistical Association, 78(382):403–407, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Finner:1990:SNI

[Fin90] Helmut Finner. Some new inequalities for the range distribution, with application to the determination of optimum significance levels of multiple range tests. Journal of the American Statistical Association, 85(409):191–194, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Finner:1993:MPS

[Fin93] H. Finner. On a monotonicity problem in step-down multiple test procedures. Journal of the American Statistical Association, 88 (423):920–923, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fishman:1979:SBD

[Fis79] G. S. Fishman. Sampling from the binomial distribution on a computer. Journal of the American Statistical Association, 74 (366):418–423, ???? 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Faraway:1990:BCB

[FJ90] Julian J. Faraway and Myoungshic Jhun. Bootstrap choice of bandwidth for density estimation. Journal of the American Statistical Association, 85(412):1119–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Field:1989:IAI

[FJV89] J. B. F. Field, A. T. James, and W. N. Venables. The influence of assumptions implicit in a model on parametric inference. Journal of the American Statistical Association, 84(405):101–106, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fu:1994:DTR

[FK94] J. C. Fu and M. V. Koutras. Distribution theory of runs: A Markov chain approach. Journal of the American Statistical Association, 89(427):1050–1058, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fienberg:1999:CCL

[FK99] Stephen E. Fienberg and Sung-Ho Kim. Combining conditional log-linear structures. Journal of the American Statistical Association, 94(445):229–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ferretti:1999:CLG

[FKYZ99] Nelida Ferretti, Diana Kelmansky, Victor J. Yohai, and Reuben H. Zamar. A class of locally and globally robust regression estimates. Journal of the American Statistical Association, 94(445):174–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fitzmaurice:1995:RMB

[FL95] Garrett M. Fitzmaurice and Nan M. Laird. Regression models for a bivariate discrete and continuous outcome with clustering. Journal of the American Statistical Association, 90(431):845–852, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:1998:TSW

[FL98] Jianqing Fan and Sheng-Kuei Lin. Test of significance when data are curves. Journal of the American Statistical Association, 93 (443):1007–1021, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/sept/fan.html.

Fligner:1984:NTS

[Fli84] Michael A. Fligner. A note on two-sided distribution-free treatment versus control multiple comparisons. Journal of the American Statistical Association, 79(385):208–211, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Flury:1984:CPC

[Flu84] Bernhard N. Flury. Common principal components in k groups. Journal of the American Statistical Association, 79(388):892–898, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Flury:1985:ALC

[Flu85] Bernhard N. Flury. Analysis of linear combinations with extreme ratios of variance. Journal of the American Statistical Association, 80(392):915–922, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fitzmaurice:1996:MLM

[FLZ96] Garrett M. Fitzmaurice, Nan M. Laird, and Gwendolyn E. P. Zahner. Multivariate logistic models for incomplete binary responses. Journal of the American Statistical Association, 91(433):99–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feuerverger:1981:SFM

[FM81] Andrey Feuerverger and Philip McDunnough. On some Fourier methods for inference. Journal of the American Statistical Association, 76(374):379–387, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fishman:1982:SEM

[FM82] George S. Fishman and Louis R. Moore. A statistical evaluation of multiplicative congruential random number generators with modulus 2^{31} − 1. Journal of the American Statistical Association, 77(377):129–136, March 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://links.jstor.org/sici?sici=0162-1459%
28198203%2977%3A377%3C129%3AASEOMC%3E2.0.CO%3B2-Q.

Faltin:1983:SSP

[FM83] Frederick W. Faltin and Charles E. McCulloch. On the smallsample properties of the Olkin-Sobel-Tong estimator of the probability of correct selection. Journal of the American Statistical Association, 78(382):464–467, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fox:1992:GCD

[FM92] John Fox and Georges Monette. Generalized collinearity diagnostics. Journal of the American Statistical Association, 87(417): 178–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fairbanks:1982:CIE

[FMD82] Kenneth Fairbanks, Richard Madsen, and Richard Dykstra. A confidence interval for an exponential parameter from a hybrid life test. Journal of the American Statistical Association, 77(377): 137–140, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:1962:DSP

[FMR62] C. T. Fan, M. E. Muller, and I. Rezucha. Development of sampling plans (item by item) selection techniques and digital computers. Journal of the American Statistical Association, ??(??):??, June 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Findley:1990:SSD

[FMSP90] David F. Findley, Brian C. Monsell, Holly B. Shulman, and Marian G. Pugh. Sliding-spans diagnostics for seasonal and related adjustments. Journal of the American Statistical Association, 85 (410):345–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Flury:1995:SDS

[FNP95] Bernard D. Flury, Daan G. Nel, and Inet Pienaar. Simultaneous detection of shift in means and variances. Journal of the American Statistical Association, 90(432):1474–1481, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Follmann:1996:SMT

[Fol96] Dean Follmann. A simple multivariate test for one-sided alternatives. Journal of the American Statistical Association, 91(434): 854–861, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fong:1992:RER

[Fon92] Duncan K. H. Fong. Ranking and estimation of related means in the presence of a covariate — A Bayesian approach. Journal of the American Statistical Association, 87(420):1128–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Formann:1992:LLL

[For92] Anton K. Formann. Linear logistic latent class analysis for polytomous data. Journal of the American Statistical Association, 87(418):476–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Formann:1994:MCL

[For94] Anton K. Formann. Measuring change in latent subgroups using dichotomous data: Unconditional, conditional, and semiparametric maximum likelihood estimation. Journal of the American Statistical Association, 89(427):1027–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fernandez:1995:MIS

[FOS95] Carmen Fernández, Jacek Osiewalski, and Mark F. J. Steel. Modeling and inference with v-spherical distributions. Journal of the American Statistical Association, 90(432):1331–1340, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fernandez:1997:CBI

[FOS97] Carmen Fernandez, Jacek Osiewalski, and Mark F. J. Steel. Classical and Bayesian inference robustness in multivariate regression models. Journal of the American Statistical Association, 92(440): 1434–1444, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/FERNANDE.HTM.

Fox:1974:FHB

[Fox74] B. L. Fox. Finite horizon behavior of policies for two-arm bandits. Journal of the American Statistical Association, 69(348):963–965, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Freedman:1984:BRE

[FP84] David A. Freedman and Stephen C. Peters. Bootstrapping a regression equation: some empirical results. Journal of the American Statistical Association, 79(385):97–106, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Foresi:1995:CDE

[FP95] Silverio Foresi and Franco Peracchi. The conditional distribution of excess returns: An empirical analysis. Journal of the American Statistical Association, 90(430):451–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frees:1990:EFS

[FR90] Edward W. Frees and David Ruppert. Estimation following a sequentially designed experiment. Journal of the American Statistical Association, 85(412):1123–1129, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Forbes:1999:BMF

[FR99] Florence Forbes and Adrian E. Raftery. Bayesian morphology: Fast unsupervised Bayesian image analysis. Journal of the American Statistical Association, 94(446):555–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Franck:1984:LRT

[Fra84] Wallace E. Franck. A likelihood ratio test for stochastic ordering. Journal of the American Statistical Association, 79(387):686–691, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fraser:1991:SIL

[Fra91] D. A. S. Fraser. Statistical inference: Likelihood to significance. Journal of the American Statistical Association, 86(414):258–265, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frees:1994:EDF

[Fre94] Edward W. Frees. Estimating densities of functions of observations. Journal of the American Statistical Association, 89(426): 517–525, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Friedman:1987:EPP

[Fri87] Jerome H. Friedman. Exploratory projection pursuit. Journal of the American Statistical Association, 82(397):249–266, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Friedman:1989:RDA

[Fri89] Jerome H. Friedman. Regularized discriminant analysis. Journal of the American Statistical Association, 84(405):165–175, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Friendly:1994:MDM

[Fri94] Michael Friendly. Mosaic displays for multi-way contingency tables. Journal of the American Statistical Association, 89(425): 190–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frydman:1984:MLE

[Fry84] Halina Frydman. Maximum likelihood estimation in the moverstayer model. Journal of the American Statistical Association, 79 (387):632–638, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Friedman:1981:PPR

[FS81] Jerome H. Friedman and Werner Stuetzle. Projection pursuit regression. Journal of the American Statistical Association, 76(376): 817–823, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frank:1986:MG

[FS86] Ove Frank and David Strauss. Markov graphs. Journal of the American Statistical Association, 81(395):832–842, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fitzpatrick:1987:QSC

[FS87] Simon Fitzpatrick and Alastair Scott. Quick simultaneous confidence intervals for multinomial proportions. Journal of the American Statistical Association, 82(399):875–878, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frigessi:1994:IPB

[FS94] Arnoldo Frigessi and Julian Stander. Informative priors for the Bayesian classification of satellite images. Journal of the American Statistical Association, 89(426):703–709, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Faraway:1995:SCB

[FS95a] Julian J. Faraway and Jiayang Sun. Simultaneous confidence bands for linear regression with heteroscedastic errors. Journal of the American Statistical Association, 90(431):1094–1098, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Forbes:1995:EOR

[FS95b] Andrew B. Forbes and Thomas J. Santner. Estimators of odds ratio regression parameters in matched case- control studies with covariate measurement error. Journal of the American Statistical Association, 90(431):1075–1084, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fay:1998:PTU

[FS98a] Michael P. Fay and Joanna H. Shih. Permutation tests using estimated distribution functions. Journal of the American Statistical Association, 93(441):387–396, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/abstracts_98/FAY.HTM.

Fernandez:1998:BMF

[FS98b] Carmen Fernández and Mark F. J. Steel. On Bayesian modeling of fat tails and skewness. Journal of the American Statistical Association, 93(441):359–371, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/FERNANDC.HTM.

Faucett:1998:ACS

[FSE98] Cheryl L. Faucett, Nathaniel Schenker, and Robert M. Elashoff. Analysis of censored survival data with intermittently observed time-dependent binary covariates. Journal of the American Statistical Association, 93(442):427–437, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/june/FAUCETT.HTM.

Friedman:1984:PPD

[FSS84] Jerome H. Friedman, Werner Stuetzle, and Anne Schroeder. Projection pursuit density estimation. Journal of the American Statistical Association, 79(387):599–608, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fahrmeir:1994:DSM

[FT94] Ludwig Fahrmeir and Gerhard Tutz. Dynamic stochastic models for time-dependent ordered paired comparison systems. Journal of the American Statistical Association, 89(428):1438–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fung:1993:UOL

[Fun93] Wing-Kam Fung. Unmasking outliers and leverage points: A confirmation. Journal of the American Statistical Association, 88 (422):515–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fung:1995:DLD

[Fun95] Wing K. Fung. Diagnostics in linear discriminant analysis. Journal of the American Statistical Association, 90(431):952–956, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Furlan:1946:HSU

[Fur46] Luigi Vladimir Furlan. Das Harmoniegesetz der Statistik: Eine Untersuchung über die metrische Interdependenz der sozialen Erscheinungen. (German) [The Law of Harmony in statistics: an investigation of the metrical interdependence of social phenomena], volume 2 of Beihefte zum Assekuranzjahrbuch. Verlag für Recht und Gesellschaft A.-G., Basel, Switzerland, 1946. xii + 504 pp. LCCN HA29 .F85.

Fligner:1988:MRM

[FV88] Michael A. Fligner and Joseph S. Verducci. Multistage ranking models. Journal of the American Statistical Association, 83(403): 892–901, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fenech:1988:PFC

[FW88] Alan P. Fenech and Peter H. Westfall. The power function of conditional log-linear model tests. Journal of the American Statistical Association, 83(401):198–203, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fahrmeir:1996:SHF

[FW96] Ludwig Fahrmeir and Stefan Wagenpfeil. Smoothing hazard functions and time-varying effects in discrete duration and competing risks models. Journal of the American Statistical Association, 91 (436):1584–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fygenson:1997:NAM

[Fyg97] Mendel Fygenson. A new approach in modeling a categorical response. I. binary case. Journal of the American Statistical Association, 92(437):322–332, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/fygenson.htm.

Fan:1998:CSS

[FZ98] Jianqing Fan and Jin-Ting Zhang. Comment: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443):980–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Griffiths:1982:UTS

[GA82] W. E. Griffiths and J. R. Anderson. Using time-series and crosssection data to estimate a production function with positive and negative marginal risks. Journal of the American Statistical Association, 77(379):529–536, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gage:1943:CTR

[Gag43] Robert Gage. Contents of Tippett’s “Random Sampling Numbers”. Journal of the American Statistical Association, 38(??): 223–227, ???? 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gail:1996:SA

[Gai96] Mitchell H. Gail. Statistics in action. Journal of the American Statistical Association, 91(433):1–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Galbraith:1994:SAR

[Gal94] Rex F. Galbraith. Some applications of radial plots. Journal of the American Statistical Association, 89(428):1232–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Garthwaite:1994:IPL

[Gar94] Paul H. Garthwaite. An interpretation of partial least squares. Journal of the American Statistical Association, 89(425):122–127, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gastwirth:1985:UME

[Gas85] Joseph L. Gastwirth. The use of maximin efficiency robust tests in combining contingency tables and survival analysis. Journal of the American Statistical Association, 80(390):381–384, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gaver:1978:PNA

[Gav78] Donald P. Gaver. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments: Comment. Journal of the American Statistical Association, 73 (363):522, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/pdfplus/2286593.pdf. See [SM78a].

Gelman:1997:CAN

[GB97] Andrew Gelman and Frederic Y. Bois. Comment: Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials. Journal of the American Statistical Association, 92(440):1248–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gopalan:1998:BMC

[GB98] Ramanan Gopalan and Donald A. Berry. Bayesian multiple comparisons using Dirichlet process priors. Journal of the American Statistical Association, 93(443):1130–1139, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/gopalan.html.

Gelman:1996:PPA

[GBJ96] Andrew Gelman, Frederic Bois, and Jiming Jiang. Physiological pharmacokinetic analysis using population modeling and informative prior distributions. Journal of the American Statistical Association, 91(436):1400–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geronimus:1996:VUC

[GBN96] Arline T. Geronimus, John Bound, and Lisa J. Neidert. On the validity of using census geocode characteristics to proxy individual socioeconomic characteristics. Journal of the American Statistical Association, 91(434):529–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goldstein:1982:SSS

[GC82] Michael Goldstein and David Collett. On sample-size selection and the evaluation of discriminability in the model choice problem. Journal of the American Statistical Association, 77(380):922–928, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gentleman:1991:GMC

[GC91] R. Gentleman and J. Crowley. Graphical methods for censored data. Journal of the American Statistical Association, 86(415): 678–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Godtliebsen:1995:ENT

[GC95] Fred Godtliebsen and Chih Kang Chu. Estimation of the number of true gray levels, their values, and relative frequencies in a noisy image. Journal of the American Statistical Association, 90(431): 890–899, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goutis:1996:AEM

[GCW96] Constantinos Goutis, George Casella, and Martin T. Wells. Assessing evidence in multiple hypotheses. Journal of the American Statistical Association, 91(435):1268–1277, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gebhardt:1971:IBI

[Geb71] Friedrich Gebhardt. Incomplete beta-integral B(x; 2/3, 2/3) and [p(1 − p)]^{−1/6} for use with Borges’ approximation of the binomial distribution. Journal of the American Statistical Association, 66 (??):189–191, ???? 1971. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geiringer:1948:RVF

[Gei48] Hilda Geiringer. Review of L. V. Furlan’s book: Das Harmoniegesetz der Statistik: Eine Untersuchung über die metrische Interdependenz der sozialen Erscheinungen. (German) [The Law of Harmony in statistics: an investigation of the metrical interdependence of social phenomena]. Journal of the American Statistical Association, 43(242):325–328, June 1948. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2280379. See [Fur46].

Geisser:1992:SSI

[Gei92] Seymour Geisser. Some statistical issues in medicine and forensics. Journal of the American Statistical Association, 87(419):607–??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelman:1994:DPM

[Gel94] Andrew Gelman. Discussion: A probabilistic model for the spatial distribution of party support in multiparty electorates. Journal of the American Statistical Association, 89(428):1198–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelfand:1996:CEB

[Gel96] Alan E. Gelfand. Comment: Empirical Bayes methods for combining likelihoods. Journal of the American Statistical Association, 91(434):538–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

George:1986:CMS

[Geo86] Edward I. George. Combining minimax shrinkage estimators. Journal of the American Statistical Association, 81(394):437–445, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

George:1996:CEB

[Geo96] Edward I. George. Comment: Empirical Bayes methods for combining likelihoods. Journal of the American Statistical Association, 91(434):538–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gertsbakh:1982:CLH

[Ger82] I. Gertsbakh. Confidence limits for highly reliable coherent systems with exponentially distributed component life. Journal of the American Statistical Association, 77(379):673–678, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geweke:1982:MLD

[Gew82] John Geweke. Measurement of linear dependence and feedback between multiple time series. with discussion and with a reply by the author. Journal of the American Statistical Association, 77 (378):304–324, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geweke:1984:MCL

[Gew84] John F. Geweke. Measures of conditional linear dependence and feedback between time series. Journal of the American Statistical Association, 79(388):907–915, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geyer:1991:CML

[Gey91] Charles J. Geyer. Constrained maximum likelihood exemplified by isotonic convex logistic regression. Journal of the American Statistical Association, 86(415):717–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goria:1996:CCV

[GF96a] Mohammed N. Goria and Bernard D. Flury. Common canonical variates in k independent groups. Journal of the American Statistical Association, 91(436):1735–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goutis:1996:PLS

[GF96b] Constantinos Goutis and Tom Fearn. Partial least squares regression on smooth factors. Journal of the American Statistical Association, 91(434):627–632, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gaynor:1993:UCS

[GFT^{+}93] Jeffrey J. Gaynor, Eric J. Feuer, Claire C. Tan, Danny H. Wu, Claudia R. Little, David J. Straus, Bayard D. Clarkson, and Murray F. Brennan. On the use of cause-specific failure and conditional failure probabilities: Examples from clinical oncology data. Journal of the American Statistical Association, 88(422):400–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gastwirth:1987:ECR

[GG87] Joseph L. Gastwirth and Samuel W. Greenhouse. Estimating a common relative risk: application in equal employment. Journal of the American Statistical Association, 82(397):38–45, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gilula:1986:CAC

[GH86] Zvi Gilula and Shelby J. Haberman. Canonical analysis of contingency tables by maximum likelihood. Journal of the American Statistical Association, 81(395):780–788, ???? 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gideon:1987:RCC

[GH87a] Rudy A. Gideon and Robert A. Hollister. A rank correlation coefficient resistant to outliers. Journal of the American Statistical Association, 82(398):656–666, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Greenacre:1987:GIC

[GH87b] Michael Greenacre and Trevor Hastie. The geometric interpretation of correspondence analysis. Journal of the American Statistical Association, 82(398):437–447, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gilula:1988:AMC

[GH88] Zvi Gilula and Shelby J. Haberman. The analysis of multivariate contingency tables by restricted canonical and restricted association models. Journal of the American Statistical Association, 83 (403):760–771, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goodman:1990:ANT

[GH90] Leo A. Goodman and Shelby J. Haberman. The analysis of nonadditivity in two-way analysis of variance. Journal of the American Statistical Association, 85(409):139–145, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gilula:1994:CLL

[GH94] Zvi Gilula and Shelby J. Haberman. Conditional log-linear models for analyzing categorical panel data. Journal of the American Statistical Association, 89(426):645–656, June 1994. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gilula:1995:DCV

[GH95] Zvi Gilula and Shelby J. Haberman. Dispersion of categorical variables and penalty functions: Derivation, estimation, and comparability. Journal of the American Statistical Association, 90(432): 1447–1462, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gibbons:1994:RPM

[GHCF94] Robert D. Gibbons, Donald Hedeker, Sara C. Charles, and Paul Frisch. A random-effects probit model for predicting medical malpractice claims. Journal of the American Statistical Association, 89(427):760–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghysels:1996:PST

[GHL96] Eric Ghysels, Alastair Hall, and Hahn Shik Lee. On periodic structures and testing for seasonal unit roots. Journal of the American Statistical Association, 91(436):1551–??, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gentleman:1994:SQC

[GHMW94] R. C. Gentleman, M. S. Hamada, D. E. Matthews, and A. R. Wilson. Statistical quality control of HIV-1 ELISA test performance. Journal of the American Statistical Association, 89(428):1200–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:1992:CBE

[Gho92] Malay Ghosh. Constrained Bayes estimation with applications. Journal of the American Statistical Association, 87(418):533–540, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Griffiths:1987:SSP

[GHP87] William E. Griffiths, R. Carter Hill, and Peter J. Pope. Small sample properties of probit model estimators. Journal of the American Statistical Association, 82(399):929–937, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelfand:1990:IBI

[GHRPS90] Alan E. Gelfand, Susan E. Hills, Amy Racine-Poon, and Adrian F. M. Smith. Illustration of Bayesian inference in normal data models using Gibbs sampling. Journal of the American Statistical Association, 85(412):972–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gill:1984:UCR

[Gil84] Richard D. Gill. Understanding Cox’s regression model: a martingale approach. Journal of the American Statistical Association, 79(386):441–447, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gilula:1986:GAC

[Gil86] Zvi Gilula. Grouping and association in contingency tables: an exploratory canonical correlation approach. Journal of the American Statistical Association, 81(395):773–779, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Glaz:1984:PIM

[GJ84] Joseph Glaz and Bruce McK. Johnson. Probability inequalities for multivariate distributions with dependence structures. Journal of the American Statistical Association, 79(386):436–440, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gastwirth:1994:SCE

[GJ94] J. L. Gastwirth and Wesley O. Johnson. Screening with costeffective quality control: Potential applications to HIV and drug testing. Journal of the American Statistical Association, 89(427): 972–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gatto:1999:CSA

[GJ99] Riccardo Gatto and S. Rao Jammalamadaka. A conditional saddlepoint approximation for testing problems. Journal of the American Statistical Association, 94(446):533–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gersovitz:1998:ACS

[GJDT98] Mark Gersovitz, Hanan G. Jacoby, F. Seri Dedy, and A. Gozé Tapé. Applications and case studies — the balance of self-reported heterosexual activity in KAP surveys and the AIDS epidemic in Africa. Journal of the American Statistical Association, 93(443): 875–883, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/gersovitz.html.

Golan:1996:MEA

[GJP96] Amos Golan, George Judge, and Jeffrey M. Perloff. A maximum entropy approach to recovering information from multinomial response data. Journal of the American Statistical Association, 91 (434):841–853, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guihenneuc-Jouyaux:1998:DCM

[GJR98] Chantal Guihenneuc-Jouyaux and Christian P. Robert. Discretization of continuous Markov chains and Markov chain Monte Carlo convergence assessment. Journal of the American Statistical Association, 93(443):1055–1067, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/guihenneuc-joyaux.html.

Gelman:1990:EEC

[GK90] Andrew Gelman and Gary King. Estimating the electoral consequences of legislative redistricting. Journal of the American Statistical Association, 85(410):274–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Graubard:1993:HTC

[GK93] Barry I. Graubard and Edward L. Korn. Hypothesis testing with complex survey data: The use of classical quadratic test statistics with particular reference to regression problems. Journal of the American Statistical Association, 88(422):629–641, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gasser:1995:SSC

[GK95] Theo Gasser and Alois Kneip. Searching for structure in curve samples. Journal of the American Statistical Association, 90(432):

1179–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelman:1998:EPE

[GKB98] Andrew Gelman, Gary King, and W. John Boscardin. Estimating the probability of events that have never occurred: When is your vote decisive? Journal of the American Statistical Association, 93(441):1–9, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/GELMAN.HTM.

Gasser:1991:FFM

[GKK91] Theo Gasser, Alois Kneip, and Walter Köhler. A flexible and fast method for automatic smoothing. Journal of the American Statistical Association, 86(415):643–652, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelman:1998:ACSa

[GKL98a] Andrew Gelman, Gary King, and Chuanhai Liu. Applications and case studies — not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443):846–857, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/gelman.html.

Gelman:1998:ACSb

[GKL98b] Andrew Gelman, Gary King, and Chuanhai Liu. Applications and case studies — rejoinder: Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443):869–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gilula:1988:OAC

[GKR88] Zvi Gilula, Abba M. Krieger, and Yaakov Ritov. Ordinal association in contingency tables: some interpretive aspects. Journal of the American Statistical Association, 83(402):540–545, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gilstein:1983:SWR

[GL83] C. Zachary Gilstein and Edward E. Leamer. The set of weighted regression estimates. Journal of the American Statistical Association, 78(384):942–948, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:1987:REB

[GL87] Malay Ghosh and Parthasarathi Lahiri. Robust empirical Bayes estimation of means from stratified samples. Journal of the American Statistical Association, 82(400):1153–1162, ???? 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gill:1992:TRD

[GL92] Len Gill and Arthur Lewbel. Testing the rank and definiteness of estimated matrices with applications to factor, state-space, and ARMA models. Journal of the American Statistical Association, 87(419):766–776, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gross:1996:NER

[GL96] Shulamith T. Gross and Tze Leung Lai. Nonparametric estimation and regression analysis with left- truncated and right-censored data. Journal of the American Statistical Association, 91(435): 1166–1180, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Glaser:1980:BRF

[Gla80] Ronald E. Glaser. Bathtub and related failure rate characterizations. Journal of the American Statistical Association, 75(371): 667–672, ???? 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Glaz:1989:ABD

[Gla89] Joseph Glaz. Approximations and bounds for the distribution of the scan statistic. Journal of the American Statistical Association, 84(406):560–566, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goldman:1998:CCA

[GLB98] Dana P. Goldman, Arleen Leibowitz, and Joan L. Buchanan. Cost containment and adverse selection in Medicaid HMOs. Journal of the American Statistical Association, 93(441):54–62, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/GOLDMAN.HTM.

Gleser:1985:EPG

[Gle85] Leon Jay Gleser. Exact power of goodness-of-fit tests of Kolmogorov type for discontinuous distributions. Journal of the American Statistical Association, 80(392):954–958, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gleser:1992:IAM

[Gle92] Leon Jay Gleser. The importance of assessing measurement reliability in multivariate regression. Journal of the American Statistical Association, 87(419):696–707, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gleason:1993:UES

[Gle93] John R. Gleason. Understanding elongation: The scale contaminated normal family. Journal of the American Statistical Association, 88(421):327–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Giles:1992:OSP

[GLG92] David E. A. Giles, Offer Lieberman, and Judith A. Giles. The optimal size of a preliminary test of linear restrictions in a misspecified regression model. Journal of the American Statistical Association, 87(420):1153–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gragg:1979:NSM

[GLT79] W. B. Gragg, R. J. Leveque, and J. A. Trangenstein. Numerically stable methods for updating regressions. Journal of the American Statistical Association, 74(365):161–168, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:1986:EBE

[GM86a] Malay Ghosh and Glen Meeden. Empirical Bayes estimation in finite population sampling. Journal of the American Statistical Association, 81(396):1058–1062, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guttman:1986:CWA

[GM86b] Irwin Guttman and Ulrich Menzefricke. Counting by weighing: an approach using renewal theory. Journal of the American Statistical Association, 81(393):129–131, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gomez:1994:EPI

[GM94] Víctor Gómez and Agustín Maravall. Estimation, prediction, and interpolation for nonstationary series with the Kalman filter. Journal of the American Statistical Association, 89(426):611–624, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goetghebeur:1996:CIP

[GM96] Els Goetghebeur and Geert Molenberghs. Causal inference in a placebo-controlled clinical trial with binary outcome and ordered compliance. Journal of the American Statistical Association, 91 (435):928–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gonzalez-Manteiga:1996:BSS

[GMCM96] W. González-Manteiga, R. Cao, and J. S. Marron. Bootstrap selection of the smoothing parameter in nonparametric hazard rate estimation. Journal of the American Statistical Association, 91 (435):1130–1140, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelfand:1995:MEO

[GMD95] Alan E. Gelfand, Bani K. Mallick, and Dipak K. Dey. Modeling expert opinion arising as a partial probabilistic specification. Journal of the American Statistical Association, 90(430):598–604, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelfand:1997:BBS

[GMP97] Alan E. Gelfand, Bani K. Mallick, and Wolfgang Polasek. Broken biological size relationships: A truncated semiparametric regression approach with measurement error. Journal of the American Statistical Association, 92(439):836–845, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/gelfand.htm.

Gijbels:1999:EMC

[GMPS99] Irène Gijbels, Enno Mammen, Byeong U. Park, and Léopold Simar. On estimation of monotone and concave frontier functions. Journal of the American Statistical Association, 94(445): 220–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guttorp:1997:COE

[GMS97] Peter Guttorp, Wendy Meiring, and Paul D. Sampson. Comment: Ozone exposure and population density in Harris County, Texas. Journal of the American Statistical Association, 92(438):405–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Greenwood:1999:ISP

[GMWZ99] Michael J. Greenwood, John M. McDowell, Donald M. Waldman, and Steven S. Zahniser. The influence of social programs in source countries on various classes of U.S. immigration. Journal of the American Statistical Association, 94(445):64–??, March 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Golub:1982:NAV

[GN82] Gene H. Golub and Stephen G. Nash. Nonorthogonal analysis of variance using a generalized conjugate-gradient algorithm. Journal of the American Statistical Association, 77(377):109–116, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:1996:EMI

[GNK96] Malay Ghosh, Narinder Nangia, and Dal Ho Kim. Estimation of median income of four-person families: A Bayesian time series approach. Journal of the American Statistical Association, 91(436): 1423–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:1998:GLM

[GNSC98] Malay Ghosh, Kannan Natarajan, T. W. F. Stroud, and Bradley P. Carlin. Generalized linear models for small area estimation. Journal of the American Statistical Association, 93 (441):273–282, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/GHOSH.HTM.

Godambe:1982:APS

[God82a] V. P. Godambe. Ancillarity principle and a statistical paradox. Journal of the American Statistical Association, 77(380):931–933, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Godambe:1982:ESS

[God82b] V. P. Godambe. Estimation in survey sampling: robustness and optimality. with comments by W. A. Ericson and C.-M. Cassel and with a reply by the author. Journal of the American Statistical Association, 77(378):393–406, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goel:1983:IMB

[Goe83] Prem K. Goel. Information measures and Bayesian hierarchical models. Journal of the American Statistical Association, 78(382): 408–410, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Golub:1963:RGA

[Gol63] Gene H. Golub. Review: A Guide to ALGOL Programming by Daniel D. McCracken. Journal of the American Statistical Association, 58(304):1202, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283384.

Goldstein:1983:GVM

[Gol83a] Michael Goldstein. General variance modifications for linear Bayes estimators. Journal of the American Statistical Association, 78 (383):616–618, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goldstein:1983:PP

[Gol83b] Michael Goldstein. The prevision of a prevision. Journal of the American Statistical Association, 78(384):817–819, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goldstein:1986:EBS

[Gol86] Michael Goldstein. Exchangeable belief structures. Journal of the American Statistical Association, 81(396):971–976, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goodman:1991:MMG

[Goo91] Leo A. Goodman. Measures, models, and graphical displays in the analysis of cross-classified data. with comments and a rejoinder by the author. Journal of the American Statistical Association, 86 (416):1085–1138, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Good:1992:BNB

[Goo92] I. J. Good. The Bayes/non-Bayes compromise: A brief review. Journal of the American Statistical Association, 87(419):597–606, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goodman:1994:QIQ

[Goo94] Leo A. Goodman. On quasi-independence and quasi-dependence in contingency tables, with special reference to ordinal triangular contingency tables. Journal of the American Statistical Association, 89(427):1059–1063, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goodman:1996:SGM

[Goo96] Leo A. Goodman. A single general method for the analysis of crossclassified data: Reconciliation and synthesis of some methods of Pearson, Yule, and Fisher, and also some methods of correspondence analysis and association analysis. Journal of the American Statistical Association, 91(433):408–428, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goutis:1997:NEM

[Gou97] Constantinos Goutis. Nonparametric estimation of a mixing density via the kernel method. Journal of the American Statistical Association, 92(440):1445–1450, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/GOUTIS.HTM.

Gumpertz:1992:NRV

[GP92] Marcia L. Gumpertz and Sastry G. Pantula. Nonlinear regression with variance components. Journal of the American Statistical Association, 87(417):201–209, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gutierrez-Pena:1995:CPN

[GPS95] E. Gutiérrez-Peña and A. F. M. Smith. Conjugate parameterizations for natural exponential families. Journal of the American Statistical Association, 90(432):1347–1356, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gans:1981:DGK

[GR81] Lydia P. Gans and C. A. Robertson. Distributions of Goodman and Kruskal’s gamma and Spearman’s rho in 2×2 tables for small and moderate sample sizes. Journal of the American Statistical Association, 76(376):942–946, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guang:1992:EMP

[GR92a] Guo Guang and German Rodriguez. Estimating a multivariate proportional hazards model for clustered data using the EM algorithm, with an application to child survival in Guatemala. Journal of the American Statistical Association, 87(420):969–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gwet:1992:ORA

[GR92b] Jean-Philippe Gwet and Louis-Paul Rivest. Outlier resistant alternatives to the ratio estimator. Journal of the American Statistical Association, 87(420):1174–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Genest:1993:SIP

[GR93] Christian Genest and Louis-Paul Rivest. Statistical inference procedures for bivariate Archimedean copulas. Journal of the American Statistical Association, 88(423):1034–1043, ???? 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gatto:1996:GSA

[GR96a] Riccardo Gatto and Elvezio Ronchetti. General saddlepoint approximations of marginal densities and tail probabilities. Journal of the American Statistical Association, 91(434):666–673, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Givens:1996:LAI

[GR96b] Geof H. Givens and Adrian E. Raftery. Local adaptive importance sampling for multivariate densities with strong nonlinear relationships. Journal of the American Statistical Association, 91(433): 132–141, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geiser:1997:NTI

[GR97a] Peter W. Geiser and Ronald H. Randles. A nonparametric test of independence between two vectors. Journal of the American Statistical Association, 92(438):561–??, June 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gieser:1997:NTI

[GR97b] Peter W. Gieser and Ronald H. Randles. A nonparametric test of independence between two vectors. Journal of the American Statistical Association, 92(438):561–??, June 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gray:1992:FMA

[Gra92] Robert J. Gray. Flexible methods for analyzing survival data using splines, with applications to breast cancer prognosis. Journal of the American Statistical Association, 87(420):942–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gray:1995:TVG

[Gra95] Robert J. Gray. Tests for variation over groups in survival data. Journal of the American Statistical Association, 90(429):198–203, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Grambsch:1995:MLI

[GRB^{+}95] Patricia M. Grambsch, Bryan L. Randall, Roberd M. Bostick, John D. Potter, and Thomas A. Louis. Modeling the labeling index distribution: An application of functional data analysis. Journal of the American Statistical Association, 90(431):813–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Grillenzoni:1990:MTV

[Gri90] Carlo Grillenzoni. Modeling time-varying dynamical systems. Journal of the American Statistical Association, 85(410):499–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gripenberg:1992:CIP

[Gri92] G. Gripenberg. Confidence intervals for partial rank correlations. Journal of the American Statistical Association, 87(418):546–551, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Grillenzoni:1994:ORE

[Gri94] Carlo Grillenzoni. Optimal recursive estimation of dynamic models. Journal of the American Statistical Association, 89(427):777– ??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gross:1981:APE

[Gro81] Shulamith T. Gross. On asymptotic power and efficiency of tests of independence in contingency tables with ordered classifications. Journal of the American Statistical Association, 76(376):935–941, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Grogger:1990:DEC

[Gro90] Jeffrey Grogger. The deterrent effect of capital punishment: An analysis of daily homicide counts. Journal of the American Statistical Association, 85(410):295–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gilks:1998:AMC

[GRS98] Walter R. Gilks, Gareth O. Roberts, and Sujit K. Sahu. Adaptive Markov chain Monte Carlo through regeneration. Journal of the American Statistical Association, 93(443):1045–1054, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/sept/gilks.html.

Ghosh:1989:MUP

[GS89] Malay Ghosh and Pranab Kumar Sen. Median unbiasedness and Pitman closeness. Journal of the American Statistical Association, 84(408):1089–1091, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelfand:1990:SBA

[GS90a] Alan E. Gelfand and Adrian F. M. Smith. Sampling-based approaches to calculating marginal densities. Journal of the American Statistical Association, 85(410):398–409, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gordon:1990:MMB

[GS90b] K. Gordon and A. F. M. Smith. Modeling and monitoring biomedical time series. Journal of the American Statistical Association, 85(410):328–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Giles:1991:UEC

[GS91a] D. E. A. Giles and V. K. Srivastava. An unbiased estimator of the covariance matrix of the mixed regression estimator. Journal of the American Statistical Association, 86(414):441–444, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Green:1991:JST

[GS91b] Edwin J. Green and William E. Strawderman. A James–Stein type estimator for combining unbiased and possibly biased estimators. Journal of the American Statistical Association, 86(416): 1001–1006, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ganio:1992:DO

[GS92] Lisa M. Ganio and Daniel W. Schafer. Diagnostics for overdispersion. Journal of the American Statistical Association, 87(419):795– ??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gonul:1993:CPB

[GS93] Fusun Gonul and Kannan Srinivasan. Consumer purchase behavior in a frequently bought product category: Estimation issues and managerial insights from a hazard function model with heterogeneity. Journal of the American Statistical Association, 88 (424):1219–??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Giani:1994:TSE

[GS94] Guido Giani and Klaus Straßburger. Testing and selecting for equivalence with respect to a control. Journal of the American Statistical Association, 89(425):320–??, March 1994. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gurland:1995:HPF

[GS95] John Gurland and Jayaram Sethuraman. How pooling failure data may reverse increasing failure rates. Journal of the American Statistical Association, 90(432):1416–1423, ???? 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Giani:1997:OPP

[GS97] Guido Giani and Klaus Straßburger. Optimum partition procedures for separating good and bad treatments. Journal of the American Statistical Association, 92(437):291–298, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/giani.htm.

Ge:1998:CHD

[GS98a] Nanxiang Ge and Douglas G. Simpson. Correlation and high dimensional consistency in pattern recognition. Journal of the American Statistical Association, 93(443):995–1006, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/sept/ge.html.

Glickman:1998:SSM

[GS98b] Mark E. Glickman and Hal S. Stern. A state-space model for National Football League scores. Journal of the American Statistical Association, 93(441):25–35, March 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/GLICKMAN.HTM.

Gelfand:1999:IIP

[GS99] Alan E. Gelfand and Sujit K. Sahu. Identifiability, improper priors, and Gibbs sampling for generalized linear models. Journal of the American Statistical Association, 94(445):247–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gupta:1989:IEC

[GSS89] Arjun K. Gupta, A. K. Md. Ehsanes Saleh, and Pranab K. Sen. Improved estimation in a contingency table: independence structure. Journal of the American Statistical Association, 84(406): 525–532, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelfand:1992:BAC

[GSTM92] Alan E. Gelfand, Adrian F. M. Smith, and Lee Tai-Ming. Bayesian analysis of constrained parameter and truncated data problems using Gibbs sampling. Journal of the American Statistical Association, 87(418):523–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guttorp:1991:ESO

[GT91] Peter Guttorp and Mary Lou Thompson. Estimating secondorder parameters of volcanicity from historical data. Journal of the American Statistical Association, 86(415):578–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geyer:1995:AMC

[GT95] Charles J. Geyer and Elizabeth A. Thompson. Annealing Markov chain Monte Carlo with applications to ancestral inference. Journal of the American Statistical Association, 90(431):909–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gallant:1998:RPO

[GT98] A. Ronald Gallant and George Tauchen. Reprojecting partially observed systems with application to interest rate diffusions. Journal of the American Statistical Association, 93 (441):10–24, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/GALLANT.HTM.

Gu:1990:ASS

[Gu90] Chong Gu. Adaptive spline smoothing in Non-Gaussian regression models. Journal of the American Statistical Association, 85 (411):801–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guilbaud:1983:NPI

[Gui83] Olivier Guilbaud. Nonparametric prediction intervals for sample medians in the general case. Journal of the American Statistical Association, 78(384):937–941, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guilbaud:1988:EKT

[Gui88] Olivier Guilbaud. Exact Kolmogorov-type tests for left-truncated and/or right-censored data. Journal of the American Statistical Association, 83(401):213–221, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guilbaud:1993:EIA

[Gui93] Olivier Guilbaud. Exact inference about the within-subject variability in 2×2 crossover trials. Journal of the American Statistical Association, 88(423):939–946, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gurler:1996:BER

[Gür96] Ülkü Gürler. Bivariate estimation with right-truncated data. Journal of the American Statistical Association, 91(435):1152–1165, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gustafson:1996:LSI

[Gus96] Paul Gustafson. Local sensitivity of inferences to prior marginals. Journal of the American Statistical Association, 91(434):774–781, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guthery:1974:PR

[Gut74] S. B. Guthery. Partition regression. Journal of the American Statistical Association, 69(348):945–947, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gutmann:1982:SPI

[Gut82] Sam Gutmann. Stein’s paradox is impossible in the nonanticipative context. Journal of the American Statistical Association, 77 (380):934–935, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gupta:1986:LOT

[GV86] Ashis Sen Gupta and Lea Vermeire. Locally optimal tests for multiparameter hypotheses. Journal of the American Statistical Association, 81(395):819–825, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Godambe:1996:OER

[GV96] V. P. Godambe and K. Vijayan. Optimal estimation for responsedependent retrospective sampling. Journal of the American Statistical Association, 91(436):1724–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gates:1984:DSS

[GW84] David J. Gates and Mark Westcott. On the distributions of scan statistics. Journal of the American Statistical Association, 79 (386):423–429, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gray:1986:NAS

[GW86] H. L. Gray and Wayne A. Woodward. A new ARMA spectral estimator. Journal of the American Statistical Association, 81 (396):1100–1108, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gay:1988:MLQ

[GW88] David M. Gay and Roy E. Welsch. Maximum likelihood and quasi-likelihood for nonlinear exponential family regression models. Journal of the American Statistical Association, 83(404):990– 998, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gray:1991:GMA

[GW91] H. L. Gray and Suojin Wang. A general method for approximating tail probabilities. Journal of the American Statistical Association, 86(413):159–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gong:1990:CSD

[GWG90] Gail Gong, Alice S. Whittemore, and Stella Grosser. Censored survival data with misclassified covariates: a case study of breastcancer mortality. Journal of the American Statistical Association, 85(409):20–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gibbons:1998:HSU

[GWG98] Robert D. Gibbons and Virginia Wilcox-Gok. Health service utilization and insurance coverage: A multivariate probit analysis. Journal of the American Statistical Association, 93 (441):63–72, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/GIBBONS.HTM.

Gu:1993:SAC

[GY93] Minggao Gu and Zhiliang Ying. Sequential analysis for censored regression data. Journal of the American Statistical Association, 88(423):890–898, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Helland:1994:CPM

[HA94] Inge S. Helland and Trygve Almøy. Comparison of prediction methods when only a few components are relevant. Journal of the American Statistical Association, 89(426):583–591, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [HA95].

Helland:1995:CCP

[HA95] Inge S. Helland and Trygve Almøy. Correction: “Comparison of prediction methods when only a few components are relevant”. Journal of the American Statistical Association, 90(429):399, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [HA94].

Haas:1990:LMW

[Haa90] Timothy C. Haas. Lognormal and moving window methods of estimating acid deposition. Journal of the American Statistical Association, 85(412):950–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Haas:1995:LPS

[Haa95] Timothy C. Haas. Local prediction of a spatio-temporal process with an application to wet sulfate deposition. Journal of the American Statistical Association, 90(432):1189–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Haberman:1982:ADM

[Hab82] Shelby J. Haberman. Analysis of dispersion of multinomial responses. Journal of the American Statistical Association, 77(379): 568–580, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Haberman:1988:WUC

[Hab88] Shelby J. Haberman. A warning on the use of chi-squared statistics with frequency tables with small expected cell counts. Journal of the American Statistical Association, 83(402):555–560, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Haberman:1995:CML

[Hab95] Shelby J. Haberman. Computation of maximum likelihood estimates in association models. Journal of the American Statistical Association, 90(432):1438–1446, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Halpern:1973:PRB

[Hal73] E. F. Halpern. Polynomial regression from a Bayesian approach. Journal of the American Statistical Association, 68(??):137–143, ???? 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hall:1994:EBE

[Hal94] Peter Hall. On the erratic behavior of estimators of N in the binomial N, p distribution. Journal of the American Statistical Association, 89(425):344–352, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Handcock:1999:CPS

[Han99] Mark S. Handcock. Comment: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):100–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hartigan:1981:CSL

[Har81] J. A. Hartigan. Consistency of single linkage for high-density clusters. Journal of the American Statistical Association, 76(374): 388–394, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hart:1984:EKD

[Har84] Jeffrey D. Hart. Efficiency of a kernel density estimator under an autoregressive dependence model. Journal of the American Statistical Association, 79(385):110–117, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harville:1985:DPE

[Har85] David A. Harville. Decomposition of prediction error. Journal of the American Statistical Association, 80(389):132–138, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hartigan:1987:ECD

[Har87] J. A. Hartigan. Estimation of a convex density contour in two dimensions. Journal of the American Statistical Association, 82 (397):267–270, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harder:1990:ECV

[Har90a] Melinda H. Harder. Effects of censoring on the validity of confidence intervals. Journal of the American Statistical Association, 85(409):90–96, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harris:1990:RDI

[Har90b] Jeffrey E. Harris. Reporting delays and the incidence of AIDS. Journal of the American Statistical Association, 85(412):915–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harder:1992:MDE

[Har92] Melinda H. Harder. Minimum distance estimators of scale with censored data. Journal of the American Statistical Association, 87(419):832–843, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hasselblad:1969:EFM

[Has69] V. Hasselblad. Estimation of finite mixtures of distributions from the exponential family. Journal of the American Statistical Association, 64(??):1459–1471, ???? 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hobert:1997:AFS

[HAS97] James P. Hobert, N. S. Altman, and Carl L. Schofield. Analyses of fish species richness with spatial covariate. Journal of the American Statistical Association, 92(439):846–854, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/hobert.htm.

Haunsperger:1992:DPS

[Hau92] Deanna B. Haunsperger. Dictionaries of paradoxes for statistical tests on k samples. Journal of the American Statistical Association, 87(417):149–155, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hawkins:1993:AES

[Haw93] Douglas M. Hawkins. The accuracy of elemental set approximations for regression. Journal of the American Statistical Association, 88(422):580–589, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hayter:1986:MFE

[Hay86] Anthony J. Hayter. The maximum familywise error rate of Fisher’s least significant difference test. Journal of the American Statistical Association, 81(396):1000–1004, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hayter:1989:PCG

[Hay89] A. J. Hayter. Pairwise comparisons of generally correlated means. Journal of the American Statistical Association, 84(405):208–213, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hayter:1990:OSS

[Hay90] A. J. Hayter. A one-sided Studentized range test for testing against a simple ordered alternative. Journal of the American Statistical Association, 85(411):778–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hayakawa:1994:CNB

[Hay94] Yu Hayakawa. The construction of new bivariate exponential distributions from a Bayesian perspective. Journal of the American Statistical Association, 89(427):1044–1049, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Haber:1986:MLM

[HB86] Michael Haber and Morton B. Brown. Maximum likelihood methods for log-linear models when expected frequencies are subject to linear constraints. Journal of the American Statistical Association, 81(394):477–482, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hardle:1988:BNR

[HB88] Wolfgang Härdle and Adrian W. Bowman. Bootstrapping in nonparametric regression: local adaptive smoothing and confidence bands. Journal of the American Statistical Association, 83(401): 102–110, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hsu:1999:SCI

[HB99] Jason C. Hsu and Roger L. Bergier. Stepwise confidence intervals without multiplicity adjustment for dose response and toxicity studies. Journal of the American Statistical Association, 94(446): 468–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hora:1984:STW

[HC84] Stephen C. Hora and W. J. Conover. The F statistic in the two-way layout with rank-score transformed data. Journal of the American Statistical Association, 79(387):668–673, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hobert:1996:EIP

[HC96] James P. Hobert and George Casella. The effect of improper priors on Gibbs sampling in hierarchical linear mixed models. Journal of the American Statistical Association, 91(436):1461–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hong:1988:GE

[HD88] Chong Sun Hong and H. T. David. Granularity and efficiency. Journal of the American Statistical Association, 83(403):880–885, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hwang:1997:PIA

[HD97] J. T. Gene Hwang and A. Adam Ding. Prediction intervals for artificial neural networks. Journal of the American Statistical Association, 92(438):747–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Higgins:1997:TSA

[HDG97] Karen M. Higgins, Marie Davidian, and David M. Giltinan. A two-step approach to measurement error in time-dependent covariates in nonlinear mixed-effects models, with application to IGF-I pharmacokinetics. Journal of the American Statistical Association, 92(438):436–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hsu:1983:SMC

[HE83] Jason C. Hsu and Donald G. Edwards. Sequential multiple comparisons with the best. Journal of the American Statistical Association, 78(384):958–964, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:1990:APE

[He90] Kun He. An ancillarity paradox in the estimation of multinomial probabilities. Journal of the American Statistical Association, 85 (411):824–828, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:1999:CRD

[He99] Xuming He. Comment: Regression depth. Journal of the American Statistical Association, 94(446):403–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heckman:1988:MES

[Hec88] Nancy E. Heckman. Minimax estimates in a semiparametric model. Journal of the American Statistical Association, 83(404): 1090–1096, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heckman:1996:CIC

[Hec96] James J. Heckman. Comment: Identification of causal effects using instrumental variables. Journal of the American Statistical Association, 91(434):444–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heitjan:1991:NMS

[Hei91] Daniel F. Heitjan. Nonlinear modeling of serial immunologic data: A case study. Journal of the American Statistical Association, 86 (416):891–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heitjan:1995:BAP

[Hei95] Daniel F. Heitjan. Bayesian analysis of a protocol for sampling preserved tumor specimens. Journal of the American Statistical Association, 90(429):38–??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hemmerle:1979:BHU

[Hem79] W. J. Hemmerle. Balanced hypotheses and unbalanced data. Journal of the American Statistical Association, 74(368):794–798, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hayre:1981:SSL

[HG81] Lakhbir S. Hayre and John C. Gittins. Sequential selection of the larger of two normal means. Journal of the American Statistical Association, 76(375):696–700, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:1982:NSU

[HG82] J. S. Huang and Malay Ghosh. A note on strong unimodality of order statistics. Journal of the American Statistical Association, 77(380):929–930, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See corrigenda [HG83].

Huang:1983:CNS

[HG83] J. S. Huang and M. Ghosh. Corrigenda: “A note on strong unimodality of order statistics” [J. Amer. Statist. Assoc. 77 (1982), no. 380, 929–930, MR 84d:62098]. Journal of the American Statistical Association, 78(384):1008, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [HG82].

Hall:1990:BTD

[HH90] Peter Hall and Jeffrey D. Hart. Bootstrap test for difference between means in nonparametric regression. Journal of the American Statistical Association, 85(412):1039–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hulting:1991:SBN

[HH91] Frederick L. Hulting and David A. Harville. Some Bayesian and non-Bayesian procedures for the analysis of comparative experiments and for small-area estimation: Computational aspects, frequentist properties, and relationships. Journal of the American Statistical Association, 86(415):557–568, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hayter:1994:RBS

[HH94] Anthony J. Hayter and Jason C. Hsu. On the relationship between stepwise decision procedures and confidence sets. Journal of the American Statistical Association, 89(425):128–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Horowitz:1996:DSE

[HH96] Joel L. Horowitz and Wolfgang Hardle. Direct semiparametric estimation of single-index models with discrete covariates. Journal of the American Statistical Association, 91(436):1632–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hougaard:1992:MSB

[HHH92] Philip Hougaard, Bent Harvald, and Niels V. Holm. Measuring the similarities between the lifetimes of adult Danish twins born between 1881–1930. Journal of the American Statistical Association, 87(417):17–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hardle:1988:HFA

[HHM88] Wolfgang Härdle, Peter Hall, and J. S. Marron. How far are automatically chosen regression smoothing parameters from their optimum? with comments by David W. Scott and Iain Johnstone and a reply by the authors. Journal of the American Statistical Association, 83(401):86–101, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hardle:1992:RSP

[HHM92] W. Härdle, P. Hall, and J. S. Marron. Regression smoothing parameters that are not far from their optimum. Journal of the American Statistical Association, 87(417):227–233, March 1992.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hansen:1988:EBI

[HHO88] Lars Peter Hansen, John C. Heaton, and Masao Ogaki. Efficiency bounds implied by multiperiod conditional moment restrictions. Journal of the American Statistical Association, 83(403):863–871, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hall:1997:CMO

[HHS97] Peter Hall, Catherine Huber, and Paul L. Speckman. Covariatematched one-sided tests for the difference between functional means. Journal of the American Statistical Association, 92(439): 1074–1083, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/sept/hall_p.htm.

Higdon:1998:AVM

[Hig98] David M. Higdon. Auxiliary variable methods for Markov chain Monte Carlo with applications. Journal of the American Statistical Association, 93(442):585–595, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/june/HIGDON.HTM.

Hill:1970:ZLP

[Hil70] Bruce M. Hill. Zipf’s law and prior distributions for the composition of a population. Journal of the American Statistical Association, 65(331):1220–1232, September 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2284288.

Hill:1974:RFF

[Hil74] Bruce M. Hill. The rank-frequency form of Zipf’s law. Journal of the American Statistical Association, 69(348):1017–1026, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hinkley:1971:ITP

[Hin71] D. V. Hinkley. Inference in two-phase regression. Journal of the American Statistical Association, 66(??):736–743, ???? 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hirji:1992:CED

[Hir92] Karim F. Hirji. Computing exact distributions for polytomous response data. Journal of the American Statistical Association, 87(418):487–492, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoaglin:1986:PSR

[HIT86] David C. Hoaglin, Boris Iglewicz, and John W. Tukey. Performance of some resistant rules for outlier labeling. Journal of the American Statistical Association, 81(396):991–999, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comment [Sim87a].

Harville:1992:MSE

[HJ92] David A. Harville and Daniel R. Jeske. Mean squared error of estimation or prediction under a general linear model. Journal of the American Statistical Association, 87(419):724–731, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hjort:1994:EAN

[Hjo94] Nils Lid Hjort. The exact amount of t-ness that the normal model can tolerate. Journal of the American Statistical Association, 89 (426):665–675, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hendricks:1992:HSM

[HK92] Wallace Hendricks and Roger Koenker. Hierarchical spline models for conditional quantiles and the demand for electricity. Journal of the American Statistical Association, 87(417):58–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harvey:1993:FHE

[HK93] Andrew Harvey and Siem Jan Koopman. Forecasting hourly electricity demand using time-varying splines. Journal of the American Statistical Association, 88(424):1228–??, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hansen:1998:TM

[HKS98] Mark Hansen, Charles Kooperberg, and Sylvain Sardy. Triogram models. Journal of the American Statistical Association, 93 (441):101–119, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/HANSEN.HTM.

Holland:1981:EFP

[HL81] Paul W. Holland and Samuel Leinhardt. An exponential family of probability distributions for directed graphs. with comments by Ronald L. Breiger, Stephen E. Fienberg, Stanley Wasserman, Ove Frank and Shelby J. Haberman and a reply by the authors. Journal of the American Statistical Association, 76(373):33–65, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heitjan:1994:AST

[HL94] Daniel F. Heitjan and J. Richard Landis. Assessing secular trends in blood pressure: A multiple- imputation approach. Journal of the American Statistical Association, 89(427):750–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:1996:ERM

[HL96] X. Joan Hu and Jerald F. Lawless. Estimation of rate and mean functions from truncated recurrent event data. Journal of the American Statistical Association, 91(433):300–310, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huffer:1997:ADS

[HL97] Fred W. Huffer and Chien-Tai Lin. Approximating the distribution of the scan statistic using moments of the number of clumps. Journal of the American Statistical Association, 92(440): 1466–1475, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/HUFFER.HTM.

Heagerty:1998:CLA

[HL98] Patrick J. Heagerty and Subhash R. Lele. A composite likelihood approach to binary spatial data. Journal of the American Statistical Association, 93(443):1099–1111, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/heagerty.html.

Hines:1992:DCM

[HLC92] R. J. O’Hara Hines, J. F. Lawless, and E. M. Carter. Diagnostics for a cumulative multinomial generalized linear model, with applications to grouped toxicological mortality data. Journal of the American Statistical Association, 87(420):1059–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hettmansperger:1983:GII

[HM83] Thomas P. Hettmansperger and Joseph W. McKean. A geometric interpretation of inferences based on ranks in the linear model. Journal of the American Statistical Association, 78(384):885–893, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hedayat:1985:CEU

[HM85] A. S. Hedayat and Dibyen Majumdar. Combining experiments under Gauss–Markov models. Journal of the American Statistical Association, 80(391):698–703, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hallin:1988:RBT

[HM88a] Marc Hallin and Guy Mélard. Rank-based tests for randomness against first-order serial dependence. Journal of the American Statistical Association, 83(404):1117–1128, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hannan:1988:RPA

[HM88b] E. J. Hannan and A. J. McDougall. Regression procedures for ARMA estimation. Journal of the American Statistical Association, 83(402):490–498, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [HM89].

Hannan:1989:CRP

[HM89] E. J. Hannan and A. J. McDougall. Correction: “Regression procedures for ARMA estimation”. Journal of the American Statistical Association, 84(406):636, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [HM88b].

Huffer:1991:WLS

[HM91] Fred W. Huffer and Ian W. McKeague. Weighted least squares estimation for Aalen’s additive risk model. Journal of the American Statistical Association, 86(413):114–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hawkins:1997:HBL

[HM97] Douglas M. Hawkins and Geoffrey J. McLachlan. High-breakdown linear discriminant analysis. Journal of the American Statistical Association, 92(437):136–143, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/hawkins.htm.

Hoover:1993:UED

[HMC^{+}93] D. R. Hoover, A. Munoz, V. Carey, J. M. G. Taylor, M. VanRaden, J. S. Chmiel, L. Kingsley, and The Multicenter Aids Cohort Study. Using events from dropouts in nonparametric survival function estimation with application to incubation of AIDS. Journal of the American Statistical Association, 88(421):37–??, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hardle:1998:TPV

[HMM98] Wolfgang Härdle, Enno Mammen, and Marlene Müller. Testing parametric versus semiparametric modeling in generalized linear models. Journal of the American Statistical Association, 93(444): 1461–1474, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/dec/hardle.html.

Hettmansparger:1997:AIM

[HMO97a] Thomas P. Hettmansparger, Jyrki Mottonen, and Hannu Oja. Affine-invariant multivariate one-sample signed-rank tests. Journal of the American Statistical Association, 92(440):1591– 1600, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/HETTMANS.HTM.

Hettmansperger:1997:AIM

[HMO97b] Thomas P. Hettmansperger, Jyrki Mottonen, and Hanne Oja. Affine-invariant multivariate one-sample signed-rank tests. Journal of the American Statistical Association, 92(440):1591– 1600, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/HETTMANS.HTM.

Hirji:1987:CDE

[HMP87] Karim F. Hirji, Cyrus R. Mehta, and Nitin R. Patel. Computing distributions for exact logistic regression. Journal of the American Statistical Association, 82(400):1110–1117, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hardle:1992:BCA

[HMT92] W. Härdle, J. S. Marron, and A. B. Tsybakov. Bandwidth choice for average derivative estimation. Journal of the American Statistical Association, 87(417):218–226, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hollander:1997:LRB

[HMY97] Myles Hollander, Ian W. McKeague, and Jie Yang. Likelihood ratio — based confidence bands for survival functions. Journal of the American Statistical Association, 92(437):215–226, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/hollander.htm.

Hettmansperger:1987:TPA

[HN87] Thomas P. Hettmansperger and Robert M. Norton. Tests for patterned alternatives in k-sample problems. Journal of the American Statistical Association, 82(397):292–299, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Horowitz:1992:GMS

[HN92] Joel L. Horowitz and George R. Neumann. A generalized moments specification test of the proportional hazards model. Journal of the American Statistical Association, 87(417):234–240, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoaglin:1973:RRE

[Hoa73] David C. Hoaglin. Review: Robust Estimates of Location: Survey and Advances, by D. F. Andrews, P. J. Bickel, F. R. Hampel, P. J. Huber, W. H. Rogers, and J. W. Tukey. Journal of the American Statistical Association, 69(345):272–274, March 1973. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2285544.

Hodges:1987:AAN

[Hod87] James S. Hodges. Assessing the accuracy of normal approximations. Journal of the American Statistical Association, 82(397): 149–154, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hodoshima:1988:ETI

[Hod88] Jiro Hodoshima. The effect of truncation on the identifiability of regression coefficients. Journal of the American Statistical Association, 83(404):1055–1056, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hodges:1994:AUD

[Hod94] James S. Hodges. Analytical use of data from army training exercise: a case study of tactical reconnaissance. Journal of the American Statistical Association, 89(426):444–??, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoeffding:1963:PIS

[Hoe63] W. Hoeffding. Probability inequalities for sums of bounded random variables. Journal of the American Statistical Association, 58(301):13–30, March 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoeffding:1984:RPU

[Hoe84] Wassily Hoeffding. Range preserving unbiased estimators in the multinomial case. Journal of the American Statistical Association, 79(387):712–714, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoffman:1992:CLB

[Hof92] Lorrie L. Hoffman. Characterizing linear birth and death processes. Journal of the American Statistical Association, 87(420): 1183–1187, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hogg:1974:ARP

[Hog74] R. V. Hogg. Adaptive robust procedures: a partial review and some suggestions for future applications and theory. Journal of the American Statistical Association, 69(348):909–923, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Holland:1986:SCI

[Hol86] Paul W. Holland. Statistics and causal inference. with discussion and a reply by the author. Journal of the American Statistical Association, 81(396):945–970, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Holden:1987:TSA

[Hol87] Robert T. Holden. Time series analysis of a contagious process. Journal of the American Statistical Association, 82(400):1019– 1026, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Holmgren:1995:PMC

[Hol95] Eric B. Holmgren. The P − P plot as a method for comparing treatment effects. Journal of the American Statistical Association, 90(429):360–??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Holm:1999:MCS

[Hol99] Sture Holm. Multiple confidence sets based on stagewise tests. Journal of the American Statistical Association, 94(446):489–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hughes-Oliver:1998:AUS

[HOLDG98] Jacqueline M. Hughes-Oliver, Jye-Chyi Lu, Joseph C. Davis, and Ronald S. Gyurcsik. Achieving uniformity in a semiconductor fabrication process using spatial modeling. Journal of the American Statistical Association, 93(441):36–45, March 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/HUGHES-O.HTM.

Hooper:1989:ERV

[Hoo89] Peter M. Hooper. Experimental randomization and the validity of normal-theory inference. Journal of the American Statistical Association, 84(406):576–586, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hooper:1993:IWL

[Hoo93] Peter M. Hooper. Iterative weighted least squares estimation in heteroscedastic linear models. Journal of the American Statistical Association, 88(421):179–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hossjer:1994:RBE

[Hös94a] Ola Hössjer. Rank-based estimates in the linear model with high breakdown point. Journal of the American Statistical Association, 89(425):149–158, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hughes-Oliver:1994:TSA

[HOS94b] Jacqueline M. Hughes-Oliver and William H. Swallow. A two-stage adaptive group-testing procedure for estimating small proportions. Journal of the American Statistical Association, 89(427):982–993, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Host:1995:SIE

[HOS95] Gudmund Høst, Henning Omre, and Paul Switzer. Spatial interpolation errors for monitoring data. Journal of the American Statistical Association, 90(431):853–861, ???? 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Houdre:1995:SAC

[Hou95] Christian Houdré. Some applications of covariance identities and inequalities to functions of multivariate normal variables. Journal of the American Statistical Association, 90(431):965–968, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hollander:1988:NTU

[HP88] Myles Hollander and Edsel Peña. Nonparametric tests under restricted treatment-assignment rules. Journal of the American Statistical Association, 83(404):1144–1151, ???? 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hollander:1992:CSG

[HP92] Myles Hollander and Edsel A. Peña. A chi-squared goodness-of-fit test for randomly censored data. Journal of the American Statistical Association, 87(418):458–463, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hwang:1981:OHP

[HPE81] Frank K. Hwang, Charles G. Pfeifer, and Peter Enis. An optimal hierarchical procedure for a modified binomial group-testing problem. Journal of the American Statistical Association, 76(376): 947–949, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hollander:1986:CLD

[HPP86] Myles Hollander, Dong Ho Park, and Frank Proschan. A class of life distributions for aging. Journal of the American Statistical Association, 81(393):91–95, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hardwick:1998:SDB

[HPS98] Janis Hardwick, Connie Page, and Quentin F. Stout. Sequentially deciding between two experiments for estimating a common success probability. Journal of the American Statistical Association, 93(444):1502–1511, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/dec/hardwick.html.

Hinkley:1984:ATD

[HR84] D. V. Hinkley and G. Runger. The analysis of transformed data. with discussion. Journal of the American Statistical Association, 79(386):302–320, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heitjan:1990:ICD

[HR90] Daniel F. Heitjan and Donald B. Rubin. Inference from coarse data via multiple imputation with application to age heaping. Journal of the American Statistical Association, 85(410):304–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heritier:1994:RBI

[HR94] Stéphane Heritier and Elvezio Ronchetti. Robust boundedinfluence tests in general parametric models. Journal of the American Statistical Association, 89(427):897–904, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:1997:SEP

[HR97] Jian Huang and A. J. Rossini. Sieve estimation for the proportional-odds failure-TIme regression model with interval censoring. Journal of the American Statistical Association, 92(439): 960–967, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/sept/huang.htm.

Hampel:1981:CVC

[HRR81] Frank R. Hampel, Peter J. Rousseeuw, and Elvezio Ronchetti. The change-of-variance curve and optimal redescending M-estimators. Journal of the American Statistical Association, 76(375):643–648, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heckman:1990:TME

[HRW90] James J. Heckman, Richard Robb, and James R. Walker. Testing the mixture of exponentials hypothesis and estimating the mixing distribution by the method of moments. Journal of the American Statistical Association, 85(410):582–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hardle:1989:ISM

[HS89a] Wolfgang Härdle and Thomas M. Stoker. Investigating smooth multiple regression by the method of average derivatives. Journal of the American Statistical Association, 84(408):986–995, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hastie:1989:PC

[HS89b] Trevor Hastie and Werner Stuetzle. Principal curves. Journal of the American Statistical Association, 84(406):502–516, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hedayat:1989:RBP

[HS89c] A. Hedayat and J. Stufken. A relation between pairwise balanced and variance balanced block designs. Journal of the American Statistical Association, 84(407):753–755, ???? 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hadi:1993:PIM

[HS93] Ali S. Hadi and Jeffrey S. Simonoff. Procedures for the identification of multiple outliers in linear models. Journal of the American Statistical Association, 88(424):1264–??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huggins:1994:VCM

[HS94] Richard M. Huggins and Robert G. Staudte. Variance components models for dependent cell populations. Journal of the American Statistical Association, 89(425):19–??, March 1994. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Haas:1998:ENC

[HS98a] Peter J. Haas and Lynne Stokes. Estimating the number of classes in a finite population. Journal of the American Statistical Association, 93(444):1475–1487, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/dec/haas.html.

Hall:1998:EGE

[HS98b] Daniel B. Hall and Thomas A. Severini. Extended generalized estimating equations for clustered data. Journal of the American Statistical Association, 93(444):1365–1375, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/dec/hall.html.

He:1998:MBS

[HS98c] Xuming He and Peide Shi. Monotone B-spline smoothing. Journal of the American Statistical Association, 93(442):643–650, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/june/HE.HTM.

Hsiao:1997:ABF

[Hsi97] Chuhsing Kate Hsiao. Approximate Bayes factors when a mode occurs on the boundary. Journal of the American Statistical Association, 92(438):656–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:1990:BRT

[HSP90] Xuming He, Douglas G. Simpson, and Stephen L. Portnoy. Breakdown robustness of tests. Journal of the American Statistical Association, 85(410):446–452, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hsu:1982:SIR

[Hsu82] Jason C. Hsu. Simultaneous inference with respect to the best treatment in block designs. Journal of the American Statistical Association, 77(378):461–467, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hudson:1981:SPE

[HT81] H. Malcolm Hudson and Kam Wah Tsui. Simultaneous Poisson estimators for a priori hypotheses about means. Journal of the American Statistical Association, 76(373):182–187, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hillmer:1982:AMB

[HT82] C. Hillmer and G. C. Tiao. An ARIMA-model-based approach to seasonal adjustment. Journal of the American Statistical Association, 77(377):63–70, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hillmer:1987:BET

[HT87] Steven C. Hillmer and Abdelwahed Trabelsi. Benchmarking of economic time series. Journal of the American Statistical Association, 82(400):1064–1071, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hansen:1990:REF

[HT90] Morris H. Hansen and Benjamin J. Tepping. Regression estimates in Federal welfare quality control programs. Journal of the American Statistical Association, 85(411):856–??, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hall:1997:IMA

[HT97a] Peter Hall and Berwin A. Turlach. Interpolation methods for adapting to sparse design in nonparametric regression. Journal of the American Statistical Association, 92(438):466–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/june/hall.htm.

Hall:1997:RIM

[HT97b] Peter Hall and Berwin A. Turlach. Rejoinder: Interpolation methods for adapting to sparse design in nonparametric regression. Journal of the American Statistical Association, 92(438):476–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hsiao:1997:PAL

[HT97c] Cheng Hsiao and A. Kamil Tahmiscioğlu. A panel analysis of liquidity constraints and firm investment. Journal of the American Statistical Association, 92(438):455–??, June 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hastie:1994:FDA

[HTB94] Trevor Hastie, Robert Tibshirani, and Andreas Buja. Flexible discriminant analysis by optimal scoring. Journal of the American Statistical Association, 89(428):1255–1270, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hughes:1995:OSS

[HTP95] Michael D. Hughes, Simon G. Thompson, and Stuart J. Pocock. Optimal sequential screening guidelines for quantitative risk factors measured with error. Journal of the American Statistical Association, 90(429):19–??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:1997:TGF

[Hua97] Li-Shan Huang. Testing goodness-of-fit based on a roughness measure. Journal of the American Statistical Association, 92(440): 1399–1402, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/HUANG_LS.HTM.

Huber:1983:MAB

[Hub83] Peter J. Huber. Minimax aspects of bounded-influence regression. with discussion. Journal of the American Statistical Association, 78(381):66–80, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hunsberger:1994:SRL

[Hun94a] Sally Hunsberger. Semiparametric regression in likelihood-based models. Journal of the American Statistical Association, 89(428): 1354–1365, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hunter:1994:SP

[Hun94b] J. Stuart Hunter. Statistics as a profession. Journal of the American Statistical Association, 89(425):1–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hurd:1985:NEM

[Hur85a] Cuthbert C. Hurd. A note on early Monte Carlo computations and scientific meetings. Annals of the History of Computing, 7(2):141–155, April/June 1985. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1985/pdf/a2141.pdf;   http://www.computer.org/
annals/an1985/a2141abs.htm. See [MU49].

Hurvich:1985:DDC

[Hur85b] Clifford M. Hurvich. Data-driven choice of a spectrum estimate: extending the applicability of cross-validation methods. Journal of the American Statistical Association, 80(392):933–940, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hartung:1986:BIU

[HV86] Joachim Hartung and Bernard Voet. Best invariant unbiased estimators for the mean squared error of variance component estimators. Journal of the American Statistical Association, 81(395): 689–691, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hochberg:1984:SPC

[HVS84] Yosef Hochberg and Yetty Varon-Salomon. On simultaneous pairwise comparisons in analysis of covariance. Journal of the American Statistical Association, 79(388):863–866, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heiberger:1983:GTD

[HVY83] Richard M. Heiberger, Paul F. Velleman, and M. Agelia Ypelaar. Generating test data with independently controllable features for multivariate general linear forms. Journal of the American Statistical Association, 78(383):585–595, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hill:1975:SFZ

[HW75] Bruce M. Hill and Michael Woodroofe. Stronger forms of Zipf’s law. Journal of the American Statistical Association, 70(349):212– 219, March 1975. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hart:1986:KRE

[HW86] Jeffrey D. Hart and Thomas E. Wehrly. Kernel regression estimation using repeated measurements data. Journal of the American Statistical Association, 81(396):1080–1088, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hinich:1988:TAU

[HW88] Melvin J. Hinich and Murray A. Wolinsky. A test for aliasing using bispectral analysis. Journal of the American Statistical Association, 83(402):499–502, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hadi:1990:AEM

[HW90a] Ali S. Hadi and Martin T. Wells. Assessing the effects of multiple rows on the condition number of a matrix. Journal of the American Statistical Association, 85(411):786–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heckman:1990:EFD

[HW90b] James J. Heckman and James R. Walker. Estimating fecundability from data on waiting times to first conception. Journal of the American Statistical Association, 85(410):283–??, June 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hall:1991:GMR

[HW91] Peter Hall and Thomas E. Wehrly. A geometrical method for removing edge effects from kernel- type nonparametric regression estimators. Journal of the American Statistical Association, 86 (415):665–672, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hart:1992:KRW

[HW92a] Jeffrey D. Hart and Thomas E. Wehrly. Kernel regression when the boundary region is large, with an application to testing the adequacy of polynomial models. Journal of the American Statistical Association, 87(420):1018–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hasofer:1992:TEV

[HW92b] A. M. Hasofer and Z. Wang. A test for extreme value domain of attraction. Journal of the American Statistical Association, 87 (417):171–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Handcock:1994:ASS

[HW94] Mark S. Handcock and James R. Wallis. An approach to statistical spatial-temporal modeling of meteorological fields. with comments and a rejoinder by Handcock. Journal of the American Statistical Association, 89(426):368–390, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hamada:1995:ACD

[HW95a] M. Hamada and C. F. J. Wu. Analysis of censored data from fractionated experiments: A Bayesian approach. Journal of the American Statistical Association, 90(430):467–??, ???? 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:1995:EOR

[HW95b] Ying Huang and Mei Cheng Wang. Estimating the occurrence rate for prevalent survival data in competing risks models. Journal of the American Statistical Association, 90(432):1406–1415, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hwang:1986:MEV

[Hwa86] Jiunn T. Hwang. Multiplicative errors-in-variables models with applications to recent data released by the U.S. Department of Energy. Journal of the American Statistical Association, 81(395): 680–688, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hall:1999:MEC

[HWY99] Peter Hall, Rodney C. L. Wolff, and Qiwei Yao. Methods for estimating a conditional distribution function. Journal of the American Statistical Association, 94(445):154–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoel:1986:IHC

[HY86] D. G. Hoel and T. Yanagawa. Incorporating historical controls in testing for a trend in proportions. Journal of the American Statistical Association, 81(396):1095–1099, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hart:1998:OSC

[HY98] Jeffrey D. Hart and Seongbaek Yi. One-sided cross-validation. Journal of the American Statistical Association, 93(442):620–631, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic).  URL http://www.amstat.org/publications/
jasa/abstracts_98/june/HART.HTM.

Hedayat:1997:MIO

[HYP97] A. S. Hedayat, Bo Yan, and John M. Pezzuto. Modeling and identifying optimum designs for fitting dose- response curves based on raw optical density data. Journal of the American Statistical Association, 92(439):1132–1140, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/hedayat.htm.

Hurvich:1990:FDS

[HZ90] Clifford M. Hurvich and Scott L. Zeger. A frequency domain selection criterion for regression with autocorrelated errors. Journal of the American Statistical Association, 85(411):705–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heagerty:1996:MRM

[HZ96] Patrick J. Heagerty and Scott L. Zeger. Marginal regression models for clustered ordinal measurements. Journal of the American Statistical Association, 91(435):1024–??, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heagerty:1998:LRA

[HZ98] Patrick J. Heagerty and Scott L. Zeger. Lorelogram: A regression approach to exploring dependence in longitudinal categorical responses. Journal of the American Statistical Association, 93 (441):150–162, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/HEAGERTY.HTM.

Haseman:1990:PRC

[HZS^{+}90] Joseph K. Haseman, Errol Zeiger, Michael D. Shelby, Barry H. Margolin, and Raymond W. Tennant. Predicting rodent carcinogenicity from four in vitro genetic toxicity assays: An evaluation of 114 chemicals studied by the National Toxicology Program. Journal of the American Statistical Association, 85(412):964–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ibrahim:1990:IDG

[Ibr90] Joseph G. Ibrahim. Incomplete data in generalized linear models. Journal of the American Statistical Association, 85(411):765–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Isaki:1982:SDU

[IF82] Cary T. Isaki and Wayne A. Fuller. Survey design under the regression superpopulation model. Journal of the American Statistical Association, 77(377):89–96, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Iman:1984:CAD

[IHC84] Ronald L. Iman, Stephen C. Hora, and W. J. Conover. Comparison of asymptotically distribution-free procedures for the analysis of complete blocks. Journal of the American Statistical Association, 79(387):674–685, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ibrahim:1991:BAG

[IL91] Joseph G. Ibrahim and Purushottam W. Laud. On Bayesian analysis of generalized linear models using Jeffreys’s prior. Journal of the American Statistical Association, 86(416):981–986, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ibrahim:1994:PAA

[IL94] Joseph G. Ibrahim and Purushottam W. Laud. A predictive approach to the analysis of designed experiments. Journal of the American Statistical Association, 89(425):309–319, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Iman:1995:NPS

[Ima95] Ronald L. Iman. New paradigms for the statistics profession. Journal of the American Statistical Association, 90(429):1–??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ibrahim:1998:ACS

[IRC98] Joseph G. Ibrahim, Louise M. Ryan, and Ming-Hui Chen. Applications and case studies — using historical controls to adjust for covariates in trend tests for binary data. Journal of the American Statistical Association, 93(444):1282–1293, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/dec/ibrahim.html.

Iwase:1983:UMV

[IS83] Kōsei Iwase and Noriaki Setô. Uniformly minimum variance unbiased estimation for the inverse Gaussian distribution. Journal of the American Statistical Association, 78(383):660–663, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Izenman:1987:SCR

[IS87] Alan J. Izenman and Sanat K. Sarkar. Simultaneous confidence regions for the frequency analysis of multiple time series. Journal of the American Statistical Association, 82(397):271–275, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Isaki:1974:BTC

[Isa74] C. T. Isaki. Bias of two classification procedures in a dichotomous population. Journal of the American Statistical Association, 69 (348):957–960, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Isaki:1983:VEU

[Isa83] Cary T. Isaki. Variance estimation using auxiliary information. Journal of the American Statistical Association, 78(381):117–123, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Israel:1981:SPN

[Isr81] Robert B. Israel. Stronger players need not win more knockout tournaments. Journal of the American Statistical Association, 76 (376):950–951, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Inclan:1994:UCS

[IT94] Carla Inclán and George C. Tiao. Use of cumulative sums of squares for retrospective detection of changes of variance. Journal of the American Statistical Association, 89(427):913–923, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Izenman:1991:RDN

[Ize91] Alan Julian Izenman. Recent developments in nonparametric density estimation. Journal of the American Statistical Association, 86(413):205–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jacroux:1989:OBD

[Jac89] Mike Jacroux. The A-optimality of block designs for comparing test treatments with a control. Journal of the American Statistical Association, 84(405):310–317, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jacroux:1993:CTR

[Jac93] Mike Jacroux. On the construction of trend-resistant designs for comparing a set of test treatments with a set of controls. Journal of the American Statistical Association, 88(424):1398–1403, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jaynes:1979:IMD

[Jay79] E. T. Jaynes. Inference, method and decision: Towards a Bayesian philosophy of science. Journal of the American Statistical Association, 74(367):740–741, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jayasuriya:1996:TPR

[Jay96] Bodhini R. Jayasuriya. Testing for polynomial regression using nonparametric regression techniques. Journal of the American Statistical Association, 91(436):1626–??, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ji:1992:BCV

[JB92] Zhang Ji and Dennis D. Boos. Bootstrap critical values for testing homogeneity of covariance matrices. Journal of the American Statistical Association, 87(418):425–??, June 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jae:1991:SCC

[JD91] Won Lee Jae and David L. DeMets. Sequential comparison of changes with repeated measurements data. Journal of the American Statistical Association, 86(415):757–??, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jones:1994:VKT

[JDP94] M. C. Jones, S. J. Davies, and B. U. Park. Versions of kerneltype regression estimators. Journal of the American Statistical Association, 89(427):825–832, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jennings:1986:JIA

[Jen86a] Dennis E. Jennings. Judging inference adequacy in logistic regression. Journal of the American Statistical Association, 81(394): 471–476, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jennings:1986:ORD

[Jen86b] Dennis E. Jennings. Outliers and residual distributions in logistic regression. Journal of the American Statistical Association, 81 (396):987–990, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jensen:1991:LDT

[Jen91] Jens Ledet Jensen. A large deviation-type approximation for the “Box Class” of likelihood ratio criteria. Journal of the American Statistical Association, 86(414):437–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1983:PVD

[JG83] Wesley Johnson and Seymour Geisser. A predictive view of the detection and characterization of influential observations in regression analysis. Journal of the American Statistical Association, 78 (381):137–144, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jacqmin-Gadda:1995:THG

[JGC95] Hélène Jacqmin-Gadda and Daniel Commenges. Tests of homogeneity for generalized linear models. Journal of the American Statistical Association, 90(432):1237–1246, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jensen:1982:EFT

[JH82] D. R. Jensen and Y. V. Hui. Efficiency of Friedman’s χ_{r}^{2} test under dependence. Journal of the American Statistical Association, 77 (378):468–474, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:1998:CEG

[Jia98] Jiming Jiang. Consistent estimators in generalized linear mixed models. Journal of the American Statistical Association, 93 (442):720–729, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/JIANG.HTM.

Jamshidian:1993:CGA

[JJ93] Mortaza Jamshidian and Robert I. Jennrich. Conjugate gradient acceleration of the EM algorithm. Journal of the American Statistical Association, 88(421):221–228, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Juritz:1983:ASP

[JJS83] J. M. Juritz, J. W. F. Juritz, and M. A. Stephens. On the accuracy of simulated percentage points. Journal of the American Statistical Association, 78(382):441–444, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

James:1984:EE

[JJW84] Barry R. James, Kang Ling James, and Henrique Westenberger. An efficient R-estimator for the ed50. Journal of the American Statistical Association, 79(385):164–173, ???? 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jorn:1989:CAU

[JK89] Hongsuk Jorn and Jerome Klotz. The contrast algorithm for unbalanced multifactor analysis of variance. Journal of the American Statistical Association, 84(405):318–324, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1993:EHN

[JK93] Richard Johnson and Jerome Klotz. Estimating hot numbers and testing uniformity for the lottery. Journal of the American Statistical Association, 88(422):662–668, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1996:CIS

[JK96] Darren J. Johnson and K. Krishnamoorthy. Combining independent studies in a calibration problem. Journal of the American Statistical Association, 91(436):1707–??, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jones:1996:BSB

[JMS96] M. C. Jones, J. S. Marron, and S. J. Sheather. A brief survey of bandwidth selection for density estimation. Journal of the American Statistical Association, 91(433):401–407, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jacroux:1997:DCO

[JMS97] Mike Jacroux, Dibyen Majumdar, and Kirti R. Shah. On the determination and construction of optimal block designs in the presence of linear trends. Journal of the American Statistical Association, 92(437):375–382, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/march/jacroux.htm.

Jewell:1994:NEF

[JMV94] Nicholas P. Jewell, Hina M. Malani, and Eric Vittinghoff. Nonparametric estimation for a form of doubly censored data, with application to two problems in AIDS. Journal of the American Statistical Association, 89(425):7–??, March 1994. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Joe:1989:REM

[Joe89] Harry Joe. Relative entropy measures of multivariate dependence. Journal of the American Statistical Association, 84(405):157–164, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Joe:1995:AMN

[Joe95] Harry Joe. Approximations to multivariate normal rectangle probabilities based on conditional expectations. Journal of the American Statistical Association, 90(431):957–964, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1978:MTC

[Joh78] N. J. Johnson. Modified t tests and confidence intervals for asymmetrical populations. Journal of the American Statistical Association, 73(363):536–544, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1990:MHE

[Joh90] Robert Alan Johnson. Measurement of Hispanic ethnicity in the U.S. census: An evaluation based on latent-class analysis. Journal of the American Statistical Association, 85(409):58–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1992:TEM

[Joh92] Valen E. Johnson. A technique for estimating marginal posterior densities in hierarchical models using mixtures of conditional densities. Journal of the American Statistical Association, 87(419): 852–860, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1994:MSA

[Joh94] Valen E. Johnson. A model for segmentation and analysis of noisy images. Journal of the American Statistical Association, 89(425): 230–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1996:BAM

[Joh96a] Valen E. Johnson. On Bayesian analysis of multirater ordinal data: An application to automated essay grading. Journal of the American Statistical Association, 91(433):42–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1996:SCM

[Joh96b] Valen E. Johnson. Studying convergence of Markov chain Monte Carlo algorithms using coupled sample paths. Journal of the American Statistical Association, 91(433):154–166, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1998:CEP

[Joh98a] Valen E. Johnson. Comment: Edge preserving smoothers for image processing. Journal of the American Statistical Association, 93(442):541–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1998:CRS

[Joh98b] Valen E. Johnson. A coupling-regeneration scheme for diagnosing convergence in Markov chain Monte Carlo algorithms. Journal of the American Statistical Association, 93(441):238–248, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/JOHNSON.HTM.

Jones:1989:DIK

[Jon89] M. C. Jones. Discretized and interpolated kernel density estimates. Journal of the American Statistical Association, 84(407):733–741, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jones:1996:ISM

[Jon96] Michael P. Jones. Indicator and stratification methods for missing explanatory variables in multiple linear regression. Journal of the American Statistical Association, 91(433):222–230, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Joe:1985:ENF

[JR85] H. Joe and N. Reid. Estimating the number of faults in a system. Journal of the American Statistical Association, 80(389):222–226, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jan:1996:ITS

[JR96] Show-Li Jan and Ronald H. Randles. Interdirection tests for simple repeated-measures designs. Journal of the American Statistical Association, 91(436):1611–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jensen:1994:AJD

[JS94] D. R. Jensen and Herbert Solomon. Approximations to joint distributions of definite quadratic forms. Journal of the American Statistical Association, 89(426):480–486, June 1994. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jones:1997:CHO

[JS97a] M. C. Jones and D. F. Signorini. A comparison of higher-order bias kernel density estimators. Journal of the American Statistical Association, 92(439):1063–1073, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/jones_mc.htm.

Juster:1997:IQE

[JS97b] F. Thomas Juster and James P. Smith. Improving the quality of economic data: Lessons from the HRS and AHEAD. Journal of the American Statistical Association, 92(440):1268– 1278, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/JUSTER.HTM.

Jennen-Steinmetz:1988:UAN

[JSG88] Christine Jennen-Steinmetz and Theo Gasser. A unifying approach to nonparametric regression estimation. Journal of the American Statistical Association, 83(404):1084–1089, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1998:MMQ

[JSKK98] Richard A. Johnson, Songyong Sim, Barbara E. Klein, and Ronald Klein. A multivariate MultiSample quantile test for ordered alternatives. Journal of the American Statistical Association, 93 (442):807–818, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/JOHNSON.HTM.

Jennison:1997:GSA

[JT97] Christopher Jennison and Bruce W. Turnbull. Group-sequential analysis incorporating covariate information. Journal of the American Statistical Association, 92(440):1330–1341, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/JENNISON.HTM.

Johnson:1980:NFP

[JTB80] M. E. Johnson, G. L. Tietjen, and R. J. Beckman. A new family of probability distributions with applications to Monte Carlo studies. Journal of the American Statistical Association, 75(370):276–279, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [JTB81].

Johnson:1981:CNF

[JTB81] Mark E. Johnson, Gary L. Tietjen, and Richard J. Beckman. Correction: “A new family of probability distributions with applications to Monte Carlo studies” [J. Amer. Statist. Assoc. 75 (1980), no. 370, 276–279, MR 81f:62032]. Journal of the American Statistical Association, 76(376):1022, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [FH81c, JTB80].

Jiang:1999:SRM

[JTC99] Wenxin Jiang, Bruce W. Turnbull, and Larry C. Clark. Semiparametric regression models for repeated events with random effects and measurement error. Journal of the American Statistical Association, 94(445):111–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Judkins:1996:CVE

[Jud96] David R. Judkins. Comment: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91(434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Judkins:1998:ACS

[Jud98] David R. Judkins. Applications and case studies — comment: Not asked and not answered: Multiple imputation for multiple surveys.

Journal of the American Statistical Association, 93(443):861–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jung:1996:QLM

[Jun96] Sin-Ho Jung. Quasi-likelihood for median regression models. Journal of the American Statistical Association, 91(433):251–257, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnstone:1985:RLR

[JV85] Iain M. Johnstone and Paul F. Velleman. The resistant line and related regression methods. with a comment by John W. Tukey. Journal of the American Statistical Association, 80(392):1041– 1059, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:1988:BSM

[JW88] Richard A. Johnson and Samaradasa Weerahandi. A Bayesian solution to the multivariate Behrens–Fisher problem. Journal of the American Statistical Association, 83(401):145–149, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:1997:RBS

[jW97] Tiee jian Wu. Root n bandwidth selectors for kernel estimation of density derivatives. Journal of the American Statistical Association, 92(438):536–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kohn:1986:EPI

[KA86] Robert Kohn and Craig F. Ansley. Estimation, prediction, and interpolation for ARIMA models with missing data. Journal of the American Statistical Association, 81(395):751–761, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kohn:1991:SEA

[KA91] Robert Kohn and Craig F. Ansley. A signal extraction approach to the estimation of treatment and control curves. Journal of the American Statistical Association, 86(416):1034–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:1995:IEI

[KA95] Donguk Kim and Alan Agresti. Improved exact inference about conditional association in three-way contingency tables. Journal of the American Statistical Association, 90(430):632–639, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kadane:1990:SAA

[Kad90] Joseph B. Kadane. A statistical analysis of adverse impact of employer decisions. Journal of the American Statistical Association, 85(412):925–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kadane:1998:ACS

[Kad98] Joseph B. Kadane. Applications and case studies — comment: Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443):867–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kaigh:1994:DFT

[Kai94] W. D. Kaigh. Distribution-free two-sample tests based on rank spacings. Journal of the American Statistical Association, 89(425): 159–167, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kanoh:1988:RWC

[Kan88] Satoru Kanoh. The reduction of the width of confidence bands in linear regression. Journal of the American Statistical Association, 83(401):116–122, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kariya:1981:BCM

[Kar81] Takeaki Kariya. Bounds for the covariance matrices of Zellner’s estimator in the SUR model and the 2SAE in a heteroscedastic model. Journal of the American Statistical Association, 76(376): 975–979, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kariya:1985:NVG

[Kar85] Takeaki Kariya. A nonlinear version of the Gauss–Markov theorem. Journal of the American Statistical Association, 80(390):

476–477, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Karrison:1987:RML

[Kar87] Theodore Karrison. Restricted mean life with adjustment for covariates. Journal of the American Statistical Association, 82(400): 1169–1176, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kohn:1991:PCV

[KAT91] Robert Kohn, Craig F. Ansley, and David Tharm. The performance of cross-validation and maximum likelihood estimators of spline smoothing parameters. Journal of the American Statistical Association, 86(416):1042–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kano:1993:SIB

[KBB93] Yutaka Kano, Maia Berkane, and Peter M. Bentler. Statistical inference based on pseudo-maximum likelihood estimators in elliptical populations. Journal of the American Statistical Association, 88(421):135–143, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kooperberg:1997:PR

[KBS97] Charles Kooperberg, Smarajit Bose, and Charles J. Stone. Polychotomous regression. Journal of the American Statistical Association, 92(437):117–??, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/kooperbe.htm.

Kim:1993:RLR

[KC93] Hyune-Ju Kim and Lujing Cai. Robustness of the likelihood ratio test for a change in simple linear regression. Journal of the American Statistical Association, 88(423):864–871, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kosorok:1996:ALO

[KC96] Michael R. Kosorok and Wei-Hsiung Chao. The analysis of longitudinal ordinal response data in continuous time. Journal of the American Statistical Association, 91(434):807–817, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:1993:SER

[KCHH93] P. T. Kim, E. M. Carter, J. J. Hubert, and K. J. Hand. Shrinkage estimators of relative potency. Journal of the American Statistical Association, 88(422):615–621, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kramer:1987:SAA

[KD87] Walter Krämer and Christian Donninger. Spatial autocorrelation among errors and the relative efficiency of OLS in the linear regression model. Journal of the American Statistical Association, 82(398):577–579, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Keenan:1982:TSA

[Kee82] Daniel MacRae Keenan. A time series analysis of binary data. Journal of the American Statistical Association, 77(380):816–821, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kettenring:1997:PAS

[Ket97] Jon R. Kettenring. Presidential address: Shaping statistics for success in the 21st Century. Journal of the American Statistical Association, 92(440):1229–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/kettenring.html.

Kackar:1984:ASE

[KH84] Raghu N. Kackar and David A. Harville. Approximations for standard errors of estimators of fixed and random effects in mixed linear models. Journal of the American Statistical Association, 79 (388):853–862, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Khuri:1981:SCI

[Khu81] A. I. Khuri. Simultaneous confidence intervals for functions of variance components in random models. Journal of the American Statistical Association, 76(376):878–885, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kiefer:1978:PNA

[Kie78] J. Kiefer. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments: Comment. Journal of the American Statistical Association, 73(363): 523–524, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/pdfplus/2286594.pdf. See [SM78a].

Kim:1986:LCB

[Kim86] Woo Chul Kim. A lower confidence bound on the probability of a correct selection. Journal of the American Statistical Association, 81(396):1012–1017, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

King:1985:SCP

[Kin85] Benjamin F. King. Surveys combining probability and quota methods of sampling. Journal of the American Statistical Association, 80(392):890–896, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kitagawa:1987:NGS

[Kit87] Genshiro Kitagawa. Non-Gaussian state-space modeling of nonstationary time series. with comments and a reply by the author. Journal of the American Statistical Association, 82(400): 1032–1063, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kitagawa:1998:SOS

[Kit98] Genshiro Kitagawa. A self-organizing state space model. Journal of the American Statistical Association, 93(443):1203– 1215, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/kitagawa.html.

Kanoh:1984:OSS

[KK84a] Satoru Kanoh and Yoshiatsu Kusunoki. One-sided simultaneous bounds in linear regression. Journal of the American Statistical Association, 79(387):715–719, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kelton:1984:HTM

[KK84b] W. David Kelton and Christina M. L. Kelton. Hypothesis tests for Markov process models estimated from aggregate frequency data. Journal of the American Statistical Association, 79(388):922–928, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Knight:1998:CBC

[KKT98] Keith Knight, Rafal Kustra, and Robert Tibshirani. Comment: Bayesian CART model search. Journal of the American Statistical Association, 93(443):950–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kalbfleisch:1985:APD

[KL85] J. D. Kalbfleisch and J. F. Lawless. The analysis of panel data under a Markov assumption. Journal of the American Statistical Association, 80(392):863–871, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Keramidas:1990:FTS

[KL90] Elaine M. Keramidas and Jack C. Lee. Forecasting technological substitutions with concurrent short time series. Journal of the American Statistical Association, 85(411):625–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Keyes:1996:GPF

[KL96] Tim K. Keyes and Martin S. Levy. Goodness of prediction fit for multivariate linear models. Journal of the American Statistical Association, 91(433):191–197, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kallenberg:1997:DDS

[KL97] Wilbert C. M. Kallenberg and Teresa Ledwina. Data-driven smooth tests when the hypothesis is composite. Journal of the American Statistical Association, 92(439):1094–1104, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/kallenberg.htm.

Kallenberg:1999:DDR

[KL99] Wilbert C. M. Kallenberg and Teresa Ledwina. Data-driven rank tests for independence. Journal of the American Statistical Association, 94(445):285–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kong:1994:SIB

[KLW94] Augustine Kong, Jun S. Liu, and Wong Wing Hung. Sequential imputations and Bayesian missing data problems. Journal of the American Statistical Association, 89(425):278–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Krantz:1983:PLR

[KM83] David H. Krantz and John Miyamoto. Priors and likelihood ratios as evidence. Journal of the American Statistical Association, 78 (382):418–423, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kao:1991:OSA

[KM91a] Tzu-Cheg Kao and George P. McCabe. Optimal sample allocation for normal discrimination and logistic regression under stratified sampling. Journal of the American Statistical Association, 86(414):432–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Karlin:1991:SSP

[KM91b] Samuel Karlin and Catherine Macken. Some statistical problems in the assessment of inhomogeneities of DNA sequence data. Journal of the American Statistical Association, 86(413):27–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kitagawa:1996:DCC

[KM96a] Genshiro Kitagawa and Norio Matsumoto. Detection of coseismic changes of underground water level. Journal of the American Statistical Association, 91(434):521–??, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koshevoy:1996:LZM

[KM96b] Gleb Koshevoy and Karl Mosler. The Lorenz zonoid of a multivariate distribution. Journal of the American Statistical Association, 91(434):873–882, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kadane:1999:YAP

[KMT99] Joseph B. Kadane, Michael M. Meyer, and John W. Tukey. Yule’s association paradox and ignored stratum heterogeneity in capturerecapture studies. Journal of the American Statistical Association, 94(447):855–859, 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kneip:1998:CSS

[Kne98] Alois Kneip. Comment: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443):984–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kadane:1995:UFA

[KO95] Joseph B. Kadane and Anthony O’Hagan. Using finitely additive probability: Uniform distributions on the natural numbers. Journal of the American Statistical Association, 90(430):626–631, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kooperberg:1996:POP

[KO96] Charles Kooperberg and Finbarr O’Sullivan. Predictive oscillation patterns: A synthesis of methods for spatial-temporal decomposition of random fields. Journal of the American Statistical Association, 91(436):1485–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kobayashi:1985:CES

[Kob85] Masahito Kobayashi. Comparison of efficiencies of several estimators for linear regressions with autocorrelated errors. Journal of the American Statistical Association, 80(392):951–953, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kobayashi:1986:BTE

[Kob86] Masahito Kobayashi. A bounds test of equality between sets of coefficients in two linear regressions when disturbance variances are unequal. Journal of the American Statistical Association, 81 (394):510–513, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koehler:1986:GFT

[Koe86] Kenneth J. Koehler. Goodness-of-fit tests for log-linear models in sparse contingency tables. Journal of the American Statistical Association, 81(394):483–493, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koenker:1999:DRD

[Koe99] Roger Koenker. Discussion: Regression depth. Journal of the American Statistical Association, 94(446):405–??, June 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kolaczyk:1996:WSA

[Kol96] Eric D. Kolaczyk. A wavelet shrinkage approach to tomographic image reconstruction. Journal of the American Statistical Association, 91(435):1079–1090, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koopman:1997:EIK

[Koo97] Siem Jan Koopman. Exact initial Kalman filtering and smoothing for nonstationary time series models. Journal of the American Statistical Association, 92(440):1630–1638, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/KOOPMAN.HTM.

Korn:1981:HLL

[Kor81] Edward L. Korn. Hierarchical log-linear models not preserved by classification error. Journal of the American Statistical Association, 76(373):110–113, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kallenberg:1985:NCC

[KOS85] W. C. M. Kallenberg, J. Oosterhoff, and B. F. Schriever. The number of classes in chi-squared goodness-of-fit tests. Journal of the American Statistical Association, 80(392):959–968, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kott:1989:CCR

[Kot89] Phillip S. Kott. Comment to: “Conditioning ratio estimates under simple random sampling” [J. Amer. Statist. Assoc. 82 (1987), no. 399, 826–831, MR 88h:62016] by J. Robinson. Journal of the American Statistical Association, 84(405):352, ???? 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Rob87].

Kott:1994:NHN

[Kot94] Phillip S. Kott. A note on handling nonresponse in sample surveys. Journal of the American Statistical Association, 89(426):693–696, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kourouklis:1981:CAG

[KP81a] S. Kourouklis and C. C. Paige. A constrained approach to the general Gauss–Markov linear model. Journal of the American Statistical Association, 76(??):620–625, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kronmal:1981:VAR

[KP81b] R. A. Kronmal and A. V. Peterson. A variant of the acceptancerejection method for computer generation of random variables. Journal of the American Statistical Association, 76(374):446–451, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See corrigendum [KP82].

Kronmal:1982:CVA

[KP82] Richard A. Kronmal and Arthur V. Peterson, Jr. Corrigendum: “A variant of the acceptance-rejection method for computer generation of random variables” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 446–451, MR 82h:62037]. Journal of the American Statistical Association, 77(380):954, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [KP81b].

Kenett:1983:SDS

[KP83] Ron Kenett and Moshe Pollak. On sequential detection of a shift in the probability of a rare event. Journal of the American Statistical Association, 78(382):389–395, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koenker:1987:ELM

[KP87] Roger Koenker and Stephen Portnoy. L-estimation for linear models. Journal of the American Statistical Association, 82(399):851– 857, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koenker:1990:EMR

[KP90] Roger Koenker and Stephen Portnoy. M estimation of multivariate regressions. Journal of the American Statistical Association, 85 (412):1060–1068, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kedem:1991:TMR

[KP91] Benjamin Kedem and Harry Pavlopoulos. On the threshold method for rainfall estimation: Choosing the optimal threshold level. Journal of the American Statistical Association, 86(415): 626–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Klein:1992:ABC

[KP92] Rubin Klein and S. James Press. Adaptive Bayesian classification of spatial data. Journal of the American Statistical Association, 87(419):844–??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kateri:1997:AMC

[KP97] Maria Kateri and Takis Papaioannou. Asymmetry models for contingency tables. Journal of the American Statistical Association, 92(439):1124–1032, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/kateri.htm.

Kleffe:1991:OEC

[KPR91] J. Kleffe, N. G. N. Prasad, and J. N. K. Rao. “optimal” estimation of correlated response variance under additive models. Journal of the American Statistical Association, 86(413):144–150, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kinderman:1976:CGN

[KR76] A. J. Kinderman and J. G. Ramage. Computer generation of normal random variables. Journal of the American Statistical Association, 71(356):893–896, December 1976. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2286857.

Kepner:1982:DUM

[KR82] James L. Kepner and Ronald H. Randles. Detecting unequal marginal scales in a bivariate population. Journal of the American Statistical Association, 77(378):475–482, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kepner:1984:DFR

[KR84] James L. Kepner and David H. Robinson. A distribution-free rank test for ordered alternatives in randomized complete block designs. Journal of the American Statistical Association, 79(385):212–217, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kleffe:1986:EAU

[KR86] J. Kleffe and J. N. K. Rao. The existence of asymptotically unbiased nonnegative quadratic estimates of variance components in ANOVA models. Journal of the American Statistical Association, 81(395):692–698, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kepner:1988:NMD

[KR88] James L. Kepner and David H. Robinson. Nonparametric methods for detecting treatment effects in repeated-measures designs. Journal of the American Statistical Association, 83(402):456–461, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Krzysztofowicz:1993:SBP

[KR93] Roman Krzysztofowicz and Sharon Reese. Stochastic bifurcation processes and distributions of fractions. Journal of the American Statistical Association, 88(421):345–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kass:1995:BF

[KR95] Robert E. Kass and Adrian E. Raftery. Bayes factors. Journal of the American Statistical Association, 90(430):773–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kahn:1996:DRM

[KR96] Michael J. Kahn and Adrian E. Raftery. Discharge rates of medicare stroke patients to skilled nursing facilities: Bayesian logistic regression with unobserved heterogeneity. Journal of the American Statistical Association, 91(433):29–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kramer:1986:LSR

[Krä86] Walter Krämer. Least squares regression when the independent variable follows an ARIMA process. Journal of the American Statistical Association, 81(393):150–154, ???? 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kramer:1990:SPW

[Kra90] Fredrica D. Kramer. Statistics and policy in welfare quality control: A basis for understanding and assessing competing views. Journal of the American Statistical Association, 85(411):850–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kremers:1986:CUE

[Kre86] Walter K. Kremers. Completeness and unbiased estimation for sum-quota sampling. Journal of the American Statistical Association, 81(396):1070–1073, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kruskal:1988:MSC

[Kru88] William Kruskal. Miracles and statistics: the casual assumption of independence. Journal of the American Statistical Association, 83(404):929–940, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Krzanowski:1979:BGC

[Krz79] W. J. Krzanowski. Between-groups comparison of principal components. Journal of the American Statistical Association, 74(367): 703–707, ???? 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [Krz81].

Krzanowski:1981:CBG

[Krz81] W. J. Krzanowski. Correction: “Between-groups comparison of principal components” [J. Amer. Statist. Assoc. 74 (1979), no. 367, 703–707, MR 81a:62049]. Journal of the American Statistical Association, 76(376):1022, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Krz79].

Krzysztofowicz:1987:MFP

[Krz87] Roman Krzysztofowicz. Markovian forecast processes. Journal of the American Statistical Association, 82(397):31–37, ???? 1987.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kvam:1993:EDF

[KS93] Paul H. Kvam and Francisco J. Samaniego. On estimating distribution functions using nomination samples. Journal of the American Statistical Association, 88(424):1317–1322, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kvam:1994:NML

[KS94] Paul H. Kvam and Francisco J. Samaniego. Nonparametric maximum likelihood estimation based on ranked set samples. Journal of the American Statistical Association, 89(426):526–537, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kianifard:1996:RDA

[KS96] Farid Kianifard and William H. Swallow. A review of the development and application of recursive residuals in linear models. Journal of the American Statistical Association, 91(433):391–400, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kouassi:1997:SAH

[KS97] Djokouri A. Kouassi and Jagbir Singh. A semiparametric approach to hazard estimation with randomly censored observations. Journal of the American Statistical Association, 92(440): 1351–1355, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/KOUASSI.HTM.

Kunsch:1989:CUB

[KSC89] Hans R. Künsch, Leonard A. Stefanski, and Raymond J. Carroll. Conditionally unbiased bounded-influence estimation in general regression models, with applications to generalized linear models. Journal of the American Statistical Association, 84(406):460–466, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kaiser:1994:SML

[KSJ94] Mark S. Kaiser, Paul L. Speckman, and John R. Jones. Statistical models for limiting nutrient relations in inland waters. Journal of the American Statistical Association, 89(426):410–??, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kadane:1996:RFC

[KSS96] Joseph B. Kadane, Mark J. Schervish, and Teddy Seidenfeld. Reasoning to a foregone conclusion. Journal of the American Statistical Association, 91(435):1228–1235, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kooperberg:1995:HR

[KST95] Charles Kooperberg, Charles J. Stone, and Young K. Truong. Hazard regression. Journal of the American Statistical Association, 90(429):78–94, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kakizawa:1998:DCM

[KST98] Yoshihide Kakizawa, Robert H. Shumway, and Masanobu Taniguchi. Discrimination and clustering for multivariate time series. Journal of the American Statistical Association, 93(441): 328–340, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/KAKIZAWA.HTM.

Kimeldorf:1992:MMS

[KSW92] George Kimeldorf, Allan R. Sampson, and Lyn R. Whitaker. Min and max scorings for two-sample ordinal data. Journal of the American Statistical Association, 87(417):241–247, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Knafl:1985:CBR

[KSY85] George Knafl, Jerome Sacks, and Donald Ylvisaker. Confidence bands for regression functions. Journal of the American Statistical Association, 80(391):683–691, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kafadar:1988:BT

[KT88] Karen Kafadar and John W. Tukey. A bidec t table. Journal of the American Statistical Association, 83(402):532–539, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:1994:TBS

[KT94a] Dong K. Kim and Jeremy M. G. Taylor. Transform-both-sides approach for overdispersed binomial data when N is unobserved. Journal of the American Statistical Association, 89(427):833–845, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kolassa:1994:ACI

[KT94b] John E. Kolassa and Martin A. Tanner. Approximate conditional inference in exponential families via the Gibbs sampler. Journal of the American Statistical Association, 89(426):697–702, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [KT94c].

Kolassa:1994:CAC

[KT94c] John E. Kolassa and Martin A. Tanner. Correction to: “Approximate conditional inference in exponential families via the Gibbs sampler”. Journal of the American Statistical Association, 89 (428):1572, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [KT94b].

Kim:1995:REA

[KT95] Dong K. Kim and Jeremy M. G. Taylor. The restricted EM algorithm for maximum likelihood estimation under linear restrictions on the parameters. Journal of the American Statistical Association, 90(430):708–716, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kulasekera:1995:CRC

[Kul95] K. B. Kulasekera. Comparison of regression curves using quasiresiduals. Journal of the American Statistical Association, 90 (431):1085–1093, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kunitomo:1980:AED

[Kun80] Naoto Kunitomo. Asymptotic expansions of the distributions of estimators in a linear functional relationship and simultaneous equations. Journal of the American Statistical Association, 75(371): 693–700, ???? 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kunert:1990:PTM

[Kun90] Joachim Kunert. On the power of tests for multiple comparisons of three normal means. Journal of the American Statistical Association, 85(411):808–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kushner:1998:OER

[Kus98] H. B. Kushner. Optimal and efficient repeated measurements designs for uncorrelated observations. Journal of the American Statistical Association, 93(443):1176–1187, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/kushner.html.

Krasker:1982:EBI

[KW82] William S. Krasker and Roy E. Welsch. Efficient boundedinfluence regression estimation. Journal of the American Statistical Association, 77(379):595–604, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Keener:1985:MLR

[KW85] Robert W. Keener and Donald M. Waldman. Maximum likelihood regression of rank-censored data. Journal of the American Statistical Association, 80(390):385–392, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kadane:1988:SPE

[KW88] Joseph B. Kadane and Robert L. Winkler. Separating probability elicitation from utilities. Journal of the American Statistical Association, 83(402):357–363, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kass:1995:RBT

[KW95] Robert E. Kass and Larry Wasserman. A reference Bayesian test for nested hypotheses and its relationship to the Schwarz criterion. Journal of the American Statistical Association, 90(431):928–934, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kass:1996:SPD

[KW96a] Robert E. Kass and Larry Wasserman. The selection of prior distributions by formal rules. Journal of the American Statistical Association, 91(435):1343–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kepner:1996:RTT

[KW96b] James L. Kepner and Dennis D. Wackerly. On rank transformation techniques for balanced incomplete repeated-measures designs. Journal of the American Statistical Association, 91(436): 1619–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kulasekera:1997:SPS

[KW97] K. B. Kulasekera and J. Wang. Smoothing parameter selection for power optimality in testing of regression curves. Journal of the American Statistical Association, 92(438):500–??, June 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kunitomo:1985:PPM

[KY85] Naoto Kunitomo and Taku Yamamoto. Properties of predictors in misspecified autoregressive time series models. Journal of the American Statistical Association, 80(392):941–950, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koch:1986:MTI

[KY86] Paul D. Koch and Shie-Shien Yang. A method for testing the independence of two time series that accounts for a potential pattern in the cross-correlation function. Journal of the American Statistical Association, 81(394):533–544, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kuo:1996:BCN

[KY96] Lynn Kuo and Tae Young Yang. Bayesian computation for nonhomogeneous Poisson processes in software reliability. Journal of the American Statistical Association, 91(434):763–773, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kosorok:1999:VFI

[KY99] Michael R. Kosorok and Chin-Lu Yin. The versatility of functionindexed weighted log-rank statistics. Journal of the American Statistical Association, 94(445):320–??, March 1999. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Leger:1993:AIV

[LA93] Christian Léger and Naomi Altman. Assessing influence in variable selection problems. Journal of the American Statistical Association, 88(422):547–556, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lang:1994:SMJ

[LA94] Joseph B. Lang and Alan Agresti. Simultaneously modeling joint and marginal distributions of multivariate categorical responses. Journal of the American Statistical Association, 89(426):625–??, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lambert:1982:QRT

[Lam82] Diane Lambert. Qualitative robustness of tests. Journal of the American Statistical Association, 77(378):352–357, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lang:1996:PGF

[Lan96] Joseph B. Lang. On the partitioning of goodness-of-fit statistics for multivariate categorical response models. Journal of the American Statistical Association, 91(435):1017–1023, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lange:1998:CSS

[Lan98] Nicholas Lange. Comment: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443):986–??, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

LaRiccia:1986:ACS

[LaR86] Vincent N. LaRiccia. Asymptotically chi-squared distributed tests of normality for type ii censored samples. Journal of the American Statistical Association, 81(396):1026–1031, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

LaRiccia:1991:SGF

[LaR91] V. N. LaRiccia. Smooth goodness-of-fit tests: A quantile function approach. Journal of the American Statistical Association, 86(414):427–431, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laslett:1994:KSE

[Las94] Geoffrey M. Laslett. Kriging and splines: an empirical comparison of their predictive performance in some applications. with comments and a rejoinder by the author. Journal of the American Statistical Association, 89(426):391–409, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lauritzen:1992:PPM

[Lau92] Steffen L. Lauritzen. Propagation of probabilities, means, and variances in mixed graphical association models. Journal of the American Statistical Association, 87(420):1098–1108, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lavine:1991:ARB

[Lav91a] Michael Lavine. An approach to robust Bayesian analysis for multidimensional parameter spaces. Journal of the American Statistical Association, 86(414):400–403, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lavine:1991:PEI

[Lav91b] Michael Lavine. Problems in extrapolation illustrated with Space Shuttle O-ring data. Journal of the American Statistical Association, 86(416):919–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lavine:1991:SBS

[Lav91c] Michael Lavine. Sensitivity in Bayesian statistics: The prior and the likelihood. Journal of the American Statistical Association, 86(414):396–399, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lawless:1981:MSE

[Law81] J. F. Lawless. Mean squared error properties of generalized ridge estimators. Journal of the American Statistical Association, 76 (374):462–466, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See corrigendum [Law82].

Lawless:1982:CMS

[Law82] J. F. Lawless. Corrigendum: “Mean squared error properties of generalized ridge estimators” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 462–466, MR 82h:62113]. Journal of the American Statistical Association, 77(380):954, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Law81].

Lawless:1987:RMP

[Law87] J. F. Lawless. Regression methods for Poisson process data. Journal of the American Statistical Association, 82(399):808–815, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lindsay:1993:MNM

[LB93] Bruce G. Lindsay and Prasanta Basak. Multivariate normal mixtures: A fast consistent method of moments. Journal of the American Statistical Association, 88(422):468–476, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lewis:1994:GSC

[LB94a] Roger J. Lewis and Donald A. Berry. Group sequential clinical trials: A classical evaluation of Bayesian decision-theoretic designs. Journal of the American Statistical Association, 89(428):1528–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:1994:REM

[LB94b] Zhaohai Li and Colin B. Begg. Random effects models for combining results from controlled and uncontrolled studies in a metaanalysis. Journal of the American Statistical Association, 89(428): 1523–1527, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lindstrom:1994:CNR

[LB94c] Mary J. Lindstrom and Douglas M. Bates. Correction to: “Newton–Raphson and EM algorithms for linear mixed-effects models for repeated measures-data” [J. Amer. Statist. Assoc. 83 (1988), no. 404, 1014–1022]. Journal of the American Statistical Association, 89(428):1572, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [LB94d].

Lindstrom:1994:NRE

[LB94d] Mary J. Lindstrom and Douglas M. Bates. Newton–Raphson and EM algorithms for linear mixed-effects models for repeated measures-data. Journal of the American Statistical Association, 89(404):1014–1022, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [LB94c].

Lin:1996:BCG

[LB96] Xihong Lin and Norman E. Breslow. Bias correction in generalized linear mixed models with multiple components of dispersion. Journal of the American Statistical Association, 91(435): 1007–1016, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

LeBlanc:1993:STG

[LC93] Michael LeBlanc and John Crowley. Survival trees by goodness of split. Journal of the American Statistical Association, 88(422): 457–467, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

LeBlanc:1995:SRF

[LC95a] Michael LeBlanc and John Crowley. Semiparametric regression functionals. Journal of the American Statistical Association, 90 (429):95–105, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1995:BDS

[LC95b] Jun S. Liu and Rong Chen. Blind deconvolution via sequential imputations. Journal of the American Statistical Association, 90 (430):567–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1998:SMC

[LC98a] Jun S. Liu and Rong Chen. Sequential Monte Carlo methods for dynamic systems. Journal of the American Statistical Association, 93(443):1032–1044, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/liu2.html.

Llu:1998:SMC

[LC98b] Jun S. Llu and Rong Chen. Sequential Monte Carlo methods for dynamic systems. Journal of the American Statistical Association, 93(443):1032–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lindsay:1991:SER

[LCG91] Bruce Lindsay, Clifford C. Clogg, and John Grego. Semiparametric estimation in the Rasch model and related exponential response models, including a simple latent class model for item analysis. Journal of the American Statistical Association, 86(413):96–107, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lange:1992:HBM

[LCG^{+}92] Nicholas Lange, Bradley P. Carlin, Alan E. Gelfand, Thomas A. Louis, Jeremy M. G. Taylor, Mark R. Segal, and Bradley P. Carlin. Hierarchical Bayes models for the progression of HIV infection using longitudinal CD4 T-cell numbers. Journal of the American Statistical Association, 87(419):615–??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lambert:1997:ER

[LCK97] Diane Lambert, George Casella, and Karen Kafadar. Editors’ report for 1996. Journal of the American Statistical Association, 92(438):391–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:1992:MFU

[LCMA92] Ronald D. Lee, Lawrence R. Carter, Robert McNown, and Juha M. Alho. Modeling and forecasting U.S. mortality. Journal of the American Statistical Association, 87(419):659–??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1998:RCS

[LCW98] Jun S. Liu, Rong Chen, and Wing Hung Wong. Rejection control and sequential importance sampling. Journal of the American Statistical Association, 93(443):1022–1031, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/liu1.html.

Linton:1997:ATA

[LCWH97] Oliver B. Linton, Rong Chen, Naiysin Wang, and Wolfgang Hardle. An analysis of transformations for additive nonparametric regression. Journal of the American Statistical Association, 92(440): 1512–1521, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/LINTON.HTM.

Lee:1991:SCC

[LD91] Jae Won Lee and David L. DeMets. Sequential comparison of changes with repeated measurements data. Journal of the American Statistical Association, 86(415):757–762, ???? 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:1992:SRT

[LD92] Jae Won Lee and David L. DeMets. Sequential rank tests with repeated measurements in clinical trials. Journal of the American Statistical Association, 87(417):136–142, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Leamer:1981:CFR

[Lea81] Edward E. Leamer. Coordinate-free ridge regression bounds. Journal of the American Statistical Association, 76(376):842–849, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Leamer:1984:GSR

[Lea84] Edward E. Leamer. Global sensitivity results for generalized least squares estimates. Journal of the American Statistical Association, 79(388):867–870, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ledwina:1994:DDV

[Led94] Teresa Ledwina. Data-driven version of Neyman’s smooth test of fit. Journal of the American Statistical Association, 89(427): 1000–1005, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:1987:CST

[Lee87] Chu-In Charles Lee. Chi-squared tests for and against an order restriction on multinomial parameters. Journal of the American Statistical Association, 82(398):611–618, ???? 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:1988:PEG

[Lee88] Jack C. Lee. Prediction and estimation of growth curves with special covariance structures. Journal of the American Statistical Association, 83(402):432–440, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:1996:EMD

[Lee96] Chu-In Charles Lee. On estimation for monotone dose-response curves. Journal of the American Statistical Association, 91(435): 1110–1119, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lehmann:1981:ICB

[Leh81] E. L. Lehmann. An interpretation of completeness and Basu’s theorem. Journal of the American Statistical Association, 76(374): 335–340, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lehmann:1983:EII

[Leh83] E. L. Lehmann. Estimation with inadequate information. Journal of the American Statistical Association, 78(383):624–627, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lehmann:1993:FNP

[Leh93] E. L. Lehmann. The Fisher, Neyman–Pearson theories of testing hypotheses: One theory or two? Journal of the American Statistical Association, 88(424):1242–1249, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lele:1994:EFC

[Lel94] Subhash Lele. Estimating functions in chaotic systems. Journal of the American Statistical Association, 89(426):512–516, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lenk:1988:LND

[Len88] Peter J. Lenk. The logistic normal distribution for Bayesian, nonparametric, predictive densities. Journal of the American Statistical Association, 83(402):509–516, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lejeune:1982:SPD

[LF82] Michel Lejeune and G. David Faulkenberry. A simple predictive density function. with a comment by Tom Leonard and a rejoinder by the authors. Journal of the American Statistical Association, 77(379):654–659, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1993:TAO

[LGWC93] P. Y. Liu, Stephanie Green, Michael Wolf, and John Crowley. Testing against ordered alternatives for censored survival data. Journal of the American Statistical Association, 88(421):153–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:1991:SAW

[LH91] Tsung-Hua Lin and David A. Harville. Some alternatives to Wald’s confidence interval and test. Journal of the American Statistical Association, 86(413):179–187, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:1994:ER

[LHA94] Roderick J. A. Little, Myles Hollander, and Alan Agresti. Editor’s report for 1993. Journal of the American Statistical Association, 89(426):367–??, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lambert:1996:ER

[LHA96] Diane Lambert, Myles Hollander, and Alan Agresti. Editors’ report for 1995. Journal of the American Statistical Association, 91(434):443–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:1995:ER

[LHAL95] Roderick J. A. Little, Myles Hollander, Alan Agresti, and Diane Lambert. Editor’s report for 1994. Journal of the American Statistical Association, 90(430):401–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Leonard:1989:BMI

[LHT89] Tom Leonard, John S. J. Hsu, and Kam-Wah Tsui. Bayesian marginal inference. Journal of the American Statistical Association, 84(408):1051–1058, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:1991:SIR

[Li91] Ker-Chau Li. Sliced inverse regression for dimension reduction. with discussion and a rejoinder by the author. Journal of the American Statistical Association, 86(414):316–342, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:1992:PHD

[Li92] Ker-Chau Li. On principal Hessian directions for data visualization and dimension reduction: Another application of Stein’s lemma. Journal of the American Statistical Association, 87(420): 1025–1039, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:1995:NLR

[Li95] Gang Li. Nonparametric likelihood ratio estimation of probabilities for truncated data. Journal of the American Statistical Association, 90(431):997–1003, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:1996:DTS

[Li96] Ta-Hsin Li. Discrimination of time series by parametric filtering. Journal of the American Statistical Association, 91(433):284–293, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lancaster:1998:PDS

[LI98a] Tony Lancaster and Orna Intrator. Panel data with survival: Hospitalization of HIV patients. Journal of the American Statistical Association, 93(441):46–53, March 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/LANCASTE.HTM.

Li:1998:CPH

[Li98b] Ker-Chau Li. Comment: Principle Hessian directions revisited. Journal of the American Statistical Association, 93(441):94–??, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lieberman:1994:SAD

[Lie94] Offer Lieberman. Saddlepoint approximation for the distribution of a ratio of quadratic forms in normal variables. Journal of the American Statistical Association, 89(427):924–928, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ling:1974:CSA

[Lin74] R. F. Ling. Comparison of several algorithms for computing sample means and variances. Journal of the American Statistical Association, 69(348):859–866, ???? 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lindsay:1985:EII

[Lin85] Bruce G. Lindsay. Errors in inspection: integer parameter maximum likelihood in a finite population. Journal of the American Statistical Association, 80(392):879–885, ???? 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:1991:GFA

[Lin91] D. Y. Lin. Goodness-of-fit analysis for the Cox regression model based on a class of parameter estimators. Journal of the American Statistical Association, 86(415):725–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:1982:MNS

[Lit82] Roderick J. A. Little. Models for nonresponse in sample surveys. Journal of the American Statistical Association, 77(378):237–250, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:1992:RMX

[Lit92] Roderick J. A. Little. Regression with missing X’s: A review. Journal of the American Statistical Association, 87(420):1227–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:1993:PMM

[Lit93] Roderick J. A. Little. Pattern-mixture models for multivariate incomplete data. Journal of the American Statistical Association, 88(421):125–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:1995:MDM

[Lit95] Roderick J. A. Little. Modeling the drop-out mechanism in repeated-measures studies. Journal of the American Statistical Association, 90(431):1112–1121, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1994:CGS

[Liu94] Jun S. Liu. The collapsed Gibbs sampler in Bayesian computations with applications to a gene regulation problem. Journal of the American Statistical Association, 89(427):958–966, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1995:CCM

[Liu95] Regina Y. Liu. Control charts for multivariate processes. Journal of the American Statistical Association, 90(432):1380–1387, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1996:BRM

[Liu96] Chuanhai Liu. Bayesian robust multivariate linear regression with incomplete data. Journal of the American Statistical Association, 91(435):1219–1227, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lou:1993:COG

[LJK93] Sheldon Lou, Jiong Jiang, and Kenneth Keng. Clustering objects generated by linear regression models. Journal of the American Statistical Association, 88(424):1356–1362, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ljungovist:1993:UAM

[Lju93] Lars Ljungovist. A unified approach to measures of privacy in randomized response models: A utilitarian perspective. Journal of the American Statistical Association, 88(421):97–??, March 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Langholz:1991:TDH

[LK91] Bryan Langholz and Richard A. Kronmal. Tests of distributional hypotheses with nuisance parameters using Fourier series methods. Journal of the American Statistical Association, 86(416): 1077–1084, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lesaffre:1992:EUM

[LK92a] Emmanuel Lesaffre and Heinz Kaufmann. Existence and uniqueness of the maximum likelihood estimator for a multivariate probit model. Journal of the American Statistical Association, 87(419): 805–811, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lesperance:1992:ACN

[LK92b] Mary L. Lesperance and John D. Kalbfleisch. An algorithm for computing the nonparametric MLE of a mixing distribution. Journal of the American Statistical Association, 87(417):120–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Long:1995:FBD

[LK95] Dou Long and Roman Krzysztofowicz. A family of bivariate densities constructed from marginals. Journal of the American Statistical Association, 90(430):739–746, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lam:1997:MLA

[LK97] K. F. Lam and Anthony Y. C. Kuk. A marginal likelihood approach to estimation in frailty models. Journal of the American Statistical Association, 92(439):985–990, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/lam.htm.

Lu:1997:SIM

[LKC97] Jiandong Lu, Daijin Ko, and Ted Chang. The standardized influence matrix and its applications. Journal of the American Statistical Association, 92(440):1572–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lahiri:1999:PSC

[LKCH99a] Soumendra N. Lahiri, Mark S. Kaiser, Noel Cressie, and Nan-Jung Hsu. Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94 (445):86–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lahiri:1999:RPS

[LKCH99b] Soumendra N. Lahiri, Mark S. Kaiser, Noel Cressie, and Nan-Jung Hsu. Rejoinder: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):107–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lancaster:1998:RCS

[LKM98] Vicki A. Lancaster and Sallie Keller-McNulty. A review of composite sampling methods. Journal of the American Statistical Association, 93(443):1216–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Littell:1981:CRB

[LL81] Ramon C. Littell and William C. Louv. Confidence regions based on methods of combining test statistics. Journal of the American Statistical Association, 76(373):125–130, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Louv:1986:COS

[LL86] William C. Louv and Ramon C. Littell. Combining one-sided binomial tests. Journal of the American Statistical Association, 81(394):550–554, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laird:1987:EBC

[LL87] Nan M. Laird and Thomas A. Louis. Empirical Bayes confidence intervals based on bootstrap samples. with discussion and with a reply by the authors. Journal of the American Statistical Association, 82(399):739–757, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:1997:CAN

[LL97a] Xihong Lin and Roderick J. Little. Comment: Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials. Journal of the American Statistical Association, 92 (440):1245–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ling:1997:FIA

[LL97b] Shiqing Ling and W. K. Li. On fractionally integrated autoregressive moving-average time series models with conditional heteroscedasticity. Journal of the American Statistical Association, 92 (439):1184–1194, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/ling.htm.

Lee:1993:RCP

[LLE93] J. Jack Lee, Ker-Chau Li, and Robert M. Elashoff. On recensoring for censored paired data. Journal of the American Statistical Association, 88(421):104–118, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lloyd:1990:CID

[Llo90] Christopher J. Lloyd. Confidence intervals from the difference between two correlated proportions. Journal of the American Statistical Association, 85(412):1154–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lloyd:1998:USR

[Llo98] Chris J. Lloyd. Using smoothed receiver operating characteristic curves to summarize and compare diagnostic systems. Journal of the American Statistical Association, 93(444):1356– 1364, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/dec/lloyd.html.

Legler:1995:EPT

[LLR95] Julie M. Legler, Myrto Lefkopoulou, and Louise M. Ryan. Efficiency and power of tests for multiple binary outcomes. Journal of the American Statistical Association, 90(430):680–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laird:1987:MLC

[LLS87] Nan Laird, Nicholas Lange, and Daniel Stram. Maximum likelihood computations with repeated measures: application of the EM algorithm. Journal of the American Statistical Association, 82(397):97–105, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laderman:1954:EIP

[LLT54] J. Laderman, S. B. Littauer, and John W. Tukey. Errata: The inventory problem. Journal of the American Statistical Association, 49(268):906, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281550.

Lange:1989:RSM

[LLT89] Kenneth L. Lange, Roderick J. A. Little, and Jeremy M. G. Taylor. Robust statistical modeling using the t distribution. Journal of the American Statistical Association, 84(408):881–896, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lye:1993:REN

[LM93] Jenny N. Lye and Vance L. Martin. Robust estimation, nonnormalities, and generalized exponential distributions. Journal of the American Statistical Association, 88(421):261–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Le:1996:MFS

[LMR96] Nhu D. Le, R. Douglas Martin, and Adrian E. Raftery. Modeling flat stretches, bursts, and outliers in time series using mixture transition distribution models. Journal of the American Statistical Association, 91(436):1504–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1995:BMM

[LNL95] Jun S. Liu, Andrew F. Neuwald, and Charles E. Lawrence. Bayesian models for multiple local sequence alignment and Gibbs sampling strategies. Journal of the American Statistical Association, 90(432):1156–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1999:MSB

[LNL99] Jun S. Liu, Andrew F. Neuwald, and Charles E. Lawrence. Markovian structures in biological sequence alignment. Journal of the American Statistical Association, 94(445):1–??, March 1999. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Loh:1985:NMT

[Loh85] Wei-Yin Loh. A new method for testing separate families of hypotheses. Journal of the American Statistical Association, 80(390): 362–368, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Loh:1986:IER

[Loh86] Wei-Yin Loh. Improved estimators for ratios of variance components. Journal of the American Statistical Association, 81(395): 699–702, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Loh:1987:CCC

[Loh87] Wei-Yin Loh. Calibrating confidence coefficients. Journal of the American Statistical Association, 82(397):155–162, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Longley:1967:ALS

[Lon67] J. W. Longley. An appraisal of least squares programs for the electronic computer from the point of view of the user. Journal of the American Statistical Association, 62(??):819–841, ???? 1967. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Louis:1984:EPP

[Lou84] Thomas A. Louis. Estimating a population of parameter values using Bayes and empirical Bayes methods. Journal of the American Statistical Association, 79(386):393–398, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lou:1996:RLR

[Lou96] W. Y. Wendy Lou. On runs and longest run tests: A method of finite Markov chain imbedding. Journal of the American Statistical Association, 91(436):1595–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Levy:1986:OPF

[LP86] Martin S. Levy and S. K. Perng. An optimal prediction function for the normal linear model. Journal of the American Statistical Association, 81(393):196–198, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lambert:1991:NDL

[LPT91] Diane Lambert, Bruce Peterson, and Irma Terpenning. Nondetects, detection limits, and the probability of detection. Journal of the American Statistical Association, 86(414):266–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lancaster:1985:NTL

[LQ85] J. F. Lancaster and Dana Quade. A nonparametric test for linear regression based on combining Kendall’s tau with the sign test. Journal of the American Statistical Association, 80(390):393–397, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:1997:SLR

[LQT97] Gang Li, Jing Qin, and Ram C. Tiwari. Semiparametric likelihood ratio — based inferences for truncated data. Journal of the American Statistical Association, 92(437):236–245, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/li.htm.

Lindsay:1987:UTI

[LR87] Bruce G. Lindsay and Kathryn Roeder. A unified treatment of integer parameter models. Journal of the American Statistical Association, 82(399):758–764, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lindsay:1992:RDM

[LR92] Bruce G. Lindsay and Kathryn Roeder. Residual diagnostics for mixture models. Journal of the American Statistical Association, 87(419):785–794, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lahiri:1995:REM

[LR95a] P. Lahiri and J. N. K. Rao. Robust estimation of mean squared error of small area estimators. Journal of the American Statistical Association, 90(430):758–766, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lambert:1995:ODG

[LR95b] Diane Lambert and Kathryn Roeder. Overdispersion diagnostics for generalized linear models. Journal of the American Statistical Association, 90(432):1225–1236, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lawrence:1996:LIP

[LR96] Charles Lawrence and Andrew Reilly. Likelihood inference for permuted data with application to gene regulation. Journal of the American Statistical Association, 91(433):76–??, ???? 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Legler:1997:LVM

[LR97a] Julie M. Legler and Louise M. Ryan. Latent variable models for teratogenesis using multiple binary outcomes. Journal of the American Statistical Association, 92(437):13–??, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lewis:1997:MLR

[LR97b] Peter A. W. Lewis and Bonnie K. Ray. Modeling long-range dependence, nonlinearity, and periodic phenomena in sea surface temperatures using TSMARS. Journal of the American Statistical Association, 92(439):881–893, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/lewis_aw.htm.

Lewis:1997:EBF

[LR97c] Steven M. Lewis and Adrian E. Raftery. Estimating Bayes factors via posterior simulation with the Laplace-Metropolis estimator. Journal of the American Statistical Association, 92(438):648–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Le:1996:RBM

[LRM96] Nhu D. Le, Adrian E. Raftery, and R. Douglas Martin. Rohust Bayesian model selection for autoregressive processes with additive outliers. Journal of the American Statistical Association, 91(433): 123–131, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:1991:LSS

[LRR91] K. H. Li, T. E. Raghunathan, and D. B. Rubin. Large-sample significance levels from multiply imputed data using moment-based statistics and an F reference distribution. Journal of the American Statistical Association, 86(416):1065–1073, December 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lindley:1986:RFT

[LS86] Dennis V. Lindley and Nozer D. Singpurwalla. Reliability (and fault tree) analysis using expert opinions. Journal of the American Statistical Association, 81(393):87–90, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:1987:EIQ

[LS87] Roderick J. A. Little and Philip J. Smith. Editing and imputation for quantitative survey data. Journal of the American Statistical Association, 82(397):58–68, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lewis:1991:NMT

[LS91a] P. A. W. Lewis and J. G. Stevens. Nonlinear modeling of time series using multivariate adaptive regression splines (MARS). Journal of the American Statistical Association, 86(416):864–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lindley:1991:ENR

[LS91b] Dennis V. Lindley and Nozer D. Singpurwalla. On the evidence needed to reach agreed action between adversaries, with application to acceptance sampling. Journal of the American Statistical Association, 86(416):933–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1993:QIB

[LS93] Regina Y. Liu and Kesar Singh. A quality index based on data depth and multivariate rank tests. Journal of the American Statistical Association, 88(421):252–260, March 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:1995:ELN

[LS95] Lung Fei Lee and Jungsywan H. Sepanski. Estimation of linear and nonlinear errors-in-variables models using validation data. Journal of the American Statistical Association, 90(429):130–140, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1997:NLV

[LS97] Regina Y. Liu and Kesar Singh. Notions of limiting P values based on data depth and bootstrap. Journal of the American Statistical Association, 92(437):266–277, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/liu.htm.

Liu:1999:DRD

[LS99] Regina Liu and Kesar Singh. Discussion: Regression depth. Journal of the American Statistical Association, 94(446):407–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Limam:1988:STI

[LT88] Mohamed M. T. Limam and David R. Thomas. Simultaneous tolerance intervals for the linear regression model. Journal of the American Statistical Association, 83(403):801–804, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

LeBlanc:1994:APS

[LT94a] Michael LeBlanc and Robert Tibshirani. Adaptive principal surfaces. Journal of the American Statistical Association, 89(425): 53–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:1994:SPF

[LT94b] Ronald D. Lee and Shripad Tuljapurkar. Stochastic population forecasts for the united states: Beyond high, medium, and low. Journal of the American Statistical Association, 89(428):1175–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

LeBlanc:1996:CER

[LT96a] Michael LeBlanc and Robert Tibshirani. Combining estimates in regression and classification. Journal of the American Statistical Association, 91(436):1641–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:1996:CCD

[LT96b] Regina Y. Liu and Jen Tang. Control charts for dependent and independent measurements based on bootstrap methods. Journal of the American Statistical Association, 91(436):1694–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lambert:1997:NML

[LT97] Diane Lambert and Luke Tierney. Nonparametric maximum likelihood estimation from samples with irrelevant data and verification bias. Journal of the American Statistical Association, 92(439): 937–944, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/sept/lambert.htm.

Li:1998:UAI

[LT98] Hong Li and Ruey S. Tsay. A unified approach to identifying multivariate time series models. Journal of the American Statistical Association, 93(442):770–782, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

amstat.org/publications/jasa/abstracts_98/june/LI.HTM.

Laska:1992:THA

[LTM92] Eugene M. Laska, Dei-In Tang, and Morris J. Meisner. Testing hypotheses about an identified treatment when there are multiple endpoints. Journal of the American Statistical Association, 87 (419):825–831, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Litvak:1994:SPD

[LTP94] Eugene Litvak, Xin M. Tu, and Marcello Pagano. Screening for the presence of a disease by pooling sera samples. Journal of the American Statistical Association, 89(426):424–??, June 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:1996:QCF

[LTW96] Gang Li, Ram C. Tiwari, and Martin T. Wells. Quantile comparison functions in two-sample problems, with application to comparisons of diagnostic markers. Journal of the American Statistical Association, 91(434):689–698, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lucas:1993:WCN

[Luc93a] T. W. Lucas. When is conflict normal? Journal of the American Statistical Association, 88(424):1433–1437, December 1993. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luceno:1993:FAR

[Luc93b] Alberto Luceno. A fast algorithm for the repeated evaluation of the likelihood of a general linear process for long series. Journal of the American Statistical Association, 88(421):229–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Loh:1988:TSC

[LV88] Wei-Yin Loh and Nunta Vanichsetakul. Tree-structured classification via generalized discriminant analysis. with a comment by Leo Breiman and Jerome H. Friedman and with a reply by the authors. Journal of the American Statistical Association, 83(403): 715–728, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:1989:RIC

[LW89] D. Y. Lin and L. J. Wei. The robust inference for the Cox proportional hazards model. Journal of the American Statistical Association, 84(408):1074–1078, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:1990:ESE

[LW90] Kung-Yee Liang and Myron A. Waclawiw. Extension of the Stein estimating procedure through the use of estimating functions. Journal of the American Statistical Association, 85(410):435–440, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:1991:MCT

[LW91] Roderick J. A. Little and Mei-Miau Wu. Models for contingency tables with known margins when target and sampled populations differ. Journal of the American Statistical Association, 86(413): 87–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:1992:LRA

[LW92] J. S. Lin and L. J. Wei. Linear regression analysis for multivariate failure time observations. Journal of the American Statistical Association, 87(420):1091–1097, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luo:1997:HAS

[LW97] Zhen Luo and Grace Wahba. Hybrid adaptive splines. Journal of the American Statistical Association, 92(437):107–116, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/luo.htm.

Little:1993:ER

[LWG93] Roderick J. A. Little, Edward J. Wegman, and Donald Guthrie. Editors’ report for 1992. Journal of the American Statistical Association, 88(422):389–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lu:1994:ISF

[LWT94] Henry H. S. Lu, Martin T. Wells, and Ram C. Tiwari. Inference for shift functions in the two-sample problem with right-censored data: With applications. Journal of the American Statistical Association, 89(427):1017–1026, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lavine:1991:BIS

[LWW91] Michael Lavine, Larry Wasserman, and Robert L. Wolpert. Bayesian inference with specified prior marginals. Journal of the American Statistical Association, 86(416):964–971, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:1993:LRA

[LWY93] Eric W. Lee, L. J. Wei, and Z. Ying. Linear regression analysis for highly stratified failure time data. Journal of the American Statistical Association, 88(422):557–565, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:1993:CRI

[LY93] D. Y. Lin and Z. Ying. Cox regression with incomplete covariate measurements. Journal of the American Statistical Association, 88 (424):1341–1349, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:1999:GTS

[LYY99] D. Y. Lin, Q. Yao, and Zhiliang Ying. A general theory on stochastic curtailment for censored survival data. Journal of the American Statistical Association, 94(446):510–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:1989:CLR

[LZ89] Kung-Yee Liang and Scott L. Zeger. A class of logistic regression models for multivariate binary time series. Journal of the American Statistical Association, 84(406):447–451, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:1998:ACS

[LZ98] Sandra J. Lee and Marvin Zelen. Applications and case studies — scheduling periodic examinations for the early detection of disease: Applications to breast cancer. Journal of the American Statistical Association, 93(444):1271–1281, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/dec/lee.html.

Mack:1986:EWA

[Mac86] Michael E. Mack. Efficiencies of weighted averages in stationary autoregressive processes. Journal of the American Statistical Association, 81(395):730–735, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [Mac87].

Mack:1987:CEW

[Mac87] Michael E. Mack. Correction to: “Efficiencies of weighted averages in stationary autoregressive processes” [J. Amer. Statist. Assoc. 81 (1986), no. 395, 730–735, MR 87j:62174]. Journal of the American Statistical Association, 82(398):704, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Mac86].

Madansky:1959:FSL

[Mad59] A. Madansky. The fitting of straight lines when both variables are subject to error. Journal of the American Statistical Association, 54(??):173–205, ???? 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mallows:1978:PNA

[Mal78] Colin L. Mallows. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments: Comment. Journal of the American Statistical Association, 73 (363):520, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/pdfplus/2286591.pdf. See [SM78a].

Malley:1982:SCI

[Mal82] James D. Malley. Simultaneous confidence intervals for ratios of normal means. Journal of the American Statistical Association, 77(377):170–176, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Manski:1988:IBR

[Man88] Charles F. Manski. Identification of binary response models. Journal of the American Statistical Association, 83(403):729–738, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Manski:1990:UID

[Man90] Charles F. Manski. The use of intentions data to predict behavior: A best-case analysis. Journal of the American Statistical Association, 85(412):934–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mann:1994:MWE

[Man94] Nancy R. Mann. In memoriam: W. Edwards Deming 1900–1993. Journal of the American Statistical Association, 89(426):365–366, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marsaglia:1964:CMC

[Mar64] G. Marsaglia. Conditional means and covariances of normal variables with singular covariance matrix. Journal of the American Statistical Association, 59:1203–1204, 1964. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marsaglia:1965:RNV

[Mar65] George Marsaglia. Ratios of normal variables and ratios of sums of uniform variables. Journal of the American Statistical Association, 60(309):193–204, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283145.

Marquardt:1980:CSC

[Mar80] D. W. Marquardt. Comment on Smith and Campbell (1980). Journal of the American Statistical Association, 75(??):87–91, ???? 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Martinsek:1981:NVH

[Mar81] Adam T. Martinsek. A note on the variance and higher central moments of the stopping time of an SPRT. Journal of the American Statistical Association, 76(375):701–703, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marsaglia:1984:EAM

[Mar84] George Marsaglia. The exact-approximation method for generating random variables in a computer. Journal of the American Statistical Association, 79(385):218–221, ???? 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Maravall:1986:RAS

[Mar86] Agustín Maravall. Revisions in ARIMA signal extraction. Journal of the American Statistical Association, 81(395):736–740, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Martinsek:1988:NRO

[Mar88] Adam T. Martinsek. Negative regret, optional stopping, and the elimination of outliers. Journal of the American Statistical Association, 83(401):160–163, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Martin:1990:BIC

[Mar90] Michael A. Martin. On bootstrap iteration for coverage correction in confidence intervals. Journal of the American Statistical Association, 85(412):1105–1118, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marden:1992:UNO

[Mar92] John I. Marden. Use of nested orthogonal contrasts in analyzing rank data. Journal of the American Statistical Association, 87(418):307–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Massey:1951:KSG

[Mas51] Frank J. Massey, Jr. The Kolmogorov–Smirnov for goodness of fit. Journal of the American Statistical Association, 46(253):68– 78, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2280095.

Mathew:1983:LEI

[Mat83] Thomas Mathew. Linear estimation with an incorrect dispersion matrix in linear models with a common linear part. Journal of the American Statistical Association, 78(382):468–471, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mauricio:1995:EML

[Mau95] José Alberto Mauricio. Exact maximum likelihood estimation of stationary vector ARMA models. Journal of the American Statistical Association, 90(429):282–291, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mehrotra:1982:CIJ

[MB82] Kishan G. Mehrotra and Gouri K. Bhattacharyya. Confidence intervals with jointly type-ii censored samples from two exponential distributions. Journal of the American Statistical Association, 77 (378):441–446, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mitchell:1988:BVS

[MB88a] T. J. Mitchell and J. J. Beauchamp. Bayesian variable selection in linear regression. with comments by James Berger and C. L. Mallows and with a reply by the authors. Journal of the American Statistical Association, 83(404):1023–1036, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Monlezun:1988:GNS

[MB88b] Charles J. Monlezun and David C. Blouin. A general nested splitplot analysis of covariance. Journal of the American Statistical Association, 83(403):818–823, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Markatou:1998:WLE

[MBL98] Marianthi Markatou, Ayanendranath Basu, and Bruce G. Lindsay. Weighted likelihood equations with bootstrap root search. Journal of the American Statistical Association, 93(442):740–750, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/june/MARKATOU.HTM.

Magee:1991:CKS

[MBR91] Lonnie Magee, John B. Burbidge, and A. Leslie Robb. Computing kernel-smoothed conditional quantiles from many observations. Journal of the American Statistical Association, 86(415):673–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Moreno:1998:ILP

[MBR98] Elias Moreno, Francesco Bertolino, and Walter Racugno. An intrinsic limiting procedure for model selection and hypotheses testing. Journal of the American Statistical Association, 93(444): 1451–1460, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/dec/moreno.html.

McKnight:1984:TDT

[MC84] Barbara McKnight and John Crowley. Tests for differences in tumor incidence based on animal carcinogenesis experiments. Journal of the American Statistical Association, 79(387):639–648, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McCulloch:1982:SMD

[McC82] Charles E. McCulloch. Symmetric matrix derivatives with applications. Journal of the American Statistical Association, 77(379): 679–682, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McCulloch:1994:MLV

[McC94] Charles E. McCulloch. Maximum likelihood variance components estimation for binary data. Journal of the American Statistical Association, 89(425):330–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McCulloch:1997:MLA

[McC97] Charles E. McCulloch. Maximum likelihood algorithms for generalized linear mixed models. Journal of the American Statistical Association, 92(437):162–170, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/mcculloch.htm.

McCulloch:1998:CBI

[McC98] Robert McCulloch. Comment: Bayesian inference on network traffic using link count data. Journal of the American Statistical Association, 93(442):575–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McPherson:1942:MOP

[McP42] J. C. McPherson. Mathematical operations with punched cards. Journal of the American Statistical Association, 37(??):275–281, June 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). Reprinted in Annals of the History of Computing, 7(4), October 1985, 368–371.

McWilliams:1990:DFT

[McW90] Thomas P. McWilliams. A distribution-free test for symmetry based on a runs statistic. Journal of the American Statistical Association, 85(412):1130–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McCann:1996:PLI

[ME96] Melinda McCann and Don Edwards. A path length inequality for the multivariate-t distribution, with applications to multiple comparisons. Journal of the American Statistical Association, 91 (433):211–216, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mee:1990:CIP

[Mee90] Robert W. Mee. Confidence intervals for probabilities and tolerance regions based on a generalization of the Mann–Whitney statistic. Journal of the American Statistical Association, 85(411): 793–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McCaffrey:1992:ELE

[MEGN92] Daniel F. McCaffrey, Stephen Ellner, A. Ronald Gallant, and Douglas W. Nychka. Estimating the Lyapunov exponent of a chaotic system with nonparametric regression. Journal of the American Statistical Association, 87(419):682–695, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Meilijson:1994:CRC

[Mei94] Isaac Meilijson. Competing risks on coherent reliability systems: Estimation in the parametric case. Journal of the American Statistical Association, 89(428):1459–1464, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Merrill:1994:PMS

[Mer94a] Samuel Merrill, III. A probabilistic model for the spatial distribution of party support in multiparty electorates. Journal of the American Statistical Association, 89(428):1190–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Merrill:1994:RPM

[Mer94b] Samuel Merrill, III. Rejoinder: A probabilistic model for the spatial distribution of party support in multiparty electorates. Journal of the American Statistical Association, 89(428):1199–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Metropolis:1950:BRG

[Met50] N. Metropolis. Book review: Giant Brains or Machines That Think by Edmund C. Berkeley. Journal of the American Statistical Association, 45(252):573–574, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://links.jstor.org/sici?sici=0162-1459%
28195012%2945%3A252%3C573%3AGBOMTT%3E2.0.CO%3B2-U.

Meyer:1987:OIC

[Mey87] John S. Meyer. Outer and inner confidence intervals for finite population quantile intervals. Journal of the American Statistical Association, 82(397):201–204, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mezrich:1984:DRM

[MFS84] J. J. Mezrich, S. P. Frysinger, and R. Slivjanovski. Dynamic representation of multivariate time series data. Journal of the American Statistical Association, 79(385):34–40, July 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Morton:1984:SE

[MG84] M. J. Morton and H. L. Gray. The G-spectral estimator. Journal of the American Statistical Association, 79(387):692–701, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Monahan:1997:SRI

[MG97a] John Monahan and Alan Genz. Spherical-radial integration rules for Bayesian computation. Journal of the American Statistical Association, 92(438):664–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mukhopadhyay:1997:DPM

[MG97b] Saurabh Mukhopadhyay and Alan E. Gelfand. Dirichlet process mixed generalized linear models. Journal of the American Statistical Association, 92(438):633–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Markowski:1982:IBS

[MH82] Edward P. Markowski and Thomas P. Hettmansperger. Inference based on simple rank step score statistics for the location model. Journal of the American Statistical Association, 77(380):901–907, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Markatou:1990:RBI

[MH90] Marianthi Markatou and Thomas P. Hettmansperger. Robust bounded-influence test in linear models. Journal of the American Statistical Association, 85(409):187–190, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Markatou:1994:BIH

[MH94a] Marianthi Markatou and Xuming He. Bounded influence and high breakdown point testing procedures in linear models. Journal of the American Statistical Association, 89(426):543–549, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McArdle:1994:LML

[MH94b] John J. McArdle and Fumiaki Hamagami. Logit and multilevel logit modeling of college graduation for 1984–1985 freshman student-athletes. Journal of the American Statistical Association, 89(427):1107–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Miwa:1999:CAO

[MH99] Tetsuhisa Miwa and Tony Hayter. Combining the advantages of one-sided and two-sided test procedures for comparing several treatment effects. Journal of the American Statistical Association, 94(445):302–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Millar:1992:ESS

[Mil92] Russell B. Millar. Estimating the size-selectivity of fishing gear by conditioning on the total catch. Journal of the American Statistical Association, 87(420):962–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Minkin:1987:ODB

[Min87] Salomon Minkin. Optimal designs for binary data. Journal of the American Statistical Association, 82(400):1098–1103, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Minkin:1993:EDC

[Min93] Salomon Minkin. Experimental design for clonogenic assays in chemotherapy. Journal of the American Statistical Association, 88(422):410–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Minnotte:1998:AHO

[Min98] Michael C. Minnotte. Achieving higher-order convergence rates for density estimation with binned data. Journal of the American Statistical Association, 93(442):663–672, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/june/MINNOTTE.HTM.

Mislevy:1985:ELG

[Mis85] Robert J. Mislevy. Estimation of latent group effects. Journal of the American Statistical Association, 80(392):993–997, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mittal:1991:HSS

[Mit91] Yashaswini Mittal. Homogeneity of subpopulations and Simpson’s paradox. Journal of the American Statistical Association, 86(413): 167–172, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marasinghe:1982:TIA

[MJ82] Mervyn G. Marasinghe and Dallas E. Johnson. A test of incomplete additivity in the multiplicative interaction model. Journal of the American Statistical Association, 77(380):869–877, ????

1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mason:1990:CLE

[MKSB90] Robert L. Mason, Jerome P. Keating, Pranab K. Sen, and Neil W. Blaylock, Jr. Comparison of linear estimators using Pitman’s measure of closeness. Journal of the American Statistical Association, 85(410):579–581, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Molenberghs:1994:MMC

[ML94] Geert Molenberghs and Emmanuel Lesaffre. Marginal modeling of correlated ordinal data using a multivariate Plackett distribution. Journal of the American Statistical Association, 89(426):633–??, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mesenbrink:1994:COW

[MLMT94] Peter Mesenbrink, Jye-Chyi Lu, Richard McKenzie, and Javad Taheri. Characterization and optimization of a wave-soldering process. Journal of the American Statistical Association, 89(428): 1209–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:1992:PCG

[MLRR92] Keith E. Muller, Lisa M. La Vange, Sharon Landesman Ramey, and Craig T. Ramey. Power calculations for general linear multivariate models including repeated measures applications. Journal of the American Statistical Association, 87(420):1209–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McDermott:1993:SAT

[MM93] Michael P. McDermott and Govind S. Mudholkar. A simple approach to testing homogeneity of order-constrained means. Journal of the American Statistical Association, 88(424):1371–??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marden:1995:RTM

[MM95] John I. Marden and Maria Emerita T. Muyot. Rank tests for main and interaction effects in analysis of variance. Journal of the American Statistical Association, 90(432):1388–1398, ???? 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Meier:1993:NER

[MN93] Kristen Meier and Douglas Nychka. Nonparametric estimation of rate equations for nutrient uptake. Journal of the American Statistical Association, 88(422):602–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Miller:1999:IMP

[MNS99] Forrest R. Miller, James W. Neill, and Brian W. Sherfey. Implementation of a maximum power clustering criterion to select near replicates for regression lack of fit tests. Journal of the American Statistical Association, 94(446):610–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marshall:1985:FBD

[MO85] Albert W. Marshall and Ingram Olkin. A family of bivariate distributions generated by the bivariate Bernoulli distribution. Journal of the American Statistical Association, 80(390):332–338, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marshall:1988:FMD

[MO88] Albert W. Marshall and Ingram Olkin. Families of multivariate distributions. Journal of the American Statistical Association, 83 (403):834–841, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Maitra:1998:VAP

[MO98] Ranjan Maitra and Finbarr O’Sullivan. Variability assessment in positron emission tomography and related generalized deconvolution models. Journal of the American Statistical Association, 93 (444):1340–1355, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/dec/maitra.html.

Moffitt:1996:CIC

[Mof96] Robert A. Moffitt. Comment: Identification of causal effects using instrumental variables. Journal of the American Statistical Association, 91(434):444–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mojirsheibani:1999:CCD

[Moj99] Majid Mojirsheibani. Combining classifiers via discretization. Journal of the American Statistical Association, 94(446):600–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Moors:1981:ILI

[Moo81] J. J. A. Moors. Inadmissibility of linearly invariant estimators in truncated parameter spaces. Journal of the American Statistical Association, 76(376):910–915, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Moore:1998:PAS

[Moo98] David S. Moore. Presidential address — statistics among the liberal arts. Journal of the American Statistical Association, 93(444): 1253–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/moore.html.

Morris:1983:PEB

[Mor83] Carl N. Morris. Parametric empirical Bayes inference: theory and applications. with discussion. Journal of the American Statistical Association, 78(381):47–65, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Morgenthaler:1986:RCI

[Mor86] Stephan Morgenthaler. Robust confidence intervals for a location parameter: the configural approach. Journal of the American Statistical Association, 81(394):518–525, ???? 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Morris:1996:CEB

[Mor96] Carl N. Morris. Comment: Empirical Bayes methods for combining likelihoods. Journal of the American Statistical Association, 91(434):538–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Moynihan:1999:DDP

[Moy99] Daniel Patrick Moynihan. Data and dogma in public policy. Journal of the American Statistical Association, 94(446):359–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
moynihan.pdf.

Mehta:1983:NAP

[MP83] Cyrus R. Mehta and Nitin R. Patel. A network algorithm for performing Fisher’s exact test in r × c contingency tables. Journal of the American Statistical Association, 78(382):427–434, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:1995:ODC

[MP95] Peter Müller and Giovanni Parmigiani. Optimal design via curve fitting of Monte Carlo experiments. Journal of the American Statistical Association, 90(432):1322–1330, ???? 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mohanty:1996:EGP

[MP96] R. Mohanty and M. Pourahmadi. Estimation of the generalized prediction error variance of a multiple time series. Journal of the American Statistical Association, 91(433):294–299, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Morrell:1995:EUT

[MPCB95] Christopher H. Morrell, Jay D. Pearson, H. Ballentine Carter, and Larry J. Brant. Estimating unknown transition times using a piecewise nonlinear mixed-effects model in men with prostate cancer. Journal of the American Statistical Association, 90(429): 45–??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mehta:1985:CEC

[MPG85] Cyrus R. Mehta, Nitin R. Patel, and Robert Gray. Computing an exact confidence interval for the common odds ratio in several 2 × 2 contingency tables. Journal of the American Statistical Association, 80(392):969–973, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [MPG86].

Mehta:1986:CCE

[MPG86] C. R. Mehta, N. R. Patel, and R. Gray. Correction: “Computing an exact confidence interval for the common odds ratio in several 2 × 2 contingency tables” [J. Amer. Statist. Assoc. 80 (1985), no. 392, 969–973]. Journal of the American Statistical Association, 81 (396):1132, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MPG85].

Mehta:1988:ISE

[MPS88] Cyrus R. Mehta, Nitin R. Patel, and Pralay Senchaudhuri. Importance sampling for estimating exact probabilities in permutational inference. Journal of the American Statistical Association, 83(404):999–1005, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mehta:1998:CAE

[MPS98] Cyrus R. Mehta, Nitin R. Patel, and Pralay Senchaudhuri. Comment: Approximately exact inference for the common odds ratio in several 2 × 2 tables. Journal of the American Statistical Association, 93(444):1313–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Madigan:1994:MSA

[MR94] David Madigan and Adrian E. Raftery. Model selection and accounting for model uncertainty in graphical models using Occam’s Window. Journal of the American Statistical Association, 89(428): 1535–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Moller:1997:BPM

[MR97a] Peter Möller and Gary Rosner. A Bayesian population model with hierarchical mixture priors applied to blood count data. Journal of the American Statistical Association, 92(440):1279– 1292, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/MULLER.HTM.

Muller:1997:BPM

[MR97b] Peter Muller and Gary L. Rosner. A Bayesian population model with hierarchical mixture priors applied to blood count data. Journal of the American Statistical Association, 92(440):1279– 1292, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/MULLER.HTM.

Murphy:1997:MLE

[MRvdV97] S. A. Murphy, A. J. Rossini, and A. W. van der Vaart. Maximum likelihood estimation in the proportional odds model. Journal of the American Statistical Association, 92(439):968–976, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/murphy.htm.

Mukerjee:1987:CST

[MRW87] Hari Mukerjee, Tim Robertson, and F. T. Wright. Comparison of several treatments with a control using multiple contrasts. Journal of the American Statistical Association, 82(399):902–910, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Malec:1985:BIF

[MS85a] Donald Malec and J. Sedransk. Bayesian inference for finite population parameters in multistage cluster sampling. Journal of the American Statistical Association, 80(392):897–902, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Montagne:1985:RSE

[MS85b] Ernest R. Montagne and Nozer D. Singpurwalla. Robustness of sequential exponential life-testing procedures. Journal of the American Statistical Association, 80(391):715–179, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mathew:1988:OTF

[MS88a] Thomas Mathew and Bimal Kumar Sinha. Optimum tests for fixed effects and variance components in balanced models. Journal of the American Statistical Association, 83(401):133–135, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:1988:KPE

[MS88b] Hans-Georg Müller and Thomas Schmitt. Kernel and probit estimates in quantal bioassay. Journal of the American Statistical Association, 83(403):750–759, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:1991:EME

[MS91a] D. W. Müller and G. Sawitzki. Excess mass estimates and tests for multimodality. Journal of the American Statistical Association, 86(415):738–746, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See acknowledgement of priority [MS95].

Mulry:1991:TEPa

[MS91b] Mary H. Mulry and Bruce D. Spencer. Total error in PES estimates of population. Journal of the American Statistical Association, 86(416):839–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mulry:1991:TEPb

[MS91c] Mary H. Mulry and Bruce D. Spencer. Total error in PES estimates of population rejoinder. Journal of the American Statistical Association, 86(416):861–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mathew:1992:EOT

[MS92] Thomas Mathew and Bimal Kumar Sinha. Exact and optimum tests in unbalanced split-plot designs under mixed and random models. Journal of the American Statistical Association, 87(417): 192–200, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mukarjee:1994:FNE

[MS94] Hari Mukarjee and Steven Stern. Feasible nonparametric estimation of multiargument monotone functions. Journal of the American Statistical Association, 89(425):77–80, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:1995:APE

[MS95] D. W. Müller and G. Sawitzki. Acknowledgement of priority: “Excess mass estimates and tests for multimodality” [J. Amer. Statist. Assoc. 86 (1991), no. 415, 738–746, MR 92i:62091]. Journal of the American Statistical Association, 90(429):399, ???? 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MS91a].

Meng:1996:FFI

[MS96] Xiao-Li Meng and Stephen Schilling. Fitting full-information item factor models and an empirical investigation of bridge sampling. Journal of the American Statistical Association, 91(435):1254–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McCabe:1998:EPS

[MS98a] Brendan P. M. McCabe and Richard J. Smith. Errata: “The power of some tests for difference stationarity under local heteroscedastic integration” [J. Amer. Statist. Assoc. 93 (1998), no. 442, 751–761; 1 631 382]. Journal of the American Statistical Association, 93 (443):1252, 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MS98b].

McCabe:1998:PST

[MS98b] Brendan P. M. McCabe and Richard J. Smith. The power of some tests for difference stationarity under local heteroscedastic integration. Journal of the American Statistical Association, 93 (442):751–761, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/MCCABE.HTM. See er-

rata [MS98a].

McKean:1999:DRD

[MS99] Joseph W. McKean and Simon J. Sheather. Discussion: Regression depth. Journal of the American Statistical Association, 94(446): 411–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McKean:1990:RDR

[MSH90] Joseph W. McKean, Simon J. Sheather, and Thomas P. Hettmansperger. Regression diagnostics for rank-based methods. Journal of the American Statistical Association, 85(412):1018–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McKean:1993:UIR

[MSH93] Joseph W. McKean, Simon J. Sheather, and Thomas P. Hettmansperger. The use and interpretation of residuals based on robust estimation. Journal of the American Statistical Association, 88(424):1254–1263, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mudholkar:1996:GWD

[MSK96] Govind S. Mudholkar, Deo Kumar Srivastava, and Georgia D. Kollia. A generalization of the Weibull distribution with application to the analysis of survival data. Journal of the American Statistical Association, 91(436):1575–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Murry:1993:EAD

[MSL93] John P. Murry, Antonie Stam, and John L. Lastovicka. Evaluating an anti-drinking and driving advertising campaign with a sample survey and time series intervention analysis. Journal of the American Statistical Association, 88(421):50–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Malec:1997:SAI

[MSML97] Donald Malec, J. Sedransk, Christopher L. Moriarity, and Felicia B. LeClere. Small area inference for binary variables in the National Health Interview Survey. Journal of the American Statistical Association, 92(439):815–826, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/malec.htm.

Manski:1992:AEE

[MSMP92] Charles F. Manski, Gary D. Sandefur, Sara Mclanahan, and Daniel Powers. Alternative estimates of the effect of family structure during adolescence on high school graduation. Journal of the American Statistical Association, 87(417):25–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:1997:SSG

[MST97] Hans-Georg Müller, Ulrich Stadtmüller, and Farzaneh Tabnak. Spatial smoothing of geographically aggregated data, with application to the construction of incidence maps. Journal of the American Statistical Association, 92(437):61–71, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McDonald:1983:TPE

[MSW^{+}83] J. F. McDonald, M. A. Selby, C. S. Wong, L. D. Favro, and P. K. Kuo. Tests, point estimations, and confidence sets for a capturerecapture model. Journal of the American Statistical Association, 78(384):913–919, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mosteller:1949:UUB

[MT49] F. Mosteller and John W. Tukey. The uses and usefulness of binomial probability paper. Journal of the American Statistical Association, 44(??):174–212, ???? 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mudholkar:1981:GAD

[MT81] Govind S. Mudholkar and Madhusudan C. Trivedi. A Gaussian approximation to the distribution of the sample variance for nonnormal populations. Journal of the American Statistical Association, 76(374):479–485, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See corrigendum [MT82].

Mudholkar:1982:CGA

[MT82] Govind S. Mudholkar and Madhusudan C. Trivedi. Corrigendum: “A Gaussian approximation to the distribution of the sample variance for nonnormal populations” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 479–485, MR 82h:62034]. Journal of the American Statistical Association, 77(380):954–955, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MT81].

Marron:1995:VEC

[MT95a] J. S. Marron and A. B. Tsybakov. Visual error criteria for qualitative smoothing. Journal of the American Statistical Association, 90(430):499–507, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mukerjee:1995:ORI

[MT95b] Hari Mukerjee and Renjin Tu. Order-restricted inferences in linear regression. Journal of the American Statistical Association, 90 (430):717–728, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Meer:1998:CEP

[MT98] Peter Meer and David Tyler. Comment: Edge preserving smoothers for image processing. Journal of the American Statistical Association, 93(442):548–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mykland:1995:RMC

[MTY95] Per Mykland, Luke Tierney, and Bin Yu. Regeneration in Markov chain samplers. Journal of the American Statistical Association, 90(429):233–241, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Metropolis:1949:MCM

[MU49] N. Metropolis and S. Ulam. The Monte Carlo method. Journal of the American Statistical Association, 44(247):335–341, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://links.jstor.org/sici?sici=0162-
1459%28194909%2944%3A247%3C335%3ATMCM%3E2.0.CO%3B2-3.

Muirhead:1985:EPF

[Mui85] Robb J. Muirhead. Estimating a particular function of the multiple correlation coefficient. Journal of the American Statistical Association, 80(392):923–925, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mukerjee:1996:ESF

[Muk96] Hari Mukerjee. Estimation of survival functions under uniform stochastic ordering. Journal of the American Statistical Association, 91(436):1684–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:1987:WLR

[Mül87] Hans-Georg Müller. Weighted local regression and kernel methods for nonparametric curve fitting. Journal of the American Statistical Association, 82(397):231–238, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Munk:1996:EIT

[Mun96] Axel Munk. Equivalence and interval testing for Lehmann’s alternative. Journal of the American Statistical Association, 91(435): 1187–1196, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Murphy:1995:LRB

[Mur95] S. A. Murphy. Likelihood ratio-based confidence intervals in survival analysis. Journal of the American Statistical Association, 90 (432):1399–1405, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Meeden:1985:BAS

[MV85] Glen Meeden and Stephen Vardeman. Bayes and admissible set estimation. Journal of the American Statistical Association, 80 (390):465–471, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Malani:1988:CTT

[MV88] Hina Mehta Malani and John Van Ryzin. Comparison of two treatments in animal carcinogenicity experiments. Journal of the American Statistical Association, 83(404):1171–1177, ???? 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Meeden:1991:NBA

[MV91] Glen Meeden and Stephen Vardeman. A noninformative Bayesian approach to interval estimation in finite population sampling. Journal of the American Statistical Association, 86(416):972–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mack:1981:SRT

[MW81] Gregory A. Mack and Douglas A. Wolfe. k-sample rank tests for umbrella alternatives. Journal of the American Statistical Association, 76(373):175–181, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See corrigendum [MW82].

Mack:1982:CSR

[MW82] Gregory A. Mack and Douglas A. Wolfe. Corrigendum: “k-sample rank tests for umbrella alternatives” [J. Amer. Statist. Assoc. 76 (1981), no. 373, 175–181, MR 82d:62084]. Journal of the American Statistical Association, 77(380):954, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MW81].

Marais:1998:COV

[MW98a] M. Laurentius Marais and William E. Wecker. Correcting for omitted-variables and measurement-error bias in regression with an application to the effect of lead on IQ. Journal of the American Statistical Association, 93(442):494–504, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/june/MARAIS.HTM.

Marais:1998:RCO

[MW98b] M. Laurentius Marais and William E. Wecker. Rejoinder: Correcting for omitted-variables and measurement-error bias in regression with an application to the effect of lead on IQ. Journal of the American Statistical Association, 93(442):515–??, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Maronna:1991:BPS

[MY91] Ricardo A. Maronna and Víctor J. Yohai. The breakdown point of simultaneous general M estimates of regression and scale. Journal of the American Statistical Association, 86(415):699–703, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Maronna:1995:BSD

[MY95] Ricardo A. Maronna and Víctor J. Yohai. The behavior of the Stahel-Donoho robust multivariate estimator. Journal of the American Statistical Association, 90(429):330–341, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mosteller:1967:DSR

[MYZ67] Frederick Mosteller, Cleo Youtz, and Douglas Zahn. The distribution of sums of rounded percentages. Demography, 4(2):850–858, June 1967. CODEN ???? ISSN 0070-3370 (print), 1533-7790 (elec-

tronic). URL http://muse.jhu.edu/journals/dem/;   http:
//www.biomedsearch.com/nih/Distribution-Sums-Rounded-
Percentages/21318695.html; http://www.jstor.org/stable/

2060324. See [DF79] for further work.

Martin:1989:AMM

[MZ89] R. Douglas Martin and Ruben H. Zamar. Asymptotically min-max bias robust M-estimates of scale for positive random variables. Journal of the American Statistical Association, 84(406):494–501, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Maller:1994:TSF

[MZ94] R. A. Maller and S. Zhou. Testing for sufficient follow-up and outliers in survival data. Journal of the American Statistical Association, 89(428):1499–1506, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Magder:1996:SNE

[MZ96] Laurence S. Magder and Scott L. Zeger. A smooth nonparametric estimate of a mixing distribution using mixtures of Gaussians. Journal of the American Statistical Association, 91(435):1141– 1151, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mathew:1997:MUC

[MZ97] Thomas Mathew and Wenxing Zha. Multiple use confidence regions in multivariate calibration. Journal of the American Statistical Association, 92(439):1141–1150, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/mathew.htm.

Montgomery:1998:FUU

[MZTT98] Alan L. Montgomery, Victor Zarnowitz, Ruey S. Tsay, and George C. Tiao. Forecasting the U.S. unemployment rate. Journal of the American Statistical Association, 93(442):478–493, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/june/MONTGOME.HTM.

Naiman:1983:CST

[Nai83] Daniel Q. Naiman. Comparing Scheffé-type to constant-width confidence bounds in regression. Journal of the American Statistical Association, 78(384):906–912, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nair:1984:BSE

[Nai84] Vijayan N. Nair. On the behavior of some estimators from probability plots. Journal of the American Statistical Association, 79 (388):823–831, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Naiman:1987:MRS

[Nai87a] Daniel Q. Naiman. Minimax regret simultaneous confidence bands for multiple regression functions. Journal of the American Statistical Association, 82(399):894–901, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Naiman:1987:SCB

[Nai87b] Daniel Q. Naiman. Simultaneous confidence bounds in multiple regression using predictor variable constraints. Journal of the American Statistical Association, 82(397):214–219, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nair:1987:CST

[Nai87c] Vijayan N. Nair. Chi-squared-type tests for ordered alternatives in contingency tables. Journal of the American Statistical Association, 82(397):283–291, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Naus:1965:DSM

[Nau65] J. I. Naus. The distribution of the size of the maximum cluster of points on a line. Journal of the American Statistical Association, 60(??):532–538, ???? 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Naus:1982:ADS

[Nau82] Joseph I. Naus. Approximations for distributions of scan statistics. Journal of the American Statistical Association, 77(377):177–183, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ng:1999:AST

[NC99] Edmund T. M. Ng and Richard J. Cook. Adjusted score tests of homogeneity for Poisson processes. Journal of the American Statistical Association, 94(445):308–??, March 1999. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Newton:1996:BIS

[NCC96] Michael A. Newton, Claudia Czado, and Rick Chappell. Bayesian inference for semiparametric binary regression. Journal of the American Statistical Association, 91(433):142–153, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nusser:1996:STA

[NCDF96] S. M. Nusser, A. L. Carriquiry, K. W. Dodd, and W. A. Fuller. A semiparametric transformation approach to estimating usual daily intake distributions. Journal of the American Statistical Association, 91(436):1440–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Neftci:1982:SET

[Nef82] S. N. Neftci. Specification of economic time series models using akaike’s criterion. Journal of the American Statistical Association, 77(379):537–540, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Neuhaus:1998:EEO

[Neu98] John M. Neuhaus. Estimation efficiency with omitted covariates in generalized linear models. Journal of the American Statistical Association, 93(443):1124–1129, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/neuhaus.html.

Newcombe:1992:UNL

[NFL92] Howard B. Newcombe, Martha E. Fair, and Pierre Lalonde. The use of names for linking personal records. Journal of the American Statistical Association, 87(420):1193–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Newton:1994:BRM

[NG94] Michael A. Newton and Charles J. Geyer. Bootstrap recycling: A Monte Carlo alternative to the nested bootstrap. Journal of the American Statistical Association, 89(427):905–912, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Newton:1995:SME

[NGC^{+}95] Michael A. Newton, Peter Guttorp, Sandra Catlin, Renato Assuncao, and Janis L. Abkowitz. Stochastic modeling of early hematopoiesis. Journal of the American Statistical Association, 90(432):1146–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Normand:1997:SMP

[NGG97] Sharon-Lise T. Normand, Mark E. Glickman, and Constantine A. Gatsonis. Statistical methods for profiling providers of medical care: Issues and applications. Journal of the American Statistical Association, 92(439):803–??, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nychka:1995:NRA

[NGH^{+}95] Doug Nychka, Gerry Gray, Perry Haaland, David Martin, and Michael O’Connell. A nonparametric regression approach to syringe grading for quality improvement. Journal of the American Statistical Association, 90(432):1171–??, ???? 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Niu:1996:NAM

[Niu96] Xu-Feng Niu. Nonlinear additive models for environmental time series, with applications to ground-level ozone data analysis. Journal of the American Statistical Association, 91(435):1310–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Noorbaloochi:1983:UDB

[NM83a] Siamak Noorbaloochi and Glen Meeden. Unbiasedness as the dual of being Bayes. Journal of the American Statistical Association, 78(383):619–623, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nyblom:1983:CTP

[NM83b] Jukka Nyblom and Timo Mäkeläinen. Comparisons of tests for the presence of random walk coefficients in a simple linear model. Journal of the American Statistical Association, 78(384):856–864, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Neerchal:1998:LCR

[NM98] Nagaraj K. Neerchal and Jorge G. Morel. Large cluster results for TWO parametric multinomial extra variation models. Journal of the American Statistical Association, 93(443):1078– 1087, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/neerchal.html.

Naik-Nimbalkar:1995:FSE

[NNR95] U. V. Naik-Nimbalkar and M. B. Rajarshi. Filtering and smoothing via estimating functions. Journal of the American Statistical Association, 90(429):301–306, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Noether:1987:SSD

[Noe87] Gottfried E. Noether. Sample size determination for some common nonparametric tests. Journal of the American Statistical Association, 82(398):645–647, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Norwood:1990:SPP

[Nor90] Janet L. Norwood. Statistics and public policy: Reflections of a changing world. Journal of the American Statistical Association, 85(409):1–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nowak:1992:IMD

[Now92] Eugen Nowak. Identifiability in multivariate dynamic linear errorsin- variables models. Journal of the American Statistical Association, 87(419):714–723, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Newton:1983:MDP

[NP83] H. Joseph Newton and Marcello Pagano. A method for determining periods in time series. Journal of the American Statistical Association, 78(381):152–161, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ng:1995:URT

[NP95] Serena Ng and Pierre Perron. Unit root tests in ARMA models with data-dependent methods for the selection of the truncation lag. Journal of the American Statistical Association, 90(429):268– 281, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nagarsenker:1983:DCS

[NS83] B. N. Nagarsenker and J. Suniaga. Distributions of a class of statistics useful in multivariate analysis. Journal of the American Statistical Association, 78(382):472–475, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [NS86].

Nagarsenker:1986:CDC

[NS86] B. N. Nagarsenker and J. Suniaga. Correction: “Distributions of a class of statistics useful in multivariate analysis” [J. Amer. Statist. Assoc. 78 (1983), no. 382, 472–475, MR 84i:62070]. Journal of the American Statistical Association, 81(396):1132, ???? 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [NS83].

Nandram:1993:EBE

[NS93] B. Nandram and J. Sedransk. Empirical Bayes estimation of the finite population mean on the current occasion. Journal of the American Statistical Association, 88(423):994–1000, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nandram:1997:ORB

[NSS97] Balgobin Nandram, J. Sedransk, and Stephen J. Smith. Orderrestricted Bayesian estimation of the age composition of a population of Atlantic cod. Journal of the American Statistical Association, 92(437):33–??, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/nandram.htm.

Normand:1992:PUB

[NT92] Sharon-Lise Normand and David Tritchler. Parameter updating in a Bayes network. Journal of the American Statistical Association, 87(420):1109–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Niu:1995:MSO

[NT95] Xu Feng Niu and George C. Tiao. Modeling satellite ozone data. Journal of the American Statistical Association, 90(431):969–983, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Naus:1997:DSS

[NW97] Joseph I. Naus and Dan Wartenberg. A double-scan statistic for clusters of two types of events. Journal of the American Statistical Association, 92(439):1105–1113, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/naus.htm.

Nychka:1984:CVS

[NWGP84] Douglas Nychka, Grace Wahba, Stanley Goldfarb, and Thomas Pugh. Cross-validated spline methods for the estimation of threedimensional tumor size distributions from observations on twodimensional cross sections. Journal of the American Statistical Association, 79(388):832–846, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nyblom:1986:TDL

[Nyb86] Jukka Nyblom. Testing for deterministic linear trend in time series. Journal of the American Statistical Association, 81(394):545–549, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nyblom:1989:TCP

[Nyb89] Jukka Nyblom. Testing for the constancy of parameters over time. Journal of the American Statistical Association, 84(405):223–230, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nychka:1988:BCI

[Nyc88] Douglas Nychka. Bayesian confidence intervals for smoothing splines. Journal of the American Statistical Association, 83(404): 1134–1143, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nychka:1991:CRA

[Nyc91] Douglas Nychka. Choosing a range for the amount of smoothing in nonparametric regression. Journal of the American Statistical Association, 86(415):653–664, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oakes:1985:SCP

[Oak85] David Oakes. Self-calibrating priors do not exist. with comments by A. P. Dawid and Mark J. Schervish. Journal of the American Statistical Association, 80(390):339–342, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oakes:1989:BSM

[Oak89] David Oakes. Bivariate survival models induced by frailties. Journal of the American Statistical Association, 84(406):487–493, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

OHagan:1988:RPP

[OB88] Anthony O’Hagan and James O. Berger. Ranges of posterior probabilities for quasiunimodal priors with specified quantiles. Journal of the American Statistical Association, 83(402):503–508, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oh:1993:IMF

[OB93] Man-Suk Oh and James O. Berger. Integration of multimodal functions by Monte Carlo importance sampling. Journal of the American Statistical Association, 88(422):450–456, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oden:1991:AEM

[Ode91] Neal L. Oden. Allocation of effort in Monte Carlo simulation for power of permutation tests. Journal of the American Statistical Association, 86(416):1074–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oehlert:1993:RTS

[Oeh93] Gary W. Oehlert. Regional trends in sulfate wet deposition. Journal of the American Statistical Association, 88(422):390–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

OGorman:1982:DSA

[O’G82] T. W. O’Gorman. On the design of seasonal adjustment methods using linear programming techniques. Journal of the American Statistical Association, 77(380):739–742, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

OHagan:1987:BLE

[O’H87] A. O’Hagan. Bayes linear estimators for randomized response models. Journal of the American Statistical Association, 82(398): 580–585, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

OHagan:1990:OCL

[O’H90] A. O’Hagan. Outliers and credence for location parameter inference. Journal of the American Statistical Association, 85(409):

172–176, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Olive:1999:DRD

[OH99] David Olive and Douglas M. Hawkins. Discussion: Regression depth. Journal of the American Statistical Association, 94(446): 416–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

OlusegunGeorge:1996:TIT

[OK96] E. Olusegun George and Ralph L. Kodell. Tests of independence, treatment heterogeneity, and dose-related trend with exchangeable binary data. Journal of the American Statistical Association, 91 (436):1602–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ord:1997:EPC

[OKS97] J. K. Ord, A. B. Koehler, and R. D. Snyder. Estimation and prediction for a class of dynamic nonlinear statistical models. Journal of the American Statistical Association, 92(440):1621– 1629, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/ORD.HTM.

Oler:1985:NPC

[Ole85] Jacqueline Oler. Noncentrality parameters in chi-squared goodness-of-fit analyses with an application to log-linear procedures. Journal of the American Statistical Association, 80(389): 181–189, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Olson:1974:CRS

[Ols74] C. L. Olson. Comparative robustness of six tests in multivariate analysis of variance. Journal of the American Statistical Association, 69(348):894–908, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oman:1981:CBA

[Oma81] Samuel D. Oman. A confidence bound approach to choosing the biasing parameter in ridge regression. Journal of the American Statistical Association, 76(374):452–461, ???? 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oakes:1993:UCM

[OMF^{+}93] D. Oakes, A. J. Moss, J. L. Fleiss, J. T. Bigger, T. Therneau, S. W. Eberly, M. P. McDermott, A. Manatunga, E. Carleen, Benhorin, J., et al. Use of compliance measures in an analysis of the effect of diltiazem on mortality and reinfarction after myocardial infarction. Journal of the American Statistical Association, 88(421):44–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

ONeill:1984:GFT

[O’N84] Terence J. O’Neill. A goodness-of-fit test for one-sample life table data. Journal of the American Statistical Association, 79(385): 194–199, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oja:1989:BST

[ON89] Hannu Oja and Jukka Nyblom. Bivariate sign tests. Journal of the American Statistical Association, 84(405):249–259, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

ONeill:1994:BEE

[O’N94] Terence J. O’Neill. The bias of estimating equations with application to the error RATE of logistic discrimination. Journal of the American Statistical Association, 89(428):1492–1498, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Olshen:1997:CDA

[OO97] Adam B. Olshen and Finbarr O’Sullivan. Camouflaged deconvolution with application to blood curve modeling in FDG PET studies. Journal of the American Statistical Association, 92(440): 1293–1303, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/OLSHEN.HTM.

OSullivan:1996:BSI

[OP96] Finbarr O’Sullivan and Yudi Pawitan. Bandwidth selection for indirect density estimation based on corrupted histogram data. Journal of the American Statistical Association, 91(434):610–626, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Osius:1992:NGF

[OR92] Gerhard Osius and Dieter Rojek. Normal goodness-of-fit tests for multinomial models with large degrees of freedom. Journal of the American Statistical Association, 87(420):1145–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Olsson:1996:QER

[OR96] J. Olsson and H. Rootzen. Quantile estimation from repeated measurements. Journal of the American Statistical Association, 91(436):1560–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Opsomer:1998:FAB

[OR98] Jean D. Opsomer and David Ruppert. A fully automated bandwidth selection method for fitting additive models. Journal of the American Statistical Association, 93(442):605–619, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/june/OPSOMER.HTM.

Olkin:1987:SAD

[OS87] Ingram Olkin and Clifford H. Spiegelman. A semiparametric approach to density estimation. Journal of the American Statistical Association, 82(399):858–865, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

OSullivan:1990:IAT

[O’S90] Finbarr O’Sullivan. An iterative approach to two-dimensional Laplacian smoothing with application to image restoration. Journal of the American Statistical Association, 85(409):213–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

OSullivan:1991:DLS

[O’S91] Finbarr O’Sullivan. Discretized Laplacian smoothing by Fourier methods. Journal of the American Statistical Association, 86(415): 634–642, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

OSullivan:1993:MEM

[O’S93] Finbarr O’Sullivan. Mixture estimation from multichannel image data. Journal of the American Statistical Association, 88(421): 209–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Olkin:1995:CAE

[OV95] Ingram Olkin and Marlos Viana. Correlation analysis of extreme observations from a multivariate normal distribution. Journal of the American Statistical Association, 90(432):1373–1379, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Orban:1982:CDF

[OW82] John Orban and Douglas A. Wolfe. A class of distribution-free two-sample tests based on placements. Journal of the American Statistical Association, 77(379):666–672, ???? 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Owen:1994:CCL

[Owe94] Art B. Owen. Controlling correlations in Latin hypercube samples. Journal of the American Statistical Association, 89(428):1517–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Owen:1995:NLC

[Owe95] Art B. Owen. Nonparametric likelihood confidence bands for a distribution function. Journal of the American Statistical Association, 90(430):516–521, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

OSullivan:1986:ASR

[OYR86] Finbarr O’Sullivan, Brian S. Yandell, and William J. Raynor, Jr. Automatic smoothing of regression functions in generalized linear models. Journal of the American Statistical Association, 81(393): 96–103, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Padgett:1986:KTE

[Pad86] W. J. Padgett. A kernel-type estimator of a quantile function from right-censored data. Journal of the American Statistical Association, 81(393):215–222, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Page:1990:APC

[Pag90] Connie Faye Page. Allocation proportional to coefficients of variation when estimating the product of parameters. Journal of the American Statistical Association, 85(412):1134–1139, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Paik:1997:GEE

[Pai97] Myunghee Cho Paik. The generalized estimating equation approach when data are not missing completely at random. Journal of the American Statistical Association, 92(440):1320– 1329, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/PAIK.HTM.

Pan:1997:CSC

[Pan97] Guohua Pan. Confidence subset containing the unknown peaks of an umbrella ordering. Journal of the American Statistical Association, 92(437):307–314, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/pan.htm.

Paparoditis:1996:FDB

[Pap96] Efstathios Paparoditis. A frequency domain bootstrap-based method for checking the fit of a transfer function model. Journal of the American Statistical Association, 91(436):1535–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:1993:CVC

[Par93a] Byeong U. Park. A cross-validatory choice of smoothing parameter in adaptive location estimation. Journal of the American Statistical Association, 88(423):848–854, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Parmigiani:1993:OSA

[Par93b] Giovanni Parmigiani. On optimal screening ages. Journal of the American Statistical Association, 88(422):622–628, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:1996:FIO

[Par96] Sangun Park. Fisher information in order statistics. Journal of the American Statistical Association, 91(433):385–390, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pena:1987:ISS

[PB87] Daniel Peña and George E. P. Box. Identifying a simplifying structure in time series. Journal of the American Statistical Association, 82(399):836–843, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Picard:1992:CSV

[PB92] Richard Picard and Maurice Bryson. Calibrated seismic verification of the threshold test ban treaty. Journal of the American Statistical Association, 87(418):293–??, June 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:1994:MCD

[PB94] Taesung Park and Morton B. Brown. Models for categorical data with nonignorable nonresponse. Journal of the American Statistical Association, 89(425):44–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Paul:1998:ATW

[PB98] S. R. Paul and T. Banerjee. Analysis of two-way layout of count data involving multiple counts in each cell. Journal of the American Statistical Association, 93(444):1419–1429, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/dec/paul.html.

Picard:1984:CVR

[PC84] Richard R. Picard and R. Dennis Cook. Cross-validation of regression models. Journal of the American Statistical Association, 79(387):575–583, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peddada:1996:BCR

[PC96] Shyamal Das Peddada and Ted Chang. Bootstrap confidence region estimation of the motion of rigid bodies. Journal of the American Statistical Association, 91(433):231–241, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pepe:1997:MPC

[PC97] Margaret Sullivan Pepe and David Couper. Modeling partly conditional means with longitudinal data. Journal of the American Statistical Association, 92(439):991–998, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/sullivan.htm.

Peach:1961:BPR

[Pea61] Paul Peach. Bias in pseudo-random numbers. Journal of the American Statistical Association, 56(??):610–618, 1961. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pearson:1991:SSA

[Pea91] Robert W. Pearson. Social statistics and an American urban underclass: Improving the knowledge base. Journal of the American Statistical Association, 86(414):504–??, June 1991. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pennello:1997:RMC

[Pen97] Gene Pennello. The k-ratio multiple comparisons Bayes rule for the balanced two-way design. Journal of the American Statistical Association, 92(438):675–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pena:1998:SGF

[Peñ98] Edsel A. Peña. Smooth goodness-of-fit tests for the baseline hazard in Cox’s proportional hazards model. Journal of the American Statistical Association, 93(442):673–692, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/june/PENA.HTM.

Pepe:1991:IED

[Pep91] Margaret Sullivan Pepe. Inference for events with dependent risks in multiple endpoint studies. Journal of the American Statistical Association, 86(415):770–778, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peruggia:1997:VCD

[Per97] Mario Peruggia. On the variability of case-deletion importance sampling weights in the Bayesian linear model. Journal of the American Statistical Association, 92(437):199–207, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/peruggia.htm.

Peskun:1993:NCI

[Pes93] Peter H. Peskun. A new confidence interval method based on the normal approximation for the difference of two binomial probabilities. Journal of the American Statistical Association, 88(422): 656–661, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pettit:1992:BFO

[Pet92] Lawrence I. Pettit. Bayes factors for outlier models using the device of imaginary observations. Journal of the American Statistical Association, 87(418):541–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pepe:1991:NMD

[PF91] Margaret Sullivan Pepe and Thomas R. Fleming. A nonparametric method for dealing with mismeasured covariate data. Journal of the American Statistical Association, 86(413):108–113, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Proschan:1997:RTC

[PF97] Michael A. Proschan and Dean A. Follmann. A restricted test of circadian rhythm. Journal of the American Statistical Association, 92(438):716–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peck:1989:ACI

[PFV89] Roger Peck, Lloyd Fisher, and John Van Ness. Approximate confidence intervals for the number of clusters. Journal of the American Statistical Association, 84(405):184–191, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Phillips:1974:STS

[PH74] G. D. A. Phillips and A. C. Harvey. A simple test for serial correlation in regression analysis. Journal of the American Statistical Association, 69(348):935–939, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pagano:1981:AFE

[PH81] Marcello Pagano and Katherine Taylor Halvorsen. An algorithm for finding the exact significance levels of r × c contingency tables. Journal of the American Statistical Association, 76(376):931–934, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peixoto:1986:CAP

[PH86] Julio L. Peixoto and David A. Harville. Comparisons of alternative predictors under the balanced one-way random model. Journal of the American Statistical Association, 81(394):431–436, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Porter-Hudak:1990:ASF

[PH90a] Susan Porter-Hudak. An application of the seasonal fractionally difference model to the monetary aggregates. Journal of the American Statistical Association, 85(410):338–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Puma:1990:FSP

[PH90b] Michael J. Puma and David C. Hoaglin. Food stamp payment error rates: Can state-specific performance standards be developed? Journal of the American Statistical Association, 85(411):891–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pulkstenis:1998:MAB

[PHL98] Erik P. Pulkstenis, Thomas R. Ten Have, and J. Richard Landis. Model for the analysis of binary longitudinal pain data subject to informative dropout through remedication. Journal of the American Statistical Association, 93(442):438–450, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/june/PULKSTEN.HTM.

Presnell:1994:TMR

[PHS94] Brett Presnell, Myles Hollander, and Jayaram Sethuraman. Testing the minimal repair assumption in an imperfect repair model. Journal of the American Statistical Association, 89(425):289–297, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pickard:1987:IDM

[Pic87] David K. Pickard. Inference for discrete Markov fields: the simplest nontrivial case. Journal of the American Statistical Association, 82(397):90–96, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Piegorsch:1987:MRS

[Pie87] Walter W. Piegorsch. Model robustness for simultaneous confidence bands. Journal of the American Statistical Association, 82 (399):879–885, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Poli:1994:NNM

[PJ94] I. Poli and R. D. Jones. A neural net model for prediction. Journal of the American Statistical Association, 89(425):117–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peng:1996:BIM

[PJT96] Fengchun Peng, Robert A. Jacobs, and Martin A. Tanner. Bayesian inference in mixtures-of-experts and hierarchical mixtures-of-experts models with an application to speech recognition. Journal of the American Statistical Association, 91(435): 953–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:1990:CDD

[PM90] Byeong U. Park and J. S. Marron. Comparison of data-driven bandwidth selectors. Journal of the American Statistical Association, 85(409):66–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Percival:1997:ASC

[PM97] Donald B. Percival and Harold O. Mofjeld. Analysis of subtidal coastal sea level fluctuations using wavelets. Journal of the American Statistical Association, 92(439):868–880, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/percival.htm.

Portnoy:1999:DRD

[PM99] Stephen Portnoy and Ivan Mizera. Discussion: Regression depth. Journal of the American Statistical Association, 94(446):417–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Patel:1988:SAT

[PMF88] Kartik R. Patel, Govind S. Mudholkar, and J. L. Indrasiri Fernando. Student’s t approximations for three simple robust estimators. Journal of the American Statistical Association, 83(404): 1203–1210, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Presnell:1998:PML

[PML98] Brett Presnell, Scott P. Morrison, and Ramon C. Littell. Projected multivariate linear models for directional data. Journal of the American Statistical Association, 93(443):1068–1077, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/sept/presnell.html.

Potthoff:1993:CNB

[PMW93] Richard F. Potthoff, Kenneth G. Manton, and Max A. Woodbury. Correcting for nonavailability bias in surveys by weighting based on number of callbacks. Journal of the American Statistical Association, 88(424):1197–??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pawitan:1994:NSD

[PO94] Yudi Pawitan and Finbarr O’Sullivan. Nonparametric spectral density estimation using penalized Whittle likelihood. Journal of the American Statistical Association, 89(426):600–610, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Priebe:1997:SSS

[POH97] Carey E. Priebe, Tim Olson, and Dennis M. Healy, Jr. A spatial scan statistic for stochastic scan partitions. Journal of the American Statistical Association, 92(440):1476–1484, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/PRIEBE.HTM.

Polasek:1984:RDG

[Pol84] Wolfgang Polasek. Regression diagnostics for general linear regression models. Journal of the American Statistical Association, 79(386):336–340, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pollock:1991:MCR

[Pol91] Kenneth H. Pollock. Modeling capture, recapture, and removal statistics for estimation of demographic parameters for fish and wildlife populations: Past, present, and future. Journal of the American Statistical Association, 86(413):225–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Poole:1974:EDF

[Poo74] W. K. Poole. Estimation of the distribution function of a continuous type random variable through randomized response. Journal of the American Statistical Association, 69(348):1002–1005, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Portnoy:1997:CIM

[Por97] Stephen Portnoy. Comment: Interpolation methods for adapting to sparse design in nonparametric regression. Journal of the American Statistical Association, 92(438):473–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Potthoff:1987:SGM

[Pot87] Richard F. Potthoff. Some generalizations of the Mitofsky-Waksberg technique for random digit dialing. Journal of the American Statistical Association, 82(398):409–418, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Phillips:1988:AEO

[PP88] P. C. B. Phillips and Joon Y. Park. Asymptotic equivalence of ordinary least squares and generalized least squares in regressions with integrated regressors. Journal of the American Statistical Association, 83(401):111–115, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peters:1990:MSR

[PR90a] Dawn Peters and Ronald H. Randles. A multivariate signed-rank test for the one-sample location problem. Journal of the American Statistical Association, 85(410):552–557, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Prasad:1990:EMS

[PR90b] N. G. N. Prasad and J. N. K. Rao. The estimation of the mean squared error of small-area estimators. Journal of the American Statistical Association, 85(409):163–171, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pukelsheim:1993:EDM

[PR93] Friedrich Pukelsheim and James L. Rosenberger. Experimental designs for model discrimination. Journal of the American Statistical Association, 88(422):642–649, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Politis:1994:SB

[PR94] Dimitris N. Politis and Joseph P. Romano. The stationary bootstrap. Journal of the American Statistical Association, 89(428): 1303–1313, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pratt:1981:CLL

[Pra81] John W. Pratt. Concavity of the log likelihood. Journal of the American Statistical Association, 76(373):103–106, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See corrigendum [Pra82].

Pratt:1982:CCL

[Pra82] John W. Pratt. Corrigendum: “Concavity of the log likelihood” [J. Amer. Statist. Assoc. 76 (1981), no. 373, 103–106, MR 83h:62098]. Journal of the American Statistical Association, 77(380):954, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Pra81].

Press:1996:DNA

[Pre96] S. James Press. The directional neighborhoods approach to contextual classification of images from noisy data. Journal of the American Statistical Association, 91(435):1091–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pringle:1974:SRE

[Pri74] R. M. Pringle. Some results on the estimation of variance components by MINQUE. Journal of the American Statistical Association, 69(348):987–989, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Priebe:1994:AM

[Pri94] Carey E. Priebe. Adaptive mixtures. Journal of the American Statistical Association, 89(427):796–806, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Priebe:1996:NAU

[Pri96] Carey E. Priebe. Nonhomogeneity analysis using borrowed strength. Journal of the American Statistical Association, 91(436): 1497–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pruitt:1993:IBS

[Pru93] Ronald C. Pruitt. Identifiability of bivariate survival curves from censored data. Journal of the American Statistical Association, 88(422):573–579, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Parr:1982:JSS

[PS82] William C. Parr and William R. Schucany. Jackknifing L-statistics with smooth weight functions. Journal of the American Statistical Association, 77(379):629–638, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pierce:1986:RGL

[PS86] Donald A. Pierce and Daniel W. Schafer. Residuals in generalized linear models. Journal of the American Statistical Association, 81 (396):977–986, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:1994:ESE

[PS94a] B. U. Park and L. Simar. Efficient semiparametric estimation in a stochastic frontier model. Journal of the American Statistical Association, 89(427):929–936, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Phillips:1994:BFH

[PS94b] D. B. Phillips and A. F. M. Smith. Bayesian faces via hierarchical template modeling. Journal of the American Statistical Association, 89(428):1151–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peruggia:1996:BAT

[PS96] Mario Peruggia and Thomas Santner. Bayesian analysis of time evolution of earthquakes. Journal of the American Statistical Association, 91(435):1209–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pitt:1999:FSA

[PS99] Michael K. Pitt and Neil Shephard. Filtering via simulation: Auxiliary particle filters. Journal of the American Statistical Association, 94(446):590–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pericchi:1993:PCR

[PSS93] L. R. Pericchi, B. Sansó, and A. F. M. Smith. Posterior cumulant relationships in Bayesian inference involving the exponential family. Journal of the American Statistical Association, 88(424): 1419–1426, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pesaran:1999:PMG

[PSS99] M. Hashem Pesaran, Yongcheol Shin, and Ron P. Smith. Pooled mean group estimation of dynamic heterogeneous panels. Journal of the American Statistical Association, 94(446):621–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pierce:1992:EVP

[PSVS92] Donald A. Pierce, Daniel O. Stram, Michael Vaeth, and Daniel W. Schafer. The errors-in-variables problem: Considerations provided by radiation dose-response analyses of the A-bomb survivor data. Journal of the American Statistical Association, 87(418):351–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Patil:1982:DCM

[PT82] G. P. Patil and C. Taillie. Diversity as a concept and its measurement. with comments by I. J. Good and George Sugihara and a rejoinder by the authors. Journal of the American Statistical Association, 77(379):548–567, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pagano:1983:OPD

[PT83] Marcello Pagano and David Tritchler. On obtaining permutation distributions in polynomial time. Journal of the American Statistical Association, 78(382):435–440, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pudney:1982:ELV

[Pud82] S. E. Pudney. Estimating latent variable systems when specification is uncertain: generalized component analysis and the eliminant method. Journal of the American Statistical Association, 77 (380):883–889, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pan:1996:CGU

[PW96] Guohua Pan and Douglas A. Wolfe. Comparing groups with umbrella orderings. Journal of the American Statistical Association, 91(433):311–317, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Potthoff:1992:ESS

[PWM92] Richard F. Potthoff, Max A. Woodbury, and Kenneth G. Manton. “equivalent sample size” and “equivalent degrees of freedom” refinements for inference using survey weights under super- population models. Journal of the American Statistical Association, 87(418):383–396, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pena:1999:FPO

[PY99] Daniel Peña and Victor Yohai. A fast procedure for outlier diagnostics in large regression problems. Journal of the American Statistical Association, 94(446):434–??, June 1999. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.amstat.org/publications/jasa/pena.pdf.

Qu:1998:ACS

[QH98] Yinsheng Qu and Alula Hadgu. Applications and case studies — A model for evaluating sensitivity and specificity for correlated diagnostic tests in efficacy studies with an imperfect reference test. Journal of the American Statistical Association, 93(443): 920–928, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/qu.html.

Quintana:1998:AOD

[QN98] Fernando A. Quintana and Michael A. Newton. Assessing the order of dependence for partially exchangeable binary data. Journal of the American Statistical Association, 93(441):194–202, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/QUINTANA.HTM.

Qian:1996:IAP

[QTC96] W. Qian, D. M. Titterington, and J. N. Chapman. An image analysis problem in electron microscopy. Journal of the American Statistical Association, 91(435):944–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Quintana:1998:NBA

[Qui98a] Fernando A. Quintana. Nonparametric Bayesian analysis for assessing homogeneity in k × l contingency tables with fixed right margin totals. Journal of the American Statistical Association, 93 (443):1140–1149, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_98/sept/quintana.html.

Quintos:1998:FMV

[Qui98b] Carmela E. Quintos. Fully modified vector autoregressive inference in partially nonstationary models. Journal of the American Statistical Association, 93(442):783–795, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/june/QUINTOS.HTM.

Ruppert:1989:TSH

[RA89] David Ruppert and Brian Aldershof. Transformations to symmetry and homoscedasticity. Journal of the American Statistical Association, 84(406):437–446, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raftery:1987:IPG

[Raf87] Adrian E. Raftery. Inference and prediction for a general order statistic model with unknown population size. Journal of the American Statistical Association, 82(400):1163–1168, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raghunathan:1993:QEB

[Rag93] T. E. Raghunathan. A quasi-empirical Bayes method for small area estimation. Journal of the American Statistical Association, 88(424):1444–??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Randles:1984:TAR

[Ran84] Ronald H. Randles. On tests applied to residuals. Journal of the American Statistical Association, 79(386):349–354, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Randles:1989:DFM

[Ran89] Ronald H. Randles. A distribution-free multivariate sign test based on interdirections. Journal of the American Statistical Association, 84(408):1045–1050, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rao:1996:RVE

[Rao96] J. N. K. Rao. Rejoinder: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91 (434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raveh:1989:NAL

[Rav89] Adi Raveh. A nonmetric approach to linear discriminant analysis. Journal of the American Statistical Association, 84(405):176–183, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raz:1990:TNE

[Raz90] Jonathan Raz. Testing for no effect when estimating a smooth function by nonparametric regression: A randomization approach.

Journal of the American Statistical Association, 85(409):132–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rousseeuw:1990:RRA

[RB90] Peter J. Rousseeuw and Gilbert W. Bassett, Jr. The remedian: A robust averaging method for large data sets. Journal of the American Statistical Association, 85(409):97–104, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rodgers:1993:ESR

[RBD93] Willard L. Rodgers, Charles Brown, and Greg J. Duncan. Errors in survey reports of earnings, hours worked, and hourly wages. Journal of the American Statistical Association, 88(424):1208–??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robertson:1974:IPE

[RC74] T. Robertson and J. D. Cryer. An iterative procedure for estimating the mode. Journal of the American Statistical Association, 69 (348):1012–1016, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ruppert:1980:TLS

[RC80] David Ruppert and Raymond J. Carroll. Trimmed least squares estimation in the linear model. Journal of the American Statistical Association, 75(372):828–838, ???? 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See corrigendum [RC82].

Royall:1981:FPL

[RC81] Richard M. Royall and William G. Cumberland. The finitepopulation linear regression estimator and estimators of its variance—an empirical study. Journal of the American Statistical Association, 76(376):924–930, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ruppert:1982:CTL

[RC82] David Ruppert and Raymond J. Carroll. Corrigendum: “Trimmed least squares estimation in the linear model” [J. Amer. Statist. Assoc. 75 (1980), no. 372, 828–838, MR 82d:62073]. Journal of the American Statistical Association, 77(380):954, ???? 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [RC80].

Rousseeuw:1993:AMA

[RC93] Peter J. Rousseeuw and Christophe Croux. Alternatives to the median absolute deviation. Journal of the American Statistical Association, 88(424):1273–1283, December 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rathburn:1994:STS

[RC94] Stephen L. Rathburn and Noel Cressie. A space-time survival point process for a longleaf pine forest in Southern Georgia. Journal of the American Statistical Association, 89(428):1164–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ridder:1994:TEE

[RCH94] G. Ridder, J. S. Cramer, and P. Hopstaken. Down to ten: Estimating the effect of a red card in soccer. Journal of the American Statistical Association, 89(427):1124–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Roeder:1996:SMA

[RCL96] Kathryn Roeder, Raymond J. Carroll, and Bruce G. Lindsay. A semiparametric mixture approach to case-control studies with errors in covariables. Journal of the American Statistical Association, 91(434):722–732, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Read:1984:SSC

[Rea84] Timothy R. C. Read. Small-sample comparisons for the power divergence goodness-of-fit statistics. Journal of the American Statistical Association, 79(388):929–935, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reinsel:1982:MRM

[Rei82] Greg Reinsel. Multivariate repeated-measurement or growth curve models with multivariate random-effects covariance structure. Journal of the American Statistical Association, 77(377): 190–195, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reinsel:1984:EPM

[Rei84] Gregory Reinsel. Estimation and prediction in a multivariate random effects generalized linear model. Journal of the American Statistical Association, 79(386):406–414, ???? 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reid:1996:CEB

[Rei96] N. Reid. Comment: Empirical Bayes methods for combining likelihoods. Journal of the American Statistical Association, 91(434): 538–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Relles:1972:SAG

[Rel72] D. A. Relles. A simple algorithm for generating binomial random variables when N is large. Journal of the American Statistical Association, 67(339):612–613, ???? 1972. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ronchetti:1997:RLM

[RFB97] Elvezio Ronchetti, Christopher Field, and Wade Blanchard. Robust linear model selection by cross-validation. Journal of the American Statistical Association, 92(439):1017–1023, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/ronchetti.htm.

Ritov:1993:ACT

[RG93] Ya’acov Ritov and Zvi Gilula. Analysis of contingency tables by correspondence models subject to order constraints. Journal of the American Statistical Association, 88(424):1380–1387, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robins:1994:ADR

[RG94] James M. Robins and Sander Greenland. Adjusting for differential rates of prophylaxis therapy for PCP in high- versus lowdose AZT treatment arms in an Aids randomized trial. Journal of the American Statistical Association, 89(427):737–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raghunathan:1995:SQS

[RG95] Trivellore E. Raghunathan and James E. Grizzle. A split questionnaire survey design. Journal of the American Statistical Association, 90(429):54–??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robins:1996:CIC

[RG96] James M. Robins and Sander Greenland. Comment: Identification of causal effects using instrumental variables. Journal of the American Statistical Association, 91(434):444–??, ???? 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raftery:1995:IDP

[RGZ95a] Adrian E. Raftery, Geof H. Givens, and Judith E. Zeh. Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90(430): 402–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raftery:1995:RID

[RGZ95b] Adrian E. Raftery, Geof H. Givens, and Judith E. Zeh. Rejoinder: Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90(430):427–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robert:1996:MLE

[RH96] Christian P. Robert and J. T. Gene Hwang. Maximum likelihood estimation under order restrictions by the prior feedback method. Journal of the American Statistical Association, 91(433):167–172, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rousseeuw:1999:RD

[RH99] Peter J. Rousseeuw and Mia Hubert. Regression depth. Journal of the American Statistical Association, 94(446):388–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
rousseeuw.pdf.

Robert:1993:PCR

[RHS93] Christian P. Robert, J. T. Gene Hwang, and William E. Strawderman. Is Pitman closeness a reasonable criterion? with discussion and a reply by the authors. Journal of the American Statistical Association, 88(421):57–76, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rousseeuw:1999:RRD

[RHV99] Peter J. Rousseeuw, Mia Hubert, and Stefan Van Aelst. Rejoinder: Regression depth. Journal of the American Statistical Association, 94(446):419–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Richmond:1982:GMC

[Ric82] J. Richmond. A general method for constructing simultaneous confidence intervals. Journal of the American Statistical Association, 77(378):455–460, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Richardson:1997:BIE

[Ric97] Alice M. Richardson. Bounded influence estimation in the mixed linear model. Journal of the American Statistical Association, 92 (437):154–161, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/march/richardson.htm.

Rigby:1997:BDB

[Rig97] Robert A. Rigby. Bayesian discrimination between TWO multivariate normal populations with equal covariance matrices. Journal of the American Statistical Association, 92(439):1151– 1154, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/sept/rigby.htm.

Rinaman:1983:DFR

[Rin83a] William C. Rinaman, Jr. On distribution-free rank tests for twoway layouts. Journal of the American Statistical Association, 78 (383):655–659, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See corrections [Rin85].

Ringland:1983:RMC

[Rin83b] James T. Ringland. Robust multiple comparisons. Journal of the American Statistical Association, 78(381):145–151, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rinaman:1985:CDF

[Rin85] William C. Rinaman, Jr. Corrections: “On distribution-free rank tests for two-way layouts” [J. Amer. Statist. Assoc. 78 (1983), no. 383, 655–659, MR 85a:62120]. Journal of the American Statistical Association, 80(392):1084, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Rin83a].

Rivest:1986:BCH

[Riv86] Louis-Paul Rivest. Bartlett’s, Cochran’s, and Hartley’s tests on variances are liberal when the underlying distribution is longtailed. Journal of the American Statistical Association, 81(393): 124–128, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rivest:1998:SLM

[Riv98] Louis-Paul Rivest. Some linear models for estimating the motion of rigid bodies with applications to geometric quality assurance. Journal of the American Statistical Association, 93 (442):632–642, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/RIVEST.HTM.

Rizzo:1992:CCE

[Riz92] Louis Rizzo. Conditionally consistent estimators using only probabilities of selection in complex sample surveys. Journal of the American Statistical Association, 87(420):1166–1173, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:1990:STS

[RK90] Paul R. Rosenbaum and Abba M. Krieger. Sensitivity of twosample permutation inferences in observational studies. Journal of the American Statistical Association, 85(410):493–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ronning:1996:EEO

[RK96] Gerd Ronning and Martin Kukuk. Efficient estimation of ordered probit models. Journal of the American Statistical Association, 91 (435):1120–1129, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raftery:1997:BMA

[RMH97] Adrian E. Raftery, David Madigan, and Jennifer A. Hoeting. Bayesian model averaging for linear regression models. Journal of the American Statistical Association, 92(437):179–191, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/raftery.htm.

Rubin:1990:MRT

[RMS90] Gail Rubin, Charles E. McCulloch, and Michael A. Shapiro. Multinomial runs tests to detect clustering in constrained free recall. Journal of the American Statistical Association, 85(410):315–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Renssen:1997:AEC

[RN97] Robbert H. Renssen and Nico J. Nieuwenbroek. Aligning estimates for common variables in two or more sample surveys. Journal of the American Statistical Association, 92(437):368–374, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/march/renssen.htm.

Robinson:1987:CRE

[Rob87] J. Robinson. Conditioning ratio estimates under simple random sampling. Journal of the American Statistical Association, 82 (399):826–831, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comment [Kot89].

Robinson:1994:ETN

[Rob94] P. M. Robinson. Efficient tests of nonstationary hypotheses. Journal of the American Statistical Association, 89(428):1420–1437, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rochon:1992:ACS

[Roc92] James Rochon. ARMA covariance structures with time heteroscedasticity for repeated measures experiments. Journal of the American Statistical Association, 87(419):777–??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rochon:1995:SIA

[Roc95] James Rochon. Supplementing the intent-to-treat analysis: Accounting for covariates observed postrandomization in clinical trials. Journal of the American Statistical Association, 90(429):292– ??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Roeder:1990:DEC

[Roe90] Kathryn Roeder. Density estimation with confidence sets exemplified by superclusters and voids in the galaxies. Journal of the American Statistical Association, 85(411):617–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Roeder:1994:GTD

[Roe94] Kathryn Roeder. A graphical technique for determining the number of components in a mixture of normals. Journal of the American Statistical Association, 89(426):487–495, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rogers:1988:BET

[Rog88] Alan J. Rogers. The Bahadur efficiency of tests of some joint hypotheses. Journal of the American Statistical Association, 83 (401):257–264, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rojo:1996:TCP

[Roj96] Javier Rojo. On tail categorization of probability laws. Journal of the American Statistical Association, 91(433):378–384, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rolle:1994:BNI

[Rol94] Jean-Daniel Rolle. Best nonnegative invariant partially orthogonal quadratic estimation in normal regression. Journal of the American Statistical Association, 89(428):1378–1385, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Romano:1988:BRS

[Rom88] Joseph P. Romano. A bootstrap revival of some nonparametric distance tests. Journal of the American Statistical Association, 83 (403):698–708, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Romano:1990:BRT

[Rom90] Joseph P. Romano. On the behavior of randomization tests without a group invariance assumption. Journal of the American Statistical Association, 85(411):686–692, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosander:1942:UIT

[Ros42] A. C. Rosander. The use of inversions as a tsst of random order. Journal of the American Statistical Association, 37(??):352– 358, 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rose:1984:OSS

[Ros84a] John S. Rose. Optimal sequential selection based on relative ranks with renewable call options. Journal of the American Statistical Association, 79(386):430–435, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:1984:CPT

[Ros84b] Paul R. Rosenbaum. Conditional permutation tests and the propensity score in observational studies. Journal of the American Statistical Association, 79(387):565–574, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosner:1989:MMC

[Ros89] Bernard Rosner. Multivariate methods for clustered binary data with more than one level of nesting. Journal of the American Statistical Association, 84(406):373–380, ???? 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:1993:HLP

[Ros93a] Paul R. Rosenbaum. Hodges-Lehmann point estimates of treatment effect in observational studies. Journal of the American Statistical Association, 88(424):1250–1253, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:1993:SLT

[Ros93b] Paul R. Rosenbaum. Sampling the leaves of a tree with equal probabilities. Journal of the American Statistical Association, 88 (424):1455–1457, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:1995:QNS

[Ros95a] Paul R. Rosenbaum. Quantiles in nonrandom samples and observational studies. Journal of the American Statistical Association, 90(432):1424–1431, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenthal:1995:CMC

[Ros95b] Jeffrey Rosenthal. Correction: “Minorization conditions and convergence rates for Markov chain Monte Carlo”. Journal of the American Statistical Association, 90(431):1136, ???? 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Ros95c].

Rosenthal:1995:MCC

[Ros95c] Jeffrey S. Rosenthal. Minorization conditions and convergence rates for Markov chain Monte Carlo. Journal of the American Statistical Association, 90(430):558–566, ???? 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [Ros95b].

Rosenbaum:1996:CIC

[Ros96] Paul R. Rosenbaum. Comment: Identification of causal effects using instrumental variables. Journal of the American Statistical Association, 91(434):444–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Roth:1983:RTT

[Rot83] Arthur J. Roth. Robust trend tests derived and simulated: analogs of the Welch and Brown-Forsythe tests. Journal of the American Statistical Association, 78(384):972–980, ???? 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rousseeuw:1984:LMS

[Rou84] Peter J. Rousseeuw. Least median of squares regression. Journal of the American Statistical Association, 79(388):871–880, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Royall:1986:PAR

[Roy86] Richard M. Royall. The prediction approach to robust variance estimation in two-stage cluster sampling. Journal of the American Statistical Association, 81(393):119–123, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Runger:1996:MCM

[RP96] George C. Runger and Sharad S. Prabhu. A Markov chain model for the multivariate exponentially weighted moving averages control chart. Journal of the American Statistical Association, 91 (436):1701–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robins:1995:SEM

[RR95] James M. Robins and Andrea Rotnitzky. Semiparametric efficiency in multivariate regression models with missing data. Journal of the American Statistical Association, 90(429):122–129, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rotnitzky:1998:SRR

[RRS98] Andrea Rotnitzky, James M. Robins, and Daniel O. Scharfstein. Semiparametric regression for repeated outcomes with non-ignorable non-response. Journal of the American Statistical Association, 93(444):1321–1339, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/dec/rotnitzky.html.

Robins:1994:ERC

[RRZ94] James M. Robins, Andrea Rotnitzky, and Lue Ping Zhao. Estimation of regression coefficients when some regressors are not always observed. Journal of the American Statistical Association, 89(427):846–866, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robins:1995:ASR

[RRZ95] James M. Robins, Andrea Rotnitzky, and Lue Ping Zhao. Analysis of semiparametric regression models for repeated outcomes in the presence of missing data. Journal of the American Statistical Association, 90(429):106–121, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rukhin:1982:EQE

[RS82] Andrew L. Rukhin and William E. Strawderman. Estimating a quantile of an exponential distribution. Journal of the American Statistical Association, 77(377):159–162, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rubin:1986:MII

[RS86] Donald B. Rubin and Nathaniel Schenker. Multiple imputation for interval estimation from simple random samples with ignorable nonresponse. Journal of the American Statistical Association, 81 (394):366–374, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reneau:1990:ESC

[RS90] D. M. Reneau and F. J. Samaniego. Estimating the survival curve when new is better than used of a specified age. Journal of the American Statistical Association, 85(409):123–131, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rojo:1993:ESC

[RS93] Javier Rojo and Francisco J. Samaniego. On estimating a survival curve subject to a uniform stochastic ordering constraint. Journal of the American Statistical Association, 88(422):566–572, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rayens:1994:DPG

[RS94a] William S. Rayens and Cidambi Srinivasan. Dependence properties of generalized Liouville distributions on the simplex. Journal of the American Statistical Association, 89(428):1465–1470, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ronchetti:1994:RVM

[RS94b] Elvezio Ronchetti and Robert G. Staudte. A robust version of Mallow’s C_{p}. Journal of the American Statistical Association, 89 (426):550–559, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rao:1996:BHS

[RS96] J. N. K. Rao and J. Shao. On balanced half-sample variance estimation in stratified random sampling. Journal of the American Statistical Association, 91(433):343–348, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ramsay:1998:CSS

[RS98] James O. Ramsay and Bernard W. Silverman. Comment: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93 (443):988–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rubin:1995:HDK

[RSV95] Donald B. Rubin, Hal S. Stern, and Vasja Vehovar. Handling “don’t know” survey responses: The case of the Slovenian plebiscite. Journal of the American Statistical Association, 90 (431):822–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ruppert:1995:EBS

[RSW95] D. Ruppert, S. J. Sheather, and M. P. Wand. An effective bandwidth selector for local least squares regression. Journal of the American Statistical Association, 90(432):1257–1270, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rao:1993:GLS

[RSY93] J. N. K. Rao, Brajendra C. Sutradhar, and Kim Yue. Generalized least squares F test in regression analysis with two-stage cluster samples. Journal of the American Statistical Association, 88(424): 1388–??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ritter:1992:FGS

[RT92] Christian Ritter and Martin A. Tanner. Facilitating the Gibbs sampler: The Gibbs stopper and the Griddy-Gibbs sampler. Journal of the American Statistical Association, 87(419):861–??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Romano:1996:IAU

[RT96a] Joseph P. Romano and Lori A. Thombs. Inference for autocorrelations under weak assumptions. Journal of the American Statistical Association, 91(434):590–600, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rossini:1996:SPO

[RT96b] A. J. Rossini and A. A. Tsiatis. A semiparametric proportional odds regression model for the analysis of current status data. Journal of the American Statistical Association, 91(434):713–721, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raz:1992:FDE

[RTF92] Jonathan Raz, Bruce Turetsky, and George Fein. Frequency domain estimation of the parameters of human brain electrical dipoles. Journal of the American Statistical Association, 87(417): 69–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rubin:1996:MIA

[Rub96a] Donald B. Rubin. Multiple imputation after 18+ years. Journal of the American Statistical Association, 91(434):473–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rubin:1996:RVE

[Rub96b] Donald B. Rubin. Rejoinder: On variance estimation with imputed survey data. Journal of the American Statistical Association, 91 (434):499–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rue:1995:NLF

[Rue95] Håvard Rue. New loss functions in Bayesian imaging. Journal of the American Statistical Association, 90(431):900–908, ????

1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rukhin:1984:ACP

[Ruk84] Andrew L. Rukhin. Adaptive classification procedures. Journal of the American Statistical Association, 79(386):415–422, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rukhin:1986:IEL

[Ruk86] Andrew L. Rukhin. Improved estimation in lognormal models. Journal of the American Statistical Association, 81(396):1046– 1049, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rukhin:1987:HMB

[Ruk87] Andrew L. Rukhin. How much better are better estimators of a normal variance. Journal of the American Statistical Association, 82(399):925–928, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rukhin:1993:BET

[Ruk93] Andrew L. Rukhin. Bahadur efficiency of tests of separate hypotheses and adaptive test statistics. Journal of the American Statistical Association, 88(421):161–165, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ruppert:1985:BIR

[Rup85] David Ruppert. On the bounded-influence regression estimator of Krasker and Welsch. Journal of the American Statistical Association, 80(389):205–208, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ruppert:1997:EBB

[Rup97] David Ruppert. Empirical-bias bandwidths for local polynomial nonparametric regression and density estimation. Journal of the American Statistical Association, 92(439):1049–1062, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/ruppert.htm.

Rousseeuw:1990:RUM

[Rv90a] Peter J. Rousseeuw and Bert C. van Zomeren. Rejoinder: Unmasking multivariate outliers and leverage points. Journal of the American Statistical Association, 85(411):648–??, September 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rousseeuw:1990:UMO

[Rv90b] Peter J. Rousseeuw and Bert C. van Zomeren. Unmasking multivariate outliers and leverage points. Journal of the American Statistical Association, 85(411):633–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rust:1990:SDC

[RV90c] Philip F. Rust and Eberhard O. Voit. Statistical densities, cumulatives, quantiles, and power obtained by S-system differential equations. Journal of the American Statistical Association, 85(410):572–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rukhin:1998:ECM

[RV98] Andrew L. Rukhin and Mark G. Vangel. Estimation of a common mean and weighted means statistics. Journal of the American Statistical Association, 93(441):303–308, March 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/RUKHIN.HTM.

Rao:1985:ISS

[RW85] J. N. K. Rao and C.-F. J. Wu. Inference from stratified samples: second-order analysis of three methods for nonlinear statistics. Journal of the American Statistical Association, 80(391):620–630, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rao:1988:RIC

[RW88] J. N. K. Rao and C.-F. J. Wu. Resampling inference with complex survey data. Journal of the American Statistical Association, 83 (401):231–241, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rocke:1996:IOM

[RW96] David M. Rocke and David L. Woodruff. Identification of outliers in multivariate data. Journal of the American Statistical Association, 91(435):1047–1061, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Roeder:1997:PBD

[RW97] Kathryn Roeder and Larry Wasserman. Practical Bayesian density estimation using mixtures of normals. Journal of the American Statistical Association, 92(439):894–??, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rogan:1998:CCO

[RW98] Walter J. Rogan and James H. Ware. Comment: Correcting for omitted-variables and measurement-error bias in regression with an application to the effect of lead on IQ. Journal of the American Statistical Association, 93(442):513–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ryu:1993:EMO

[Ryu93] Keunkwan Ryu. An extension of Marshall and Olkin’s bivariate exponential distribution. Journal of the American Statistical Association, 88(424):1458–1465, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ryu:1994:GDA

[Ryu94] Keunkwan Ryu. Group duration analysis of the proportional hazard model: Minimum chi-squared estimators and specification tests. Journal of the American Statistical Association, 89(428): 1386–??, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raftery:1998:EBW

[RZ98] Adrian E. Raftery and Judith E. Zeh. Estimating bowhead whale, Balaena mysticetus, population size and rate of increase from the 1993 census. Journal of the American Statistical Association, 93 (442):451–463, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/RAFTERY.HTM.

Sadooghi-Alvandi:1986:CSS

[SA86] Mohammad Sadooghi-Alvandi. The choice of subsample size in two-stage sampling. Journal of the American Statistical Association, 81(394):555–558, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sobel:1992:MHF

[SA92] Michael E. Sobel and Gerhard Arminger. Modeling household fertility decisions: A nonlinear simultaneous probit model. Journal of the American Statistical Association, 87(417):38–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shively:1990:FED

[SAK90] Thomas S. Shively, Craig F. Ansley, and Robert Kohn. Fast evaluation of the distribution of the Durbin-Watson and other invariant test statistics in time series regression. Journal of the American Statistical Association, 85(411):676–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Samuels:1981:MSR

[Sam81] Stephen M. Samuels. Minimax stopping rules when the underlying distribution is uniform. Journal of the American Statistical Association, 76(373):188–197, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Samarov:1985:BIR

[Sam85a] Alexander M. Samarov. Bounded-influence regression via local minimax mean squared error. Journal of the American Statistical Association, 80(392):1032–1040, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Samuels:1985:BCP

[Sam85b] Stephen M. Samuels. A best-choice problem with linear travel cost. Journal of the American Statistical Association, 80(390):461–464, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sampson:1990:MCM

[Sam90] Michael Sampson. A Markov chain model for unskilled workers and the highly mobile. Journal of the American Statistical Association, 85(409):177–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Samarov:1993:ERS

[Sam93a] Alexander M. Samarov. Exploring regression structure using nonparametric functional estimation. Journal of the American Statistical Association, 88(423):836–847, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Samuels:1993:SPR

[Sam93b] Myra L. Samuels. Simpson’s paradox and related phenomena. Journal of the American Statistical Association, 88(421):81–88, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Santos:1998:ACS

[San98] Robert L. Santos. Applications and case studies — comment: Not asked and not answered: Multiple imputation for multiple surveys. Journal of the American Statistical Association, 93(443):865–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sarkar:1987:CBE

[Sar87] Sanat K. Sarkar. A continuous bivariate exponential distribution. Journal of the American Statistical Association, 82(398):667–675, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sarkar:1989:QIO

[Sar89] Sanat K. Sarkar. Quasi-independence in ordinal triangular contingency tables. Journal of the American Statistical Association, 84(406):592–597, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sarndal:1996:EES

[Sär96] Carl-Erik Särndal. Efficient estimators with simple variance in unequal probability sampling. Journal of the American Statistical Association, 91(435):1289–1300, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sasieni:1993:MWP

[Sas93] Peter Sasieni. Maximum weighted partial likelihood estimators for the Cox model. Journal of the American Statistical Association, 88(421):144–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sastry:1997:NFM

[Sas97] Narayan Sastry. A nested frailty model for survival data, with an application to the study of child survival in northeast Brazil.

Journal of the American Statistical Association, 92(438):426–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Satten:1995:ULB

[Sat95] Glen A. Satten. Upper and lower bound distributions that give simultaneous confidence intervals for quantiles. Journal of the American Statistical Association, 90(430):747–752, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sihota:1981:ASC

[SB81] S. S. Sihota and K. S. Banerjee. On the algebraic structures in the construction of confounding plans in mixed factorial designs on the lines of White and Hultquist. Journal of the American Statistical Association, 76(376):996–1001, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shapiro:1987:ACS

[SB87] Alexander Shapiro and Michael W. Browne. Analysis of covariance structures under elliptical distributions. Journal of the American Statistical Association, 82(400):1092–1097, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stasny:1990:DRC

[SB90] Elizabeth A. Stasny and Harold R. Bauer. Deviations from randomness in children’s early speech. Journal of the American Statistical Association, 85(409):46–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stefanski:1995:IVE

[SB95] L. A. Stefanski and J. S. Buzas. Instrumental variable estimation in binary regression measurement error models. Journal of the American Statistical Association, 90(430):541–550, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sengupta:1997:ROC

[SB97] Debasis Sengupta and Pochiraju Bhimasankaram. On the roles of observations in collinearity in the linear model. Journal of the American Statistical Association, 92(439):1024–1032, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/sengupta.htm.

Sheiner:1997:ANC

[SBD97a] Lewis B. Sheiner, Stuart L. Beal, and Adrian Dunne. Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials. Journal of the American Statistical Association, 92(440):1235–1244, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/SHEINER.HTM.

Sheiner:1997:RAN

[SBD97b] Lewis B. Sheiner, Stuart L. Beal, and Adrian Dunne. Rejoinder: Analysis of nonrandomly censored ordered categorical longitudinal data from analgesic trials. Journal of the American Statistical Association, 92(440):1253–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sain:1994:CVM

[SBS94] Stephan R. Sain, Keith A. Baggerly, and David W. Scott. Crossvalidation of multivariate densities. Journal of the American Statistical Association, 89(427):807–817, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smith:1980:CSR

[SC80] G. Smith and F. Campbell. A critique of some ridge regression methods. Journal of the American Statistical Association, 75(??): 74–81, ???? 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Storer:1985:DCR

[SC85] Barry E. Storer and John Crowley. A diagnostic for Cox regression and general conditional likelihoods. Journal of the American Statistical Association, 80(389):139–147, ???? 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schlax:1992:FDD

[SC92a] Michael G. Schlax and Dudley B. Chelton. Frequency domain diagnostics for linear smoothers. Journal of the American Statistical Association, 87(420):1070–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

StLaurent:1992:LSN

[SC92b] Roy T. St. Laurent and R. Dennis Cook. Leverage and superleverage in nonlinear regression. Journal of the American Statistical Association, 87(420):985–990, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sherman:1994:NEM

[SC94] Michael Sherman and Edward Carlstein. Nonparametric estimation of the moments of a general statistic computed from spatial data. Journal of the American Statistical Association, 89(426): 496–500, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stefanski:1995:SEM

[SC95] L. A. Stefanski and J. R. Cook. Simulation-extrapolation: The measurement error jackknife. Journal of the American Statistical Association, 90(432):1247–1256, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Samuel-Cahn:1996:OSR

[SC96a] Ester Samuel-Cahn. Optimal stopping with random horizon with application to the full-information best-choice problem with random freeze. Journal of the American Statistical Association, 91 (433):357–364, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sherman:1996:RH

[SC96b] Michael Sherman and Edward Carlstein. Replicate histograms. Journal of the American Statistical Association, 91(434):566–576, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sakkar:1997:SMM

[SC97a] Sanat K. Sakkar and Chung-Kuei Chang. The Simes method for multiple hypothesis testing with positively dependent test statistics. Journal of the American Statistical Association, 92(440): 1601–1608, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/SARKAR.HTM.

Sarkar:1997:SMM

[SC97b] Sanat K. Sarkar and Cung-Kuei Chang. The Simes method for multiple hypothesis testing with positively dependent test statistics. Journal of the American Statistical Association, 92(440): 1601–1608, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/december/SARKAR.HTM.

SullivanPepe:1997:MPC

[SC97c] Margaret Sullivan Pepe and David Couper. Modeling partly conditional means with longitudinal data. Journal of the American Statistical Association, 92(439):991–??, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sherman:1999:CPS

[SC99] Michael Sherman and Edward Carlstein. Comment: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):102–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shao:1998:BRR

[SCC98] Jun Shao, Ying Chen, and Yinzhong Chen. Balanced repeated replication for stratified multistage survey data under imputation. Journal of the American Statistical Association, 93 (442):819–831, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/SHAO.HTM.

Symons:1993:BNC

[SCCF93] Michael J. Symons, Chen Chu-Chih, and Michael R. Flynn. Bayesian nonparametrics for compliance to exposure standards. Journal of the American Statistical Association, 88(424):1237–??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schlicht:1981:SAP

[Sch81] Ekkehart Schlicht. A seasonal adjustment principle and a seasonal adjustment method derived from this principle. Journal of the American Statistical Association, 76(374):374–378, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See corrigendum [Sch82].

Schlicht:1982:CSA

[Sch82] Ekkehart Schlicht. Corrigendum: “A seasonal adjustment principle and a seasonal adjustment method derived from this principle” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 374–378, MR 82h:62162]. Journal of the American Statistical Association, 77 (380):954, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Sch81].

Schervish:1983:UOI

[Sch83a] Mark J. Schervish. User-oriented inference. Journal of the American Statistical Association, 78(383):611–615, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schwager:1983:RPS

[Sch83b] Steven J. Schwager. Run probabilities in sequences of Markovdependent trials. Journal of the American Statistical Association, 78(381):168–180, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schmoyer:1984:SCM

[Sch84a] Richard L. Schmoyer. Sigmoidally constrained maximum likelihood estimation in quantal bioassay. Journal of the American Statistical Association, 79(386):448–453, ???? 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schuster:1984:CPL

[Sch84b] Eugene F. Schuster. Classification of probability laws by tail behavior. Journal of the American Statistical Association, 79(388): 936–939, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schenker:1985:QAB

[Sch85] Nathaniel Schenker. Qualms about bootstrap confidence intervals. Journal of the American Statistical Association, 80(390):360–361, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schafer:1986:CIM

[Sch86a] Daniel W. Schafer. Combining information on measurement error in the errors-in-variables model. Journal of the American Statistical Association, 81(393):181–185, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schilling:1986:MTS

[Sch86b] Mark F. Schilling. Multivariate two-sample tests based on nearest neighbors. Journal of the American Statistical Association, 81 (395):799–806, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schwarz:1987:LAV

[Sch87a] Gideon Schwarz. Least-absolute-value prediction lines in closed form. Journal of the American Statistical Association, 82(400): 1150–1152, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schwarz:1987:MPL

[Sch87b] Gideon Schwarz. A minimax property of linear regression. Journal of the American Statistical Association, 82(397):220, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schmoyer:1988:LIN

[Sch88] Richard L. Schmoyer. Linear interpolation with a nonparametric accelerated failure-time model. Journal of the American Statistical Association, 83(402):441–449, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schirm:1991:ECU

[Sch91a] Allen L. Schirm. The effects of census undercount adjustment on Congressional apportionment. Journal of the American Statistical Association, 86(414):526–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schott:1991:TSP

[Sch91b] James R. Schott. A test for a specific principal component of a correlation matrix. Journal of the American Statistical Association, 86(415):747–751, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schmoyer:1994:PTC

[Sch94a] Richard L. Schmoyer. Permutation tests for correlation in regression errors. Journal of the American Statistical Association, 89 (428):1507–1516, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schott:1994:DDS

[Sch94b] James R. Schott. Determining the dimensionality in sliced inverse regression. Journal of the American Statistical Association, 89 (425):141–148, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schucany:1995:ABC

[Sch95a] William R. Schucany. Adaptive bandwidth choice for kernel regression. Journal of the American Statistical Association, 90(430): 535–540, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schweder:1995:CID

[Sch95b] Tore Schweder. Comment: Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90(430):420–??, ???? 1995. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schmid:1996:EAF

[Sch96] Christopher H. Schmid. An EM algorithm fitting first-order conditional autoregressive models to longitudinal data. Journal of the American Statistical Association, 91(435):1322–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schmertmann:1999:EMT

[Sch99] Carl P. Schmertmann. Estimating multistate transition hazards from last-move data. Journal of the American Statistical Association, 94(445):53–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Samuels:1991:IBR

[SCM91] Myra L. Samuels, George Casella, and George P. McCabe. Interpreting blocks and random factors. with comments and a rejoinder by the authors. Journal of the American Statistical Association, 86(415):798–821, ???? 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scott:1985:FPT

[Sco85] David W. Scott. Frequency polygons: theory and application. Journal of the American Statistical Association, 80(390):348–354, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Strawderman:1996:PSS

[SCW96] Robert L. Strawderman, George Casella, and Martin T. Wells. Practical small-sample asymptotics for regression problems. Journal of the American Statistical Association, 91(434):643–654, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [SCW97].

Strawderman:1997:CPS

[SCW97] Robert L. Strawderman, George Casella, and Martin T. Wells. Correction to “Practical Small-Sample Asymptotics for Regression Problems”. Journal of the American Statistical Association, 92 (440):1657–??, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [SCW96].

Said:1985:HTM

[SD85] Said E. Said and David A. Dickey. Hypothesis testing in arima(p, 1, q) models. Journal of the American Statistical Association, 80(390):369–374, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sinha:1997:SBA

[SD97] Debajyoti Sinha and Dipak K. Dey. Semiparametric Bayesian analysis of survival data. Journal of the American Statistical Association, 92(439):1195–1212, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/sinha.htm.

Satten:1998:IBI

[SDW98] Glen A. Satten, Somnath Datta, and John M. Williamson. Inference based on imputed failure times for the proportional hazards model with interval censored data. Journal of the American Statistical Association, 93(441):318–327, March 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/SATTEN.HTM.

Segal:1992:TSM

[Seg92] Mark Robert Segal. Tree-structured methods for longitudinal data. Journal of the American Statistical Association, 87(418):

407–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Soofi:1995:IDA

[SEH95] Ehsan S. Soofi, Nader Ebrahimi, and Mohamed Habibullah. Information distinguishability with application to analysis of failure data. Journal of the American Statistical Association, 90(430): 657–668, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sen:1974:UER

[Sen74] P. Kumar Sen. On unbiased estimation for randomized response models. Journal of the American Statistical Association, 69(348): 997–1001, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sen:1986:GCP

[Sen86] Pranab Kumar Sen. The Gini coefficient and poverty indexes: some reconciliations. Journal of the American Statistical Association, 81(396):1050–1057, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Severini:1991:CML

[Sev91] Thomas A. Severini. A comparison of the maximum likelihood estimator and the posterior mean in the single-parameter case. Journal of the American Statistical Association, 86(416):997–1000, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Severini:1994:ABI

[Sev94] Thomas A. Severini. Approximately Bayesian inference. Journal of the American Statistical Association, 89(425):242–249, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Severini:1995:ICI

[Sev95] Thomas A. Severini. Information and conditional inference. Journal of the American Statistical Association, 90(432):1341–1346, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Severini:1996:MSR

[Sev96] Thomas A. Severini. Measures of the sensitivity of regression estimates to the choice of estimator. Journal of the American Statistical Association, 91(436):1651–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stein:1997:COE

[SF97] Michael L. Stein and Dongping Fang. Comment: Ozone exposure and population density in Harris County, Texas. Journal of the American Statistical Association, 92(438):405–??, June 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spruill:1982:ECC

[SG82] Nancy L. Spruill and Joseph L. Gastwirth. On the estimation of the correlation coefficient from grouped data. Journal of the American Statistical Association, 77(379):614–620, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sebenius:1983:DBI

[SG83] James K. Sebenius and John Geanakoplos. Don’t bet on it: contingent agreements with asymmetric information. Journal of the American Statistical Association, 78(382):424–426, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sampson:1992:NEN

[SG92] Paul D. Sampson and Peter Guttorp. Nonparametric estimation of nonstationary spatial covariance structure. Journal of the American Statistical Association, 87(417):108–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sison:1995:CSC

[SG95a] Christina P. Sison and Joseph Glaz. Correction: “Simultaneous confidence intervals and sample size determination for multinomial proportions” [J. Amer. Statist. Assoc. 90 (1995), no. 429, 366–369, MR 96e:62039]. Journal of the American Statistical Association, 90(432):1494, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [SG95b].

Sison:1995:SCI

[SG95b] Cristina P. Sison and Joseph Glaz. Simultaneous confidence intervals and sample size determination for multinomial proportions.

Journal of the American Statistical Association, 90(429):366–369, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [SG95a].

Seifert:1996:FSV

[SG96a] Burkhardt Seifert and Theo Gasser. Finite-sample variance of local polynomials: Analysis and solutions. Journal of the American Statistical Association, 91(433):267–275, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sinclair:1996:PEE

[SG96b] Michael D. Sinclair and Joseph Gastwirth. On procedures for evaluating the effectiveness of reinterview survey methods: Applications to labor force data. Journal of the American Statistical Association, 91(435):961–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Swanson:1997:IRF

[SG97] Norman R. Swanson and Clive W. J. Granger. Impulse response functions based on a causal approach to residual orthogonalization in vector autoregressions. Journal of the American Statistical Association, 92(437):357–367, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/swanson.htm.

Schwartz:1995:TMC

[SGC95] Pamela F. Schwartz, Chris Gennings, and Vernon M. Chinchilli. Threshold models for combination data from reproductive and developmental experiments. Journal of the American Statistical Association, 90(431):862–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stute:1998:BAM

[SGP98] W. Stute, W. González Manteiga, and M. Presedo Quindimil. Bootstrap approximations in model checks for regression. Journal of the American Statistical Association, 93(441):141–149, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/STUTE.HTM.

Sallas:1981:BLR

[SH81] William M. Sallas and David A. Harville. Best linear recursive estimation for mixed linear models. Journal of the American Statistical Association, 76(376):860–869, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sleeper:1990:RSC

[SH90] Lynn A. Sleeper and David P. Harrington. Regression splines in the Cox model with application to covariate effects in liver disease. Journal of the American Statistical Association, 85(412):941–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sanchez:1992:OSR

[SH92] Susan M. Sanchez and Julia L. Higle. Observational studies of rare events: A subset selection approach. Journal of the American Statistical Association, 87(419):878–883, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shafer:1982:LPD

[Sha82] Glenn Shafer. Lindley’s paradox. with discussion and with a reply by the author. Journal of the American Statistical Association, 77(378):325–351, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shapiro:1983:SAO

[Sha83] C. P. Shapiro. Sequential allocation and optional stopping in Bayesian simultaneous estimation. Journal of the American Statistical Association, 78(382):396–402, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shapiro:1985:ASE

[Sha85] C. Page Shapiro. Allocation schemes for estimating the product of positive parameters. Journal of the American Statistical Association, 80(390):449–454, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shapiro:1986:ATO

[Sha86] Alexander Shapiro. Asymptotic theory of overparameterized structural models. Journal of the American Statistical Association, 81 (393):142–149, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shao:1989:ECA

[Sha89a] Jun Shao. The efficiency and consistency of approximations to the jackknife variance estimators. Journal of the American Statistical Association, 84(405):114–119, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shao:1989:MCA

[Sha89b] Jun Shao. Monte Carlo approximations in Bayesian decision theory. Journal of the American Statistical Association, 84(407):727– 732, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shao:1993:LMS

[Sha93] Jun Shao. Linear model selection by cross-validation. Journal of the American Statistical Association, 88(422):486–494, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shao:1996:BMS

[Sha96] Jun Shao. Bootstrap model selection. Journal of the American Statistical Association, 91(434):655–665, ???? 1996. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spiegelhalter:1994:EEP

[SHBF94] David J. Spiegelhalter, Nomi L. Harris, Kate Bull, and Rodney C. G. Franklin. Empirical evaluation of prior beliefs about frequencies: Methodology and a case study in congenital heart disease. Journal of the American Statistical Association, 89(426): 435–??, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shephard:1993:MLE

[She93] Neil Shephard. Maximum likelihood estimation of regression models with stochastic trend components. Journal of the American Statistical Association, 88(422):590–??, June 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sherman:1997:SMR

[She97] Michael Sherman. Subseries methods in regression. Journal of the American Statistical Association, 92(439):1041–1048, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/sherman.htm.

Sun:1996:BMV

[SHGL96] Li Sun, John S. J. Hsu, Irwin Guttman, and Tom Leonard. Bayesian methods for variance component models. Journal of the American Statistical Association, 91(434):743–752, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Speed:1978:MAL

[SHH78] F. M. Speed, R. R. Hocking, and D. P. Hackney. Methods of analysis of linear models with unbalanced data. Journal of the American Statistical Association, 73(??):105–112, ???? 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shiller:1984:SPN

[Shi84] Robert J. Shiller. Smoothness priors and nonlinear regression. Journal of the American Statistical Association, 79(387):609–615, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:1991:THO

[Shi91] Ning-Zhong Shi. A test of homogeneity of odds ratios against order restrictions. Journal of the American Statistical Association, 86 (413):154–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Simpson:1998:CEP

[SHL98] Douglas G. Simpson, Xuming He, and Yao-Tsorng Liu. Comment: Edge preserving smoothers for image processing. Journal of the American Statistical Association, 93(442):544–??, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Short:1993:ADM

[Sho93] Thomas H. Short. An algorithm for the detection and measurement of rail surface defects. Journal of the American Statistical Association, 88(422):436–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Skinner:1986:ESD

[SHS86] C. J. Skinner, D. J. Holmes, and T. M. F. Smith. The effect of sample design on principal component analysis. Journal of the American Statistical Association, 81(395):789–798, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shun:1997:ALS

[Shu97] Zhenming Shun. Another look at the salamander mating data: A modified Laplace approximation approach. Journal of the American Statistical Association, 92(437):341–??, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/shun.htm.

Spurrier:1985:ESC

[SI85] John D. Spurrier and Steven P. Isham. Exact simultaneous confidence intervals for pairwise comparisons of three normal means. Journal of the American Statistical Association, 80(390):438–442, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Strauss:1990:PES

[SI90] David Strauss and Michael Ikeda. Pseudolikelihood estimation for social networks. Journal of the American Statistical Association, 85(409):204–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Siegel:1993:SCI

[Sie93] Andrew F. Siegel. A surprising covariance involving the minimum of multivariate normal variables. Journal of the American Statistical Association, 88(421):77–80, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Silverman:1984:FEC

[Sil84] B. W. Silverman. A fast and efficient cross-validation method for smoothing parameter choice in spline regression. Journal of the American Statistical Association, 79(387):584–589, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Silvapulle:1992:RWT

[Sil92] Mervyn J. Silvapulle. Robust Wald-type tests of one-sided hypotheses in the linear model. Journal of the American Statistical Association, 87(417):156–161, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Silvapulle:1996:TPN

[Sil96] Mervyn J. Silvapulle. A test in the presence of nuisance parameters. Journal of the American Statistical Association, 91(436): 1690–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Silliman:1997:HSM

[Sil97] Nancy Paul Silliman. Hierarchical selection models with applications in meta- analysis. Journal of the American Statistical Association, 92(439):926–936, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/silliman.htm.

Simon:1974:AAS

[Sim74] G. Simon. Alternative analyses for the singly-ordered contingency table. Journal of the American Statistical Association, 69(348): 971–976, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Simon:1978:PNA

[Sim78] Gary A. Simon. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments: Comment. Journal of the American Statistical Association, 73 (363):520–521, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/pdfplus/2286592.pdf. See [SM78a].

Simonoff:1986:JBG

[Sim86] Jeffrey S. Simonoff. Jackknifing and bootstrapping goodness-of-fit statistics in sparse multinomials. Journal of the American Statistical Association, 81(396):1005–1011, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Simonoff:1987:CPS

[Sim87a] Jeffrey S. Simonoff. Comment on: “Performance of some resistant rules for outlier labeling” [J. Amer. Statist. Assoc. 81 (1986), no. 396, 991–999] by D. C. Hoaglin, B. Iglewicz and J. W. Tukey. With reply by Hoaglin, Iglewicz and Tukey. Journal of the American Statistical Association, 82(398):703–704, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [HIT86].

Simpson:1987:MHD

[Sim87b] Douglas G. Simpson. Minimum Hellinger distance estimation for the analysis of count data. Journal of the American Statistical Association, 82(399):802–807, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Simpson:1989:HDT

[Sim89] Douglas G. Simpson. Hellinger deviance tests: efficiency, breakdown points, and examples. Journal of the American Statistical Association, 84(405):107–113, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Singh:1986:SAO

[Sin86] N. Singh. A simple and asymptotically optimal test for the equality of normal populations. A pragmatic approach to one-way classification. Journal of the American Statistical Association, 81(395): 703–704, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sitter:1992:RPC

[Sit92] R. R. Sitter. A resampling procedure for complex survey data. Journal of the American Statistical Association, 87(419):755–765, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sitter:1997:VER

[Sit97] R. R. Sitter. Variance estimation for the regression estimator in two-phase sampling. Journal of the American Statistical Association, 92(438):779–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Suchman:1990:ITF

[SJ90] Lucy Suchman and Brigitte Jordan. Interactional troubles in faceto-face survey interviews. Journal of the American Statistical Association, 85(409):232–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shiboski:1992:SAT

[SJ92] Stephen C. Shiboski and Nicholas P. Jewell. Statistical analysis of the time dependence of HIV infectivity based on partner study data. Journal of the American Statistical Association, 87(418): 360–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Storer:1990:EPS

[SK90] Barry E. Storer and Choongrak Kim. Exact properties of some exact test statistics for comparing two binomial proportions. Journal of the American Statistical Association, 85(409):146–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shumway:1991:APM

[SK91] Robert H. Shumway and Myron J. Katzoff. Adjustment of provisional mortality series: The dynamic linear model with structured measurement errors. Journal of the American Statistical Association, 86(415):611–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Satten:1993:IAE

[SK93a] Glen A. Satten and Lawrence L. Kupper. Inferences about exposure-disease associations using probability-of-exposure information. Journal of the American Statistical Association, 88(421): 200–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:1993:ACS

[SK93b] Jianguo Sun and John D. Kalbfleisch. The analysis of current status data on point processes. Journal of the American Statistical Association, 88(424):1449–1454, December 1993. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smith:1997:BAN

[SK97] Michael Smith and Robert Kohn. A Bayesian approach to nonparametric bivariate regression. Journal of the American Statistical Association, 92(440):1522–1535, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/SMITH_M.HTM.

Sepanski:1994:SCA

[SKC94] J. H. Sepanski, R. Knickerbocker, and R. J. Carroll. A semiparametric correction for attenuation. Journal of the American Statistical Association, 89(428):1366–1373, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stasny:1998:FDP

[SKF98] Elizabeth A. Stasny, Joseph B. Kadane, and Kathleen S. Fritsch. On the fairness of death-penalty jurors: A comparison of Bayesian models with different levels of hierarchy and various missing-data mechanisms. Journal of the American Statistical Association, 93(442):464–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/3r0c0005.pdf;   http:
//www.amstat.org/publications/jasa/abstracts_98/june/
STASNY.HTM.

Skinner:1991:ERR

[Ski91] C. J. Skinner. On the efficiency of raking ratio estimation for multiple frame surveys. Journal of the American Statistical Association, 86(415):779–784, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schmeiser:1980:SMG

[SL80] B. W. Schmeiser and R. Lal. Squeeze methods for generating gamma variates. Journal of the American Statistical Association, 75(371):679–682, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Self:1987:APM

[SL87] Steven G. Self and Kung-Yee Liang. Asymptotic properties of maximum likelihood estimators and likelihood ratio tests under nonstandard conditions. Journal of the American Statistical Association, 82(398):605–610, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Saikkonen:1993:TMA

[SL93a] Pentti Saikkonen and Ritva Luukkonen. Testing for a moving average unit root in autoregressive integrated moving average models. Journal of the American Statistical Association, 88(422):596–601, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Su:1993:LCM

[SL93b] John Q. Su and Jun S. Liu. Linear combinations of multiple diagnostic markers. Journal of the American Statistical Association, 88 (424):1350–1355, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spiekerman:1998:MRM

[SL98a] C. F. Spiekerman and D. Y. Lin. Marginal regression models for multivariate failure time data. Journal of the American Statistical Association, 93(443):1164–1175, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/spiekerman.html.

Staniswalis:1998:NRA

[SL98b] Joan G. Staniswalis and J. Jack Lee. Nonparametric regression analysis of longitudinal data. Journal of the American Statistical Association, 93(444):1403–1418, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/dec/staniswalis.html; http://www.jstor.org/stable/
2670055.

Slate:1994:PNE

[Sla94] Elizabeth H. Slate. Parameterizations for natural exponential families with quadratic variance functions. Journal of the American Statistical Association, 89(428):1471–1482, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shah:1997:REM

[SLS97] Amrik Shah, Nan Laird, and David Schoenfeld. A random effects model for multiple characteristics with possibly missing data. Journal of the American Statistical Association, 92(438):774–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schruben:1978:PNAa

[SM78a] Lee W. Schruben and Barry H. Margolin. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments. Journal of the American Statistical Association, 73(363):504–520, September 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/pdfplus/2286590.pdf. See discussion

[Gav78, Kie78, Mal78, SM78b, Sim78].

Schruben:1978:PNAb

[SM78b] Lee W. Schruben and Barry H. Margolin. Pseudorandom number assignment in statistically designed simulation and distribution sampling experiments: Rejoinder. Journal of the American Statistical Association, 73(363):524–525, September 1978. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/pdfplus/2286595.pdf. See [SM78a].

Subbaiah:1981:MAS

[SM81] Perla Subbaiah and Govind S. Mudholkar. On a multivariate analog of Studentized range test. Journal of the American Statistical Association, 76(375):725–728, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Saunders:1983:MLE

[SM83] Sam C. Saunders and Janet M. Myhre. Maximum likelihood estimation for two-parameter decreasing hazard rate distributions using censored data. Journal of the American Statistical Association, 78(383):664–673, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Saunders:1984:BCM

[SM84a] Sam C. Saunders and Janet M. Myhre. On the behavior of certain maximum likelihood estimators from large, randomly censored samples. Journal of the American Statistical Association, 79(386):294–301, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smith:1984:AEM

[SM84b] David W. Smith and Leigh W. Murray. An alternative to Eisenhart’s model ii and mixed model in the case of negative variance estimates. Journal of the American Statistical Association, 79(385): 145–151, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sheather:1990:KQE

[SM90a] Simon J. Sheather and J. S. Marron. Kernel quantile estimators. Journal of the American Statistical Association, 85(410):410–416, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spencer:1990:NDE

[SM90b] Bruce D. Spencer and Lincoln E. Moses. Needed data expenditure for an ambiguous decision problem. Journal of the American Statistical Association, 85(412):1099–1104, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spall:1992:FBE

[SM92] James C. Spall and John L. Maryak. A feasible Bayesian estimator of quantiles for projectile accuracy from non-iid data. Journal of the American Statistical Association, 87(419):676–??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Seillier-Moiseiwitsch:1993:TVS

[SMD93] F. Seillier-Moiseiwitsch and A. P. Dawid. On testing the validity of sequential probability forecasts. Journal of the American Statistical Association, 88(421):355–359, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Seo:1994:GTC

[SMF94] Takashi Seo, Shigeru Mano, and Yasunori Fujikoshi. A generalized Tukey conjecture for multiple comparisons among mean vectors. Journal of the American Statistical Association, 89(426):676–679, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smouse:1982:BEF

[Smo82] Evan P. Smouse. Bayesian estimation of a finite population total using auxiliary information in the presence of nonresponse. Journal of the American Statistical Association, 77(377):97–102, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Senchaudhuri:1995:EEV

[SMP95] Pralay Senchaudhuri, Cyrus R. Mehta, and Nitin R. Patel. Estimating exact p values by the method of control variates or Monte Carlo rescue. Journal of the American Statistical Association, 90 (430):640–648, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spurrier:1990:SSA

[SN90] John D. Spurrier and Azhar Nizam. Sample size allocation for simultaneous inference in comparison with control experiments. Journal of the American Statistical Association, 85(409):181–??, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Samaniego:1996:HBB

[SN96] Francisco J. Samaniego and Andrew A. Neath. How to be a better Bayesian. Journal of the American Statistical Association, 91 (434):733–742, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shafer:1983:AVA

[SO83] Glenn Shafer and Ingram Olkin. Adjusting P values to account for selection over dichotomies. Journal of the American Statistical Association, 78(383):674–678, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stehman:1994:CVE

[SO94] Stephen V. Stehman and W. Scott Overton. Comparison of variance estimators of the Horvitz–Thompson estimator for randomized variable probability systematic sampling. Journal of the American Statistical Association, 89(425):30–??, March 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sobel:1991:BMS

[Sob91] Marc J. Sobel. Bayes and minimax sample-admissible multivariate selection procedures. Journal of the American Statistical Association, 86(416):1022–1023, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sobel:1993:BEB

[Sob93] Marc J. Sobel. Bayes and empirical Bayes procedures for comparing parameters. Journal of the American Statistical Association, 88(422):687–693, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Solo:1984:ODA

[Sol84] Victor Solo. The order of differencing in ARIMA models. Journal of the American Statistical Association, 79(388):916–921, ????

1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Solow:1991:EAO

[Sol91] Andrew R. Solow. An exploratory analysis of the occurrence of explosive volcanism in the Northern Hemisphere, 1851–1985. Journal of the American Statistical Association, 86(413):49–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Solow:1995:EAR

[Sol95] Andrew R. Solow. An exploratory analysis of a record of el nino events, 1800- 1987. Journal of the American Statistical Association, 90(429):72–??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Soofi:1992:GFC

[Soo92] Ehsan S. Soofi. A generalizable formulation of conditional logit with diagnostics. Journal of the American Statistical Association, 87(419):812–??, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Soofi:1994:CIC

[Soo94] Ehsan S. Soofi. Capturing the intangible concept of information. Journal of the American Statistical Association, 89(428): 1243–1254, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spino:1991:ECP

[SP91] Cathie Spino and Marcello Pagano. Efficient calculation of the permutation distribution of trimmed means. Journal of the American Statistical Association, 86(415):729–737, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spencer:1985:ODQ

[Spe85] Bruce D. Spencer. Optimal data quality. Journal of the American Statistical Association, 80(391):564–573, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spitzer:1982:FEA

[Spi82] J. J. Spitzer. A fast and efficient algorithm for the estimation of parameters in models with the Box-and-Cox transformation.

Journal of the American Statistical Association, 77(380):760–766, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spliid:1983:FEM

[Spl83] Henrik Spliid. A fast estimation method for the vector autoregressive moving average model with exogenous variables. Journal of the American Statistical Association, 78(384):843–849, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sposito:1982:URE

[Spo82] V. A. Sposito. On unbiased L_{p} regression estimators. Journal of the American Statistical Association, 77(379):652–653, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sprott:1983:EPC

[Spr83] David A. Sprott. Estimating the parameters of a convolution by maximum likelihood. Journal of the American Statistical Association, 78(382):457–460, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spurrier:1988:GST

[Spu88] John D. Spurrier. Generalizations of Steel’s treatments-versus-control multivariate sign test. Journal of the American Statistical Association, 83(402):471–476, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spurrier:1999:ECB

[Spu99] John D. Spurrier. Exact confidence bounds for all contrasts of three or more regression lines. Journal of the American Statistical Association, 94(446):483–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/spurrier.pdf.

Stromberg:1992:BNR

[SR92] Arnold J. Stromberg and David Ruppert. Breakdown in nonlinear regression. Journal of the American Statistical Association, 87 (420):991–997, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Samaniego:1994:TRB

[SR94] Francisco J. Samaniego and Dana M. Reneau. Toward a reconciliation of the Bayesian and frequentist approaches to point estimation. Journal of the American Statistical Association, 89(427): 947–957, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Singh:1995:AGF

[SR95] A. C. Singh and J. N. K. Rao. On the adjustment of gross flow estimates for classification error with application to data from the Canadian labour force survey. Journal of the American Statistical Association, 90(430):478–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Skinner:1996:EDF

[SR96] C. J. Skinner and J. N. K. Rao. Estimation in dual frame surveys with complex designs. Journal of the American Statistical Association, 91(433):349–356, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Simpson:1992:OSG

[SRC92] D. G. Simpson, D. Ruppert, and R. J. Carroll. On one-step GM estimates and stability of inferences in linear regression. Journal of the American Statistical Association, 87(418):439–450, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Srivastava:1972:DVE

[Sri72] V. K. Srivastava. Disturbance-variance estimation in simultaneous equations when disturbances are small. Journal of the American Statistical Association, 67(??):164–168, ???? 1972. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Silber:1995:CCG

[SRR95] Jeffrey H. Silber, Paul R. Rosenabaum, and Richard N. Ross. Comparing the contributions of groups of predictors: Which outcomes vary with hospital rather than patient characteristics? Journal of the American Statistical Association, 90(429):7–??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stahel:1994:REA

[SRSD94] Werner A. Stahel, Andreas F. Ruckstuhl, Peter Senn, and Kurt Dressler. Robust estimation in the analysis of complex molecular spectra. Journal of the American Statistical Association, 89 (427):788–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schmeiser:1980:ARM

[SS80] B. W. Schmeiser and M. A. Shalaby. Acceptance/rejection methods for beta variate generation. Journal of the American Statistical Association, 75(371):673–678, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sethuraman:1982:THD

[SS82a] J. Sethuraman and Nozer D. Singpurwalla. Testing of hypotheses for distributions in accelerated life tests. Journal of the American Statistical Association, 77(377):204–208, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smith:1982:BOE

[SS82b] Philip J. Smith and J. Sedransk. Bayesian optimization of the estimation of the age composition of a fish population. Journal of the American Statistical Association, 77(380):707–713, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sampson:1985:MSI

[SS85] Paul D. Sampson and Andrew F. Siegel. The measure of “size” independent of “shape” for multivariate lognormal populations. Journal of the American Statistical Association, 80(392):910–914, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Simon:1986:DPM

[SS86] Gary A. Simon and Jeffrey S. Simonoff. Diagnostic plots for missing data in least squares regression. Journal of the American Statistical Association, 81(394):501–509, ???? 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scholz:1987:SAD

[SS87a] F.-W. Scholz and M. A. Stephens. k-sample Anderson–Darling tests. Journal of the American Statistical Association, 82(399):

918–924, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Souvaine:1987:TSE

[SS87b] Diane L. Souvaine and J. Michael Steele. Time- and spaceefficient algorithms for least median of squares regression. Journal of the American Statistical Association, 82(399):794–801, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stokes:1988:CRS

[SS88] S. Lynne Stokes and Thomas W. Sager. Characterization of a ranked-set sample with application to estimating distribution functions. Journal of the American Statistical Association, 83 (402):374–381, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smith:1990:RBL

[SS90a] Stanley K. Smith and Terry Sincich. The relationship between the length of the base period and population forecast errors. Journal of the American Statistical Association, 85(410):367–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stine:1990:BAS

[SS90b] Robert A. Stine and Paul Shaman. Bias of autoregressive spectral estimators. Journal of the American Statistical Association, 85 (412):1091–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shumway:1991:DLM

[SS91a] R. H. Shumway and D. S. Stoffer. Dynamic linear models with switching. Journal of the American Statistical Association, 86 (415):763–??, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Staniswalis:1991:DAR

[SS91b] Joan G. Staniswalis and Thomas A. Severini. Diagnostics for assessing regression models. Journal of the American Statistical Association, 86(415):684–692, September 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Severini:1994:QLE

[SS94] Thomas A. Severini and Joan G. Staniswalis. Quasi-likelihood estimation in semiparametric models. Journal of the American Statistical Association, 89(426):501–511, June 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [Sta94].

Silvapulle:1995:STA

[SS95a] Mervyn J. Silvapulle and Paramsothy Silvapulle. A score test against one-sided alternatives. Journal of the American Statistical Association, 90(429):342–349, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smith:1995:EPP

[SS95b] Stanley K. Smith and Mohammed Shahidullah. An evaluation of population projection errors for census tracts. Journal of the American Statistical Association, 90(429):64–??, March 1995. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sain:1996:LAD

[SS96a] Stephan R. Sain and David W. Scott. On locally adaptive density estimation. Journal of the American Statistical Association, 91 (436):1525–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shao:1996:BIS

[SS96b] Jun Shao and Randy R. Sitter. Bootstrap for imputed survey data. Journal of the American Statistical Association, 91(435): 1278–1288, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Solna:1996:TTE

[SS96c] Knut Sølna and Paul Switzer. Time trend estimation for a geographic region. Journal of the American Statistical Association, 91(434):577–589, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schell:1997:RMR

[SS97] Michael J. Schell and Bahadur Singh. The reduced monotonic regression method. Journal of the American Statistical Association, 92(437):128–??, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_97/march/schell.htm.

Shao:1999:VES

[SS99] Jun Shao and Philip Steel. Variance estimation for survey data with composite imputation and nonnegligible sampling fractions. Journal of the American Statistical Association, 94(445):254–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sackrowitz:1994:RUH

[SSC94] Harold Sackrowitz and Ester Samuel-Cahn. Rationality and unbiasedness in hypothesis testing. Journal of the American Statistical Association, 89(427):967–971, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schervish:1990:SDU

[SSK90] Mark J. Schervish, Teddy Seidenfeld, and Joseph B. Kadane. State-dependent utilities. Journal of the American Statistical Association, 85(411):840–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Siegel:1960:NSR

[ST60] Sidney Siegel and John W. Tukey. A nonparametric sum of ranks procedure for relative spread in unpaired samples. Journal of the American Statistical Association, 55(??):429–445, 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction in this journal, 56, 1005–1005, 1961.

Shaked:1982:EPM

[ST82] Moshe Shaked and Lanh Tat Tran. Estimating parameters from mixed samples. Journal of the American Statistical Association, 77(377):196–203, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scott:1987:BUC

[ST87] David W. Scott and George R. Terrell. Biased and unbiased crossvalidation in density estimation. Journal of the American Statistical Association, 82(400):1131–1146, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Soper:1992:ETS

[ST92] Keith A. Soper and Philip Troilo. Exact tests for in situ hybridization experiments. Journal of the American Statistical Association, 87(417):78–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Severini:1996:CEB

[ST96] Thomas A. Severini and Martin A. Tanner. Comment: Empirical Bayes methods for combining likelihoods. Journal of the American Statistical Association, 91(434):538–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Siquiera:1999:TEL

[ST99] Arminda Lucia Siquiera and Jeremy M. G. Taylor. Treatment effects in a logistic model involving the Box–Cox transformation. Journal of the American Statistical Association, 94(445):240–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Staniswalis:1989:KER

[Sta89a] Joan G. Staniswalis. The kernel estimate of a regression function in likelihood-based models. Journal of the American Statistical Association, 84(405):276–283, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Staniswalis:1989:LBS

[Sta89b] Joan G. Staniswalis. Local bandwidth selection for kernel estimates. Journal of the American Statistical Association, 84(405): 284–288, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stasny:1991:HMP

[Sta91] Elizabeth A. Stasny. Hierarchical models for the probabilities of a survey classification and nonresponse: An example from the National Crime Survey. Journal of the American Statistical Association, 86(414):296–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Staniswalis:1994:CQL

[Sta94] Joan G. Staniswalis. Correction to: “Quasi-likelihood estimation in semiparametric models” [J. Amer. Statist. Assoc. 89 (1994), no. 426, 501–511 and 1294076] by T. A. Severini and Joan G.

Staniswalis. Journal of the American Statistical Association, 89 (428):1572, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [SS94].

Stephenson:1981:GCO

[Ste81] W. Robert Stephenson. A general class of one-sample nonparametric test statistics based on subsamples. Journal of the American Statistical Association, 76(376):960–966, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stein:1987:MNQ

[Ste87] Michael L. Stein. Minimum norm quadratic estimation of spatial variograms. Journal of the American Statistical Association, 82 (399):765–772, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stern:1990:MDP

[Ste90] Hal Stern. Models for distributions on permutations. Journal of the American Statistical Association, 85(410):558–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stern:1994:BMM

[Ste94] Hal S. Stern. A Brownian motion model for the progress of sports scores. Journal of the American Statistical Association, 89(427): 1128–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stein:1995:FDA

[Ste95] Michael L. Stein. Fixed-domain asymptotics for spatial periodograms. Journal of the American Statistical Association, 90 (432):1277–1288, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stein:1999:CPS

[Ste99] Michael L. Stein. Comment: Prediction of spatial cumulative distribution functions using subsampling. Journal of the American Statistical Association, 94(445):106–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stine:1985:BPI

[Sti85] Robert A. Stine. Bootstrap prediction intervals for regression. Journal of the American Statistical Association, 80(392):1026– 1031, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stigler:1986:JCP

[Sti86] Stephen M. Stigler. John Craig and the probability of history: from the death of Christ to the birth of Laplace. Journal of the American Statistical Association, 81(396):879–887, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stine:1987:EPA

[Sti87] Robert A. Stine. Estimating properties of autoregressive forecasts. Journal of the American Statistical Association, 82(400): 1072–1078, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stigler:1988:LBO

[Sti88] Stephen M. Stigler. A look backward on the occasion of the centenary of JASA. Journal of the American Statistical Association, 83(403):583–587, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stephan:1965:SW

[STM^{+}65] Frederick F. Stephan, John W. Tukey, Frederick Mosteller, Alex M. Mood, Morris H. Hansen, Leslie E. Simon, and W. J. Dixon. Samuel S. Wilks. Journal of the American Statistical Association, 60(312):939–966, December 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2283397.

Stoker:1982:UCS

[Sto82] Thomas M. Stoker. The use of cross-section data to characterize macrofunctions. Journal of the American Statistical Association, 77(378):369–380, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stoffer:1986:EIS

[Sto86] David S. Stoffer. Estimation and identification of space-time AR-MAX models in the presence of missing data. Journal of the American Statistical Association, 81(395):762–772, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stock:1988:ECT

[Sto88] James H. Stock. Estimating continuous-time processes subject to time deformation. Journal of the American Statistical Association, 83(401):77–85, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stock:1989:NPA

[Sto89] James H. Stock. Nonparametric policy analysis. Journal of the American Statistical Association, 84(406):567–575, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stoffer:1991:WFA

[Sto91] David S. Stoffer. Walsh-Fourier analysis and its statistical applications. Journal of the American Statistical Association, 86(414): 461–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stromberg:1993:CHB

[Str93] Arnold J. Stromberg. Computation of high breakdown nonlinear regression parameters. Journal of the American Statistical Association, 88(421):237–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scharfstein:1997:SEI

[STR97] Daniel O. Scharfstein, Anastasios A. Tsiatis, and James M. Robins. Semiparametric efficiency and its implication on the design and analysis of group- sequential studies. Journal of the American Statistical Association, 92(440):1342–1350, December 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/december/SCHARFST.HTM.

Stuart:1956:ETR

[Stu56] Alan Stuart. The efficiencies of tests of randomness, distributionfree methods vs. normal alternatives. Journal of the American Statistical Association, 51(274):285–287, June 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281348.

Studden:1982:SRT

[Stu82] W. J. Studden. Some robust-type D-optimal designs in polynomial regression. Journal of the American Statistical Association, 77 (380):916–921, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stukel:1988:GLM

[Stu88] Thérèse A. Stukel. Generalized logistic models. Journal of the American Statistical Association, 83(402):426–431, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shumway:1974:LDF

[SU74] R. H. Shumway and A. N. Unger. Linear discriminant functions for stationary time series. Journal of the American Statistical Association, 69(348):948–956, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Suzuki:1985:NEL

[Suz85] Kazuyuki Suzuki. Nonparametric estimation of lifetime distributions from a record of failures and follow-ups. Journal of the American Statistical Association, 80(389):68–72, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scarsini:1993:BDN

[SV93] Marco Scarsini and Achilles Venetoulias. Bivariate distributions with nonmonotone dependence structure. Journal of the American Statistical Association, 88(421):338–344, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shapiro:1980:BSE

[SW80] C. P. Shapiro and Robert L. Wardrop. Bayesian sequential estimation for one-parameter exponential families. Journal of the American Statistical Association, 75(372):984–988, ???? 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Slud:1982:TSR

[SW82] Eric Slud and L. J. Wei. Two-sample repeated significance tests based on the modified Wilcoxon statistic. Journal of the American Statistical Association, 77(380):862–868, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Singpurwalla:1983:KEF

[SW83] Nozer D. Singpurwalla and Man Yuen Wong. Kernel estimators of the failure-rate function and density estimation: an analogy.

Journal of the American Statistical Association, 78(382):478–481, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Srivastava:1986:LRT

[SW86] M. S. Srivastava and K. J. Worsley. Likelihood ratio tests for a change in the multivariate normal mean. Journal of the American Statistical Association, 81(393):199–204, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Silverman:1987:NEB

[SW87] B. W. Silverman and J. T. Wood. The nonparametric estimation of branching curves. Journal of the American Statistical Association, 82(398):551–558, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stock:1988:TCT

[SW88] James H. Stock and Mark W. Watson. Testing for common trends. Journal of the American Statistical Association, 83(404):1097– 1107, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sampson:1989:EMD

[SW89] Allan R. Sampson and Lyn R. Whitaker. Estimation of multivariate distributions under stochastic ordering. Journal of the American Statistical Association, 84(406):541–548, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shmueli:1990:PMC

[SW90] Uri Shmueli and George H. Weiss. Probabilistic methods in crystal structure analysis. Journal of the American Statistical Association, 85(409):6–19, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stoffer:1991:BSS

[SW91a] David S. Stoffer and Kent D. Wall. Bootstrapping state-space models: Gaussian maximum likelihood estimation and the Kalman filter. Journal of the American Statistical Association, 86(416): 1024–1033, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Su:1991:LFT

[SW91b] John Q. Su and L. J. Wei. A lack-of-fit test for the mean function in a generalized linear model. Journal of the American Statistical Association, 86(414):420–426, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sakata:1995:ADF

[SW95] Shinichi Sakata and Halbert White. An alternative definition of finite-sample breakdown point with applications to regression model estimators. Journal of the American Statistical Association, 90(431):1099–1106, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Strawderman:1997:ABC

[SW97] Robert L. Strawderman and Martin T. Wells. Accurate bootstrap confidence limits for the cumulative hazard and survivor functions under random censoring. Journal of the American Statistical Association, 92(440):1356–1374, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/STRAWDER.HTM.

Stock:1998:MUE

[SW98a] James H. Stock and Mark W. Watson. Median unbiased estimation of coefficient variance in a time varying parameter model. Journal of the American Statistical Association, 93(441): 349–358, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/STOCK.HTM.

Strawderman:1998:AEI

[SW98b] Robert L. Strawderman and Martin T. Wells. Approximately exact inference for the common odds ratio in several 2 × 2 tables. Journal of the American Statistical Association, 93(444): 1294–1306, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/dec/strawderman.html.

Strawderman:1998:RAE

[SW98c] Robert L. Strawderman and Martin T. Wells. Rejoinder: Approximately exact inference for the common odds ratio in several 2 × 2 tables. Journal of the American Statistical Association, 93(444):

1316–??, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sweeting:1984:AIL

[Swe84] Trevor J. Sweeting. Approximate inference in location-scale regression models. Journal of the American Statistical Association, 79(388):847–852, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sirvanci:1984:EWP

[SY84] Mete Sirvanci and Grace Yang. Estimation of the Weibull parameters under type i censoring. Journal of the American Statistical Association, 79(385):183–187, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:1995:RPB

[SY95] Dong Chu Sun and Keying Ye. Reference prior Bayesian analysis for normal mean products. Journal of the American Statistical Association, 90(430):589–597, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stein:1987:PES

[SZJ87] Gillian Z. Stein, Walter Zucchini, and June M. Juritz. Parameter estimation for the Sichel distribution and its multivariate extension. Journal of the American Statistical Association, 82(399): 938–944, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thompson:1989:ERT

[TA89] G. L. Thompson and L. P. Ammann. Efficacies of rank-transform statistics in two-way models with no interaction. Journal of the American Statistical Association, 84(405):325–330, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thompson:1990:EIR

[TA90] G. L. Thompson and L. P. Ammann. Efficiencies of interblock rank statistics for repeated measures designs. Journal of the American Statistical Association, 85(410):519–528, June 1990. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tjostheim:1994:NINa

[TA94a] Dag Tjøstheim and Bjørn H. Auestad. Nonparametric identification of nonlinear time series: Projections. Journal of the American Statistical Association, 89(428):1398–1409, December 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tjostheim:1994:NINb

[TA94b] Dag Tjøstheim and Bjørn H. Auestad. Nonparametric identification of nonlinear time series: selecting significant lags. Journal of the American Statistical Association, 89(428):1410–1419, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tableman:1990:BIR

[Tab90] Mara Tableman. Bounded-influence rank regression: A one-step estimator based on Wilcoxon scores. Journal of the American Statistical Association, 85(410):508–513, June 1990. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tadikamalla:1980:RSE

[Tad80] P. R. Tadikamalla. Random sampling from the exponential power distribution. Journal of the American Statistical Association, 75 (371):683–686, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Takada:1981:ISE

[Tak81] Yoshikazu Takada. Invariant sequential estimation of the exponential mean life from the ordered observations. Journal of the American Statistical Association, 76(374):470–472, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Takemura:1983:TAA

[Tak83] Akimichi Takemura. Tensor analysis of ANOVA decomposition. Journal of the American Statistical Association, 78(384):894–900, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tamhane:1981:RRT

[Tam81] Ajit C. Tamhane. Randomized response techniques for multiple sensitive attributes. Journal of the American Statistical Association, 76(376):916–923, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tamhane:1985:SSP

[Tam85] Ajit C. Tamhane. Some sequential procedures for selecting the better Bernoulli treatment by using a matched samples design. Journal of the American Statistical Association, 80(390):455–460, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tam:1988:SRR

[Tam88a] S. M. Tam. Some results on robust estimation in finite population sampling. Journal of the American Statistical Association, 83(401):242–248, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tamaki:1988:BAB

[Tam88b] Mitsushi Tamaki. A Bayesian approach to the best-choice problem. Journal of the American Statistical Association, 83(404): 1129–1133, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tam:1995:ORS

[Tam95] S. M. Tam. Optimal and robust strategies for cluster sampling. Journal of the American Statistical Association, 90(429):379–382, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tanaka:1983:ETF

[Tan83] Katsuto Tanaka. Estimation for transients in the frequency domain. Journal of the American Statistical Association, 78(383): 718–724, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tang:1993:OAB

[Tan93] Boxin Tang. Orthogonal array-based Latin hypercubes. Journal of the American Statistical Association, 88(424):1392–1397, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tang:1994:UMP

[Tan94] Dei-In Tang. Uniformly more powerful tests in a one-sided multivariate problem. Journal of the American Statistical Association, 89(427):1006–1011, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tarter:1986:MIS

[Tar86] Michael E. Tarter. Mean integrated squared error sampling. Journal of the American Statistical Association, 81(393):234–242, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tardif:1987:EOR

[Tar87] Serge Tardif. Efficiency and optimality results for tests based on weighted rankings. Journal of the American Statistical Association, 82(398):637–644, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tarpey:1999:SCP

[Tar99] Thaddeus Tarpey. Self-consistency and principle component analysis. Journal of the American Statistical Association, 94(446): 456–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Taylor:1985:MLS

[Tay85] Jeremy M. G. Taylor. Measures of location of skew distributions obtained through Box–Cox transformations. Journal of the American Statistical Association, 80(390):427–432, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tiao:1981:MMT

[TB81] G. C. Tiao and G. E. P. Box. Modeling multiple time series with applications. Journal of the American Statistical Association, 76 (376):802–816, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tryfos:1985:FR

[TB85] Peter Tryfos and Russell Blackmore. Forecasting records. Journal of the American Statistical Association, 80(389):46–50, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tamura:1986:MHD

[TB86] Roy N. Tamura and Dennis D. Boos. Minimum Hellinger distance estimation for multivariate location and covariance. Journal of the American Statistical Association, 81(393):223–229, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tan:1972:SCM

[TC72] W. Y. Tan and W. C. Chang. Some comparisons of the method of moments and the method of maximum likelihood in estimating parameters of a mixture of two normal distributions. Journal of the American Statistical Association, 67(??):702–708, ???? 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Taylor:1994:SMA

[TCS94] J. M. G. Taylor, W. G. Cumberland, and J. P. Sy. A stochastic model for analysis of longitudinal AIDS data. Journal of the American Statistical Association, 89(427):727–??, September 1994. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsiatis:1995:MRS

[TDW95] A. Tsiatis, Victor DeGruttola, and M. S. Wulfsohn. Modeling the relationship of survival to longitudinal data measured with error. applications to survival and CD4 counts in patients with AIDS. Journal of the American Statistical Association, 90(429): 27–??, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Teichroew:1965:HDS

[Tei65] D. Teichroew. A history of distribution sampling prior to the era of the computer and its relevance to simulation. Journal of the American Statistical Association, 60(309):27–49, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Telser:1967:DSM

[Tel67] L. G. Telser. Discrete samples and moving sums in stationary stochastic processes. Journal of the American Statistical Association, ??(??):484–499, June 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Terrell:1990:MSP

[Ter90] George R. Terrell. The maximal smoothing principle in density estimation. Journal of the American Statistical Association, 85 (410):470–477, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Terasvirta:1994:SEE

[Ter94] Timo Terasvirta. Specification, estimation, and evaluation of smooth transition autoregressive models. Journal of the American Statistical Association, 89(425):208–??, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tingley:1990:SSC

[TF90] Maureen Tingley and Christopher Field. Small-sample confidence intervals. Journal of the American Statistical Association, 85(410): 427–434, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tamura:1994:CSA

[TFAH94] Roy N. Tamura, Douglas E. Faries, John S. Andersen, and John H. Heiligenstein. A case study of an adaptive clinical trial in the treatment of out-patients with depressive disorder. Journal of the American Statistical Association, 89(427):768–??, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tarter:1990:MNS

[TFP90] Michael E. Tarter, William R. Freeman, and Lincoln Polissar. Modular nonparametric subsurvival estimation. Journal of the American Statistical Association, 85(409):29–??, March 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tu:1982:DAB

[TH82] Ching Tsao Tu and Chien Pai Han. Discriminant analysis based on binary and continuous variables. Journal of the American Statistical Association, 77(378):447–454, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tibshirani:1987:LLE

[TH87] Robert Tibshirani and Trevor Hastie. Local likelihood estimation. Journal of the American Statistical Association, 82(398):559–567, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Theodossiou:1993:PSM

[The93] Panayiotis T. Theodossiou. Predicting shifts in the mean of a multivariate time series process: An application in predicting business failures. Journal of the American Statistical Association, 88(422): 441–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Theberge:1999:ECE

[Thé99] Alain Théberge. Extensions of calibration estimators in survey sampling. Journal of the American Statistical Association, 94 (446):635–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thomas:1990:ICR

[Tho90a] William Thomas. Influence on confidence regions for regression coefficients in generalized linear models. Journal of the American Statistical Association, 85(410):393–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thompson:1990:ACS

[Tho90b] Steven K. Thompson. Adaptive cluster sampling. Journal of the American Statistical Association, 85(412):1050–??, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thomas:1991:IDC

[Tho91a] William Thomas. Influence diagnostics for the cross-validated smoothing parameter in spline smoothing. Journal of the American Statistical Association, 86(415):693–??, September 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thompson:1991:UAR

[Tho91b] G. L. Thompson. A unified approach to rank tests for multivariate and repeated measures designs. Journal of the American Statistical Association, 86(414):410–419, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thursby:1985:RAS

[Thu85] Jerry G. Thursby. The relationship among the specification tests of Hausman, Ramsey, and Chow. Journal of the American Statistical Association, 80(392):926–928, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tin:1965:CSR

[Tin65] M. Tin. Comparison of some ratio estimators. Journal of the American Statistical Association, 60(309):294–307, ???? 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsui:1988:NAT

[TJW88] Kwok-Leung Tsui, Nicholas P. Jewell, and C.-F. J. Wu. A nonparametric approach to the truncated regression problem. Journal of the American Statistical Association, 83(403):785–792, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tierney:1986:AAP

[TK86] Luke Tierney and Joseph B. Kadane. Accurate approximations for posterior moments and marginal densities. Journal of the American Statistical Association, 81(393):82–86, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tierney:1989:FEL

[TKK89] Luke Tierney, Robert E. Kass, and Joseph B. Kadane. Fully exponential Laplace approximations to expectations and variances of nonpositive functions. Journal of the American Statistical Association, 84(407):710–716, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tuljapurkar:1995:CID

[TL95] Shripad Tuljapurkar and Ronald Lee. Comment: Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90(430):424–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tang:1997:ALR

[TL97] Dei-In Tang and Shang P. Lin. An approximate likelihood ratio test for comparing several treatments to a control. Journal of the American Statistical Association, 92(439):1155–1162, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_97/sept/tang.htm.

Toman:1994:BOD

[Tom94] Blaza Toman. Bayes optimal designs for two- and three-level factorial experiments. Journal of the American Statistical Association, 89(427):937–946, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Toman:1996:BED

[Tom96] Blaza Toman. Bayesian experimental design for multiple hypothesis testing. Journal of the American Statistical Association, 91 (433):185–190, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thomas:1987:SSC

[TR87] D. Roland Thomas and J. N. K. Rao. Small-sample comparisons of level and power for simple goodness-of-fit statistics under cluster sampling. Journal of the American Statistical Association, 82 (398):630–636, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsou:1995:RL

[TR95] Tsung-Shan Tsou and Richard M. Royall. Robust likelihoods. Journal of the American Statistical Association, 90(429):316–320, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tam:1997:TSM

[TR97] Wing Kuen Tam and Gregory C. Reinsel. Tests for seasonal moving average unit root in ARIMA models. Journal of the American Statistical Association, 92(438):724–??, June 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tritchler:1984:IPT

[Tri84] David Tritchler. On inverting permutation tests. Journal of the American Statistical Association, 79(385):200–207, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Troendle:1995:SRM

[Tro95] James F. Troendle. A stepwise resampling method of multiple hypothesis testing. Journal of the American Statistical Association, 90(429):370–378, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Terrell:1985:OND

[TS85] George R. Terrell and David W. Scott. Oversmoothed nonparametric density estimates. Journal of the American Statistical Association, 80(389):209–214, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thombs:1990:BPI

[TS90] Lori A. Thombs and William R. Schucany. Bootstrap prediction intervals for autoregression. Journal of the American Statistical Association, 85(410):486–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Treder:1996:BST

[TS96] R. P. Treder and J. Sedransk. Bayesian sequential two-phase sampling. Journal of the American Statistical Association, 91(434): 782–790, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsay:1984:RMT

[Tsa84] Ruey S. Tsay. Regression models with time series errors. Journal of the American Statistical Association, 79(385):118–124, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsay:1987:CHT

[Tsa87] Ruey S. Tsay. Conditional heteroscedastic time series models. Journal of the American Statistical Association, 82(398):590–604, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsay:1989:TMT

[Tsa89] Ruey S. Tsay. Testing and modeling threshold autoregressive processes. Journal of the American Statistical Association, 84(405): 231–240, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsay:1998:TMM

[Tsa98] Ruey S. Tsay. Testing and modeling multivariate threshold models. Journal of the American Statistical Association, 93(443): 1188–1202, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/sept/tsay.html.

Tsiatis:1982:RST

[Tsi82] Anastasios A. Tsiatis. Repeated significance testing for a general class of statistics used in censored survival analysis. Journal of the American Statistical Association, 77(380):855–861, ????

1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsui:1984:RCZ

[Tsu84] Kam-Wah Tsui. Robustness of Clevenson-Zidek-type estimators. Journal of the American Statistical Association, 79(385):152–157, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsui:1986:FDR

[Tsu86] Kam-Wah Tsui. Further developments on the robustness of Clevenson-Zidek-type means estimators. Journal of the American Statistical Association, 81(393):176–180, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsay:1984:CEA

[TT84] Ruey S. Tsay and George C. Tiao. Consistent estimates of autoregressive parameters and extended sample autocorrelation function for stationary and nonstationary ARMA models. Journal of the American Statistical Association, 79(385):84–96, ???? 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thomas:1986:CBP

[TT86] Dana L. Thomas and David R. Thomas. Confidence bands for percentiles in the linear regression model. Journal of the American Statistical Association, 81(395):705–708, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1947:RTI

[Tuk47a] John W. Tukey. Review of A Teoria da Indução Estadistica, by J. Kingston. Journal of the American Statistical Association, 42(??): 190–191, ???? 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1947:REF

[Tuk47b] John W. Tukey. Review of The Factorial Analysis of Human Ability, 2nd ed, by G. H. Thomson. Journal of the American Statistical Association, 42(??):199–200, ???? 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1947:RFA

[Tuk47c] John W. Tukey. Review of The Factorial Analysis of Human Ability, Second Edition, by Godfrey H. Thomson. Journal of the American Statistical Association, 42(??):199–200, ???? 1947. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1949:RPT

[Tuk49a] John W. Tukey. Review of Probability Theory for Statistical Methods, by F. N. David. Journal of the American Statistical Association, 44(??):567–568, ???? 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1949:RTE

[Tuk49b] John W. Tukey. Review of Theory of Experimental Inference, by C. W. Churchman. Journal of the American Statistical Association, 44(??):136–139, ???? 1949. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1950:CCA

[Tuk50a] John W. Tukey. Chairman’s closure. acceptance sampling. A symposium. Journal of the American Statistical Association, ??(??): 150–155, ???? 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1950:SSS

[Tuk50b] John W. Tukey. Some sampling simplified. Journal of the American Statistical Association, 45(??):501–519, ???? 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1952:REI

[Tuk52] John W. Tukey. Review of The Extrapolation, Interpolation and Smoothing of Stationary Time Series with Engineering Applications, by N Wiener. Journal of the American Statistical Association, 47(??):319–321, ???? 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1954:UPE

[Tuk54] John W. Tukey. Unsolved problems of experimental statistics. Journal of the American Statistical Association, 49(??):706–731, ???? 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1960:WDW

[Tuk60] John W. Tukey. Where do we go from here? Journal of the American Statistical Association, 55(??):80–93, ???? 1960. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1979:CEP

[Tuk79a] John W. Tukey. Comments on Emanuel Parzen, Nonparametric statistical data modeling. Journal of the American Statistical Association, 74(365):121–122, March 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1979:MSR

[Tuk79b] John W. Tukey. Methodology, and the statistician’s responsibility for both accuracy and relevance. Journal of the American Statistical Association, 74(368):786–793, December 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). An expanded version of ASnumber 121.

Tukey:1985:CEP

[Tuk85] John W. Tukey. Comments on Estimating the population in a census year: 1980 and beyond. Journal of the American Statistical Association, 80(??):127–128, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1985:VSL

[TV85] John W. Tukey and P. Velleman. The variance of slopes of lines fitted by groups: An analysis of the Johnstone and Velleman Monte Carlo results. comment on “the resistant line and related regression methods,” by I. Johnstone and P. Velleman. Journal of the American Statistical Association, 80(??):1055–1059, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1946:IS

[TW46a] John W. Tukey and C. P. Winsor. Industrial statistics. Journal of the American Statistical Association, 41(??):406–411, ???? 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1946:RIE

[TW46b] John W. Tukey and J. Wolfowitz. Review: Industrial Experimentation, by K. A. Brownlee. Journal of the American Statistical Association, 41(233):125–128, March 1946. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2280164.

Tukey:1952:RSMb

[TW52] John W. Tukey and J. Wolfowitz. Review of Statistische Methoden für Naturwissenschafter, Mediziner, und Ingenieure, by A. Linder. Journal of the American Statistical Association, 47(259): 554–556, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281326.

Tanner:1987:CPD

[TW87] Martin A. Tanner and Wing Hung Wong. The calculation of posterior distributions by data augmentation. with discussion and with a reply by the authors. Journal of the American Statistical Association, 82(398):528–550, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsui:1989:GVS

[TW89] Kam-Wah Tsui and Samaradasa Weerahandi. Generalized pvalues in significance testing of hypotheses in the presence of nuisance parameters. Journal of the American Statistical Association, 84(406):602–607, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tebaldi:1998:BIN

[TW98a] Claudia Tebaldi and Mike West. Bayesian inference on network traffic using link count data. Journal of the American Statistical Association, 93(442):557–576, June 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_

98/june/TEBALDI.HTM. With comments by Y. Vardi and Robert McCulloch, and a rejoinder by the authors.

Tebaldi:1998:RBI

[TW98b] Claudia Tebaldi and Mike West. Rejoinder: Bayesian inference on network traffic using link count data. Journal of the American Statistical Association, 93(442):576–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tishler:1981:NML

[TZ81] Asher Tishler and Israel Zang. A new maximum likelihood algorithm for piecewise regression. Journal of the American Statistical Association, 76(376):980–987, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Uebersax:1993:SME

[Ueb93] John S. Uebersax. Statistical modeling of expert ratings on medical treatment appropriateness. Journal of the American Statistical Association, 88(422):421–??, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Uusipaikka:1983:ECB

[Uus83] Esa Uusipaikka. Exact confidence bands for linear regression over intervals. Journal of the American Statistical Association, 78(383): 638–644, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Uusipaikka:1985:ESC

[Uus85] Esa Uusipaikka. Exact simultaneous confidence intervals for multiple comparisons among three or four mean values. Journal of the American Statistical Association, 80(389):196–201, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Valliant:1985:NPT

[Val85] Richard Valliant. Nonlinear prediction theory and the estimation of proportions in a finite population. Journal of the American Statistical Association, 80(391):631–641, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Valliant:1987:CPS

[Val87a] Richard Valliant. Conditional properties of some estimators in stratified sampling. Journal of the American Statistical Association, 82(398):509–519, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Valliant:1987:GVF

[Val87b] Richard Valliant. Generalized variance functions in stratified twostage sampling. Journal of the American Statistical Association, 82(398):499–508, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Valliant:1993:PCV

[Val93] Richard Valliant. Poststratification and conditional variance estimation. Journal of the American Statistical Association, 88(421): 89–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vardi:1988:SMI

[Var88] Yehuda Vardi. Statistical models for intercepted data. Journal of the American Statistical Association, 83(401):183–197, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vardi:1996:NTE

[Var96] Y. Vardi. Network tomography: Estimating source-destination traffic intensities from link data. Journal of the American Statistical Association, 91(433):365–377, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vardi:1998:CBI

[Var98] Y. Vardi. Comment: Bayesian inference on network traffic using link count data. Journal of the American Statistical Association, 93(442):573–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

VanBrummelen:1997:DIH

[VB97] Glen Van Brummelen and Kenneth Butler. Determining the interdependence of historical astronomical tables. Journal of the American Statistical Association, 92(437):41–??, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/vanbrumm.htm.

vanHouwelingen:1987:MEM

[vd87] J. C. van Houwelingen and L. de Vries. Minimax estimation of the mixing proportion of two known distributions. Journal of the American Statistical Association, 82(397):300–304, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Velilla:1995:DRE

[Vel95] Santiago Velilla. Diagnostics and robust estimation in multivariate data transformations. Journal of the American Statistical Association, 90(431):945–951, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Velilla:1998:ANL

[Vel98] Santiago Velilla. Assessing the number of linear components in a general regression problem. Journal of the American Statistical Association, 93(443):1088–1098, September 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/sept/velilla.html.

Verrill:1993:PSS

[Ver93] Steve Verrill. Predictor sort sampling, tight T’s, and the analysis of covariance. Journal of the American Statistical Association, 88 (421):119–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Victoria-Feser:1997:REG

[VFR97] Maria-Pia Victoria-Feser and Elvezio Ronchetti. Robust estimation for grouped data. Journal of the American Statistical Association, 92(437):333–340, March 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/march/victoria.htm.

Vollset:1991:FCE

[VHE91] Stein E. Vollset, Karim F. Hirji, and Robert M. Elashoff. Fast computation of exact confidence limits for the common odds ratio in a series of 2 × 2 tables. Journal of the American Statistical Association, 86(414):404–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vidakovic:1998:NWS

[Vid98] Brani Vidakovic. Nonlinear wavelet shrinkage with Bayes rules and Bayes factors. Journal of the American Statistical Association, 93 (441):173–179, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/VIDAKOVI.HTM.

Villegas:1990:BIM

[Vil90] C. Villegas. Bayesian inference in models with Euclidean structures. Journal of the American Statistical Association, 85(412): 1159–1164, December 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vinod:1969:IPT

[Vin69] H. D. Vinod. Integer programming and the theory of grouping. Journal of the American Statistical Association, 64(326):506–519, June 1969. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Verrill:1988:TLS

[VJ88] Steve Verrill and Richard A. Johnson. Tables and large-sample distribution theory for censored-data correlation statistics for testing normality. Journal of the American Statistical Association, 83 (404):1192–1197, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Verdinelli:1992:BDM

[VK92] Isabella Verdinelli and Joseph B. Kadane. Bayesian designs for maximizing information and outcome. Journal of the American Statistical Association, 87(418):510–??, June 1992. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Verbeke:1996:LME

[VL96] Geert Verbeke and Emmanuel Lesaffre. A linear mixed-effects model with heterogeneity in the random-effects population. Journal of the American Statistical Association, 91(433):217–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vardeman:1983:CSD

[VM83] Stephen Vardeman and Glen Meeden. Calibration, sufficiency, and domination considerations for Bayesian probability assessors. Journal of the American Statistical Association, 78(384):808–816, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vu:1996:LRT

[VM96] H. T. V. Vu and R. A. Maller. The likelihood ratio test for Poisson versus binomial distributions. Journal of the American Statistical Association, 91(434):818–824, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Voss:1986:FOD

[Vos86] Daniel T. Voss. First-order deletion designs and the construction of efficient nearly orthogonal factorial designs in small blocks. Journal of the American Statistical Association, 81(395):813–818, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Voss:1993:CRS

[Vos93] Daniel T. Voss. Characterization of regularity for single replicate factorial designs. Journal of the American Statistical Association, 88(423):924–925, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

vanderLaan:1998:LEE

[vR98] Mark J. van der Laan and James M. Robins. Locally efficient estimation with current status data and time-dependent covariates. Journal of the American Statistical Association, 93(442):693–701, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

vanHouwelingen:1993:MEB

[VS93] Hans C. Van Houwelingen and Theo Stijnen. Monotone empirical Bayes estimators based on more informative samples. Journal of the American Statistical Association, 88(424):1438–1443, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vardi:1985:SMP

[VSK85] Y. Vardi, L. A. Shepp, and L. Kaufman. A statistical model for positron emission tomography. with discussion. Journal of the American Statistical Association, 80(389):8–37, ???? 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Villalobos:1987:ICM

[VW87] Miguel Villalobos and Grace Wahba. Inequality-constrained multivariate smoothing splines with application to the estimation of posterior probabilities. Journal of the American Statistical Association, 82(397):239–248, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Verdinelli:1995:CBF

[VW95] Isabella Verdinelli and Larry Wasserman. Computing Bayes factors using a generalization of the Savage-Dickey density ratio. Journal of the American Statistical Association, 90(430):614–618, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Velleman:1980:CRC

[VY80] P. F. Velleman and M. A. Ypelaar. Constructing regressions with controlled features: A method of probing robust regression performance. Journal of the American Statistical Association, ??(??):??, ???? 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Waugh:1967:FPM

[WA67] F. V. Waugh and M. E. Abel. On fractional powers of a matrix. Journal of the American Statistical Association, 62(??):1018–1021, ???? 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wecker:1983:SEA

[WA83] William E. Wecker and Craig F. Ansley. The signal extraction approach to nonlinear regression and spline smoothing. Journal of the American Statistical Association, 78(381):81–89, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wakefield:1996:BAP

[Wak96] Jon Wakefield. The Bayesian analysis of population pharmacokinetic models. Journal of the American Statistical Association, 91(433):62–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Walton:1986:NOC

[Wal86] Gerald S. Walton. The number of observed classes from a multiple hypergeometric distribution. Journal of the American Statistical Association, 81(393):169–171, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wallman:1993:ESL

[Wal93] Katherine K. Wallman. Enhancing statistical literacy: Enriching our society. Journal of the American Statistical Association, 88(421):1–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wampler:1970:RAS

[Wam70] R. H. Wampler. A report on the accuracy of some widely used least squares computer programs. Journal of the American Statistical Association, 65(??):549–565, ???? 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1989:SMR

[Wan89] Mei-Cheng Wang. A semiparametric model for randomly truncated data. Journal of the American Statistical Association, 84 (407):742–748, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1991:NEC

[Wan91] Mei-Cheng Wang. Nonparametric estimation from cross-sectional survival data. Journal of the American Statistical Association, 86 (413):130–143, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1992:ARA

[Wan92] Mei-Cheng Wang. The analysis of retrospectively ascertained data in the presence of reporting delays. Journal of the American Statistical Association, 87(418):397–??, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1995:SKL

[Wan95] Julian Z. Wang. Selection of the k largest order statistics for the domain of attraction of the Gumbel distribution. Journal of the American Statistical Association, 90(431):1055–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1996:LRT

[Wan96] Yazhen Wang. A likelihood ratio test against stochastic ordering in several populations. Journal of the American Statistical Association, 91(436):1676–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1997:OUT

[Wan97] Weizhen Wang. Optimal unbiased tests for equivalence in intrasubject variability. Journal of the American Statistical Association, 92 (439):1163–1170, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/wang_w.htm.

Wang:1998:CCE

[Wan98a] Yazhen Wang. Change curve estimation via wavelets. Journal of the American Statistical Association, 93(441):163–172, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/WANG.HTM.

Wang:1998:SSM

[Wan98b] Yuedong Wang. Smoothing spline models with correlated random errors. Journal of the American Statistical Association, 93 (441):341–348, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/WANG_Y.HTM.

Watson:1987:UMB

[Wat87] Mark W. Watson. Uncertainty in model-based seasonal adjustment procedures and construction of minimax filters. Journal of the American Statistical Association, 82(398):395–408, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wakefield:1996:BMC

[WB96] Jon Wakefield and James Bennett. The Bayesian modeling of covariates for population pharmacokinetic models. Journal of the American Statistical Association, 91(435):917–??, ???? 1996. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wood:1993:SAC

[WBB93] Andrew T. A. Wood, James G. Booth, and Ronald W. Butler. Saddlepoint approximations to the CDF of some statistics with nonnormal limit distributions. Journal of the American Statistical Association, 88(422):680–686, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wolter:1991:EPI

[WC91] Kirk M. Wolter and Beverley D. Causey. Evaluation of procedures for improving population estimates for small areas. Journal of the American Statistical Association, 86(414):278–??, June 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Windham:1992:IRV

[WC92] Michael P. Windham and Adele Cutler. Information ratios for validating mixture analyses. Journal of the American Statistical Association, 87(420):1188–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1999:NER

[WC99] Mei-Cheng Wang and Shu-Hui Chang. Nonparametric estimation of a recurrent survival function. Journal of the American Statistical Association, 94(445):146–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:1998:ACR

[WCH98] Colin O. Wu, Chin-Tsang Chiang, and Donald R. Hoover. Asymptotic confidence regions for kernel smoothing of a varyingcoefficient model with longitudinal data. Journal of the American Statistical Association, 93(444):1388–1402, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/dec/wu.html.

Welsh:1994:FHR

[WCR94] A. H. Welsh, R. J. Carroll, and D. Ruppert. Fitting heteroscedastic regression models. Journal of the American Statistical Association, 89(425):100–116, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Waller:1997:HST

[WCXG97] Lance A. Waller, Bradley P. Carlin, Hong Xia, and Alan E. Gelfand. Hierarchical spatio-temporal mapping of disease rates. Journal of the American Statistical Association, 92(438):607–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Weerahandi:1991:TVC

[Wee91] Samaradasa Weerahandi. Testing variance components in mixed models with generalized p values. Journal of the American Statistical Association, 86(413):151–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Weerahandi:1993:GCI

[Wee93] Samaradasa Weerahandi. Generalized confidence intervals. Journal of the American Statistical Association, 88(423):899–905, ???? 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [Wee94].

Weerahaudi:1994:CGC

[Wee94] Samaradasa Weerahaudi. Correction: “Generalized confidence intervals” [J. Amer. Statist. Assoc. 88 (1993), no. 423, 899–905, MR 94e:62031]. Journal of the American Statistical Association, 89 (426):726, ???? 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Wee93].

Wegman:1990:HDA

[Weg90] Edward J. Wegman. Hyperdimensional data analysis using parallel coordinates. Journal of the American Statistical Association, 85 (411):664–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:1981:ACE

[Wei81] L. J. Wei. Asymptotic conservativeness and efficiency of Kruskal-Wallis test for K dependent samples. Journal of the American Statistical Association, 76(376):1006–1009, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:1983:TII

[Wei83] L. J. Wei. Tests of interchangeability with incomplete paired observations. Journal of the American Statistical Association, 78 (383):725–729, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:1984:TGF

[Wei84] L. J. Wei. Testing goodness of fit for proportional hazards model with censored observations. Journal of the American Statistical Association, 79(387):649–652, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Weiss:1992:EHU

[Wei92] Andrew A. Weiss. The effects of helmet use on the severity of head injuries in motorcycle accidents. Journal of the American Statistical Association, 87(417):48–??, March 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Weiss:1995:IVS

[Wei95] Robert E. Weiss. The influence of variable selection: A Bayesian diagnostic perspective. Journal of the American Statistical Association, 90(430):619–625, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Weir:1997:FBR

[Wei97] Iain S. Weir. Fully Bayesian reconstructions from single-photon emission computed tomography data. Journal of the American Statistical Association, 92(437):49–??, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/weir.htm.

Welch:1990:CPT

[Wel90a] William J. Welch. Construction of permutation tests. Journal of the American Statistical Association, 85(411):693–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wells:1990:REG

[Wel90b] Martin T. Wells. The relative efficiency of goodness-of-fit statistics in the simple and composite hypothesis-testing problem. Journal of the American Statistical Association, 85(410):459–463, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Westfall:1987:CVC

[Wes87a] Peter H. Westfall. A comparison of variance component estimates for arbitrary underlying distributions. Journal of the American Statistical Association, 82(399):866–874, ???? 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Westfall:1987:CMT

[Wes87b] Peter H. Westfall. Computable MINQUE-type estimates of variance components. Journal of the American Statistical Association, 82(398):586–589, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

West:1995:BIC

[Wes95] Mike West. Bayesian inference in cyclical component dynamic linear models. Journal of the American Statistical Association, 90 (432):1301–1312, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

West:1997:HMM

[Wes97a] Mike West. Hierarchical mixture models in neurological transmission analysis. Journal of the American Statistical Association, 92(438):587–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Westfall:1997:MTG

[Wes97b] Peter H. Westfall. Multiple testing of general contrasts’ using logical constraints and correlations. Journal of the American Statistical Association, 92(437):299–306, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/westfall.htm.

Westfall:1998:AMT

[Wes98] Peter Westfall. Addendum: “Multiple testing of general contrasts using logical constraints and correlations” [J. Amer. Statist. Assoc. 92 (1997), no. 437, 299–306; MR 98c:62053]. Journal of the American Statistical Association, 93(441):413, 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Worsley:1991:LSC

[WEST91] K. J. Worsley, A. C. Evans, S. C. Strother, and J. L. Tyler. A linear spatial correlation model, with applications to positron emission tomography. Journal of the American Statistical Association, 86 (413):55–??, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wolfenson:1982:BLD

[WF82] Marco Wolfenson and Terrence L. Fine. Bayes-like decision making with upper and lower probabilities. Journal of the American Statistical Association, 77(377):80–88, ???? 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:1983:NES

[WG83] L. J. Wei and M. H. Gail. Nonparametric estimation for a scalechange with censored observations. Journal of the American Statistical Association, 78(382):382–388, ???? 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Welch:1988:RPT

[WG88] William J. Welch and Leopoldo G. Gutierrez. Robust permutation tests for matched-pairs designs. Journal of the American Statistical Association, 83(402):450–455, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1990:TSI

[WH90] Jane-Ling Wang and Thomas P. Hettmansperger. Two-sample inference for median survival times based on one-sample procedures for censored survival data. Journal of the American Statistical Association, 85(410):529–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Whittemore:1998:TCS

[WHG98] Alice S. Whittemore, Jerry Halpern, and Gail Gong. Testing covariance structure in multivariate models: Application to family disease data. Journal of the American Statistical Association, 93 (442):518–525, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/WHITTEMO.HTM.

Wolfe:1996:WPS

[WHGA96] Barbara Wolfe, Robert Haveman, Donna Ginther, and Chong Bum An. The “window problem” in studies of children’s attainments: A methodological exploration. Journal of the American Statistical Association, 91(435):970–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:1988:ETS

[WHH88] C.-F. J. Wu, D. Holt, and D. J. Holmes. The effect of two-stage sampling on the F statistic. Journal of the American Statistical Association, 83(401):150–159, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

White:1981:CDM

[Whi81a] Halbert White. Consequences and detection of misspecified nonlinear regression models. Journal of the American Statistical Association, 76(374):419–433, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Whittemore:1981:CSS

[Whi81b] Alice S. Whittemore. Correction: “Sample size for logistic regression”. Journal of the American Statistical Association, 76(376): 1022–1023, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Whi81c].

Whittemore:1981:SSL

[Whi81c] Alice S. Whittemore. Sample size for logistic regression with small response probability. Journal of the American Statistical Association, 76(373):27–32, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [Whi81b].

White:1989:SAR

[Whi89] Halbert White. Some asymptotic results for learning in single hidden layer feedforward network models. Journal of the American Statistical Association, 84(408):1003–1013, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comment [Fab94]. See correction [Whi92].

White:1992:SAR

[Whi92] Halbert White. “Some asymptotic results for learning in single hidden-layer feedforward network models” [J. Amer. Statist. Assoc. 84 (1989), no. 408, 1003–1013, MR 92e:62119]. Journal of the American Statistical Association, 87(420):1252, ???? 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Whi89, Fab94].

West:1985:DGL

[WHM85] Mike West, P. Jeff Harrison, and Helio S. Migon. Dynamic generalized linear models and Bayesian forecasting. with discussion. Journal of the American Statistical Association, 80(389):73–97, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wiens:1998:MRD

[Wie98] Douglas P. Wiens. Minimax robust designs and weights for approximately specified regression models with heteroscedastic errors. Journal of the American Statistical Association, 93(444): 1440–1450, December 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/dec/wiens.html.

Wijsman:1985:UIR

[Wij85] Robert A. Wijsman. A useful inequality on ratios of integrals, with application to maximum likelihood estimation. Journal of the American Statistical Association, 80(390):472–475, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Williams:1950:CNW

[Wil50] C. Arthur Williams, Jr. On the choice of the number and width of classes for the chi-square test for goodness of fit. Journal of the American Statistical Association, 45(249):77–86, March 1950. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280429.

Wilcox:1984:SBP

[Wil84] Rand R. Wilcox. Selecting the best population, provided it is better than a standard: the unequal variance case. Journal of the American Statistical Association, 79(388):887–891, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wittkowski:1988:FTS

[Wit88] Knut M. Wittkowski. Friedman-type statistics and consistent multiple comparisons for unbalanced designs with missing data. Journal of the American Statistical Association, 83(404):1163–1170, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Weerakkody:1992:EWM

[WJ92] Govinda J. Weerakkody and Dallas E. Johnson. Estimation of within model parameters in regression models with a nested error structure. Journal of the American Statistical Association, 87 (419):708–713, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wand:1993:CSP

[WJ93] M. P. Wand and M. C. Jones. Comparison of smoothing parameterizations in bivariate kernel density estimation. Journal of the American Statistical Association, 88(422):520–528, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Whittemore:1988:ARC

[WK88] Alice S. Whittemore and Joseph B. Keller. Approximations for regression with covariate measurement error. Journal of the American Statistical Association, 83(404):1057–1066, ???? 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wasserman:1992:CBE

[WK92] Larry Wasserman and Joseph B. Kadane. Computing bounds on expectations. Journal of the American Statistical Association, 87 (418):516–522, June 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1994:SVC

[WK94] Morgan C. Wang and William J. Kennedy. Self-validating computations of probabilities for selected central and noncentral univariate probability functions. Journal of the American Statistical Association, 89(427):878–887, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wood:1998:BAR

[WK98] Sally Wood and Robert Kohn. A Bayesian approach to robust binary nonparametric regression. Journal of the American Statistical Association, 93(441):203–??, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Williamson:1995:ABO

[WKL95] John M. Williamson, KyungMann Kim, and Stuart R. Lipsitz. Analyzing bivariate ordinal data using a global odds ratio. Journal of the American Statistical Association, 90(432):1432–1437, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Waclawiw:1993:PRE

[WKY93] Myron A. Waclawiw and Liang Kung-Yee. Prediction of random effects in the generalized linear model. Journal of the American Statistical Association, 88(421):171–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:1984:TSA

[WL84] L. J. Wei and J. M. Lachin. Two-sample asymptotically distribution-free tests for incomplete multivariate observations. Journal of the American Statistical Association, 79(387):653–661, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1998:BAS

[WLGC98] Naisyin Wang, Xihong Lin, Roberto G. Gutierrez, and Raymond J. Carroll. Bias analysis and SIMEX approach in generalized linear mixed measurement error models. Journal of the American Statistical Association, 93(441):249–261, March 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/WANG_N.HTM.

Wei:1989:RAM

[WLW89] L. J. Wei, D. Y. Lin, and L. Weissfeld. Regression analysis of multivariate incomplete failure time data by modeling marginal distributions. Journal of the American Statistical Association, 84 (408):1065–1073, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

White:1980:SLS

[WM80] Halbert White and Glenn M. MacDonald. Some large-sample tests for nonnormality in the linear regression model. Journal of the American Statistical Association, 75(369):16–31, ???? 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [WM81].

White:1981:CSL

[WM81] Halbert White and Glenn M. MacDonald. Correction: “Some large-sample tests for nonnormality in the linear regression model” [J. Amer. Statist. Assoc. 75 (1980), no. 369, 16–31, MR 81e:62077]. Journal of the American Statistical Association, 76(376):1022, ???? 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WM80].

Welsh:1990:RES

[WM90] A. H. Welsh and H. L. Morrison. Robust L estimation of scale with an application in astronomy. Journal of the American Statistical Association, 85(411):729–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wong:1991:CSE

[WM91] George Y. Wong and William M. Mason. Contextually specific effects and other generalizations of the hierarchical linear model for comparative analysis. Journal of the American Statistical Association, 86(414):487–??, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1998:CLR

[WM98] Yining Wang and Michael P. McDermott. Conditional likelihood ratio test for a nonnegative normal mean vector. Journal of the American Statistical Association, 93(441):380–386, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/WANG_YIN.HTM.

Wand:1991:TDE

[WMR91] M. P. Wand, J. S. Marron, and D. Ruppert. Transformations in density estimation. with discussion and a rejoinder by the authors. Journal of the American Statistical Association, 86(414):343–361, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wolter:1984:ISE

[Wol84] Kirk M. Wolter. An investigation of some estimators of variance for systematic sampling. Journal of the American Statistical Association, 79(388):781–790, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wolter:1986:SCE

[Wol86] Kirk M. Wolter. Some coverage error models for census data. Journal of the American Statistical Association, 81(394):338–346, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wolak:1987:ETM

[Wol87] Frank A. Wolak. An exact test for multiple inequality and equality constraints in the linear regression model. Journal of the American Statistical Association, 82(399):782–793, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wolpert:1995:CID

[Wol95] Robert L. Wolpert. Comment: Inference from a deterministic population dynamics model for bowhead whales. Journal of the American Statistical Association, 90(430):426–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wong:1982:RRA

[Won82a] George Y. Wong. Round robin analysis of variance via maximum likelihood. Journal of the American Statistical Association, 77 (380):714–724, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wong:1982:HCM

[Won82b] M. Anthony Wong. A hybrid clustering method for identifying high-density clusters. Journal of the American Statistical Association, 77(380):841–847, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wong:1983:NML

[Won83] Wing Hung Wong. A note on the modified likelihood for density estimation. Journal of the American Statistical Association, 78 (382):461–463, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wong:1984:ROR

[Won84] George Y. Wong. Rank-one round robin designs. Journal of the American Statistical Association, 79(385):136–144, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wong:1987:BMD

[Won87] George Y. Wong. Bayesian models for directed graphs. Journal of the American Statistical Association, 82(397):140–148, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Working:1934:RDS

[Wor34] Holbrook Working. A random-difference series for use in the analysis of time series. Journal of the American Statistical Association, 29(185):11–24, March 1934. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2278456.

Worm:1974:SAF

[Wor74] G. H. Worm. Sensitivity analysis of the fixed point vector of transition matrices. Journal of the American Statistical Association, 69(348):961–962, December 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:1985:DFM

[WP85] L. J. Wei and D. Pee. Distribution-free methods of estimating location difference with censored paired data. Journal of the American Statistical Association, 80(390):405–410, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Waternaux:1998:CPU

[WPD98] Christine Waternaux, Eva Petkova, and William DuMouchel. Comment: Problems with using auxiliary information to correct for omitted variables when estimating the effect of lead on IQ. Journal of the American Statistical Association, 93(442):505–??, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

West:1999:ECE

[WPK99] Mike West, Raquel Prado, and Andrew D. Krystal. Evaluation and comparison of EEG traces: Latent structure in nonstationary time series. Journal of the American Statistical Association, 94(446):375–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/west.pdf.

Woodward:1984:CMD

[WPSL84] Wayne A. Woodward, William C. Parr, William R. Schucany, and Hildegard Lindsey. A comparison of minimum distance and maximum likelihood estimation of a mixture proportion. Journal of the American Statistical Association, 79(387):590–598, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Warrack:1984:LRT

[WR84] Giles Warrack and Tim Robertson. A likelihood ratio test regarding two nested but oblique order-restricted hypotheses. Journal of the American Statistical Association, 79(388):881–886, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Woodruff:1994:CRE

[WR94] David L. Woodruff and David M. Rocke. Computable robust estimation of multivariate location and shape in high dimension using compound estimators. Journal of the American Statistical Association, 89(427):888–896, September 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1995:NET

[WR95] Naisyin Wang and David Ruppert. Nonparametric estimation of the transformation in the transform-both-sides regression model. Journal of the American Statistical Association, 90(430):522–534, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wright:1983:FPS

[Wri83] Roger L. Wright. Finite population sampling with multivariate auxiliary information. Journal of the American Statistical Association, 78(384):879–884, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Whitaker:1989:ERS

[WS89] Lyn R. Whitaker and Francisco J. Samaniego. Estimating the reliability of systems subject to imperfect repair. Journal of the American Statistical Association, 84(405):301–309, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1994:MRN

[WS94] Ferdinand T. Wang and David W. Scott. The L_{1} method for robust nonparametric regression. Journal of the American Statistical Association, 89(425):65–76, March 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wendell:1996:EIP

[WS96] John P. Wendell and Josef Schmee. Exact inference for proportions from a stratified finite population. Journal of the American Statistical Association, 91(434):825–830, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1992:SDA

[WSJ92] R. Wang, J. Sedransk, and J. H. Jinn. Secondary data analysis when there are missing observations. Journal of the American Statistical Association, 87(420):952–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:1990:SID

[WSLP90] L. J. Wei, R. T. Smythe, D. Y. Lin, and T. S. Park. Statistical inference with data-dependent treatment allocation rules. Journal of the American Statistical Association, 85(409):156–162, March 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:1990:PCC

[WT90a] Greg C. F. Wei and Martin A. Tanner. Posterior computations for censored regression data. Journal of the American Statistical Association, 85(411):829–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:1990:MCI

[WT90b] Greg C. G. Wei and Martin A. Tanner. A Monte Carlo implementation of the EM algorithm and the poor man’s data augmentation algorithms. Journal of the American Statistical Association, 85(411):699–??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:1985:ESD

[Wu85] C.-F. Jeff Wu. Efficient sequential designs with binary data. Journal of the American Statistical Association, 80(392):974–984, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wegman:1983:SS

[WW83] Edward J. Wegman and Ian W. Wright. Splines in statistics. Journal of the American Statistical Association, 78(382):351–365, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1987:SBD

[WW87a] Yuchung J. Wang and George Y. Wong. Stochastic blockmodels for directed graphs. Journal of the American Statistical Association, 82(397):8–19, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wittes:1987:PMH

[WW87b] Janet Wittes and Sylvan Wallenstein. The power of the Mantel– Haenszel test. Journal of the American Statistical Association, 82 (400):1104–1109, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1991:ACA

[WW91a] J. C. Wang and C. F. J. Wu. An approach to the construction of asymmetrical orthogonal arrays. Journal of the American Statistical Association, 86(414):450–456, June 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

West:1991:DBE

[WW91b] Mike West and Robert L. Winkler. Data base error trapping and prediction. Journal of the American Statistical Association, 86 (416):987–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Walden:1993:DBT

[WW93] Andrew T. Walden and Mark L. Williams. Deconvolution, bandwidth, and the trispectrum. Journal of the American Statistical Association, 88(424):1323–??, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1998:CSS

[WW98] Yuedong Wang and Grace Wahba. Comment: Smoothing spline models for the analysis of nested and crossed samples of curves. Journal of the American Statistical Association, 93(443):976–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:1997:WSE

[WWZtO97] C. Y. Wang, Suojin Wang, Lue Ping Zhao, and Shyh tyan Ou. Weighted semiparametric estimation in regression analysis with missing covariate data. Journal of the American Statistical Association, 92(438):512–??, June 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wells:1996:DEB

[WY96] Martin T. Wells and Kwee Poo Yeo. Density estimation with bivariate censored data. Journal of the American Statistical Association, 91(436):1566–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wellner:1997:HAC

[WZ97a] Jon A. Wellner and Yihui Zhan. A hybrid algorithm for computation of the nonparametric maximum likelihood estimator from censored data. Journal of the American Statistical Association, 92 (439):945–959, September 1997. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.

org/publications/jasa/abstracts_97/sept/wellner.htm.

Wiens:1997:RDB

[WZ97b] Douglas P. Wiens and Julie Zhou. Robust designs based on the infinitesimal approach. Journal of the American Statistical Association, 92(440):1503–1511, December 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/december/WIENS.HTM.

Xiong:1995:CSC

[Xio95] Xiaoping Xiong. A class of sequential conditional probability ratio tests. Journal of the American Statistical Association, 90(432): 1463–1473, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xiao-Li:1991:UEO

[XLR91] Meng Xiao-Li and Donald B. Rubin. Using EM to obtain asymptotic variance-covariance matrices: The SEM algorithm. Journal of the American Statistical Association, 86(416):899–??, December 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xin:1993:AEE

[XXP93] Ming Tu Xin, Meng Xiao- Li, and Marcello Pagano. The AIDS epidemic: Estimating survival after AIDS diagnosis from surveillance data. Journal of the American Statistical Association, 88 (421):26–??, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yakowitz:1985:NDE

[Yak85] Sidney J. Yakowitz. Nonparametric density estimation, prediction, and regression for Markov sequences. Journal of the American Statistical Association, 80(389):215–221, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yamaguchi:1990:HSD

[Yam90] Kazuo Yamaguchi. Homophily and social distance in the choice of multiple friends. Journal of the American Statistical Association, 85(410):356–??, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yamaguchi:1992:AFT

[Yam92] Kazuo Yamaguchi. Accelerated failure-time regression models with a regression model of surviving fraction: An application to the analysis of “permanent employment” in Japan. Journal of the American Statistical Association, 87(418):284–??, June 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:1984:SAM

[Yan84] Shie-Shien Yang. Sample analogs to multidimensional coverages with applications to prediction. Journal of the American Statistical Association, 79(388):804–806, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:1985:SNE

[Yan85] Shie-Shien Yang. A smooth nonparametric estimator of a quantile function. Journal of the American Statistical Association, 80(392): 1004–1011, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:1997:EWL

[Yan97] Song Yang. Extended weighted log-rank estimating functions in censored regression. Journal of the American Statistical Association, 92(439):977–984, September 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/sept/yang_s.htm.

Yang:1999:CMR

[Yan99] Song Yang. Censored median regression using weighted empirical survival and hazard functions. Journal of the American Statistical Association, 94(445):137–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yashchin:1995:LRM

[Yas95] Emmanuel Yashchin. Likelihood ratio methods for monitoring parameters of a nested random effect model. Journal of the American Statistical Association, 90(430):729–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yuan:1997:MCS

[YB97] Ke Hai Yuan and Peter M. Bentler. Mean and covariance structure analysis: Theoretical and practical improvements. Journal of the American Statistical Association, 92(438):766–??, June 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yeh:1985:NTF

[YBN85] Ching Ming Yeh, Ralph A. Bradley, and William I. Notz. Nearly trend-free block designs. Journal of the American Statistical Association, 80(392):985–992, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yeo:1984:STA

[YD84] W. B. Yeo and H. A. David. Selection through an associated characteristic, with applications to the random effects model. Journal of the American Statistical Association, 79(386):399–405, ???? 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ye:1993:RPW

[Ye93] Keying Ye. Reference priors when the stopping rule depends on the parameter of interest. Journal of the American Statistical Association, 88(421):360–363, March 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ye:1998:MCE

[Ye98a] Jianming Ye. On measuring and correcting the effects of data mining and model selection. Journal of the American Statistical Association, 93(441):120–131, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.amstat.org/publications/jasa/abstracts_98/YE.HTM.

Ye:1998:OCL

[Ye98b] Kenny Q. Ye. Orthogonal column Latin hypercubes and their application in computer experiments. Journal of the American Statistical Association, 93(444):1430–1439, December 1998. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
98/dec/ye.html.

Yanagawa:1990:HTG

[YF90] Takashi Yanagawa and Yoshinori Fujii. Homogeneity test with a generalized Mantel-Haenzel estimator for L2 × K contingency tables. Journal of the American Statistical Association, 85(411):744– ??, September 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yanagawa:1995:PMM

[YF95] Takashi Yanagawa and Yoshinori Fujii. Projection-method Mantel–Haenszel estimator for K2×J tables. Journal of the American Statistical Association, 90(430):649–656, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:1998:LLQ

[YJ98] Keming Yu and M. C. Jones. Local linear quantile regression. Journal of the American Statistical Association, 93(441): 228–237, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/YU.HTM.

Ying:1995:SAM

[YJW95] Z. Ying, S. H. Jung, and L. J. Wei. Survival analysis with median regression models. Journal of the American Statistical Association, 90(429):178–184, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yamaguchi:1996:PES

[YK96] Kazuo Yamaguchi and Denise B. Kandel. Parametric event sequence analysis: An application to an analysis of gender and racial/ethnic differences in patterns of drug-use progression. Journal of the American Statistical Association, 91(436):1388–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yanagawa:1997:NOA

[YKB97] Takashi Yanagawa, Yasuki Kikuchi, and Kenneth G. Brown. Noobserved-adverse-effect levels in severity data. Journal of the American Statistical Association, 92(438):449–??, June 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yasui:1997:RMS

[YL97] Yutaka Yasui and Subhash Lele. A regression method for spatial disease rates: An estimating function approach. Journal of the American Statistical Association, 92(437):21–??, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/yasui.htm.

Yang:1985:DFC

[YM85] Mark C. K. Yang and Tse Chin Mo. Distribution-free confidence bounds for P (X_{1} + X_{2} + ⋯ + X_{k} < z). Journal of the American Statistical Association, 80(389):227–230, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See correction [YM86].

Yang:1986:CDF

[YM86] Mark C. K. Yang and T. C. Mo. Correction: “Distribution-free confidence bounds for P (X_{1}+X_{2}+⋯+X_{k} < z)” [J. Amer. Statist. Assoc. 80 (1985), no. 389, 227–230, MR 87b:62061]. Journal of the American Statistical Association, 81(396):1132, ???? 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [YM85].

Yang:1999:ITK

[YM99] Lijan Yang and J. S. Marron. Iterated transformation-kernel density estimation. Journal of the American Statistical Association, 94(446):580–??, June 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:1992:APM

[YP92] Xie Yu and Ellen Efron Pimentel. Age patterns of marital fertility: Revising the coale- trussell method. Journal of the American Statistical Association, 87(420):977–??, December 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:1999:SIP

[YP99] Song Yang and Ross L. Prentice. Semiparametric inference in the proportional odds regression model. Journal of the American Statistical Association, 94(445):124–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yap:1995:ETU

[YR95] Sook Fwe Yap and Gregory C. Reinsel. Estimation and testing for unit roots in a partially nonstationary vector autoregressive moving average model. Journal of the American Statistical Association, 90(429):253–267, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yohai:1988:HBP

[YZ88] Víctor J. Yohai and Ruben H. Zamar. High breakdown-point estimates of regression by means of the minimization of an efficient scale. Journal of the American Statistical Association, 83(402): 406–413, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zacks:1986:ESP

[Zac86] S. Zacks. Estimating the scale parameter of an exponential distribution from a sample of time-censored rth-order statistics. Journal of the American Statistical Association, 81(393):205–209, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeger:1986:RAC

[ZB86] Scott L. Zeger and Ron Brookmeyer. Regression analysis with censored autocorrelated data. Journal of the American Statistical Association, 81(395):722–729, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zelterman:1988:HTA

[ZC88] Daniel Zelterman and Chan Fu Chen. Homogeneity tests against central-mixture alternatives. Journal of the American Statistical Association, 83(401):179–182, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeng:1997:BAC

[ZD97] Qi Zeng and Marie Davidian. Bootstrap-adjusted calibration confidence intervals for immunoassay. Journal of the American Statistical Association, 92(437):278–290, March 1997. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.amstat.org/publications/jasa/abstracts_
97/march/zeng.htm.

Zackin:1996:NME

[ZDL96] Robert Zackin, Victor De Gruttola, and Nan Laird. Nonparametric mixed-effects models for repeated binary data arising in serial dilution assays: An application to estimating viral burden in AIDS. Journal of the American Statistical Association, 91(433): 52–??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zehnwirth:1988:GKF

[Zeh88] Ben Zehnwirth. A generalization of the Kalman filter for models with state-dependent observation variance. Journal of the American Statistical Association, 83(401):164–167, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zellner:1986:BEP

[Zel86] Arnold Zellner. Bayesian estimation and prediction using asymmetric loss functions. Journal of the American Statistical Association, 81(394):446–451, ???? 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zelterman:1987:GFT

[Zel87] Daniel Zelterman. Goodness-of-fit tests for large sparse multinomial distributions. Journal of the American Statistical Association, 82(398):624–629, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zellner:1992:SSP

[Zel92] Arnold Zellner. Statistics, science and public policy. Journal of the American Statistical Association, 87(417):1–??, March 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zelterman:1993:SBT

[Zel93] Daniel Zelterman. A semiparametric bootstrap technique for simulating extreme order statistics. Journal of the American Statistical Association, 88(422):477–485, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeger:1987:FDM

[ZH87] Scott L. Zeger and Clifford M. Hurvich. A frequency-domain median time series. Journal of the American Statistical Association, 82(399):832–835, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:1992:DPM

[Zha92] Ping Zhang. On the distributional properties of model selection criteria. Journal of the American Statistical Association, 87(419): 732–737, September 1992. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:1996:NIS

[Zha96] Ping Zhang. Nonparametric importance sampling. Journal of the American Statistical Association, 91(435):1245–1253, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:1998:CTM

[Zha98a] Heping Zhang. Classification trees for multiple binary responses. Journal of the American Statistical Association, 93 (441):180–193, March 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/ZHANG.HTM.

Zhang:1998:CBC

[Zha98b] Heping Zhang. Comment: Bayesian CART model search. Journal of the American Statistical Association, 93(443):945–??, September 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:1999:IEN

[Zha99] Linda H. Zhao. Improved estimators in nonparametric regression problems. Journal of the American Statistical Association, 94(445):164–??, March 1999. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zieschang:1990:SWM

[Zie90] Kimberly D. Zieschang. Sample weighting methods and estimation of totals in the consumer expenditure survey. Journal of the American Statistical Association, 85(412):986–??, December 1990. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeger:1991:GLM

[ZK91] Scott L. Zeger and M. Rezaul Karim. Generalized linear models with random effects and a Gibbs sampling approach. Journal of the American Statistical Association, 86(413):79–86, March 1991. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zheng:1995:CVS

[ZL95] Xiaodong Zheng and Wei-Yin Loh. Consistent variable selection in linear models. Journal of the American Statistical Association, 90(429):151–156, March 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:1998:SSM

[ZLRS98] Daowen Zhang, Xihong Lin, Jonathan Raz, and MaryFran Sowers. Semiparametric stochastic mixed models for longitudinal data. Journal of the American Statistical Association, 93 (442):710–719, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.amstat.org/

publications/jasa/abstracts_98/june/ZHANG.HTM.

Zhou:1993:CIT

[ZM93] Leping Zhou and Thomas Mathew. Combining independent tests in linear models. Journal of the American Statistical Association, 88(422):650–655, June 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zellner:1995:GSC

[ZM95] Arnold Zellner and Chung-Ki Min. Gibbs sampler convergence criteria. Journal of the American Statistical Association, 90(431): 921–??, ???? 1995. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeger:1997:EMS

[ZT97] Linda M. Zeger and Neal Thomas. Efficient matrix sampling instruments for correlated latent traits: Examples from the National Assessment of Educational Progress. Journal of the American Statistical Association, 92(438):416–??, June 1997. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zucker:1998:RML

[Zuc98] David M. Zucker. Restricted mean life with covariates: Modification and extension of a useful survival analysis method. Journal of the American Statistical Association, 93(442):702–709, June 1998. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.amstat.org/publications/jasa/
abstracts_98/june/ZUCKER.HTM.

Zhou:1996:REM

[ZWWB96] Haibo Zhou, Clarice R. Weinberg, Allen J. Wilcox, and Donna D. Baird. A random-effects model for cycle viability in fertility studies. Journal of the American Statistical Association, 91(436):1413– ??, ???? 1996. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).